### v1

# Adjoint motives of modular forms and the Tamagawa number conjecture

Abstract.
Let $f$ be a newform of weight $k\geq 2$, level NN with coefficients in a number field $K$, and AA the adjoint motive of the motive $M$ associated to $f$. We carefully discuss the construction of the realisations of $M$ and AA, as well as natural integral structures in these realisations. We then use the method of Taylor and Wiles to verify the $\lambda$-part of the Tamagawa number conjecture of Bloch and Kato for $L(A,0)$ and $L(A,1)$. Here $\lambda$ is any prime of $K$ not dividing $Nk!$, and so that the mod $\lambda$ representation associated to $f$ is absolutely irreducible when restricted to the Galois group over ${\mathbb{Q}}(\sqrt{(-1)^{(\ell-1)/2}\ell})$ where $\lambda\mid\ell$. The method also establishes modularity of all lifts of the mod $\lambda$ representation which are crystalline of Hodge-Tate type (0,k1)(0,k-1).
Key words and phrases:
Modular forms, adjoint motives, Bloch-Kato conjecture
1991 Mathematics Subject Classification:
Primary 11F67, 11F80, 11G40; Secondary 14G10, 14F, 19F27.
0. Introduction
This paper concerns the Tamagawa number conjecture of Bloch and Kato [B-K] for adjoint motives of modular forms of weight $k\geq 2$. The conjecture relates the value at 0 of the associated $L$-function to arithmetic invariants of the motive. We prove that it holds up to powers of certain “bad primes.” The strategy for achieving this is essentially due to Wiles [Wi], as completed with Taylor in [T-W]. The Taylor-Wiles construction yields a formula relating the size of a certain module measuring congruences between modular forms to that of a certain Galois cohomology group. This was carried out in [Wi] and [T-W] in the context of modular forms of weight 22, where it was used to prove results in the direction of the Fontaine-Mazur conjecture [F-M]. While it was no surprise that the method could be generalized to higher weight modular forms and that the resulting formula would be related to the Bloch-Kato conjecture, there remained many technical details to verify in order to accomplish this. In particular, the very formulation of the conjecture relies on a comparison isomorphism between the $\ell$-adic and de Rham realizations of the motive provided by theorems of Faltings [Fa] or Tsuji [Ts], and verification of the conjecture requires the careful application of such a theorem. We also need to generalize results on congruences between modular forms to higher weight, and to compute certain local Tamagawa numbers.
0.1. Some history
Special values of $L$-functions have long played an important role in number theory. The underlying principle is that the values of $L$-functions at integers reflect arithmetic properties of the object used to define them. A prime example of this is Dirichlet’s class number formula; another is the Birch and Swinnerton-Dyer conjecture. The Tamagawa number conjecture of Bloch and Kato [B-K], refined by Fontaine, Kato and Perrin-Riou [Kato2, F-P, Fo4], is a vast generalization of these. Roughly speaking, they predict the precise value of the first non-vanishing derivative of the $L$-function at zero (hence any integer) for every motive over $\mathbb{Q}$}. This was already done up to a rational multiple by conjectures of Deligne and Beilinson; the additional precision of the Bloch-Kato conjecture can be thought of as a generalized class number formula, where ideal class groups are replaced by groups defined using Galois cohomology.
Dirichlet’s class number formula amounts to the conjecture for the Dedekind zeta function for a number field at $s=0$ or 11. The conjecture is also known up to a power of 22 for Dirichlet $L$-functions (including the Riemann zeta function) at any integer ([Ma-W], [B-K], [H-K]). It is known up to an explicit set of bad primes for the $L$-function of a CM elliptic curve at $s=1$ if the order of vanishing is $\leq 1$ ([C-W], [Ru1], [Ko]). There are also partial results for $L$-functions of other modular forms at the central critical value ([G-Z], [Kato3], [Ko-L], [N], [Z]) and for values of certain Hecke $L$-functions ([Hn], [Gu2], [Ki]).
Here we consider the adjoint $L$-function of a modular form of weight $k\geq 2$ at $s=0$ and 11. Special values of the $L$-function associated to the adjoint of a modular form, and more generally, twists of its symmetric square, have been studied by many mathematicians. A method of Rankin relates the values to those of Petersson inner products, and this was used by Ogg [Ogg], Shimura [Sh4], Sturm [St1, St2], Coates and Schmidt [C-S, Schm] to obtain nonvanishing results and rationality results along the lines of Deligne’s conjecture. Hida [Hi1] related the precise value to a number measuring congruences between modular forms. In the case of forms corresponding to (modular) elliptic curves, results relating the value to certain Galois cohomology groups (Selmer groups) were obtained by Coates and Schmidt in the context of Iwasawa theory, and by one of the authors, who in [Fl3] obtained results in the direction of the Bloch-Kato conjecture.
A key point of Wiles’ paper [Wi] is that for many elliptic curves, modularity could be deduced from a formula relating congruences and Galois cohomology [Wi]. This formula could be regarded as a primitive form of the Bloch-Kato conjecture for the adjoint motive of a modular form. His attempt to prove it using the Euler system method introduced in [Fl3] was not successful except in the CM case using generalizations of results in [Gu1] and [Ru2]. Wiles, in work completed with Taylor [T-W], eventually proved his formula using a new construction which could be viewed as a kind of “horizontal Iwasawa theory.”
In this paper, we refine the method of [Wi] and [T-W], generalize it to higher weight modular forms and relate the result to the Bloch-Kato conjecture. Ultimately, we prove the conjecture for the adjoint of an arbitrary newform of weight $k\geq 2$ up to an explicit finite set of bad primes. We should stress the importance of making this set as small and explicit as possible; indeed the refinements in [Di3], [C-D-T] and [BCDT] which completed the proof of the Shimura-Taniyama-Weil conjecture can be viewed as work in this direction for weight two modular forms. In this paper, we make use of some of the techniques introduced in [Di3] and [C-D-T], as well as the modification of Taylor-Wiles construction in [Di5] and [Fu]. One should be able to improve our results using current technology in the weight two case (using [C-D-T], [BCDT] and [Sav]), and in the ordinary case (using [Dic], [S-W1]); one just has to relate the results in those papers to the Bloch-Kato conjecture. Finally we remark that Wiles’ method has been generalized to the setting of Hilbert modular forms by Fujiwara [Fu] and Skinner-Wiles [S-W2], but it seems much harder to extract results on special values from their work.
0.2. The framework
The Bloch-Kato conjecture is formulated in terms of “motivic structures,” a term referring to the usual collection of cohomological data associated to a motive. This data consists of:
  • vector spaces $M_{?}$, called realizations, for $\mathop{\rm B}\nolimits$}, $\mathop{\rm dR}\nolimits$} and $\ell$ for each rational prime $\ell$, each with extra structure (involution, filtration or Galois action);
  • comparison isomorphisms relating the realizations;
  • a weight filtration.
Suppose that $f$ is a newform of weight $k\geq 2$ and level NN. Much of the paper is devoted to the construction of the motivic structure $A_{f}$ for which we prove the conjecture. This construction is not new; it is due for the most part to Eichler, Shimura, Deligne, Jannsen, Scholl and Faltings ([Sh1], [De1], [Ja], [Scho2], [Fa]) We review it however in order to collect the facts we need, provide proofs we could not find in the literature, and set things up in a way suited to the formulation of the Bloch-Kato conjecture.
Let us briefly recall here how the construction works. We start with the modular curve $X_{N}$ parametrizing elliptic curves with level NN structure. Then one takes the Betti, de Rham and $\ell$-adic cohomology of $X_{N}$ with coefficients in a sheaf defined as the (k2)(k-2)-nd symmetric power of the relative cohomology of the universal elliptic curve over $X_{N}$. These come with the additional structure and comparison isomorphisms needed to define a motivic structure $M_{N,k}$, the comparison between $\ell$-adic and de Rham cohomology being provided by a theorem of Faltings [Fa]. The structures $M_{N,k}$ can also be defined as in [Scho2] using Kuga-Sato varieties; this has the advantage of showing they arise from “motives” and provides the option of applying Tsuji’s comparison theorem [Ts]. However the construction using “coefficient sheaves” is better suited to defining and comparing lattices in the realizations which play a key role in the proof.
The structures $M_{N,k}$ also come with an action of the Hecke operators and a perfect pairing. The Hecke action is used to “cut out” a piece $M_{f}$, which corresponds to the newform $f$ and has rank two over the field generated by the coefficients of $f$. The pairing comes from Poincaré duality, is related to the Petersson inner product and restricts to a perfect pairing on $M_{f}$. We finally take the trace zero endomorphisms of $M_{f}$ to obtain the motivic structure $A_{f}={\mathop{\rm ad}\nolimits}^{0}M_{f}$. The construction also yields integral structures ${\mathcal{M}}_{f}$ and ${\mathcal{A}}_{f}$, consisting of lattices in the various realizations and integral comparison isomorphisms outside a set of bad primes.
Our presentation of the Bloch-Kato conjecture is much influenced by its reformulation and generalization due to Fontaine and Perrin-Riou. Their version assumes the existence of a category of motives with conjectural properties. Without assuming conjectures however, they define a category ${\bf SPM}_{{\mathbb{Q}}}({\mathbb{Q}})$ of premotivic structures whose objects consist of realizations with additional structure and comparison isomorphisms. The category of mixed motives is supposed to admit a fully faithful functor to it, and a motivic structure is an object of the essential image. Their version of the Bloch-Kato conjecture is then stated in terms of $\operatorname{{\mathop{\rm Ext}\nolimits}}$ groups of motivic structures, but whenever there is an explicit “motivic” construction of (conjecturally) all the relevant extensions, the conjecture can be formulated entirely in terms of premotivic structures. This happens in our case, for all the relevant $\operatorname{{\mathop{\rm Ext}\nolimits}}$’s conjecturally vanish. There will therefore be no further mention of motives in this paper.
We make some other slight modifications to the framework of [F-P].
  • We work with premotivic structures with coefficients in a number field $K$.
  • We forget about the $\ell$-adic realization and comparison isomorphisms at a finite set of “bad” primes $T$.
This yields a version of the conjecture which predicts the value of $L(A_{f},0)$ up to an $T$-unit in $K$. We make our set $T$ explicit: Let $T_{f}$ be the set of finite primes $\lambda$ in $K$ such that either:
  • $\lambda\mid Nk!$
  • the two-dimensional residual Galois representation ${\mathcal{M}}_{f,\lambda}/\lambda{\mathcal{M}}_{f,\lambda}$ is not absolutely irreducible when restricted to $G_{F}$, where $F={\mathbb{Q}}(\sqrt{(-1)^{(\ell-1)/2}\ell})$ and $\lambda\mid\ell$.
Note that since $T_{f}$ includes the set of primes dividing $Nk!$, we will only be applying Faltings’ comparison theorem in the “easy” case of crystalline representations whose associated Dieudonné module has short filtration length.
0.3. The main theorems
Our main result can be stated as follows.
Theorem 0.1.
(=Theorem 8.10) Let $f$ be a newform of weight $k\geq 2$ and level NN with coefficients in $K$. If $\lambda$ is not in $T_{f}$, then the $\lambda$-part of the Bloch-Kato conjecture holds for $A_{f}$ and $A_{f}(1)$.
The main tool in the proof is the construction of Taylor and Wiles, which we axiomatize (Theorem 7.3), and apply to higher weight forms to obtain the following generalization of their class number formula.
Theorem 0.2.
(=Theorem 7.15) Let $f$ be a newform of weight $k\geq 2$ and level NN with coefficients in $K$. Suppose $\Sigma$ is a finite set of rational primes containing those dividing NN. Suppose that $\lambda$ is a prime of $K$ which is not in $T_{f}$ and does not divide any prime in $\Sigma$. Then the ${\mathcal{O}}_{K,\lambda}$-module
$H^{1}_{\Sigma}(G_{\mathbb{Q}},A_{f,\lambda}/{\mathcal{A}}_{f,\lambda})$
has length $v_{\lambda}(\eta_{f}^{\Sigma})$.
Here $\eta_{f}^{\Sigma}$, defined in § 6.4, is a generalization of the congruence ideal of Hida and Wiles; it can also be viewed as measuring the failure of the pairing on $M_{f}$ to be perfect on ${\mathcal{M}}_{f}$.
Another consequence of Theorem 0.2, is the following result in the direction of Fontaine-Mazur conjecture [F-M].
Theorem 0.3.
(=Theorem 7.12) Suppose $\rho:G_{\mathbb{Q}}\to\operatorname{{\mathop{\rm GL}\nolimits}}_{2}(K_{\lambda})$ is a continuous geometric representation whose restriction to $G_{\ell}$ is ramified and crystalline and its associated Dieudonné module has filtration length less than $\ell-1$. If its residual representation is modular and absolutely irreducible restricted to ${\mathbb{Q}}(\sqrt{(-1)^{(\ell-1)/2}\ell})$ where $\lambda\mid\ell$, then $\rho$ is modular.
0.4. Acknowledgements
Research on this project was carried out while the first author worked at Cambridge, MIT and Rutgers, visited the IAS, IHP and Paris VII, and received support from the EPSRC, NSF and an AMS Centennial Fellowship. The second author would like to thank the IAS for its hospitality and acknowledge support from the NSF and the Sloan foundation. The third author was supported in part by an NSF grant and a research grant from University of Georgia in the early stages of this project, and thanks the IAS for its hospitality.
1. Generalities and examples of premotivic structures
$1.$ Galois representations
For a field $F$, $\bar{F}$ will denote an algebraic closure, and $G_{F}=\operatorname{{\mathop{\rm Gal}\nolimits}}(\bar{F}/F)$. We fix embeddings $\bar{{\mathbb{Q}}}\to\bar{{\mathbb{Q}}}_{p}$ for each prime $p$, and an embedding $\bar{{\mathbb{Q}}}\to{\mathbb{C}}$. If $F$ is a number field, we let ${\bf I}_{F}$ denote the set of embeddings $F\to\bar{{\mathbb{Q}}}$, which we identify with the set of embeddings $F\to{\mathbb{C}}$ via our fixed one of $\bar{{\mathbb{Q}}}$ in $\mathbb{C}$}.
Suppose that $F$ is a finite extension of ${\mathbb{Q}}_{p}$ in $\bar{{\mathbb{Q}}}_{p}$. We let $F_{0}$ denote the maximal unramified subextension of $F$ and write $F^{\rm ur}$ for the maximal unramified extension of $F$ in $\bar{{\mathbb{Q}}}_{p}$. Let $\nu$ denote the natural map $G_{F}\to G_{{\mathbb{F}}_{p}}\cong\hat{{\mathbb{Z}}}$ where we view $\hat{{\mathbb{Z}}}$ as being generated by the geometric Frobenius element $\operatorname{{\mathop{\rm Frob}\nolimits}}_{p}$, which we also view as an automorphism of ${\mathbb{Q}}_{p}^{\rm ur}$. We write $\phi_{p}$ for the arithmetic Frobenius, so $\phi_{p}=\operatorname{{\mathop{\rm Frob}\nolimits}}_{p}^{-1}$. Let $I_{F}$ denote the inertia subgroup of $G_{F}$, $W_{F}$ the Weil subgroup, $\nu^{-1}({\mathbb{Z}})$ and $^{\prime}W_{F}$ the Weil-Deligne group of $F$ (see [De2, §8]). Recall that if $K$ is a field of characteristic zero, then to give a representation of $^{\prime}W_{F}$ with coefficients in $K$ is equivalent to giving a finite-dimensional $K$-vector space $V$ (with the discrete topology), a continuous homomorphism $\rho:W_{F}\to\operatorname{\operatorname{{\mathop{\rm Aut}\nolimits}}}_{K}V$ and a nilpotent endomorphism NN of $V$ satisfying $\rho(g)N=p^{-\nu(g)}N\rho(g)$. If $F={\mathbb{Q}}_{p}$, we write simply $G_{p}$, $I_{p}$, $W_{p}$ and $^{\prime}W_{p}$ for these groups and identify $I_{p}\subset W_{p}\subset G_{p}$ with their images in $G_{\mathbb{Q}}$.
If $K$ is a number field, then $S_{\bf f}(K)$ denotes the set of finite places of $K$. Suppose that $\lambda\in S_{\bf f}(K)$ divides $\ell\in S_{\bf f}({\mathbb{Q}})$ and $F\subset\bar{{\mathbb{Q}}}_{\ell}$ is a finite extension of $\mathbb{Q}$$_{\ell}$. Let $B_{\mathop{\rm dR}\nolimits}=B_{{\mathop{\rm dR}\nolimits},\ell}$, $B_{\mbox{\scriptsize crys}}=B_{{\mbox{\scriptsize crys}},\ell}$ and $B_{\rm st}=B_{{\rm st},\ell}$ be the rings defined by Fontaine [Fo2, §2],[F-P, I.2.1]. Suppose that $V$ is a finite-dimensional $K_{\lambda}$-representation (i.e., a $\lambda$-adic representation) of $G_{F}$. Then $D_{\mathop{\rm dR}\nolimits}(V)=(B_{\mathop{\rm dR}\nolimits}\otimes_{{\mathbb{Q}}_{\ell}}V)^{G_{F}}$ is filtered free $F\otimes_{{\mathbb{Q}}_{\ell}}K_{\lambda}$-module of finite rank, and $V$ is called de Rham if $\dim_{F}D_{\mathop{\rm dR}\nolimits}(V)=\dim_{{\mathbb{Q}}_{\ell}}V$. Similarly $D_{\mbox{\scriptsize crys}}(V)=(B_{\mbox{\scriptsize crys}}\otimes_{{\mathbb{Q}}_{\ell}}V)^{G_{F}}$ is a filtered free $F_{0}\otimes_{{\mathbb{Q}}_{\ell}}K_{\lambda}$-module of finite rank with a $(\phi_{\ell}\otimes 1)$-semilinear endomorphism $\phi$, and $V$ is called crystalline if $\dim_{F_{0}}D_{\mbox{\scriptsize crys}}(V)=\dim_{{\mathbb{Q}}_{\ell}}V$. Similarly $D_{\rm st}(V)=(B_{\rm st}\otimes_{{\mathbb{Q}}_{\ell}}V)^{G_{F}}$ is a filtered free $F_{0}\otimes_{{\mathbb{Q}}_{\ell}}K_{\lambda}$-module of finite rank with a $(\phi_{\ell}\otimes 1)$-semilinear endomorphism $\phi$ and an endomorphism NN satisfying $N\phi=\ell\phi N$, and $V$ is called semistable if $\dim_{F_{0}}D_{\rm st}(V)=\dim_{{\mathbb{Q}}_{\ell}}V$. One defines $D_{\mathop{\rm pst}\nolimits}(V)$ as the direct limit of $(B_{\rm st}\otimes_{{\mathbb{Q}}_{\ell}}V)^{G_{F^{\prime}}}$ where $F^{\prime}$ runs over finite extensions of $F$ in $\bar{{\mathbb{Q}}}_{\ell}$; this is a filtered free ${\mathbb{Q}}_{\ell}^{\rm ur}\otimes_{{\mathbb{Q}}_{\ell}}K_{\lambda}$-module and one checks that $(\rho,N)$ with $\rho(g)=g\phi^{\nu(g)}$ makes $D_{\mathop{\rm pst}\nolimits}(V)$ a $K_{\lambda}$-rational representation of $^{\prime}W_{F}$, and $V$ is called potentially semistable if $\dim_{{\mathbb{Q}}_{\ell}^{\rm ur}}D_{\mathop{\rm pst}\nolimits}(V)=\dim_{{\mathbb{Q}}_{\ell}}V$.
Recall that $V$ is crystalline if and only if $V$ is semistable and $N=0$ on $D_{\rm st}(V)$, in which case $D_{\mbox{\scriptsize crys}}(V)=D_{\rm st}(V)$; $V$ is semistable if and only if $V$ is potentially semistable and $I_{F}$ acts trivially on $D_{\mathop{\rm pst}\nolimits}(V)$, in which case $D_{\mathop{\rm pst}\nolimits}(V)={\mathbb{Q}}_{\ell}^{\rm ur}\otimes_{F_{0}}D_{\rm st}(V)$ with the action of $g\in W_{F}$ given by $\phi^{\nu(g)}$; $V$ is potentially semistable if and only if $V^{\prime}={\mathop{\rm res}\nolimits}_{G_{F}}^{G_{F^{\prime}}}V$ is semistable for some finite extension $F^{\prime}$ of $F$, in which case $D_{\mathop{\rm pst}\nolimits}(V)={\mathbb{Q}}_{\ell}^{\rm ur}\otimes_{F_{0}^{\prime}}D_{\rm st}(V^{\prime})$; and if $V$ is potentially semistable, then $V$ is de Rham with $F^{\rm ur}\otimes_{F}D_{\mathop{\rm dR}\nolimits}(V)=F^{\rm ur}\otimes_{{\mathbb{Q}}_{\ell}^{\rm ur}}D_{\mathop{\rm pst}\nolimits}(V)$ (with $D_{\mathop{\rm dR}\nolimits}(V)=F\otimes_{F_{0}}D_{\rm st}(V)$ if $V$ is semistable).
A $\lambda$-adic representation $V$ of $G_{\mathbb{Q}}$ is pseudo-geometric [F-P, II.2] (resp. geometric [F-P, II.3]) if it is unramified outside of a finite number of places of $\mathbb{Q}$} and its restriction to $G_{\ell}$ is de Rham (resp. potentially semistable). The representation $V$ is said to have good reduction at $p$ if its restriction to $G_{p}$ is crystalline (resp. unramified) if $p=\ell$ (resp. $p\neq\ell$).
We also recall some of the theory of Fontaine and Laffaille [F-L]. Suppose that $F\subset\bar{{\mathbb{Q}}}_{\ell}$ is a finite unramified extension of $\mathbb{Q}$$_{\ell}$. We let ${{\mathcal{M}\mathcal{F}}}({\mathcal{O}}_{F})$ denote the category whose objects are finitely generated ${\mathcal{O}}_{F}$-modules equipped with
  • a decreasing filtration such that $\operatorname{{\mathop{\rm Fil}\nolimits}}^{a}A=A$ and $\operatorname{{\mathop{\rm Fil}\nolimits}}^{b}A=0$ for some $a,b\in{\mathbb{Z}}$, and for each $i\in{\mathbb{Z}}$, $\operatorname{{\mathop{\rm Fil}\nolimits}}^{i}A$ is a direct summand of AA;
  • $\phi_{\ell}$-semilinear maps $\phi^{i}:\operatorname{{\mathop{\rm Fil}\nolimits}}^{i}A\to A$ for $i\in{\mathbb{Z}}$ satisfying $\phi^{i}|_{\operatorname{{\mathop{\rm Fil}\nolimits}}^{i+1}A}=\ell\phi^{i+1}$ and $A=\sum{\mathop{\rm Im}\nolimits}\phi^{i}$.
It follows from [F-L, 1.8] that ${{\mathcal{M}\mathcal{F}}}({\mathcal{O}}_{F})$ is an abelian category. Let ${{\mathcal{M}\mathcal{F}}}^{a}({\mathcal{O}}_{F})$ denote the full subcategory of objects AA satisfying $\operatorname{{\mathop{\rm Fil}\nolimits}}^{a}A=A$ and $\operatorname{{\mathop{\rm Fil}\nolimits}}^{a+\ell}A=0$ and having no non-trivial quotients $A^{\prime}$ such that $\operatorname{{\mathop{\rm Fil}\nolimits}}^{a+\ell-1}A^{\prime}=A^{\prime}$, and let ${{\mathcal{M}\mathcal{F}}}^{a}_{\mathop{\rm tor}\nolimits}({\mathcal{O}}_{F})$ denote the full subcategory of ${{\mathcal{M}\mathcal{F}}}^{a}({\mathcal{O}}_{F})$ consisting of objects of finite length. So ${{\mathcal{M}\mathcal{F}}}^{0}_{\mathop{\rm tor}\nolimits}({\mathcal{O}}_{F})$ is the category denoted $\underline{MF}_{\mathop{\rm tor}\nolimits}^{f,\ell^{\prime}}$ in [F-L], and it follows from [F-L, 6.1] that ${{\mathcal{M}\mathcal{F}}}^{a}({\mathcal{O}}_{F})$ and ${{\mathcal{M}\mathcal{F}}}^{a}_{\mathop{\rm tor}\nolimits}({\mathcal{O}}_{F})$ are abelian categories, stable under taking subobjects, quotients, direct products and extensions in ${{\mathcal{M}\mathcal{F}}}({\mathcal{O}}_{F})$. We write simply $\mathcal{M}\mathcal{F}$}}, ${{\mathcal{M}\mathcal{F}}}^{a}$, etc. whenever $F={\mathbb{Q}}_{\ell}$.
[…]
  • $\operatorname{{\mathop{\rm Fil}\nolimits}}^{0}D=D$ and $\operatorname{{\mathop{\rm Fil}\nolimits}}^{\ell}D=0$, where $D=(B_{\mbox{\scriptsize crys}}\otimes_{{\mathbb{Q}}_{\ell}}V)^{G_{F}}$;
  • if $V^{\prime}$ is a nonzero quotient of $V$, then $V^{\prime}\otimes_{{\mathbb{Q}}_{\ell}}{\mathbb{Q}}_{\ell}(\ell-1)$ is ramified.
In particular, the essential image of $\mathbb{V}$} is closed under taking subobjects, quotients and finite direct sums. Furthermore, one sees from [F-L, 8.4] that the natural transformations
(1) $\begin{array}[]{l}F\otimes_{{\mathcal{O}}_{F}}A\to(B_{\mbox{\scriptsize crys}}\otimes_{{\mathbb{Z}}_{\ell}}{\mathbb{V}}(A))^{G_{F}},\\ B_{\mbox{\scriptsize crys}}\otimes_{{\mathcal{O}}_{F}}A\to B_{\mbox{\scriptsize crys}}\otimes_{{\mathbb{Z}}_{\ell}}{\mathbb{V}}(A){\rm\ and}\\ \operatorname{{\mathop{\rm Fil}\nolimits}}^{0}(B_{\mbox{\scriptsize crys}}\otimes_{{\mathbb{Z}}_{\ell}}A)^{\phi=1}\to{\mathbb{Q}}_{\ell}\otimes_{{\mathbb{Z}}_{\ell}}{\mathbb{V}}(A)\end{array}$
are isomorphisms.
If $K$ is a number field and $\lambda\in S_{\bf f}(K)$ is a prime over $\ell$, we let ${\mathcal{O}}_{\lambda}={\mathcal{O}}_{K,\lambda}$ and let $\text{${\mathcal{O}}_{\lambda}$-${{\mathcal{M}\mathcal{F}}}$}^{a}$ denote the category of ${\mathcal{O}}_{\lambda}$-modules in ${{\mathcal{M}\mathcal{F}}}^{a}$. We can regard $\mathbb{V}$} as a functor from ${\mathcal{O}}_{\lambda}$-${{\mathcal{M}\mathcal{F}}}^{0}$ to the category of ${\mathcal{O}}_{\lambda}[G_{\ell}]$-modules.
If AA and $A^{\prime}$ are objects of ${\mathcal{O}}_{\lambda}$-${{\mathcal{M}\mathcal{F}}}^{0}$ such that $A\otimes_{{\mathcal{O}}_{\lambda}}A^{\prime}$ defines an object of ${\mathcal{O}}_{\lambda}$-${{\mathcal{M}\mathcal{F}}}^{0}$, then there is a canonical isomorphism
${\mathbb{V}}(A\otimes_{{\mathcal{O}}_{\lambda}}A^{\prime})\cong{\mathbb{V}}(A)\otimes_{{\mathcal{O}}_{\lambda}}{\mathbb{V}}(A^{\prime}).$
Analogous assertions hold for ${\mathop{\rm Hom}\nolimits}_{{\mathcal{O}}_{\lambda}}(A,A^{\prime})$.
1.2. Premotivic structures
We work with categories of premotivic structures based on notions from [F-P] and [B-K].
For a number field $K$, we let $\operatorname{{\bf PM}}_{K}$ denote the category of premotivic structures over $\mathbb{Q}$} with coefficients in $K$. In the notation of [F-P, III.2.1], this is the category ${\bf SPM}_{\mathbb{Q}}({\mathbb{Q}})\otimes K$ of $K$-modules in ${\bf SPM}_{{\mathbb{Q}}}({\mathbb{Q}})$. Thus an object $M$ of $\operatorname{{\bf PM}}_{K}$ consists of the following data:
  • a finite-dimensional $K$-vector space $M_{\mathop{\rm B}\nolimits}$ with an action of $G_{{\mathbb{R}}}$,
  • a finite-dimensional $K$-vector space $M_{{\mathop{\rm dR}\nolimits}}$ with a finite decreasing filtration $\operatorname{{\mathop{\rm Fil}\nolimits}}^{i}$, called the Hodge filtration;
  • for each $\lambda\in S_{\bf f}(K)$, a finite-dimensional $K_{\lambda}$ vector space $M_{\lambda}$ with a continuous pseudo-geometric (see [F-P, II.2.1]) action of $G_{\mathbb{Q}}$;
  • a ${\mathbb{C}}\otimes K$-linear isomorphism
    $I^{\infty}:{\mathbb{C}}\otimes M_{\mathop{\rm dR}\nolimits}\to{\mathbb{C}}\otimes M_{\mathop{\rm B}\nolimits}$
    respecting the action of $G_{{\mathbb{R}}}$ (where $G_{\mathbb{R}}$ acts on ${\mathbb{C}}\otimes M_{\mathop{\rm B}\nolimits}$ diagonally and acts on ${\mathbb{C}}\otimes M_{\mathop{\rm dR}\nolimits}$ via the first factor);
  • for each $\lambda\in S_{\bf f}(K)$, a $K_{\lambda}$-linear isomorphism
    $I^{\lambda}_{\mathop{\rm B}\nolimits}:K_{\lambda}\otimes_{K}M_{\mathop{\rm B}\nolimits}\to M_{\lambda}$
    respecting the action of $G_{{\mathbb{R}}}$ (where the action on $M_{\lambda}$ is via the restriction $G_{\mathbb{R}}\to G_{\mathbb{Q}}$ determined by our choice of embedding $\bar{{\mathbb{Q}}}\to{\mathbb{C}}$);
  • for each $\lambda\in S_{\bf f}(K)$, $B_{{\mathop{\rm dR}\nolimits},\ell}\otimes_{{\mathbb{Q}}_{\ell}}K_{\lambda}$-linear isomorphism
    $I^{\lambda}:B_{{\mathop{\rm dR}\nolimits},\ell}\otimes_{{\mathbb{Q}}_{\ell}}K_{\lambda}\otimes_{K}M_{{\mathop{\rm dR}\nolimits}}\to B_{{\mathop{\rm dR}\nolimits},\ell}\otimes_{{\mathbb{Q}}_{\ell}}M_{\lambda}$
    respecting filtrations and the action of $G_{{\mathbb{Q}}_{\ell}}$ (where $\ell$ is the prime which $\lambda$ divides, $K_{\lambda}$ and $M_{\lambda}$ are given the degree-0 filtration, $K_{\lambda}$ and $M_{\mathop{\rm dR}\nolimits}$ are given the trivial $G_{{\mathbb{Q}}_{\ell}}$-action and the action on $M_{\lambda}$ is determined by our choice of embedding $\bar{{\mathbb{Q}}}\to\bar{{\mathbb{Q}}}_{\ell}$);
  • increasing weight filtrations $W^{i}$ on $M_{\mathop{\rm B}\nolimits}$, $M_{{\mathop{\rm dR}\nolimits}}$ and each $M_{\lambda}$ respecting all of the above data, and such that ${\mathbb{R}}\otimes M_{\mathop{\rm B}\nolimits}$ with its Galois action and weight filtration, together with the Hodge filtration on ${\mathbb{C}}\otimes M_{\mathop{\rm B}\nolimits}$ defined via $I_{\mathop{\rm B}\nolimits}$, defines a mixed Hodge structure over $\mathbb{R}$} (see [F-P, III.1]).
If $S\subsetneq S_{\bf f}(K)$ is a set of primes of $K$, we let $\operatorname{{\bf PM}}_{K}^{S}$ denote the category defined in exactly the same way, but with $S_{\bf f}(K)$ replaced by the complement of $S$. If $S\subseteq S^{\prime}$, we use $\cdot^{S^{\prime}}$ to denote the forgetful functor from $\operatorname{{\bf PM}}_{K}^{S}$ to $\operatorname{{\bf PM}}_{K}^{S^{\prime}}$.
The category $\operatorname{{\bf PM}}_{K}^{S}$ is equipped with a tensor product, which we denote $\otimes_{K}$, and an internal hom, which we denote ${\mathop{\rm Hom}\nolimits}_{K}$. There is also a unit object, which we denote simply by $K$. These are defined in the obvious way; for example, $(M\otimes_{K}N)_{\mathop{\rm B}\nolimits}$ is the $K[G_{\mathbb{R}}]$-module $M_{\mathop{\rm B}\nolimits}\otimes_{K}N_{\mathop{\rm B}\nolimits}$. If $K\subseteq K^{\prime}$, we let $S^{K^{\prime}}$ denote the set of primes in $S_{\bf f}(K^{\prime})$ lying over those in $S$, then $K^{\prime}\otimes_{K}\cdot$ defines a functor from $\operatorname{{\bf PM}}_{K}^{S}$ to $\operatorname{{\bf PM}}_{K^{\prime}}^{S^{K^{\prime}}}$.
Definition $1.$
An object $M$ of $\operatorname{{\bf PM}}_{K}^{S}$
  • has good reduction at $p$ if for each $\lambda$ in $S$, $M_{\lambda}|G_{p}$ is crystalline at $p$ if $\lambda|p$ and unramified at $p$ otherwise;
  • is $L$-admissible at $p$ if $M_{\lambda}$ is geometric for every $\lambda\not\in S$ dividing $p$ and the representations $D_{{\mathop{\rm pst}\nolimits}}(M_{\lambda}|G_{p})^{{\mathop{\rm ss}\nolimits}}$, for $\lambda\not\in S$, form a compatible system of $K$-rational representations of the Deligne-Weil group $^{\prime}W_{p}$ (see [De2, §8]);
  • is $L$-admissible everywhere if it is $L$-admissible at $p$ for all primes pp.
If $M$ is $L$-admissible at $p$, then the local factor associated to $D_{{\mathop{\rm pst}\nolimits}}(M_{\lambda}|G_{p})^{{\mathop{\rm ss}\nolimits}}$ is of the form $P(p^{-s})^{-1}$ for some polynomial $P(u)\in K[u]$ independent of $\lambda$ not in $S$. For an embedding $\tau:K\to{\mathbb{C}}$ we put $L_{p}(M,\tau,s)=\tau P(p^{-s})$ and we regard the collection $\{L_{p}(M,\tau,s)\}_{\tau\in{\bf I}_{K}}$ as a meromorphic function on $\mathbb{C}$} with values in ${\mathbb{C}}^{{\bf I}_{K}}\cong K\otimes{\mathbb{C}}$. If $S$ is finite and $M$ is $L$-admissible everywhere, then its $L$-function
$L(M,s):=\prod_{p}L_{p}(M,s)$
is a holomorphic $K\otimes{\mathbb{C}}$-valued function in some right half plane ${\rm Re}(s)>r$ (with components $L(M,\tau,s)=\prod_{p}L_{p}(M,\tau,s)$).
For each $\tau\in{\bf I}_{K}$ the Hodge structure ${\mathbb{C}}\otimes_{K,\tau}M_{\mathop{\rm B}\nolimits}$ over $\mathbb{R}$} has attached to it an L-function $L_{\infty}(M,\tau,s)$ [De3] and we denote by $L_{\infty}(M,s)$ the resulting function with values in $K\otimes{\mathbb{C}}$. Similarly, if $M$ is $L$-admissible everywhere we define the $K\otimes{\mathbb{C}}$-valued $\epsilon$-factor $\epsilon(M,s)=\epsilon(M)c(M)^{-s}$ where the components $\epsilon(M,\tau,s)$ are defined in [De2]. $L_{\infty}(M,\tau,s)$ is essentially a product of $\Gamma$-functions, $c(M)$ (the conductor of $M$) is a positive integer divisible by $p$ if and only if $M$ does not have good reduction at $p$, and $\epsilon(M)$ is independent of $s$. Put $\Lambda(M,s):=L_{\infty}(M,s)L(M,s)$. Combining Conjecture 1 of [F-M] with Deligne’s conjectured functional equation leads to the following conjecture (see [F-P, III.4.3]):
Conjecture 1.2.
If $M$ is $L$-admissible everywhere, then $L(M,s)$ converges to a holomorphic function on some right half-plane and extends meromorphically to $\mathbb{C}$}. Moreover $\Lambda(M,s)$ satisfies the functional equation
$\Lambda(M,s)=\epsilon(M,s)\Lambda(M^{*},1-s)$
where $M^{*}={\mathop{\rm Hom}\nolimits}_{K}(M,K)$ is the contragredient of $M$, and it is holomorphic if $M$ does not have a direct summand isomorphic to the unit premotivic structure $K$.
We also define a category ${\mathcal{P}M}^{S}_{K}$ of $S$-integral premotivic structures as follows. We let ${\mathcal{O}}_{S}={\mathcal{O}}_{K,S}$ denote the set of $x\in K$ with $v_{\lambda}(x)\geq 0$ for all $\lambda\not\in S$. An object $\mathcal{M}$} of ${\mathcal{P}M}^{S}_{K}$ consists of the following data:
  • a finitely generated ${\mathcal{O}}_{K}$-module ${\mathcal{M}}_{\mathop{\rm B}\nolimits}$ with an action of $G_{{\mathbb{R}}}$,
  • a finitely generated ${\mathcal{O}}_{S}$-module ${\mathcal{M}}_{{\mathop{\rm dR}\nolimits}}$ with a finite decreasing filtration $\operatorname{{\mathop{\rm Fil}\nolimits}}^{i}$, called the Hodge filtration;
  • for each $\lambda\in S_{\bf f}(K)$, a finitely generated ${\mathcal{O}}_{\lambda}$-module ${\mathcal{M}}_{\lambda}$ with continuous action of $G_{\mathbb{Q}}$ inducing a pseudo-geometric action on ${\mathcal{M}}_{\lambda}\otimes_{{\mathcal{O}}_{\lambda}}K_{\lambda}$;
  • for each $\lambda\not\in S$, an object ${\mathcal{M}}_{{\lambda\mbox{\scriptsize-crys}}}$ of $\text{${\mathcal{O}}_{\lambda}$-${{\mathcal{M}\mathcal{F}}}$}^{0}$;
  • an ${\mathbb{R}}\otimes{\mathcal{O}}_{K}$-linear isomorphism
    $I^{\infty}:{\mathbb{C}}\otimes{\mathcal{M}}_{\mathop{\rm dR}\nolimits}\to{\mathbb{C}}\otimes{\mathcal{M}}_{\mathop{\rm B}\nolimits}$
    respecting the action of $G_{{\mathbb{R}}}$;
  • for each $\lambda$ in $S_{\bf f}(K)$ an isomorphism
    $I^{\lambda}_{\mathop{\rm B}\nolimits}:{\mathcal{M}}_{\mathop{\rm B}\nolimits}\otimes_{{\mathcal{O}}_{K}}{\mathcal{O}}_{\lambda}\cong{\mathcal{M}}_{\lambda}$
    respecting the action of $G_{\mathbb{R}}$;
  • for each $\lambda\not\in S$ an ${\mathcal{O}}_{\lambda}$-linear isomorphism
    $I^{\lambda}_{\mathop{\rm dR}\nolimits}:{\mathcal{M}}_{\mathop{\rm dR}\nolimits}\otimes_{{\mathcal{O}}}{\mathcal{O}}_{\lambda}\cong{\mathcal{M}}_{{\lambda\mbox{\scriptsize-crys}}}$
    respecting filtrations;
  • for each $\lambda\not\in S$, an ${\mathcal{O}}_{\lambda}$-linear isomorphism
    $I^{\lambda}:{\mathbb{V}}({\mathcal{M}}_{{\lambda\mbox{\scriptsize-crys}}})\to{\mathcal{M}}_{\lambda}$
    respecting the action of $G_{{\mathbb{Q}}_{\ell}}$, where $\ell$ is the prime which $\lambda$ divides.
  • increasing weight filtrations $W^{i}$ on ${\mathbb{Q}}\otimes{\mathcal{M}}_{\mathop{\rm B}\nolimits}$, ${\mathbb{Q}}\otimes{\mathcal{M}}_{{\mathop{\rm dR}\nolimits}}$ and each ${\mathbb{Q}}\otimes{\mathcal{M}}_{\lambda}$ respecting all of the above data and giving rise to a mixed Hodge structure.
With the evident notion of morphism this becomes an ${\mathcal{O}}_{K}$-linear category. In particular, it is an abelian category. Note that the forgetful functor sending $\mathcal{M}$} to ${\mathcal{M}}_{B}$ is faithful. Note also that there is a natural functor $\mathbb{Q}$$\otimes\cdot$ from ${\mathcal{P}M}_{K}^{S}$ to $\operatorname{{\bf PM}}_{K}^{S}$, where we set $({\mathbb{Q}}\otimes{\mathcal{M}})_{?}={\mathbb{Q}}\otimes$$\mathcal{M}$$_{?}$ for $\mathop{\rm B}\nolimits$}, $\mathop{\rm dR}\nolimits$} and $\lambda$ for $\lambda\not\in S$, with induced additional structure and comparison isomorphisms. (The comparison $I^{\lambda}$ for $\mathbb{Q}$$\otimes{\mathcal{M}}$ is defined as the composite
$B_{{\mathop{\rm dR}\nolimits},\ell}\otimes_{{\mathbb{Z}}_{\ell}}{\mathcal{O}}_{\lambda}\otimes_{\mathcal{O}}{\mathcal{M}}_{{\mathop{\rm dR}\nolimits}}\cong B_{{\mathop{\rm dR}\nolimits},\ell}\otimes_{{\mathbb{Z}}_{\ell}}{\mathcal{M}}_{{\lambda\mbox{\scriptsize-crys}}}\to B_{{\mathop{\rm dR}\nolimits},\ell}\otimes_{{\mathbb{Z}}_{\ell}}{\mathbb{V}}({\mathcal{M}}_{{\lambda\mbox{\scriptsize-crys}}})\to B_{{\mathop{\rm dR}\nolimits},\ell}\otimes_{{\mathbb{Z}}_{\ell}}{\mathcal{M}}_{\lambda},$
where the maps are respectively, $I^{\lambda}_{\mathop{\rm dR}\nolimits}$, the canonical map (1) and $I^{\lambda}$, each with scalars extended to $B_{{\mathop{\rm dR}\nolimits},\ell}$.)
If $S\subset S^{\prime}$, we define a functor $\cdot^{S^{\prime}}$ from ${\mathcal{P}M}_{K}^{S}$ to ${\mathcal{P}M}_{K}^{S^{\prime}}$ in the obvious way. We say that $\mathcal{M}$} is $S^{\prime}$-flat if ${\mathcal{M}}_{\mathop{\rm dR}\nolimits}^{S^{\prime}}={\mathcal{M}}_{\mathop{\rm dR}\nolimits}\otimes_{{\mathcal{O}}_{K,S}}{\mathcal{O}}_{K,S^{\prime}}$ is flat over ${\mathcal{O}}_{K,S^{\prime}}$. Note that if $\mathcal{M}$} is $S^{\prime}$-flat, then so is any subobject of $\mathcal{M}$}. Let $K^{\prime}$ be a finite extension of $K$. We also have a natural functor ${\mathcal{O}}_{K^{\prime}}\otimes_{{\mathcal{O}}_{K}}\cdot$ from ${\mathcal{P}M}_{K}^{S}$ to ${\mathcal{P}M}_{K^{\prime}}^{S^{\prime}}$ where $S^{\prime}$ is the set of primes over those in $S$.
We say that $\mathcal{M}$} has good reduction at $p$, is $L$-admissible at $p$ or is $L$-admissible everywhere according to whether the same is true for $\mathbb{Q}$$\otimes{\mathcal{M}}$. Note that if $\mathcal{M}$} is $L$-admissible at $p$ and $p$ is not invertible in $\mathcal{O}$}, then $\mathcal{M}$} necessarily has good reduction at pp.
For objects $\mathcal{M}$} and $\mathcal{M}$$^{\prime}$ of ${\mathcal{P}M}_{K}^{S}$, we can form ${\mathcal{M}}\otimes_{{\mathcal{O}}_{K}}{\mathcal{M}}^{\prime}$ in ${\mathcal{P}M}_{K}^{S}$ provided ${\mathcal{M}}_{{\lambda\mbox{\scriptsize-crys}}}\otimes_{{\mathcal{O}}_{\lambda}}{\mathcal{M}}^{\prime}_{{\lambda\mbox{\scriptsize-crys}}}$ defines an object of ${{\mathcal{M}\mathcal{F}}}^{0}_{\lambda}$ for all $\lambda\not\in S$. In particular this holds if there exist positive integers $a$, $a^{\prime}$ such that $\operatorname{{\mathop{\rm Fil}\nolimits}}^{a}{\mathcal{M}}_{\mathop{\rm dR}\nolimits}=0$, $\operatorname{{\mathop{\rm Fil}\nolimits}}^{a^{\prime}}{\mathcal{M}}^{\prime}_{\mathop{\rm dR}\nolimits}=0$ and $a+a^{\prime}<\ell$ for all primes $\ell$ not invertible in $\mathcal{O}$}. If $\mathcal{N}$} and $\mathcal{N}$$^{\prime}$ are objects of ${\mathcal{P}M}_{K}^{S}$ such that ${\mathcal{N}}\otimes_{{\mathcal{O}}_{K}}{\mathcal{N}}^{\prime}$ is as well, and if $\alpha:{\mathcal{M}}\to{\mathcal{N}}$ and $\alpha^{\prime}:{\mathcal{M}}^{\prime}\to{\mathcal{N}}^{\prime}$ are morphisms in ${\mathcal{P}M}_{K}^{S}$, then there is a well-defined morphism $\alpha\otimes\alpha^{\prime}:{\mathcal{M}}\otimes_{{\mathcal{O}}_{K}}{\mathcal{N}}\to{\mathcal{M}}^{\prime}\otimes_{{\mathcal{O}}_{K}}{\mathcal{N}}^{\prime}$ in ${\mathcal{P}M}_{K}^{S}$. Analogous assertions hold for the formation and properties of ${\mathop{\rm Hom}\nolimits}_{{\mathcal{O}}_{K}}({\mathcal{M}},{\mathcal{M}}^{\prime})$.
Note that if $\mathcal{M}$} is an object of ${\mathcal{P}M}_{K}^{S}$, then $\operatorname{{\mathop{\rm End}\nolimits}}{\mathcal{M}}$ is a finitely generated ${\mathcal{O}}_{K}$-module. If II is an ${\mathcal{O}}_{K}$-submodule of $\operatorname{{\mathop{\rm End}\nolimits}}{\mathcal{M}}$, then we define an object ${\mathcal{M}}[I]$ of ${\mathcal{P}M}_{K}^{S}$ as the kernel of
$(x_{1},\ldots,x_{r}):{\mathcal{M}}\to{\mathcal{M}}^{r}$
where $x_{1},\ldots,x_{r}$ generate II. This is independent of the choice of generators. This applies in particular when II is the image in $\operatorname{{\mathop{\rm End}\nolimits}}{\mathcal{M}}$ of an ideal in a commutative ${\mathcal{O}}_{K}$-algebra $R$ mapping to $\operatorname{{\mathop{\rm End}\nolimits}}{\mathcal{M}}$, or the augmentation ideal in ${\mathcal{O}}_{K}[G]$ where $G$ is a group acting on $\mathcal{M}$}. In the latter case, we write ${\mathcal{M}}^{G}$ instead of ${\mathcal{M}}[I]$.
1.3. Basic examples
The object ${\mathbb{Q}}(-1)$ in $\operatorname{{\bf PM}}_{\mathbb{Q}}$ is the weight two premotivic structure defined by $H^{1}({\mathbb{G}}_{\bf m})$. To give an explicit description, let $\varepsilon$ denote the generator of ${\mathbb{Z}}_{\ell}(1)=\displaystyle{\lim_{\longleftarrow}}\,\mu_{\ell^{n}}(\bar{{\mathbb{Q}}})$ defined by $(e^{2\pi i/\ell^{n}})_{n}$ via our fixed embedding $\bar{{\mathbb{Q}}}\to{\mathbb{C}}$.
  • Let ${\mathcal{T}}_{B}=H^{1}_{B}({\mathbb{G}}_{\bf m}({\mathbb{C}}),{\mathbb{Z}})\cong(2\pi i)^{-1}{\mathbb{Z}}\subset{\mathbb{C}}$ with complex conjugation in $G_{\mathbb{R}}$ acting by 1-1, and let ${\mathbb{Q}}(-1)_{B}={\mathbb{Q}}\otimes{\mathcal{T}}_{B}$.
  • Let ${\mathcal{T}}_{{\mathop{\rm dR}\nolimits}}=H^{1}_{{\mathop{\rm dR}\nolimits}}({\mathbb{G}}_{\bf m}/{\mathbb{Z}})$, which with its Hodge filtration is isomorphic to ${\mathbb{Z}}[-1]$ (where $[n]$ denotes a shift by $n$ in the filtration, so $\operatorname{{\mathop{\rm Fil}\nolimits}}^{i}V[n]=\operatorname{{\mathop{\rm Fil}\nolimits}}^{i+n}V$). Write $\iota$ for the canonical basis of ${\mathcal{T}}_{{\mathop{\rm dR}\nolimits}}\cong{\mathbb{Z}}[-1]$ and let ${\mathbb{Q}}(-1)_{{\mathop{\rm dR}\nolimits}}={\mathbb{Q}}\otimes{\mathcal{T}}_{{\mathop{\rm dR}\nolimits}}$, so $\operatorname{{\mathop{\rm Fil}\nolimits}}^{1}{\mathbb{Q}}(-1)_{{\mathop{\rm dR}\nolimits}}={\mathbb{Q}}\iota$ and $\operatorname{{\mathop{\rm Fil}\nolimits}}^{2}{\mathbb{Q}}(-1)_{{\mathop{\rm dR}\nolimits}}=0$.
  • Let ${\mathcal{T}}_{\ell}=H^{1}_{{\mathop{\rm et}\nolimits}}({\mathbb{G}}_{{\bf m},\bar{{\mathbb{Q}}}},{\mathbb{Z}}_{\ell})\cong{\mathop{\rm Hom}\nolimits}_{{\mathbb{Z}}_{\ell}}({\mathbb{Z}}_{\ell}(1),{\mathbb{Z}}_{\ell})={\mathbb{Z}}_{\ell}\delta$ where $\delta(\varepsilon)=1$, and let ${\mathbb{Q}}(-1)_{\ell}={\mathbb{Q}}\otimes{\mathcal{T}}_{\ell}$.
  • Let ${\mathcal{T}}_{\ell\mbox{\scriptsize-crys}}$ denote the object of $\mathcal{M}\mathcal{F}$$_{\ell}$ defined by ${\mathbb{Z}}_{\ell}\otimes{\mathcal{T}}_{{\mathop{\rm dR}\nolimits}}={\mathbb{Z}}_{\ell}~\iota$ with $\phi^{1}(\iota)=\iota$.
  • $I^{\infty}:{\mathbb{C}}\otimes_{\mathbb{Z}}{\mathcal{T}}_{\mathop{\rm dR}\nolimits}\to{\mathbb{C}}\otimes_{\mathbb{Z}}{\mathcal{T}}_{\mathop{\rm B}\nolimits}$ is defined by $1\otimes\iota\mapsto 2\pi i\otimes(2\pi i)^{-1}$.
  • $I^{\ell}_{\mathop{\rm B}\nolimits}:{\mathbb{Z}}_{\ell}\otimes{\mathcal{T}}_{\mathop{\rm B}\nolimits}\to{\mathcal{T}}_{\ell}$ is defined by $1\otimes(2\pi i)^{-1}\mapsto\delta$.
  • $I^{\ell}_{\mathop{\rm dR}\nolimits}:{\mathcal{T}}_{\mathop{\rm dR}\nolimits}\otimes_{\mathbb{Z}}{\mathbb{Z}}_{\ell}\to{\mathcal{T}}_{\ell\mbox{\scriptsize-crys}}$ is given by $\iota\mapsto\iota$.
  • $I^{\ell}:B_{{\mathop{\rm dR}\nolimits},\ell}\otimes{\mathbb{Q}}[-1]\to B_{{\mathop{\rm dR}\nolimits},\ell}\otimes_{{\mathbb{Q}}_{\ell}}{\mathbb{Q}}(-1)_{\ell}$ is defined by $1\otimes\iota\mapsto t\otimes\delta$ where $t=\log[\varepsilon]$.
  • For $\ell>2$, ${\mathcal{T}}_{\ell\mbox{\scriptsize-crys}}$ is an object of ${{\mathcal{M}\mathcal{F}}}^{0}$ and $I^{\ell}$ is induced by an isomorphism ${\mathbb{V}}({\mathcal{T}}_{\ell\mbox{\scriptsize-crys}})\cong{\mathcal{T}}_{\ell}$.
The above data defines objects in $\operatorname{{\bf PM}}_{\mathbb{Q}}$ and ${\mathcal{P}M}^{\{2\}}_{\mathbb{Q}}$ which we denote by ${\mathbb{Q}}(-1)$ and $\mathcal{T}$}. These could be described equivalently by $H^{2}({\mathbb{P}}^{1})$, or indeed $H^{2}(X)$ for any smooth, proper, geometrically connected curve $X$ over $\mathbb{Q}$}.
The Tate premotivic structure ${\mathbb{Q}}(1)$ is the object of $\operatorname{{\bf PM}}_{\mathbb{Q}}$ defined by ${\mathop{\rm Hom}\nolimits}_{\mathbb{Q}}({\mathbb{Q}}(-1),$ ${\mathbb{Q}})$. More generally, for any object $M$ in $\operatorname{{\bf PM}}_{K}$ and integer $n$, $M(n)$ is defined as $M\otimes_{{\mathbb{Q}}}{\mathbb{Q}}(1)^{\otimes n}$. If $M$ has good reduction at $p$, then so does $M(n)$. If $M$ is $L$-admissible at $p$, then so is $M(n)$. If they are $L$-admissible everywhere, then $L(M(n),s)=L(M,s+n)$, $\Lambda(M(n),s)=\Lambda(M,s+n)$, $c(M(n))=c(M)$, $\epsilon(M(n),s)=\epsilon(M,s+n)$ and Conjecture 1.2 for $M$ and $M(n)$ are equivalent. In particular, the premotivic structure ${\mathbb{Q}}(n)$ has good reduction and is $L$-admissible at $p$ for all primes $p$, and $L({\mathbb{Q}}(n),s)=\zeta(s+n)$ where $\zeta$ is the Riemann $\zeta$-function. Furthermore $L_{\infty}({\mathbb{Q}}(n),s)=\Gamma_{\mathbb{R}}(s+n)$ where $\Gamma_{\mathbb{R}}(s)=L_{\infty}({\mathbb{Q}},s)=\pi^{-s/2}\Gamma(s/2)$, $\epsilon({\mathbb{Q}}(n),s)=1$ and Conjecture 1.2 for $\Lambda({\mathbb{Q}}(n),s)$ is the functional equation for $\zeta(s)$.
For any integer $n\geq 0$, ${\mathcal{T}}^{S,\otimes n}$ defines an object of ${\mathcal{P}M}^{S}_{\mathbb{Q}}$ where $S$ is any set of primes containing those dividing $(n+1)!$; note also that ${\mathbb{Q}}\otimes{\mathcal{T}}^{S,\otimes n}\cong{\mathbb{Q}}(-n)^{S}$.
For any number field $F\subset\bar{{\mathbb{Q}}}$ let $M_{F}$ denote the premotivic structure $M_{F}$ of weight zero defined by $H^{0}(\operatorname{{\mathop{\rm Spec}\nolimits}}F)$, called the Dedekind premotivic structure of $F$. To give an explicit description, let $S$ denote the set of primes dividing $D=\operatorname{{\mathop{\rm Disc}\nolimits}}(F/{\mathbb{Q}})$. We let
  • ${\mathcal{M}}_{F,{\mathop{\rm B}\nolimits}}={\mathbb{Z}}^{{\bf I}_{F}}$ with the natural action of $G_{\mathbb{R}}$, and $M_{F,{\mathop{\rm B}\nolimits}}={\mathbb{Q}}\otimes{\mathcal{M}}_{F,{\mathop{\rm B}\nolimits}}={\mathbb{Q}}^{{\bf I}_{F}}$. (Recall that we identified ${\bf I}_{F}$ with the set of embeddings $F\to{\mathbb{C}}$ via the chosen embedding $\bar{{\mathbb{Q}}}\to{\mathbb{C}}$, so for $\alpha:{\bf I}_{F}\to{\mathbb{Z}}$ and $\sigma\in G_{\mathbb{R}}$, we define $\sigma\alpha$ by $\tau\mapsto\alpha(\sigma^{-1}\circ\tau)$).
  • ${\mathcal{M}}_{F,{\mathop{\rm dR}\nolimits}}={\mathcal{O}}_{F}[1/D]$ with $\operatorname{{\mathop{\rm Fil}\nolimits}}^{0}{\mathcal{M}}_{F,{\mathop{\rm dR}\nolimits}}={\mathcal{M}}_{F,{\mathop{\rm dR}\nolimits}}$ and $\operatorname{{\mathop{\rm Fil}\nolimits}}^{1}{\mathcal{M}}_{F,{\mathop{\rm dR}\nolimits}}=0$, and $M_{F,{\mathop{\rm dR}\nolimits}}={\mathbb{Q}}\otimes{\mathcal{M}}_{F,{\mathop{\rm dR}\nolimits}}=F$.
  • ${\mathcal{M}}_{F,\ell}={\mathbb{Z}}_{\ell}\otimes{\mathcal{M}}_{F,{\mathop{\rm B}\nolimits}}={\mathbb{Z}}_{\ell}^{{\bf I}_{F}}$ with the natural action of $G_{\mathbb{Q}}$, and $M_{F,\ell}={\mathbb{Q}}\otimes{\mathcal{M}}_{F,\ell}={\mathbb{Q}}_{\ell}^{{\bf I}_{F}}$ (so for $\alpha:{\bf I}_{F}\to{\mathbb{Z}}_{\ell}$, $\sigma\in G_{\mathbb{Q}}$ and $\tau:F\to\bar{{\mathbb{Q}}}$, we have $(\sigma\alpha)(\tau)=\alpha(\sigma^{-1}\circ\tau)$).
  • ${\mathcal{M}}_{F,{\ell\mbox{\scriptsize-crys}}}={\mathbb{Z}}_{\ell}\otimes{\mathcal{M}}_{F,{\mathop{\rm dR}\nolimits}}={\mathbb{Z}}_{\ell}\otimes{\mathcal{O}}_{F}$ for $\ell\not\in S$, with the same filtration as ${\mathcal{M}}_{F,{\mathop{\rm dR}\nolimits}}$ and with $\phi^{0}=\phi_{\ell}$.
  • Identifying ${\mathbb{C}}\otimes M_{F,B}$ with ${\mathbb{C}}^{{\bf I}_{F}}$, $I^{\infty}$ is defined by $I^{\infty}(1\otimes x)(\tau)=\tau(x)$.
  • $I^{\ell}_{\mathop{\rm B}\nolimits}$ is the identity map.
  • $I^{\ell}_{\mathop{\rm dR}\nolimits}$ is the identity map for $\ell\not\in S$.
  • Identifying $B_{{\mathop{\rm dR}\nolimits},\ell}\otimes M_{F,\ell}$ with $B_{{\mathop{\rm dR}\nolimits},\ell}^{{\bf I}_{F}}$ and ${\bf I}_{F}$ with the set of embeddings of $F$ in $B_{{\mathop{\rm dR}\nolimits},\ell}$ via the chosen embedding $\bar{{\mathbb{Q}}}\to\bar{{\mathbb{Q}}}_{p}$, $I^{\ell}$, is defined by $I^{\ell}(1\otimes x)(\tau)=\tau(x)$. For $\ell\not\in S$, this is induced by an isomorphism ${\mathbb{V}}({\mathcal{M}}_{F,{\ell\mbox{\scriptsize-crys}}})\to{\mathcal{M}}_{F,\ell}$ which we also denote $I^{\ell}$.
The above data defines objects ${\mathcal{M}}_{F}$ of ${\mathcal{P}M}^{S}_{\mathbb{Q}}$ and $M_{F}$ of $\operatorname{{\bf PM}}_{\mathbb{Q}}$ with ${\mathbb{Q}}\otimes{\mathcal{M}}_{F}=M_{F}^{S}$. One finds that $M_{F}$ has good reduction at all primes $p$ not dividing $D$, is $L$-admissible everywhere, and $L(M_{F},s)=\zeta_{F}(s)$ where $\zeta_{F}$ is the $\zeta$-function of $F$. Furthermore $c(M_{F})=D$, $\epsilon(M_{F})=D^{1/2}$ and $L_{\infty}(M_{F},s)=\Gamma_{\mathbb{R}}(s)^{r_{1}}\Gamma_{\mathbb{C}}(s)^{r_{2}}$ where $r_{1}$ (resp. $r_{2}$) is the number of real (resp. complex) embeddings of $F$, $\Gamma_{\mathbb{R}}$ was defined above and $\Gamma_{\mathbb{C}}=2\cdot(2\pi)^{-s}\Gamma(s)$. Conjecture 1.2 is known for $\Lambda(M_{F},s)$ (see [La2, Ch. XIII] for example).
Note that if $F$ is Galois over $\mathbb{Q}$}, then there is a natural action of $G=\operatorname{{\mathop{\rm Gal}\nolimits}}(F/{\mathbb{Q}})$ on ${\mathcal{M}}_{F}$ defined as follows: For $\alpha\in{\mathcal{M}}_{F,B}$, $g\in G$ and $\tau\in{\bf I}_{F}$, we define $g\alpha$ by $\tau\mapsto\alpha(\tau\circ g)$. The action on ${\mathcal{M}}_{F,\ell}$ is defined similarly. The action on ${\mathcal{M}}_{F,{\mathop{\rm dR}\nolimits}}={\mathcal{O}}_{F}[1/D]$ is the obvious one, as is the action on ${\mathcal{M}}_{F,{\ell\mbox{\scriptsize-crys}}}$. One checks that the action respects all the additional structure, and so defines an action on ${\mathcal{M}}_{F}$ as an object in ${\mathcal{P}M}^{S}_{\mathbb{Q}}$. We similarly have an action of $G$ on $M_{F}\in\operatorname{{\bf PM}}_{\mathbb{Q}}$.
Suppose that we are given a character $\psi:\hat{{\mathbb{Z}}}^{\times}\to K^{\times}$. We regard $\psi$ also as a character of $\mathbb{A}$$^{\times}$ and $G_{\mathbb{Q}}$ via the isomorphisms $\hat{{\mathbb{Z}}}^{\times}\cong{\mathbb{A}}^{\times}/{\mathbb{R}}_{>0}^{\times}{\mathbb{Q}}^{\times}\cong G_{\mathbb{Q}}^{\mathop{\rm ab}\nolimits}$, where the first isomorphism is induced by the natural inclusion $\hat{{\mathbb{Z}}}^{\times}\to{\mathbb{A}}^{\times}$ and the second is given by class field theory. (Our convention is that a uniformizer in ${\mathbb{Q}}_{p}^{\times}$ maps to $\operatorname{{\mathop{\rm Frob}\nolimits}}_{p}$ in the Galois group of any abelian extension of $\mathbb{Q}$} unramified at pp.)
Suppose we are given an embedding of a number field $F\to\bar{{\mathbb{Q}}}$ such that $\psi$ is trivial on the image of $G_{F}$ in $G_{\mathbb{Q}}$. We can then regard $\psi$ as a character of $G=\operatorname{{\mathop{\rm Gal}\nolimits}}(F/{\mathbb{Q}})$ and define the Dirichlet premotivic structure $M_{\psi}$ as $(V\otimes M_{F})^{G}$ where $V=K$ with $G$ acting by $\psi$. One can check that the construction is independent of the choice of $F$ and embedding $F\to\bar{{\mathbb{Q}}}$ as above. To describe $M_{\psi}$ explicitly, we choose $F={\mathbb{Q}}(e^{2\pi i/N})\subset{\mathbb{C}}$ where $\psi$ has conductor NN. We let $\tau_{0}:F\to\bar{{\mathbb{Q}}}$ denote the embedding compatible with our fixed $\bar{{\mathbb{Q}}}\to{\mathbb{C}}$, and we regard $\psi$ as a Dirichlet character via the canonical isomorphism $({\mathbb{Z}}/N{\mathbb{Z}})^{\times}\cong G$. Let $S$ denote the set of primes in $K$ lying over those dividing NN and define an object ${\mathcal{M}}_{\psi}$ of ${\mathcal{P}M}_{K}^{S}$ by $({\mathcal{V}}\otimes{\mathcal{M}}_{F})^{G}$ where ${\mathcal{V}}={\mathcal{O}}_{K}^{S}$ with $G$ acting by $\psi$. We then have:
  • ${\mathcal{M}}_{\psi,{\mathop{\rm B}\nolimits}}$ is the ${\mathcal{O}}_{K}$-submodule of ${\mathcal{O}}_{K}^{{\bf I}_{F}}$ spanned by the map $b_{B}$ defined by $\tau_{0}\circ g\mapsto\psi^{-1}(g)$, where $\tau_{0}$ is the inclusion of $F$ in $\bar{{\mathbb{Q}}}$;
  • ${\mathcal{M}}_{\psi,{\mathop{\rm dR}\nolimits}}$ is the ${\mathcal{O}}={\mathcal{O}}_{K}[1/N]$-submodule of ${\mathcal{O}}_{K}\otimes{\mathcal{O}}_{F}[1/N]$ spanned by $b_{{\mathop{\rm dR}\nolimits}}=\displaystyle\sum_{a}\psi(a)\otimes e^{2\pi ia/N}$, where $a$ runs over $({\mathbb{Z}}/N{\mathbb{Z}})^{\times}$ with $\operatorname{{\mathop{\rm Fil}\nolimits}}^{1}{\mathcal{M}}_{\psi,{\mathop{\rm dR}\nolimits}}=0$ and $\operatorname{{\mathop{\rm Fil}\nolimits}}^{0}{\mathcal{M}}_{\psi,{\mathop{\rm dR}\nolimits}}={\mathcal{M}}_{\psi,{\mathop{\rm dR}\nolimits}}$;
  • ${\mathcal{M}}_{\psi,\lambda}={\mathcal{O}}_{\lambda}\otimes_{{\mathcal{O}}_{K}}{\mathcal{M}}_{\psi,{\mathop{\rm B}\nolimits}}$ with $G_{\mathbb{Q}}$ acting via $\psi$;
[…]
The comparison isomorphisms are induced from those of ${\mathcal{M}}_{F}$. Similarly, we get the object $M_{\psi}$ of $\operatorname{{\bf PM}}_{K}$ by setting $M_{\psi,?}={\mathbb{Q}}\otimes{\mathcal{M}}_{\psi,?}$ with comparison isomorphisms induced from those of $M_{F}$. In particular, we have $I^{\infty}(1\otimes b_{{\mathop{\rm dR}\nolimits}})=G_{\psi}(1\otimes b_{B})$ where $G_{\psi}$ is the Gauss sum $\sum_{a}e^{2\pi ia/N}\otimes\psi(a)$ in ${\mathbb{C}}\otimes K$.
We have that ${\mathbb{Q}}\otimes{\mathcal{M}}_{\psi}\cong M_{\psi}^{S}$, $M_{\psi}$ has good reduction at all primes not dividing the conductor of $\psi$ and is $L$-admissible everywhere, and $L(M_{\psi},s)$ is the Dirichlet $L$-function $L(\psi^{-1},s)$. Moreover, Conjecture 1.2 is known for $\Lambda(M_{\psi},s)$ and one has $c(M_{\psi})=N$, $\epsilon(M_{\psi})=i^{-\eta}G_{\psi^{-1}}$ and $L_{\infty}(M_{\psi},s)=\Gamma_{\mathbb{R}}(s+\eta)$, where $\psi(-1)=(-1)^{\eta}$, $\eta\in\{0,1\}$ (see [Wa, Ch. 4] for example).
2. Premotivic structures for modular forms
In this section we review the construction of premotivic structures associated to the space of modular forms of weight $k$ and level NN. More precisely, if $k\geq 2$ and $N\geq 3$, we construct objects of ${\mathcal{P}M}_{\mathbb{Q}}^{S}$ whose de Rham realization contains the space of such forms, where $S=S_{N}$ is the set of primes dividing $Nk!$.
2.1. Level NN modular curves
Let $k$ and NN be integers with $k\geq 2$ and $N\geq 3$. Let $T=\operatorname{{\mathop{\rm Spec}\nolimits}}{\mathbb{Z}}[1/Nk!]$, and consider the functor which associates to a $T$-scheme $T^{\prime}$ the set of isomorphism classes of generalized elliptic curves over $T^{\prime}$ with level NN structure [D-R, IV.6.6]. By [D-R, IV.6.7], the functor is represented by a smooth, proper curve over $T$. We denote this curve by $\bar{X}$, and we let $\bar{s}:\bar{E}\to\bar{X}$ denote the universal generalized elliptic curve with level NN structure. We let $X$ denote the open subscheme of $\bar{X}$ over which $\bar{E}$ is smooth. Then $X$ is the complement of a reduced divisor, called the cuspidal divisor, which we denote by $X^{\infty}$. We let $E=\bar{s}^{-1}X$, $s=\bar{s}|_{E}$ and $E^{\infty}=\bar{E}\times_{\bar{X}}X^{\infty}$. Using the arguments of [D-R, VII.2.4], one can check that $\bar{E}$ is smooth over $T$ and $E^{\infty}$ is a reduced divisor with strict normal crossings (in the sense of [G-M, 1.8] as well as [SGA1, XIII.2.1]).
Let us also recall the standard description of
$\bar{s}^{\mathop{\rm an}\nolimits}:\bar{E}^{\mathop{\rm an}\nolimits}\to\bar{X}^{\mathop{\rm an}\nolimits},$
where we use $^{\mathop{\rm an}\nolimits}$ to denote the associated complex analytic space. We let
$X_{N}=\coprod_{t\in({\mathbb{Z}}/N{\mathbb{Z}})^{\times}}X_{N,t},$
where for each $t$, $X_{N,t}$ denotes a copy of $\Gamma(N)\backslash{\mathfrak{H}}$, the quotient of the complex upper half-plane $\mathfrak{H}$}, by the principal congruence subgroup $\Gamma(N)$ of $\operatorname{{\mathop{\rm SL}\nolimits}}_{2}({\mathbb{Z}})$. Similarly we let
$\bar{X}_{N}=\coprod_{t\in({\mathbb{Z}}/N{\mathbb{Z}})^{\times}}\bar{X}_{N,t},$
where $\bar{X}_{N,t}$ is the compactification of $X_{N,t}$ obtained by adjoining the cusps. We write $\bar{X}_{N}^{\mathop{\rm alg}\nolimits}$ for the corresponding algebraic curve over $\mathbb{C}$}.
For each $t$, we define the complex analytic surface $E_{N,t}$ to be a copy of the quotient
[…]
where $(m,n)\in{\mathbb{Z}}\times{\mathbb{Z}}$ acts on $\mathfrak{H}$$\times{\mathbb{C}}$ via $(\tau,z)\mapsto(\tau,z+m\tau+n)$, and $\gamma={\mbox{\scriptsize$\left(\begin{array}[]{cc}{{a}}&{{b}}\\ {{c}}&{{d}}\end{array}\right)$}}\in\Gamma(N)$ acts by sending the class of $(\tau,z)$ to that of $(\gamma(\tau),(c\tau+d)^{-1}z)$. We can regard $E_{N}=\coprod E_{N,t}$ as a complex analytic family of elliptic curves over $X_{N}$. We can then extend $E_{N}$ to a family $\bar{E}_{N}$ of generalized elliptic curves over $\bar{X}_{N}$ using analytic Tate curves, as in [D-R, VII.4]. More precisely, for $\alpha>0$, we let $D_{\alpha}$ denote the disk in $D[q^{1/N}]$ defined by $|q|<e^{-2\pi\alpha}$ and let $D_{\alpha}^{0}$ denote the corresponding punctured disk. Then over a punctured neighborhood around each cusp in $\bar{X}_{N}$, we have an isomorphism between $E_{N}$ and the NN-sided analytic Tate curve $E_{q}^{\mathop{\rm an}\nolimits}$ over $D_{1}^{0}$, and we define $\bar{E}_{N}$ by gluing NN-sided analytic Tate curves over $D_{1}$ to $E_{N}$. For each $a$, we define a level NN-structure on $E_{N,t}$ by the pair of sections $X_{N,t}\to E_{N,t}$ induced by $\tau\mapsto(\tau,\tau/N)$ and $\tau\mapsto(\tau,t/N)$. The resulting level NN-structure on $E_{N}$ in fact extends to $\bar{E}_{N}$.
Using for example [Ha, App. B, 3.4], one checks that $\bar{E}_{N}$ is algebraic, so $\bar{E}_{N}\to\bar{X}_{N}$ is the analytification of a generalized elliptic curve $\bar{E}_{N}^{\mathop{\rm alg}\nolimits}\to\bar{X}_{N}^{\mathop{\rm alg}\nolimits}$. One can check the resulting morphism $\bar{X}_{N}^{\mathop{\rm alg}\nolimits}\to\bar{X}_{\mathbb{C}}$ induces a bijection $X_{N}\to X({\mathbb{C}})$, and since both are smooth, proper algebraic curves over $\mathbb{C}$}, the map is an isomorphism. The analytification of the universal generalized elliptic curve with level NN-structure is therefore isomorphic to $\bar{E}_{N}\to\bar{X}_{N}$ with the level NN-structure defined above.
$2.$ Realizations of level NN modular forms
To construct the Betti realization, define ${\mathcal{F}}_{B}$ as the locally constant sheaf $R^{1}s^{\mathop{\rm an}\nolimits}_{*}{\mathbb{Z}}$ on $X^{\mathop{\rm an}\nolimits}$. The stalk of ${\mathcal{F}}_{B}$ at a point $x=\Gamma(N)\tau\in X_{N,t}$ can be canonically identified with $H^{1}(E_{x}^{\mathop{\rm an}\nolimits},{\mathbb{Z}})\cong{\mathop{\rm Hom}\nolimits}(\Lambda_{\tau},{\mathbb{Z}})$ where $\Lambda_{\tau}={\mathbb{Z}}\tau\oplus{\mathbb{Z}}$.
Let ${\mathcal{F}}_{B}^{k}=\operatorname{{\mathop{\rm Sym}\nolimits}}^{k-2}_{{\mathbb{Z}}}{\mathcal{F}}_{B}$, ${\mathcal{M}}_{B}=H^{1}(X^{\mathop{\rm an}\nolimits},{\mathcal{F}}_{B}^{k})$, and ${\mathcal{M}}_{c,B}=H^{1}_{c}(X^{\mathop{\rm an}\nolimits},{\mathcal{F}}_{B}^{k})$. Note that if $\ell$ is a prime greater than k2k-2, then multiplication by $\ell$ is injective on ${\mathcal{F}}_{B}^{k}$. Since $H^{0}_{c}(X^{\mathop{\rm an}\nolimits},{\mathbb{F}}_{\ell}\otimes{\mathcal{F}}_{B}^{k})=0$, we see that ${\mathcal{M}}_{c,B}$ has no $\ell$-torsion if $\ell>k-2$. If $k=2$, then the map
$H^{0}(X^{\mathop{\rm an}\nolimits},{\mathcal{F}}_{B}^{k})\to H^{0}(X^{\mathop{\rm an}\nolimits},{\mathbb{F}}_{\ell}\otimes{\mathcal{F}}_{B}^{k})$
is surjective, and if $k>2$ and $\ell$ does not divide N(k2)!N(k-2)!, then this last group vanishes. We conclude that ${\mathcal{M}}_{B}$ has no $\ell$-torsion if $k=2$ or $\ell$ does not divide N(k2)!N(k-2)!.
If $F_{E}$ is complex conjugation on $E^{\mathop{\rm an}\nolimits}$ and $F_{X}$ is complex conjugation on $X^{\mathop{\rm an}\nolimits}$, then the commutative diagram
[…]
gives an isomorphism
$F_{X}^{*}{\mathcal{F}}_{B}=F_{X}^{*}R^{1}s^{\mathop{\rm an}\nolimits}_{*}{\mathbb{Z}}\to R^{1}s^{\mathop{\rm an}\nolimits}_{*}F_{E}^{*}{\mathbb{Z}}={\mathcal{F}}_{B},$
and this induces an isomorphism $F_{X}^{*}{\mathcal{F}}_{B}^{k}\to{\mathcal{F}}_{B}^{k}$. The action of complex conjugation in $G_{\mathbb{R}}$ on ${\mathcal{M}}_{B}$ is defined as the composite
$H^{1}(X^{\mathop{\rm an}\nolimits},{\mathcal{F}}_{B}^{k})\stackrel{{\scriptstyle F_{X}^{*}}}{{\longrightarrow}}H^{1}(X^{\mathop{\rm an}\nolimits},F_{X}^{*}{\mathcal{F}}_{B}^{k})\to H^{1}(X^{\mathop{\rm an}\nolimits},{\mathcal{F}}_{B}^{k}),$
and the action on ${\mathcal{M}}_{c,B}$ is defined similarly. We let $M_{B}={\mathbb{Q}}\otimes{\mathcal{M}}_{B}$ and $M_{c,B}={\mathbb{Q}}\otimes{\mathcal{M}}_{c,B}$.
For any finite prime $\ell$, we let $\mathcal{F}$$_{\ell}$ denote the $\ell$-adic sheaf $R^{1}s_{*}{\mathbb{Z}}_{\ell}$ on $X$. This is a lisse $\ell$-adic sheaf whose stalk at a geometric point $\bar{x}:\operatorname{{\mathop{\rm Spec}\nolimits}}\bar{k}\to X$ can be canonically identified with $H^{1}(E_{\bar{x}},{\mathbb{Z}}_{\ell})\cong{\mathop{\rm Hom}\nolimits}(E_{\bar{x}}[\ell^{\infty}],{\mathbb{Q}}_{\ell}/{\mathbb{Z}}_{\ell})$.
Let ${\mathcal{F}}_{\ell}^{k}=\operatorname{{\mathop{\rm Sym}\nolimits}}^{k-2}_{{\mathbb{Z}}_{\ell}}{\mathcal{F}}_{\ell}$, ${\mathcal{M}}_{\ell}=H^{1}(X_{\bar{{\mathbb{Q}}}},{\mathcal{F}}_{\ell}^{k})$ and ${\mathcal{M}}_{c,\ell}=H^{1}_{c}(X_{\bar{{\mathbb{Q}}}},{\mathcal{F}}_{\ell}^{k})$. Then $G_{\mathbb{Q}}$ acts on on $\mathcal{M}$$_{\ell}$ and ${\mathcal{M}}_{c,\ell}$ by transport of structure. We let $M_{\ell}={\mathbb{Q}}\otimes{\mathcal{M}}_{\ell}$ and $M_{c,\ell}={\mathbb{Q}}\otimes{\mathcal{M}}_{c,\ell}$.
The comparison theorem between Betti and $\ell$-adic cohomology shows that $R^{1}s^{\mathop{\rm an}\nolimits}_{*}{\mathbb{Z}}_{\ell}$ is canonically isomorphic to the sheaf on $X^{\mathop{\rm an}\nolimits}$ associated to the $\ell$-adic sheaf $R^{1}s_{*}{\mathbb{Z}}_{\ell}$ on $X_{\mathbb{C}}$ (see Theorem. I.11.6 and p. 130 of [F-K]). Identifying $x\in X^{\mathop{\rm an}\nolimits}$ with a geometric point $\operatorname{{\mathop{\rm Spec}\nolimits}}{\mathbb{C}}\to X$, the isomorphism on the corresponding stalks is the canonical one between ${\mathop{\rm Hom}\nolimits}(E_{x}[\ell^{\infty}],{\mathbb{Q}}_{\ell}/{\mathbb{Z}}_{\ell})$ and $H^{1}(E_{x},{\mathbb{Z}}_{\ell})$.
Taking symmetric powers and applying the comparison theorem again gives isomorphisms
${\mathbb{Z}}_{\ell}\otimes{\mathcal{M}}_{B}\cong{\mathcal{M}}_{\ell}\quad\mbox{and}\quad{\mathbb{Z}}_{\ell}\otimes{\mathcal{M}}_{c,B}\cong{\mathcal{M}}_{c,\ell}$
respecting the action of $G_{\mathbb{R}}$.
The construction of the de Rham realizations is similar to the one given in [Scho1] except we use the language of log schemes [Kato1]. We let ${\mathcal{N}}_{E}$ denote the log structure on $\bar{E}$ associated to $E^{\infty}$, and let ${\mathcal{N}}_{X}$ denote the log structure associated to $X^{\infty}$ [Kato1, 1.5]. By [Kato1, 3.5], the morphisms $(\bar{E},{\mathcal{N}}_{E})\to(\bar{X},{\mathcal{N}}_{X})$ and $(\bar{X},{\mathcal{N}}_{X})\to(T,{\mathcal{O}}_{T}^{*})$ are smooth, so by [Kato1, 3.12], we have an exact sequence of coherent locally free ${\mathcal{O}}_{\bar{E}}$-modules
(2) $0\to\bar{s}^{*}\omega^{1}_{\bar{X}/T}\to\omega^{1}_{\bar{E}/T}\to\omega^{1}_{\bar{E}/\bar{X}}\to 0,$
where $\omega^{1}$ denotes the sheaf of logarithmic relative differentials defined in [Kato1, 1.7]. The sheaves $\omega^{1}_{\bar{X}/T}$ and $\omega^{1}_{\bar{E}/\bar{X}}$ are invertible, and can be identified, respectively, with $\Omega^{1}_{\bar{X}/T}(X^{\infty})$ and the sheaf of regular differentials for $\bar{s}$ (denoted $\omega_{\bar{E}/\bar{X}}$ in [D-R, I.2.1]).
Define ${\mathcal{F}}_{{\mathop{\rm dR}\nolimits}}$ as the locally free sheaf ${\bf R}^{1}\bar{s}_{*}\omega^{\bullet}_{\bar{E}/\bar{X}}$ of ${\mathcal{O}}_{\bar{X}}$-modules on $\bar{X}$, where $\omega^{\bullet}_{\bar{E}/\bar{X}}$ is the complex $d:{\mathcal{O}}_{\bar{E}}\to\omega^{1}_{\bar{E}/\bar{X}}$. This has a decreasing filtration with $\operatorname{{\mathop{\rm Fil}\nolimits}}^{2}{\mathcal{F}}_{{\mathop{\rm dR}\nolimits}}=0$, $\operatorname{{\mathop{\rm Fil}\nolimits}}^{1}{\mathcal{F}}_{{\mathop{\rm dR}\nolimits}}=\bar{s}_{*}\omega^{1}_{\bar{E}/\bar{X}}$, and $\operatorname{{\mathop{\rm Fil}\nolimits}}^{0}{\mathcal{F}}_{{\mathop{\rm dR}\nolimits}}={\mathcal{F}}_{{\mathop{\rm dR}\nolimits}}$. We denote $\operatorname{{\mathop{\rm Fil}\nolimits}}^{1}{\mathcal{F}}_{{\mathop{\rm dR}\nolimits}}$ simply as $\omega$. The sheaf $\omega$ is invertible and can be identified with $e^{*}\Omega^{1}_{\bar{E}/\bar{X}}$ where $e:\bar{X}\to\bar{E}$ is the zero section. We also have a canonical isomorphism ${\mathop{\rm gr}\nolimits}^{0}{\mathcal{F}}_{{\mathop{\rm dR}\nolimits}}\cong R^{1}\bar{s}_{*}{\mathcal{O}}_{\bar{E}}\cong\omega^{-1}$ given by Grothendieck-Serre duality. The fiber $x^{*}{\mathcal{F}}_{\mathop{\rm dR}\nolimits}$ of ${\mathcal{F}}_{\mathop{\rm dR}\nolimits}$ at a point $x:\operatorname{{\mathop{\rm Spec}\nolimits}}k\to\bar{X}$ can be identified with ${\bf H}^{1}(\bar{E}_{x},\omega^{\bullet}_{\bar{E}_{x}/k})$ where $\omega^{\bullet}_{\bar{E}_{x}/k}$ is the complex $d:{\mathcal{O}}_{\bar{E}_{x}}\to\omega^{1}_{\bar{E}_{x}/k}$ and $\omega^{1}_{\bar{E}_{x}/k}$ is the sheaf of regular differentials on $\bar{E}_{x}$. The fiber of $\operatorname{{\mathop{\rm Fil}\nolimits}}^{1}{\mathcal{F}}_{\mathop{\rm dR}\nolimits}$ is then identified with $H^{0}(\bar{E}_{x},\omega^{1}_{\bar{E}_{x}/k})\cong e_{x}^{*}\Omega^{1}_{\bar{E}_{x}/k}$ where $e_{x}:\operatorname{{\mathop{\rm Spec}\nolimits}}k\to\bar{E}_{x}$ is the zero section, and that of ${\mathop{\rm gr}\nolimits}^{0}{\mathcal{F}}_{\mathop{\rm dR}\nolimits}$ is identified with $H^{1}(\bar{E}_{x},{\mathcal{O}}_{\bar{E}_{x}})$.
We define ${\mathcal{F}}^{k}_{{\mathop{\rm dR}\nolimits}}$ as the filtered sheaf of ${\mathcal{O}}_{\bar{X}}$-modules $\operatorname{{\mathop{\rm Sym}\nolimits}}^{k-2}_{{\mathcal{O}}_{\bar{X}}}{\mathcal{F}}_{{\mathop{\rm dR}\nolimits}}$, and we let ${\mathcal{F}}_{c,{\mathop{\rm dR}\nolimits}}^{k}={\mathcal{F}}_{\mathop{\rm dR}\nolimits}^{k}(-X^{\infty})$.
The exact sequence (2) gives rise to an exact sequence of complexes
(3) $0\to\omega^{\bullet}_{\bar{E}/\bar{X}}\otimes_{{\mathcal{O}}_{\bar{E}}}\bar{s}^{*}\omega^{1}_{\bar{X}/T}\to\omega^{\bullet}_{\bar{E}/T}\to\omega^{\bullet}_{\bar{E}/\bar{X}}\to 0,$
where $\bar{s}^{*}\omega^{1}_{\bar{X}/T}$ has degree 1, and $\omega^{\bullet}_{\bar{E}/T}$ denotes the complex $\wedge^{\bullet}_{{\mathcal{O}}_{\bar{E}}}\omega^{1}_{\bar{E}/T}$ [Kato1, 1.9]. From the long exact sequence for ${\bf R}^{i}\bar{s}_{*}$ and the projection formula, we obtain the (logarithmic) Gauss-Manin connection
$\nabla:{\mathcal{F}}_{{\mathop{\rm dR}\nolimits}}\to{\mathcal{F}}_{{\mathop{\rm dR}\nolimits}}\otimes_{{\mathcal{O}}_{\bar{X}}}\omega^{1}_{\bar{X}/T}.$
This induces logarithmic connections on ${\mathcal{F}}_{{\mathop{\rm dR}\nolimits}}^{k}$ and ${\mathcal{F}}_{c,{\mathop{\rm dR}\nolimits}}^{k}$ satisfying Griffiths transversality. We set ${\mathcal{M}}_{{\mathop{\rm dR}\nolimits}}={\bf H}^{1}(\bar{X},\omega^{\bullet}({\mathcal{F}}_{{\mathop{\rm dR}\nolimits}}^{k}))$ and ${\mathcal{M}}_{c,{\mathop{\rm dR}\nolimits}}={\bf H}^{1}(\bar{X},\omega^{\bullet}({\mathcal{F}}_{c,{\mathop{\rm dR}\nolimits}}^{k}))$, where we write $\omega^{\bullet}({\mathcal{G}})$ for the complex associated to the module $\mathcal{G}$} with its connection. The filtrations on ${\mathcal{M}}_{{\mathop{\rm dR}\nolimits}}$ and ${\mathcal{M}}_{c,{\mathop{\rm dR}\nolimits}}$ are defined by those on ${\mathcal{F}}_{{\mathop{\rm dR}\nolimits}}^{k}$ and ${\mathcal{F}}_{c,{\mathop{\rm dR}\nolimits}}^{k}$. We let $M_{\mathop{\rm dR}\nolimits}={\mathbb{Q}}\otimes{\mathcal{M}}_{\mathop{\rm dR}\nolimits}$ and $M_{c,{\mathop{\rm dR}\nolimits}}={\mathbb{Q}}\otimes{\mathcal{M}}_{c,{\mathop{\rm dR}\nolimits}}$.
Recall that the composite
$\omega\to{\mathcal{F}}_{\mathop{\rm dR}\nolimits}\stackrel{{\scriptstyle\nabla}}{{\to}}{\mathcal{F}}_{{\mathop{\rm dR}\nolimits}}\otimes_{{\mathcal{O}}_{\bar{X}}}\omega^{1}_{\bar{X}/T}\to\omega^{-1}\otimes_{{\mathcal{O}}_{\bar{X}}}\omega^{1}_{\bar{X}/T}$
is an isomorphism (the Kodaira-Spencer isomorphism, [D-R, VI.4.5.2]). Hence $\omega^{2}\cong\omega^{1}_{\bar{X}/T}$, and one deduces that
${\mathop{\rm gr}\nolimits}^{i}{\mathcal{M}}_{{\mathop{\rm dR}\nolimits}}\cong\left\{\begin{array}[]{ll}H^{0}(\bar{X},\omega^{k-2}\otimes\omega^{1}_{\bar{X}/T}),&\mbox{if $i=k-1$;}\\ H^{1}(\bar{X},\omega^{2-k}),&\mbox{if $i=0$;}\\ 0,&\mbox{otherwise}.\end{array}\right$.
Similarly one finds
${\mathop{\rm gr}\nolimits}^{i}{\mathcal{M}}_{c,{\mathop{\rm dR}\nolimits}}\cong\left\{\begin{array}[]{ll}H^{0}(\bar{X},\omega^{k-2}\otimes\Omega^{1}_{\bar{X}/T}),&\mbox{if $i=k-1$;}\\ H^{1}(\bar{X},\omega^{2-k}(-X^{\infty})),&\mbox{if $i=0$;}\\ 0,&\mbox{otherwise}.\end{array}\right$.
In particular, it follows that ${\mathcal{M}}_{{\mathop{\rm dR}\nolimits}}$ and ${\mathcal{M}}_{c,{\mathop{\rm dR}\nolimits}}$ are torsion-free.
For lack of a reference, we now sketch the construction of the comparison isomorphism $I^{\infty}$.
For $\bar{Y}=\bar{X}$ or $\bar{E}$, let $h_{\bar{Y}}$ denote the canonical morphism $\bar{Y}^{\mathop{\rm an}\nolimits}\to\bar{Y}$ of ringed spaces, let $\omega^{1}_{\bar{Y}^{\mathop{\rm an}\nolimits}}$ denote $h_{\bar{Y}}^{*}\omega^{1}_{\bar{Y}/S}$, and let $\omega^{\bullet}_{\bar{Y}^{\mathop{\rm an}\nolimits}}$ denote the complex $\wedge^{\bullet}_{{\mathcal{O}}_{\bar{Y}^{\mathop{\rm an}\nolimits}}}\omega^{1}_{\bar{Y}^{\mathop{\rm an}\nolimits}}$. We also consider the complex
$\omega^{\bullet}_{\bar{E}^{\mathop{\rm an}\nolimits}/\bar{X}^{\mathop{\rm an}\nolimits}}=\wedge^{\bullet}_{{\mathcal{O}}_{\bar{E}^{\mathop{\rm an}\nolimits}}}\omega^{1}_{\bar{E}^{\mathop{\rm an}\nolimits}/\bar{X}^{\mathop{\rm an}\nolimits}},$
where $\omega^{1}_{\bar{E}^{\mathop{\rm an}\nolimits}/\bar{X}^{\mathop{\rm an}\nolimits}}$ denotes $h_{\bar{E}}^{*}\omega^{1}_{\bar{E}/\bar{X}}$.
We let
${\mathcal{F}}_{\mathop{\rm dR}\nolimits}^{\mathop{\rm an}\nolimits}={\bf R}^{1}\bar{s}^{\mathop{\rm an}\nolimits}_{*}\omega^{\bullet}_{\bar{E}^{\mathop{\rm an}\nolimits}/\bar{X}^{\mathop{\rm an}\nolimits}},$
${\mathcal{F}}_{\mathop{\rm dR}\nolimits}^{{\mathop{\rm an}\nolimits},k}=\operatorname{{\mathop{\rm Sym}\nolimits}}^{k-2}_{{\mathcal{O}}_{\bar{X}^{\mathop{\rm an}\nolimits}}}{\mathcal{F}}_{\mathop{\rm dR}\nolimits}^{\mathop{\rm an}\nolimits}$ and ${\mathcal{F}}_{c,{\mathop{\rm dR}\nolimits}}^{{\mathop{\rm an}\nolimits},k}={\mathcal{F}}_{\mathop{\rm dR}\nolimits}^{{\mathop{\rm an}\nolimits},k}(-X^{\infty,{\mathop{\rm an}\nolimits}})$. These are locally free sheaves of ${\mathcal{O}}_{\bar{X}^{\mathop{\rm an}\nolimits}}$-modules, and the fiber ${\mathbb{C}}\otimes_{{\mathcal{O}}_{\bar{X},x}^{\mathop{\rm an}\nolimits}}{\mathcal{F}}_{{\mathop{\rm dR}\nolimits},x}^{\mathop{\rm an}\nolimits}$ of ${\mathcal{F}}_{\mathop{\rm dR}\nolimits}^{\mathop{\rm an}\nolimits}$ at a point $x\in X^{\mathop{\rm an}\nolimits}$ can be identified with the analytic de Rham cohomology ${\bf H}^{1}(E_{x}^{\mathop{\rm an}\nolimits},\Omega^{\bullet}_{E_{x}^{\mathop{\rm an}\nolimits}})$ of $E_{x}^{\mathop{\rm an}\nolimits}$. By (2) we have an exact sequence
$0\to\bar{s}^{{\mathop{\rm an}\nolimits},*}\omega^{1}_{\bar{X}^{\mathop{\rm an}\nolimits}}\to\omega^{1}_{\bar{E}^{\mathop{\rm an}\nolimits}}\to\omega^{1}_{\bar{E}^{\mathop{\rm an}\nolimits}/\bar{X}^{\mathop{\rm an}\nolimits}}\to 0,$
and a construction like the one in the de Rham realization gives logarithmic connections
${\mathcal{G}}\to{\mathcal{G}}\otimes_{{\mathcal{O}}_{\bar{X}^{\mathop{\rm an}\nolimits}}}\omega^{1}_{\bar{X}^{\mathop{\rm an}\nolimits}}$
and complexes $\omega^{\bullet}({\mathcal{G}})$ for ${\mathcal{G}}={\mathcal{F}}_{\mathop{\rm dR}\nolimits}^{{\mathop{\rm an}\nolimits},k}$ and ${\mathcal{F}}_{c,{\mathop{\rm dR}\nolimits}}^{{\mathop{\rm an}\nolimits},k}$. We let $M_{\mathop{\rm dR}\nolimits}^{\mathop{\rm an}\nolimits}={\bf H}^{1}(\bar{X}^{\mathop{\rm an}\nolimits},\omega^{\bullet}({\mathcal{F}}_{\mathop{\rm dR}\nolimits}^{{\mathop{\rm an}\nolimits},k}))$ and $M_{c,{\mathop{\rm dR}\nolimits}}^{\mathop{\rm an}\nolimits}={\bf H}^{1}(\bar{X}^{\mathop{\rm an}\nolimits},\omega^{\bullet}({\mathcal{F}}_{{\mathop{\rm dR}\nolimits},c}^{{\mathop{\rm an}\nolimits},k}))$.
Now we compare $M_{\mathop{\rm dR}\nolimits}^{\mathop{\rm an}\nolimits}$ with ${\mathbb{C}}\otimes M_{\mathop{\rm dR}\nolimits}$, and $M_{c,{\mathop{\rm dR}\nolimits}}^{\mathop{\rm an}\nolimits}$ with ${\mathbb{C}}\otimes M_{c,{\mathop{\rm dR}\nolimits}}$. The natural map
$h_{\bar{E}}^{-1}\omega^{\bullet}_{\bar{E}/\bar{X}}\to\omega^{\bullet}_{\bar{E}^{\mathop{\rm an}\nolimits}/\bar{X}^{\mathop{\rm an}\nolimits}}$
is compatible with differentiation, and so gives rise to an $h_{\bar{X}}^{-1}{\mathcal{O}}_{\bar{X}}$-linear map
$h_{\bar{X}}^{-1}{\mathcal{F}}_{\mathop{\rm dR}\nolimits}\to{\bf R}^{1}\bar{s}^{\mathop{\rm an}\nolimits}_{*}h_{\bar{E}}^{-1}\omega^{\bullet}_{\bar{E}/\bar{X}}\to{\mathcal{F}}_{\mathop{\rm dR}\nolimits}^{\mathop{\rm an}\nolimits}.$
This map is compatible with connections and induces an isomorphism
$h_{\bar{X}}^{*}{\mathcal{F}}_{\mathop{\rm dR}\nolimits}\to{\mathcal{F}}_{\mathop{\rm dR}\nolimits}^{\mathop{\rm an}\nolimits}$
which on the fiber at $x\in X({\mathbb{C}})=X^{\mathop{\rm an}\nolimits}$ becomes the isomorphism between the algebraic de Rham cohomology of $E_{x}$ and the analytic de Rham cohomology of $E_{x}^{\mathop{\rm an}\nolimits}$ provided by GAGA [Se]. From this one obtains $h_{\bar{X}}^{-1}{\mathcal{O}}_{\bar{X}}$-linear maps
$h_{\bar{X}}^{-1}{\mathcal{F}}_{\mathop{\rm dR}\nolimits}^{k}\to{\mathcal{F}}_{\mathop{\rm dR}\nolimits}^{k,{\mathop{\rm an}\nolimits}},\quad\mbox{and}\quad h_{\bar{X}}^{-1}{\mathcal{F}}_{c,{\mathop{\rm dR}\nolimits}}^{k}\to{\mathcal{F}}_{c,{\mathop{\rm dR}\nolimits}}^{k,{\mathop{\rm an}\nolimits}}$
compatible with connections and inducing isomorphisms on replacing $h_{\bar{X}}^{-1}$ with $h_{\bar{X}}^{*}$. We thus obtain maps
$\begin{array}[]{lclcl}M_{\mathop{\rm dR}\nolimits}&\to&{\bf H}^{1}(\bar{X}^{\mathop{\rm an}\nolimits},h_{\bar{X}}^{-1}\omega^{\bullet}({\mathcal{F}}_{\mathop{\rm dR}\nolimits}^{k}))&\to&M_{\mathop{\rm dR}\nolimits}^{\mathop{\rm an}\nolimits}\quad\mbox{and}\\ M_{c,{\mathop{\rm dR}\nolimits}}&\to&{\bf H}^{1}(\bar{X}^{\mathop{\rm an}\nolimits},h_{\bar{X}}^{-1}\omega^{\bullet}({\mathcal{F}}_{c,{\mathop{\rm dR}\nolimits}}^{k}))&\to&M_{c,{\mathop{\rm dR}\nolimits}}^{\mathop{\rm an}\nolimits}.\end{array}$
Applying GAGA [Se] again, we find that these induce isomorphisms
${\mathbb{C}}\otimes M_{\mathop{\rm dR}\nolimits}\cong M_{\mathop{\rm dR}\nolimits}^{\mathop{\rm an}\nolimits}\quad\mbox{and}\quad{\mathbb{C}}\otimes M_{c,{\mathop{\rm dR}\nolimits}}\cong M_{c,{\mathop{\rm dR}\nolimits}}^{\mathop{\rm an}\nolimits}.$
Now we compare $M_{\mathop{\rm dR}\nolimits}^{\mathop{\rm an}\nolimits}$ with ${\mathbb{C}}\otimes M_{B}$ and $M_{c,{\mathop{\rm dR}\nolimits}}^{\mathop{\rm an}\nolimits}$ with ${\mathbb{C}}\otimes M_{c,{\mathop{\rm dR}\nolimits}}$. By the Poincaré Lemma, we have the resolution
${\mathbb{C}}\to\Omega^{\bullet}_{E^{\mathop{\rm an}\nolimits}}$
of the constant sheaf $\mathbb{C}$} on $E^{\mathop{\rm an}\nolimits}$. We thus obtain an isomorphism
${\mathbb{C}}\otimes{\mathcal{F}}_{B}\cong R^{1}s^{\mathop{\rm an}\nolimits}_{*}{\mathbb{C}}\cong{\bf R}^{1}s^{\mathop{\rm an}\nolimits}_{*}\Omega^{\bullet}_{E^{\mathop{\rm an}\nolimits}}.$
Applying ${\bf R}^{\bullet}s^{\mathop{\rm an}\nolimits}_{*}$ to the exact sequence
$0\to\Omega^{\bullet}_{E^{\mathop{\rm an}\nolimits}/X^{\mathop{\rm an}\nolimits}}\otimes_{{\mathcal{O}}_{E^{\mathop{\rm an}\nolimits}}}s^{{\mathop{\rm an}\nolimits},*}\Omega^{1}_{X^{\mathop{\rm an}\nolimits}}\to\Omega^{\bullet}_{E^{\mathop{\rm an}\nolimits}}\to\Omega^{\bullet}_{E^{\mathop{\rm an}\nolimits}/X^{\mathop{\rm an}\nolimits}}\to 0$
we obtain an exact sequence
${\mathcal{O}}_{X^{\mathop{\rm an}\nolimits}}\to\Omega^{1}_{X^{\mathop{\rm an}\nolimits}}\to{\bf R}^{1}s^{\mathop{\rm an}\nolimits}_{*}\Omega^{\bullet}_{E^{\mathop{\rm an}\nolimits}}\to{\mathcal{G}}\to{\mathcal{G}}\otimes_{{\mathcal{O}}_{X^{\mathop{\rm an}\nolimits}}}\Omega^{1}_{X^{\mathop{\rm an}\nolimits}},$
where $\mathcal{G}$} is the restriction of ${\mathcal{F}}_{\mathop{\rm dR}\nolimits}^{\mathop{\rm an}\nolimits}$ to $X^{\mathop{\rm an}\nolimits}$. The first map (differentiation) is surjective, as is the last map (the Gauss-Manin connection) by the local solvability of linear systems of differential equations. We thus obtain an exact sequence
(4) $0\to{\mathbb{C}}\otimes{\mathcal{F}}_{B}\to{\mathcal{G}}\to{\mathcal{G}}\otimes_{{\mathcal{O}}_{X^{\mathop{\rm an}\nolimits}}}\Omega^{1}_{X^{\mathop{\rm an}\nolimits}}\to 0.$
Composing the first map on the stalk at $x\in X^{\mathop{\rm an}\nolimits}$ with the map arising from the evaluation homomorphism ${\mathcal{O}}_{X^{\mathop{\rm an}\nolimits},x}\to{\mathbb{C}}$, we obtain the isomorphism $H^{1}(E_{x}^{\mathop{\rm an}\nolimits},{\mathbb{C}})\cong{\bf H}^{1}(E_{x}^{\mathop{\rm an}\nolimits},\Omega^{\bullet}_{E_{x}^{\mathop{\rm an}\nolimits}})$ given by the Poincaré lemma for $E_{x}^{\mathop{\rm an}\nolimits}$.
Writing ${\mathcal{G}}^{k}$ for the restriction of ${\mathcal{F}}_{\mathop{\rm dR}\nolimits}^{k,{\mathop{\rm an}\nolimits}}$ to $X^{\mathop{\rm an}\nolimits}$, we get a sequence
(5) $0\to{\mathbb{C}}\otimes{\mathcal{F}}_{B}^{k}\to{\mathcal{G}}^{k}\to{\mathcal{G}}^{k}\otimes_{{\mathcal{O}}_{X^{\mathop{\rm an}\nolimits}}}\Omega^{1}_{X^{\mathop{\rm an}\nolimits}}\to 0$
whose exactness follows from the local existence and uniqueness of solutions with initial conditions. Extending the connection on ${\mathcal{G}}^{k}$ to our logarithmic connection on ${\mathcal{F}}_{c,{\mathop{\rm dR}\nolimits}}^{k,{\mathop{\rm an}\nolimits}}$ and letting $j$ denote the inclusion of $X^{\mathop{\rm an}\nolimits}$ in $\bar{X}^{\mathop{\rm an}\nolimits}$, we claim that
$0\to{\mathbb{C}}\otimes j_{!}{\mathcal{F}}_{B}^{k}\to{\mathcal{F}}_{c,{\mathop{\rm dR}\nolimits}}^{k,{\mathop{\rm an}\nolimits}}\to{\mathcal{F}}_{c,{\mathop{\rm dR}\nolimits}}^{k,{\mathop{\rm an}\nolimits}}\otimes_{{\mathcal{O}}_{\bar{X}^{\mathop{\rm an}\nolimits}}}\omega^{1}_{\bar{X}^{\mathop{\rm an}\nolimits}}\to 0$
is exact. This amounts to the following:
Lemma 2.1.
The connection
${\mathcal{F}}_{c,{\mathop{\rm dR}\nolimits}}^{k,{\mathop{\rm an}\nolimits}}\to{\mathcal{F}}_{c,{\mathop{\rm dR}\nolimits}}^{k,{\mathop{\rm an}\nolimits}}\otimes_{{\mathcal{O}}_{\bar{X}^{\mathop{\rm an}\nolimits}}}\omega^{1}_{\bar{X}^{\mathop{\rm an}\nolimits}}$
defines an isomorphism on stalks at the cusps.
Proof.
Over a neighborhood of each cusp in $\bar{X}^{\mathop{\rm an}\nolimits}$, $\bar{E}^{\mathop{\rm an}\nolimits}$ is isomorphic to the NN-sided analytic Tate curve $E_{q}^{\mathop{\rm an}\nolimits}$ over $D_{1}$. We claim that there is an isomorphism between the pull-back of ${\mathcal{F}}_{{\mathop{\rm dR}\nolimits}}^{k,{\mathop{\rm an}\nolimits}}$ to $D_{1}$ and ${\mathcal{O}}_{D_{1}}^{k-1}$ such that the corresponding connection on ${\mathcal{O}}_{D_{1}}^{k-1}$ is given by $\nabla(f)=(A(t)f+tf^{\prime})\dfrac{dt}{t}$ where $A(t)\in M_{k-1}({\mathcal{O}}_{D_{1}})$ is such that $A(0)$ is nilpotent. Indeed this follows from the case $k=3$ which follows from the calculation in Appendix 1.3 of [Katz2] for the 1-sided Tate curve.
Now observe that the desired bijectivity follows from the local existence of a unique solution of the differential equation $(A(t)+I)f+tf^{\prime}=g$ for any $g\in{\mathcal{O}}_{D_{1}}^{k-1}$, which is guaranteed by the fact that $A(0)+I$ has no nonpositive integer eigenvalues. ∎
We can now conclude that ${\mathbb{C}}\otimes M_{c,B}\cong H^{1}(\bar{X}^{\mathop{\rm an}\nolimits},{\mathbb{C}}\otimes j_{!}{\mathcal{F}}_{B}^{k})$ is isomorphic to $M_{c,{\mathop{\rm dR}\nolimits}}^{\mathop{\rm an}\nolimits}\cong{\mathbb{C}}\otimes M_{c,{\mathop{\rm dR}\nolimits}}$.
To define $I^{\infty}$ for $M_{B}$, note that we have a restriction map
${\mathbb{C}}\otimes M_{\mathop{\rm dR}\nolimits}\cong M_{\mathop{\rm dR}\nolimits}^{\mathop{\rm an}\nolimits}\to{\bf H}^{1}(X^{\mathop{\rm an}\nolimits},\omega^{\bullet}({\mathcal{G}}^{k}))\cong{\mathbb{C}}\otimes M_{B}.$
We postpone until §2.4 the proof that it is an isomorphism. It is straightforward to check that the maps respect the action of $G_{\mathbb{R}}$.
Before defining the crystalline realization and comparison isomorphisms, we review Faltings’ generalization of the theory of Fontaine and Laffaille (see [Fa]).
Suppose that $\bar{Y}$ is a smooth, proper scheme over $\operatorname{{\mathop{\rm Spec}\nolimits}}{\mathbb{Z}}_{\ell}$ with a relative divisor $D$ with strict normal crossings, and let $Y=\bar{Y}-D$. For each integer $a$ with $0\leq a\leq\ell-2$, Faltings defines a category ${\mathcal{M}\mathcal{F}}_{[0,a]}^{\nabla}(Y)$ and a fully faithful contravariant functor $\mathbb{D}$} to the category of finite locally constant étale sheaves on $Y_{{\mathbb{Q}}_{\ell}}$ ([Fa, Thm. $2.6^{*}$]). We write $\mathbb{V}$} for the covariant functor defined by
${\mathbb{V}}({\mathcal{A}})={\mathbb{D}}({\mathcal{A}})^{*}={\mathop{\rm Hom}\nolimits}({\mathbb{D}}({\mathcal{A}}),{\mathbb{Q}}_{\ell}/{\mathbb{Z}}_{\ell}).$
In the case of $Y=\bar{Y}=\operatorname{{\mathop{\rm Spec}\nolimits}}{\mathcal{O}}_{F}$ where $F$ is an unramified extension of $\mathbb{Q}$$_{\ell}$, Faltings’ category can be identified with the full subcategory of ${{\mathcal{M}\mathcal{F}}}^{0}_{\mathop{\rm tor}\nolimits}({\mathcal{O}}_{F})$ whose objects AA satisfy $\operatorname{{\mathop{\rm Fil}\nolimits}}^{a+1}A=0$. Furthermore, identifying the category of étale sheaves on $\operatorname{{\mathop{\rm Spec}\nolimits}}F$ with that of continuous $G_{F}$-modules, one finds that the functor $\mathbb{D}$} agrees with $\underline{U}_{S}$ (see [Fo3, 3.8]).
We let ${\mathcal{O}}_{Y,{\mbox{\scriptsize crys}}}$ denote the inverse system in ${\mathcal{M}\mathcal{F}}_{[0,0]}^{\nabla}(Y)$ (indexed by integers $n\geq 1$) defined by reduction mod $\ell^{n}$ of the following data:
  • the structure sheaf ${\mathcal{O}}_{\bar{Y}}$;
  • the logarithmic connection $d:{\mathcal{O}}_{\bar{Y}}\to\omega^{1}_{\bar{Y}/{\mathbb{Z}}_{\ell}}$;
  • an affine covering $\{\operatorname{{\mathop{\rm Spec}\nolimits}}A_{\alpha}\}$ of $\bar{Y}$ with arithmetic Frobenius-lifts $\phi_{\alpha}$ of $A_{\alpha}$ as in [Fa, §2i].
Then ${\mathbb{V}}({\mathcal{O}}_{Y,{\mbox{\scriptsize crys}}})$ is canonically identified with the constant $\ell$-adic sheaf $\mathbb{Z}$$_{\ell}$ on $Y$.
Assuming $\ell$ does not divide $2N$ and applying [Fa, Thm. 6.2] to the morphism $\bar{s}$, we obtain a canonical isomorphism
${\mathbb{V}}({\mathcal{F}}_{\ell\mbox{\scriptsize-crys}})\cong{\mathcal{F}}_{\ell},$
where -${\mathcal{F}}_{\ell\mbox{\scriptsize-crys}}$ is the inverse system in ${\mathcal{M}\mathcal{F}}^{\nabla}_{[0,1]}(X_{{\mathbb{Z}}_{\ell}})$ defined by reduction mod $\ell^{n}$ of ${\mathcal{F}}_{\mathop{\rm dR}\nolimits}$ with its filtration, logarithmic Gauss-Manin connection and locally defined Frobenius maps whose precise description we shall not need. For $x:\operatorname{{\mathop{\rm Spec}\nolimits}}{\mathcal{O}}_{F}\to X$ with $F$ an unramified extension of $\mathbb{Q}$$_{\ell}$, we can identify $x^{*}{\mathcal{F}}_{\ell\mbox{\scriptsize-crys}}$ with the object $H^{1}_{\mbox{\scriptsize crys}}(E_{x},{\mathcal{O}}_{E_{x},{\mbox{\scriptsize crys}}})$ of ${\mathcal{M}\mathcal{F}}_{[0,1]}({\mathcal{O}}_{F})$. Furthermore ${\mathbb{V}}(x^{*}{\mathcal{F}}_{\ell\mbox{\scriptsize-crys}})$ is canonically identified with the stalk of ${\mathbb{V}}({\mathcal{F}}_{\ell\mbox{\scriptsize-crys}})$ at $\bar{x}:\bar{F}\to X$ and the isomorphism ${\mathbb{V}}({\mathcal{F}}_{\ell\mbox{\scriptsize-crys}})\cong{\mathcal{F}}_{\ell}$ on the stalk is the one between $H^{1}_{\mbox{\scriptsize crys}}(E_{x},{\mathcal{O}}_{E_{x},{\ell\mbox{\scriptsize-crys}}})$ and $H^{1}(E_{\bar{x}},{\mathbb{Z}}_{\ell})$ provided by [Fa, Thm. 5.3].
Using the results in [Fa, IIh)], one finds that the functor $\mathbb{V}$} respects tensor products. So assuming further that $\ell>k-1$, we obtain a canonical isomorphism
${\mathbb{V}}({\mathcal{F}}_{\ell\mbox{\scriptsize-crys}}^{k})\cong{\mathcal{F}}_{\ell}^{k},$
where -${\mathcal{F}}_{\ell\mbox{\scriptsize-crys}}^{k}$ is the inverse system in ${\mathcal{M}\mathcal{F}}^{\nabla}_{[0,k-2]}(X_{{\mathbb{Z}}_{\ell}})$ defined by $\operatorname{{\mathop{\rm Sym}\nolimits}}^{k-2}_{{\mathcal{O}}_{\bar{Y}}}{\mathcal{F}}_{\ell\mbox{\scriptsize-crys}}$. If $\ell>k$, we obtain an object
-${\mathcal{M}}_{\ell\mbox{\scriptsize-crys}}=H^{1}_{\mbox{\scriptsize crys}}(X_{{\mathbb{Z}}_{\ell}},{\mathcal{F}}_{\ell\mbox{\scriptsize-crys}}^{k})$
of ${{\mathcal{M}\mathcal{F}}}^{0}$ whose underlying filtered module is isomorphic to ${\mathbb{Z}}_{\ell}\otimes{\mathcal{M}}_{\mathop{\rm dR}\nolimits}$, and so we have the comparison $I^{\ell}_{\mathop{\rm dR}\nolimits}$. Furthermore we conclude from [Fa, Thm. 5.3] that there is a canonical isomorphism
$I^{\ell}:{\mathbb{V}}({\mathcal{M}}_{\ell\mbox{\scriptsize-crys}})\cong{\mathcal{M}}_{\ell}.$
For such primes, we similarly obtain an isomorphism
$I^{\ell}:{\mathbb{V}}({\mathcal{M}}_{c,{\ell\mbox{\scriptsize-crys}}})\cong{\mathcal{M}}_{c,\ell}$
where ${\mathcal{M}}_{c,{\ell\mbox{\scriptsize-crys}}}$ has underlying filtered module ${\mathcal{M}}_{c,{\mathop{\rm dR}\nolimits}}$.
2.3. Level NN modular forms
We recall the relation between $M_{{\mathop{\rm dR}\nolimits}}$ and the space of modular forms of weight $k$ with respect to $\Gamma(N)$.
Let $\mathcal{L}$} denote the line bundle $\omega^{k-2}\otimes\omega^{1}_{\bar{X}/S}$ on $\bar{X}$. Then we have
${\mathbb{C}}\otimes\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}{\mathcal{M}}_{{\mathop{\rm dR}\nolimits}}\cong H^{0}(\bar{X}^{\mathop{\rm an}\nolimits},{\mathcal{L}}^{\mathop{\rm an}\nolimits})\cong\bigoplus_{t\in({\mathbb{Z}}/N{\mathbb{Z}})^{\times}}H^{0}(\bar{X}_{N,t},i_{t}^{*}{\mathcal{L}}^{\mathop{\rm an}\nolimits}),$
where
${\mathcal{L}}^{\mathop{\rm an}\nolimits}=\left(e^{{\mathop{\rm an}\nolimits},*}\Omega^{1}_{\bar{E}^{\mathop{\rm an}\nolimits}/\bar{X}^{\mathop{\rm an}\nolimits}}\right)^{k-2}\otimes_{{\mathcal{O}}_{\bar{X}^{\mathop{\rm an}\nolimits}}}\Omega^{1}_{\bar{X}^{\mathop{\rm an}\nolimits}}(-X^{\infty,{\mathop{\rm an}\nolimits}})$
and $i_{t}$ is the inclusion $\bar{X}_{N,t}\to\bar{X}$ defined in §2.1. For each $t$, we have the natural projection $\pi_{t}:{\mathfrak{H}}\to\bar{X}_{N,t}$ giving an inclusion
$H^{0}(\bar{X}_{N,t},i_{t}^{*}{\mathcal{L}}^{\mathop{\rm an}\nolimits})\to H^{0}({\mathfrak{H}},\pi_{t}^{*}i_{t}^{*}{\mathcal{L}}^{\mathop{\rm an}\nolimits})$
whose image is precisely the space of sections of the form
$f(\tau)(2\pi i)^{k-1}(dz)^{\otimes(k-2)}\otimes d\tau$
with $f(\tau)$ in $M_{k}(\Gamma(N))$, the space of modular forms of weight $k$ with respect to $\Gamma(N)$. We thus obtain an isomorphism
${\mathbb{C}}\otimes\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}{\mathcal{M}}_{{\mathop{\rm dR}\nolimits}}\cong\bigoplus_{t\in({\mathbb{Z}}/N{\mathbb{Z}})^{\times}}M_{k}(\Gamma(N)).$
Note that we have chosen our normalization so that if $f(\tau)=g(e^{2\pi i\tau/N})$ is in $M_{k}(\Gamma(N))$, then the restriction of the corresponding element of $H^{0}(\bar{X}_{N,t},i_{t}^{*}{\mathcal{L}}^{\mathop{\rm an}\nolimits})$ to $H^{0}(D_{1},j_{t}^{*}i_{t}^{*}{\mathcal{L}}^{\mathop{\rm an}\nolimits})$ is
$g(q^{1/N})\omega_{\mathop{\rm can}\nolimits}^{\otimes(k-2)}\otimes\dfrac{dq}{q},$
where $j_{t}:D_{1}\to\bar{X}_{N,t}$ is the inclusion of a disk around the cusp $\infty$ used in §2.1.
By the $q$-expansion principle [D-R, VII.3.9,VII.3.13], the map
${\mathbb{C}}\otimes\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}{\mathcal{M}}_{\mathop{\rm dR}\nolimits}\cong\bigoplus_{t\in({\mathbb{Z}}/N{\mathbb{Z}})^{\times}}M_{k}(\Gamma(N))\to\bigoplus_{t\in({\mathbb{Z}}/N{\mathbb{Z}})^{\times}}{\mathbb{C}}[[q^{1/N}]],$
where the second map is given by sending $f(\tau)$ to $g(q^{1/N})$, gives rise to a pullback diagram
[…]
where $R={\mathbb{Z}}[1/Nk!,\mu_{N}]$. Here the bottom right isomorphism is induced by the isomorphism $R\otimes{\mathbb{C}}\to\oplus_{t}{\mathbb{C}}$ defined by $(e^{2\pi it/N})_{t}$ and commutativity of the right square follows from [D-R, VII.4.7].
Similarly one obtains an isomorphism
${\mathbb{C}}\otimes\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}{\mathcal{M}}_{c,{\mathop{\rm dR}\nolimits}}\cong\bigoplus_{t\in({\mathbb{Z}}/N{\mathbb{Z}})^{\times}}S_{k}(\Gamma(N)),$
where $S_{k}(\Gamma(N))$ is the space of cusp forms of weight $k$ with respect to $\Gamma(N)$. The $q$-expansion principle identifies $\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}{\mathcal{M}}_{c,{\mathop{\rm dR}\nolimits}}$ as the subset of $\bigoplus_{t\in({\mathbb{Z}}/N{\mathbb{Z}})^{\times}}S_{k}(\Gamma(N))$ whose $q$-expansion at $\infty$ has coefficients in $R$.
2.4. Pairings
We shall use Poincaré duality to construct perfect pairings $M_{c,?}\otimes M_{?}\to{\mathbb{Q}}(1-k)_{?}$ for ?=B?=B, $\mathop{\rm dR}\nolimits$} and $\ell\nmid Nk!$, compatible with the comparison isomorphisms. These pairings arise from a duality morphism ${\mathcal{M}}_{c}\to{\mathop{\rm Hom}\nolimits}({\mathcal{M}},{\mathcal{T}}^{\otimes(k-1)})$.
We begin with the pairing on Betti realizations. The cup product defines a morphism
${\mathcal{F}}_{B}\otimes{\mathcal{F}}_{B}\to R^{2}s^{\mathop{\rm an}\nolimits}_{*}{\mathbb{Z}}\cong(2\pi i)^{-1}{\mathbb{Z}}$
inducing an isomorphism
${\mathcal{F}}_{B}\to{\mathop{{\mathcal{H}}om}\nolimits}({\mathcal{F}}_{B},(2\pi i)^{-1}{\mathbb{Z}})$
of sheaves on $X^{\mathop{\rm an}\nolimits}$. We then consider the morphism
${\mathcal{F}}_{B}^{k}\otimes{\mathcal{F}}_{B}^{k}\to(2\pi i)^{2-k}{\mathbb{Z}}$
defined on sections by
(7) $x_{1}\otimes\cdots\otimes x_{k-2}\otimes y_{1}\otimes\cdots\otimes y_{k-2}\mapsto\sum_{\sigma\in\Sigma_{k-2}}\prod_{i=1}^{k-2}x_{i}\cup y_{\sigma(i)},$
where $\Sigma_{k-2}$ is the symmetric group on $\{1,\ldots,k-2\}$. The resulting map on cohomology, preceded by the cup product and followed by the trace, yields a pairing
${\mathcal{M}}_{c,B}\otimes{\mathcal{M}}_{B}\to H^{2}_{c}(X^{\mathop{\rm an}\nolimits},{\mathcal{F}}_{B}^{k}\otimes{\mathcal{F}}_{B}^{k})\to H^{2}_{c}(X^{\mathop{\rm an}\nolimits},(2\pi i)^{2-k}{\mathbb{Z}})\to(2\pi i)^{1-k}{\mathbb{Z}}={\mathcal{T}}_{B}^{\otimes(k-1)},$
respecting the action of $G_{\mathbb{R}}$. After tensoring with ${\mathbb{Z}}[1/(k-2)!]$, we have that ${\mathcal{M}}_{c,B}$ is torsion-free and
${\mathcal{F}}_{B}^{k}\to{\mathop{{\mathcal{H}}om}\nolimits}({\mathcal{F}}_{B}^{k},(2\pi i)^{2-k}{\mathbb{Z}})$
becomes an isomorphism, and it follows that
$\delta^{L}_{B}:{\mathcal{M}}_{c,B}\to{\mathop{\rm Hom}\nolimits}({\mathcal{M}}_{B},{\mathcal{T}}_{B}^{\otimes(k-1)})$
also becomes an isomorphism.
The pairing on $\ell$-adic realizations is constructed similarly. In particular we have morphisms
${\mathcal{F}}_{\ell}^{k}\to{\mathop{{\mathcal{H}}om}\nolimits}({\mathcal{F}}_{\ell}^{k},{\mathbb{Z}}_{\ell}(2-k))\quad\mbox{and}\quad\delta^{L}_{\ell}:{\mathcal{M}}_{c,\ell}\to{\mathop{\rm Hom}\nolimits}({\mathcal{M}}_{\ell},{\mathcal{T}}_{\ell}^{\otimes(k-1)})$
which are isomorphisms if $\ell>k-2$, or for any $\ell$ after tensoring with $\mathbb{Q}$$_{\ell}$.
To define the pairing on de Rham realizations, we begin with the cup product
${\bf R}^{1}\bar{s}_{*}\omega^{\bullet}_{\bar{E}/\bar{X}}\otimes_{{\mathcal{O}}_{\bar{X}}}{\bf R}^{1}\bar{s}_{*}\omega^{\bullet}_{\bar{E}/\bar{X}}\to{\bf R}^{2}\bar{s}_{*}\left(\omega^{\bullet}_{\bar{E}/\bar{X}}\otimes_{\bar{s}^{-1}{\mathcal{O}}_{\bar{X}}}\omega^{\bullet}_{\bar{E}/\bar{X}}\right).$
Combining this with the morphism induced by the wedge product on $\omega^{\bullet}_{\bar{E}/\bar{X}}$, we obtain a morphism
${\mathcal{F}}_{\mathop{\rm dR}\nolimits}\otimes_{{\mathcal{O}}_{\bar{X}}}{\mathcal{F}}_{\mathop{\rm dR}\nolimits}\to{\bf R}^{2}\bar{s}_{*}\omega^{\bullet}_{\bar{E}/\bar{X}}\cong R^{1}\bar{s}_{*}\omega^{\mathop{\rm reg}\nolimits}_{\bar{E}/\bar{X}}[-1]\cong{\mathcal{O}}_{\bar{X}}[-1]$
respecting filtrations and connections. On graded pieces, the morphism
(8) ${\mathcal{F}}_{\mathop{\rm dR}\nolimits}\to{\mathop{{\mathcal{H}}om}\nolimits}_{{\mathcal{O}}_{\bar{X}}}({\mathcal{F}}_{\mathop{\rm dR}\nolimits},{\mathcal{O}}_{\bar{X}}[-1])$
coincides with the isomorphism given by Grothendieck-Serre duality. Using the same formula on sections as for the Betti realization, we obtain a pairing
${\mathcal{F}}_{\mathop{\rm dR}\nolimits}^{k}\otimes_{{\mathcal{O}}_{\bar{X}}}{\mathcal{F}}_{\mathop{\rm dR}\nolimits}^{k}\to{\mathcal{O}}_{\bar{X}}[2-k]$
respecting filtrations and connections and inducing an isomorphism
${\mathcal{F}}_{\mathop{\rm dR}\nolimits}^{k}\to{\mathop{{\mathcal{H}}om}\nolimits}_{{\mathcal{O}}_{\bar{X}}}({\mathcal{F}}_{\mathop{\rm dR}\nolimits}^{k},{\mathcal{O}}_{\bar{X}}[2-k]).$
The pairing on ${\mathcal{F}}_{\mathop{\rm dR}\nolimits}^{k}$ gives rise to a morphism of filtered complexes
$\omega^{\bullet}({\mathcal{F}}_{c,{\mathop{\rm dR}\nolimits}}^{k})\otimes\omega^{\bullet}({\mathcal{F}}_{\mathop{\rm dR}\nolimits}^{k})\to\omega^{\bullet}({\mathcal{F}}_{c,{\mathop{\rm dR}\nolimits}}^{k}\otimes_{{\mathcal{O}}_{\bar{X}}}{\mathcal{F}}_{\mathop{\rm dR}\nolimits}^{k})\to\omega^{\bullet}({\mathcal{I}}_{X^{\infty}})[2-k].$
Composing the resulting map on cohomology with the cup product, we obtain
${\mathcal{M}}_{c,{\mathop{\rm dR}\nolimits}}\otimes{\mathcal{M}}_{\mathop{\rm dR}\nolimits}\to{\bf H}^{2}(\bar{X},\omega^{\bullet}({\mathcal{F}}_{c,{\mathop{\rm dR}\nolimits}}^{k})\otimes\omega^{\bullet}({\mathcal{F}}_{\mathop{\rm dR}\nolimits}^{k}))\to{\bf H}^{2}(\bar{X},\omega^{\bullet}({\mathcal{I}}_{X^{\infty}})[2-k]).$
Since differentiation induces the zero map $H^{1}(\bar{X},{\mathcal{I}}_{X^{\infty}})\to H^{1}(\bar{X},\Omega^{1}_{\bar{X}/S})$, we conclude that
${\bf H}^{2}(\bar{X},\omega^{\bullet}({\mathcal{I}}_{X^{\infty}})[2-k])\cong H^{1}(\bar{X},\Omega^{1}_{\bar{X}/S})[1-k].$
Composing with the trace map, we obtain a pairing
${\mathcal{M}}_{c,{\mathop{\rm dR}\nolimits}}\otimes{\mathcal{M}}_{\mathop{\rm dR}\nolimits}\to H^{1}(\bar{X},\Omega^{1}_{\bar{X}/T})[1-k]\to{\mathcal{T}}_{\mathop{\rm dR}\nolimits}^{S,\otimes(k-1)},$
respecting filtrations and inducing an isomorphism
$\delta^{L}_{\mathop{\rm dR}\nolimits}:{\mathcal{M}}_{c,{\mathop{\rm dR}\nolimits}}\to{\mathop{\rm Hom}\nolimits}({\mathcal{M}}_{\mathop{\rm dR}\nolimits},{\mathcal{T}}_{\mathop{\rm dR}\nolimits}^{S,\otimes(k-1)}).$
Finally, for $\ell\not\in S$, one checks that (8), pulled back to $\bar{X}_{{\mathbb{Z}}_{\ell}}$, arises from an isomorphism
(9) -${\mathcal{F}}_{\ell\mbox{\scriptsize-crys}}\to{\mathcal{F}}_{\ell\mbox{\scriptsize-crys}}^{t}={\mathop{{\mathcal{H}}om}\nolimits}_{{\mathcal{O}}_{\bar{X}_{{\mathbb{Z}}_{\ell}}}}({\mathcal{F}}_{\ell\mbox{\scriptsize-crys}},{\mathcal{O}}_{\bar{X}_{{\mathbb{Z}}_{\ell}}}[-1])$
in the category ${{\mathcal{M}\mathcal{F}}}^{\nabla}_{[0,1]}(X_{{\mathbb{Z}}_{\ell}})$, giving an isomorphism
-${\mathcal{F}}_{\ell\mbox{\scriptsize-crys}}^{k}\to{\mathop{{\mathcal{H}}om}\nolimits}_{\bar{X}_{{\mathbb{Z}}_{\ell}}}({\mathcal{F}}_{\ell\mbox{\scriptsize-crys}}^{k},{\mathcal{O}}_{\bar{X}_{{\mathbb{Z}}_{\ell}}}[2-k])$
in ${{\mathcal{M}\mathcal{F}}}^{\nabla}_{[0,k-1]}$ and ultimately an isomorphism
$\delta^{L}_{\ell\mbox{\scriptsize-crys}}:{\mathcal{M}}_{c,{\ell\mbox{\scriptsize-crys}}}\to{\mathcal{M}}_{\ell\mbox{\scriptsize-crys}}^{t}={\mathop{\rm Hom}\nolimits}_{{\mathbb{Z}}_{\ell}}({\mathcal{M}}_{\ell\mbox{\scriptsize-crys}},{\mathcal{T}}_{\ell\mbox{\scriptsize-crys}}^{\otimes(k-1)}).$
Now one has to check that the pairings respect the comparison isomorphisms. This is straightforward for the maps $^{\ell}_{B}, II^{\infty}$ and $I^{\ell}_{\mathop{\rm dR}\nolimits}$ using the compatibility of the comparison isomorphisms with cup products and Poincaré duality. We find that the same holds for $I^{\ell}$ for $\ell\nmid Nk!$, but we must take some care with the integral formulation since the filtration length of $M_{c,{\ell\mbox{\scriptsize-crys}}}\otimes M_{{\ell\mbox{\scriptsize-crys}}}$ is 2(k1)2(k-1), which may be greater than $\ell-2$. We proceed as follows. The isomorphism (9) is compatible with ${\mathcal{F}}_{\ell}\to{\mathcal{F}}_{\ell}^{t}={\mathop{{\mathcal{H}}om}\nolimits}_{{\mathbb{Z}}_{\ell}}({\mathcal{F}}_{\ell},{\mathbb{Z}}_{\ell}(-1))$ in the sense that the diagram
$\begin{array}[]{ccc}{\mathbb{V}}({\mathcal{F}}_{\ell\mbox{\scriptsize-crys}})&\to&{\mathbb{V}}({\mathcal{F}}_{\ell\mbox{\scriptsize-crys}}^{t})\\ \downarrow&&\downarrow\\ {\mathcal{F}}_{\ell}&\to&{\mathcal{F}}_{\ell}^{t}\end{array}$
commutes, where the vertical arrows are given by Faltings’ comparison isomorphisms (and on the right, we are also using the canonical isomorphism between ${\mathbb{V}}({\mathcal{F}}_{\ell\mbox{\scriptsize-crys}}^{t})$ and ${\mathbb{V}}({\mathcal{F}}_{\ell\mbox{\scriptsize-crys}})^{t}$ given by the discussion in [Fa, IIh)]). We deduce a similar compatibility for the dualities on ${\mathcal{F}}_{\lambda\mbox{\scriptsize-crys}}^{k}$ and ${\mathcal{F}}_{\ell}^{k}$, and eventually a commutative diagram
$\begin{array}[]{ccc}{\mathbb{V}}({\mathcal{M}}_{c,{\ell\mbox{\scriptsize-crys}}})&\to&{\mathbb{V}}({\mathcal{M}}_{\ell\mbox{\scriptsize-crys}}^{t})\\ \downarrow&&\downarrow\\ {\mathcal{M}}_{c,\ell}&\to&{\mathcal{M}}_{\ell}^{t},\end{array}$
where ${\mathcal{M}}_{\ell}^{t}={\mathop{\rm Hom}\nolimits}_{{\mathbb{Z}}_{\ell}}({\mathcal{M}}_{\ell},{\mathcal{T}}_{\ell}^{\otimes(k-1)})$.
Recall that we have shown ${\mathbb{C}}\otimes M_{c,B}\to{\mathbb{C}}\otimes M_{c,{\mathop{\rm dR}\nolimits}}$ is an isomorphism. The same now follows for ${\mathbb{C}}\otimes M_{B}\to{\mathbb{C}}\otimes M_{\mathop{\rm dR}\nolimits}$ from the perfectness of the pairings and their compatibility with these comparisons. We use their inverses to define the comparison isomorphisms $I^{\infty}$.
Finally, we need the relation with the Petersson inner product. Recall that the Petersson inner product for $g\in S_{k}(\Gamma(N))$, $h\in M_{k}(\Gamma(N))$ is defined by
[…]
For ?=B?=B, $\mathop{\rm dR}\nolimits$} or $\ell$, $g\in M_{c,?}$ and $h\in M_{?}$, we write ${(}g,h{)}_{?}$ for the image of $g\otimes h$ in ${\mathbb{Q}}(1-k)_{?}$. For $g\in{\mathbb{C}}\otimes\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}M_{c,{\mathop{\rm dR}\nolimits}}$ and $t\in({\mathbb{Z}}/N{\mathbb{Z}})^{\times}$, write $g_{t}$ for the image of $g$ in the corresponding component of $S_{k}(\Gamma(N))$, and similarly for $M_{\mathop{\rm dR}\nolimits}$ and $M_{k}(\Gamma(N))$.
Lemma $2.$
If $g\in{\mathbb{C}}\otimes\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}M_{c,{\mathop{\rm dR}\nolimits}}$ and $h\in{\mathbb{C}}\otimes\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}M_{\mathop{\rm dR}\nolimits}$, then we have
${(}g,(I^{\infty})^{-1}(F_{\infty}\otimes 1)I^{\infty}h{)}_{{\mathop{\rm dR}\nolimits}}=(k-2)!(4\pi)^{k-1}\sum_{t\in({\mathbb{Z}}/N{\mathbb{Z}})^{\times}}(g_{t},h_{t})_{\Gamma(N)}\otimes\iota^{k-1}$
in ${\mathbb{C}}\otimes{\mathbb{Q}}(1-k)_{\mathop{\rm dR}\nolimits}$.
Proof.
Let $\bar{\omega}^{\bullet}_{\bar{E}^{\mathop{\rm an}\nolimits}/\bar{X}^{\mathop{\rm an}\nolimits}}$ denote the complex $F_{\bar{E},*}\omega^{\bullet}_{\bar{E}^{\mathop{\rm an}\nolimits}/\bar{X}^{\mathop{\rm an}\nolimits}}$, and let $\bar{{\mathcal{F}}}_{\mathop{\rm dR}\nolimits}^{\mathop{\rm an}\nolimits}$ denote the sheaf $F_{\bar{X},*}{\mathcal{F}}_{\mathop{\rm dR}\nolimits}^{\mathop{\rm an}\nolimits}\cong{\bf R}^{1}s_{*}\bar{\omega}^{\bullet}_{E^{\mathop{\rm an}\nolimits}/X^{\mathop{\rm an}\nolimits}}$ of modules over $\bar{{\mathcal{O}}}_{\bar{X}^{\mathop{\rm an}\nolimits}}=F_{\bar{X},*}{\mathcal{O}}_{\bar{X}^{\mathop{\rm an}\nolimits}}$. We then have an antiholomorphic resolution of ${\mathbb{C}}\otimes{\mathcal{F}}_{B}^{k}$ analogous to (5), and complex conjugation of functions gives rise to an isomorphism $\omega^{\bullet}_{\bar{E}^{\mathop{\rm an}\nolimits}/\bar{X}^{\mathop{\rm an}\nolimits}}\cong\bar{\omega}^{\bullet}_{\bar{E}^{\mathop{\rm an}\nolimits}/\bar{X}^{\mathop{\rm an}\nolimits}}$, which induces an isomorphism
$\omega^{\bullet}({\mathcal{F}}_{\mathop{\rm dR}\nolimits}^{k,{\mathop{\rm an}\nolimits}})\cong\bar{\omega}^{\bullet}(\bar{{\mathcal{F}}}_{\mathop{\rm dR}\nolimits}^{k,{\mathop{\rm an}\nolimits}})=F_{\bar{X},*}\omega^{\bullet}({\mathcal{F}}_{\mathop{\rm dR}\nolimits}^{k,{\mathop{\rm an}\nolimits}})$
compatible with $F_{\infty}\otimes 1$ on ${\mathbb{C}}\otimes{\mathcal{F}}_{B}$.
Recall that $g$ corresponds to the class in ${\bf H}^{1}(\bar{X}^{\mathop{\rm an}\nolimits},\omega^{\bullet}({\mathcal{F}}_{c,{\mathop{\rm dR}\nolimits}}^{k,{\mathop{\rm an}\nolimits}}))$ that arises from the section of $H^{0}(\bar{X},{\mathcal{C}}^{1}_{\bar{X}}\otimes_{{\mathcal{O}}_{\bar{X}^{\mathop{\rm an}\nolimits}}}{\mathcal{F}}_{c,{\mathop{\rm dR}\nolimits}}^{k,{\mathop{\rm an}\nolimits}})$ whose pull-back by $\pi_{t}:{\mathfrak{H}}\to X_{N,t}$ is defined by
$\tau\mapsto 2\pi ig_{t}(\tau)d\tau\otimes\pi_{t}^{*}\omega_{\mathop{\rm can}\nolimits}^{\otimes(k-2)}.$
Here ${\mathcal{C}}^{n}$ is the sheaf of smooth $\mathbb{C}$}-valued $n$-forms and $\omega_{\mathop{\rm can}\nolimits}$ is the canonical section of ${\mathcal{F}}_{{\mathop{\rm dR}\nolimits}}^{\mathop{\rm an}\nolimits}$ (i.e., $dx/x$ on the Tate curve ${\mathbb{G}}_{m}/q^{\mathbb{Z}}$). Similarly $(I^{\infty})^{-1}(F_{\infty}\otimes 1)I^{\infty}h$ arise from the section of $H^{0}(\bar{X},{\mathcal{C}}^{1}_{\bar{X}}\otimes_{\bar{{\mathcal{O}}}_{\bar{X}^{\mathop{\rm an}\nolimits}}}\bar{{\mathcal{F}}}_{{\mathop{\rm dR}\nolimits}}^{k,{\mathop{\rm an}\nolimits}})$ whose pull-back is
$\tau\mapsto-2\pi i\bar{h}_{t}(\tau)d(\bar{\tau})\otimes\pi_{t}^{*}(\bar{\omega}_{\mathop{\rm can}\nolimits})^{\otimes(k-2)}.$
Note that $\omega^{\mathop{\rm can}\nolimits}$ and $\bar{\omega}^{\mathop{\rm can}\nolimits}$ both define sections of ${\mathcal{C}}^{0}_{\bar{X}}\otimes{\mathcal{F}}_{B}$, and under the pairing ${\mathcal{F}}_{B}\otimes{\mathcal{F}}_{B}\to 2\pi i^{-1}{\mathbb{Q}}$ (pulled back to $\mathfrak{H}$} and tensored with ${\mathcal{C}}^{0}_{\mathfrak{H}}$), we have
$\omega_{\mathop{\rm can}\nolimits}\otimes\bar{\omega}_{\mathop{\rm can}\nolimits}\mapsto(2\pi i)^{2}(-2i{\mathop{\rm Im}\nolimits}\tau)\otimes(2\pi i)^{-1}.$
It follows that ${(}g,(I^{\infty})^{-1}(F_{\infty}\otimes 1)I^{\infty}h{)}_{\mathop{\rm dR}\nolimits}$ is given by the class in ${\bf H}^{2}(\bar{X},\omega^{\bullet}({\mathcal{I}}_{X^{\infty}}^{\mathop{\rm an}\nolimits})[2-k])$ arising from the section of $\Gamma(\bar{X},{\mathcal{C}}^{2}_{\bar{X}}[2-k])$ whose pull-back by $\pi_{t}$ is defined by
$\tau\mapsto(k-2)!(2\pi i)^{k}g_{t}(\tau)\overline{h_{t}(\tau)}(-2i{\mathop{\rm Im}\nolimits}\tau)^{k-2}d\tau\wedge d\bar{\tau}\otimes\iota^{k-2},$
where the pairing on $\operatorname{{\mathop{\rm Sym}\nolimits}}^{k-2}$ is defined in (7). Under the trace map to ${\mathbb{C}}\otimes{\mathbb{Q}}(1-k)_{\mathop{\rm dR}\nolimits}$, this goes to $(k-2)!(4\pi)^{k-1}(g_{t},h_{t})_{\Gamma(N)}\otimes\iota^{k-1}$. ∎
2.5. Weight filtrations and summary
There is a natural map ${\mathcal{M}}_{c,?}\to{\mathcal{M}}_{?}$ for each realization respecting all of the data and comparison isomorphisms. To make $\mathcal{M}$} and ${\mathcal{M}}_{c}$ premotivic structures, one checks that setting
$W_{i}M_{?}$ $=$ $\left\{\begin{array}[]{ll}0,&\mbox{if $i<k-1$;}\\ \operatorname{{\mathop{\rm im}\nolimits}}(M_{c,?}\to M_{?}),&\mbox{if $k-1\leq i<2(k-1)$;}\\ M_{?},&\mbox{if $2(k-1)\leq i$;}\end{array}\right$.
$W_{i}M_{c,?}$ $=$ $\left\{\begin{array}[]{ll}0,&\mbox{if $i<0$;}\\ \ker(M_{c,?}\to M_{?}),&\mbox{if $0\leq i<k-1$;}\\ M_{c,?},&\mbox{if $k-1\leq i$.}\end{array}\right$.
defines weight filtrations. In each case, this amounts to the assertion that
$\left(\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}V\right)\,\cap\,\left((I^{\infty})^{-1}(F_{\infty}\otimes 1)I^{\infty}\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}V\right)=0,$
where $V=\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}({\mathbb{C}}\otimes M_{c,{\mathop{\rm dR}\nolimits}})\cong\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}W_{k-1}({\mathbb{C}}\otimes M_{\mathop{\rm dR}\nolimits})$, and this is immediate from Lemma $2.$
We can now regard $\mathcal{M}$} and ${\mathcal{M}}_{c}$ as objects of ${\mathcal{P}M}_{\mathbb{Q}}^{S}$, and $M={\mathbb{Q}}\otimes{\mathcal{M}}$ and $M_{c}={\mathbb{Q}}\otimes{\mathcal{M}}_{c}$ as objects of $\operatorname{{\bf PM}}_{\mathbb{Q}}^{S}$, where $S$ contains the set of primes dividing $Nk!$. We have also defined a (left) duality morphism
$\delta^{L}:{\mathcal{M}}_{c}\to{\mathop{\rm Hom}\nolimits}({\mathcal{M}},{\mathcal{T}}^{S,\otimes(k-1)})$
whose kernel and cokernel are torsion objects $\mathcal{N}$} satisfying ${\mathcal{N}}_{\ell}=0$ for $\ell>k-2$, and which therefore induces an isomorphism $M_{c}\to{\mathop{\rm Hom}\nolimits}(M,{\mathbb{Q}}(1-k)^{S})$. Similarly we have a (right) duality morphism $\delta^{R}=\delta^{R}_{N}:{\mathcal{M}}\to{\mathop{\rm Hom}\nolimits}({\mathcal{M}}_{c},{\mathcal{T}}^{S,\otimes(k-1)})$ whose kernel and cokernel satisfy ${\mathcal{N}}_{\ell}=0$ for $\ell\nmid(k-2)!$. Moreover the resulting diagram
[…]
is commutative, where the right-hand vertical arrow is induced by $(-1)^{k-1}$ times the map ${\mathcal{M}}_{c}\to{\mathcal{M}}$.
We let ${\mathcal{M}}_{\mathop{\rm tf}\nolimits}$ denote the maximal torsion-free quotient of $\mathcal{M}$} (i.e., ${\mathcal{M}}/{\mathcal{M}}[r]$ where $r\in{\mathbb{Z}}_{>0}$ is chosen to annihilate the the torsion in ${\mathcal{M}}_{B}$ and ${\mathcal{M}}[r]$ denotes the kernel of multiplication by $r$ on $\mathcal{M}$}.) Finally, we let $\mathcal{M}$$_{!}$ denote the premotivic structure $\operatorname{{\mathop{\rm im}\nolimits}}({\mathcal{M}}_{c}\to{\mathcal{M}}_{\mathop{\rm tf}\nolimits})$ in ${\mathcal{P}M}_{\mathbb{Q}}^{S}$, pure of weight k1k-1. We let $M_{!}={\mathbb{Q}}\otimes{\mathcal{M}}_{!}$. From the commutative diagram above we obtain morphisms
$\begin{array}[]{ccc}\delta^{L}_{!}:{\mathcal{M}}_{!}&\to&{\mathop{\rm Hom}\nolimits}({\mathcal{M}}_{!},{\mathcal{T}}^{S,\otimes(k-1)}){\rm\ and}\\ \delta^{L}_{!}:M_{!}&\to&{\mathop{\rm Hom}\nolimits}(M_{!},{\mathbb{Q}}(1-k)^{S})\end{array}$
of sign $(-1)^{k-1}$. The first map is injective and the second is an isomorphism; moreover the following lemma shows that the cokernel of the first map satisfies ${\mathcal{C}}_{\ell}=0$ unless $k>2$ and $\ell|N(k-2)!$.
Lemma 2.3.
Let $\mathcal{N}$} denote the cokernel of ${\mathcal{M}}_{c}\to{\mathcal{M}}$. Then ${\mathcal{N}}[\ell]=0$ if $k=2$ or $\ell$ does not divide N(k2)!N(k-2)!.
Proof.
It suffices to prove the assertion for the Betti realization, and for this it suffices (by [Br, II.17(2)], for example) to check that for any cusp and any $\alpha>1$, $H^{1}(D_{\alpha}^{0},j^{*}{\mathcal{F}}_{B}^{k})$ has no $\ell$-torsion, where $j$ is the inclusion $D_{\alpha}^{0}\to X^{\mathop{\rm an}\nolimits}$ in a punctured neighborhood of the cusp. This amounts to the surjectivity of
$H^{0}(D_{\alpha}^{0},j^{*}{\mathcal{F}}_{B}^{k})\to H^{0}(D_{\alpha}^{0},{\mathbb{F}}_{\ell}\otimes j^{*}{\mathcal{F}}_{B}^{k}),$
which holds if $k=2$ or $\ell$ does not divide N(k2)!N(k-2)!. ∎
We now summarize main properties of the premotivic structures ${\mathcal{M}}_{N}$ and $M_{N}$ for level NN modular forms.
Theorem 2.4.
Fix integers $N\geq 3$ and $k\geq 2$. Let $S_{N}=\{\ell\nmid Nk!\}$.
  1. (a)
    Let $S\supseteq S_{N}$. Together with the induced comparison isomorphisms and weight filtrations, ${\mathcal{M}}_{N}^{S},\ {\mathcal{M}}_{N,c}^{S}$ and ${\mathcal{M}}_{N,!}^{S}$ are in ${\mathcal{P}M}^{S}_{\mathbb{Q}}$, and ${\mathcal{M}}_{N,c}^{S}\to{\mathcal{M}}_{N,!}^{S}\to{\mathcal{M}}_{N,{\mathop{\rm tf}\nolimits}}^{S}$ are morphisms in ${\mathcal{P}M}^{S}_{\mathbb{Q}}$.
  2. (b)
    Let $S\supseteq S_{N}$. There are duality homomorphisms
    $\delta^{L}_{N}:{\mathcal{M}}_{N,c}^{S}\to{\mathop{\rm Hom}\nolimits}_{{\mathbb{Z}}}({\mathcal{M}}_{N}^{S},{\mathcal{T}}^{S,\otimes(k-1)})$
    and
    $\delta^{L}_{N,!}:{\mathcal{M}}_{N,!}^{S}\to{\mathop{\rm Hom}\nolimits}_{{\mathbb{Z}}}({\mathcal{M}}_{N,!}^{S},{\mathcal{T}}^{S,\otimes(k-1)})$
    in ${\mathcal{P}M}^{S}_{\mathbb{Q}}$, giving rise to perfect pairings
    $(\ ,\ ):M_{N,c}^{S}\otimes M_{N}^{S}\to{\mathbb{Q}}(1-k)^{S}$
    and
    $(\ ,\ )_{!}:M_{N,!}^{S}\otimes M_{N,!}^{S}\to{\mathbb{Q}}(1-k)^{S}$
    in $\operatorname{{\bf PM}}^{S}_{\mathbb{Q}}$.
3. The action of $\operatorname{{\mathop{\rm GL}\nolimits}}_{2}({\mathbb{A}}_{\bf f})$
In this section we define the adelic action on premotivic structures associated to modular forms.
3.1. Adelic modular forms
We recall the adelic definition of modular curves and forms.
Suppose that $U$ is an open compact subgroup of $\operatorname{{\mathop{\rm GL}\nolimits}}_{2}({\mathbb{A}}_{\bf f})$ where ${\mathbb{A}}_{\bf f}$ denotes the finite adeles. Let $U_{\infty}$ denote the stabilizer of $i$ in $\operatorname{{\mathop{\rm GL}\nolimits}}_{2}({\mathbb{R}})$, so $U_{\infty}={\mathbb{R}}^{\times}\operatorname{{\mathop{\rm SO}\nolimits}}_{2}({\mathbb{R}})$. The analytic modular curve $X_{U}$ of level $U$ is defined as the quotient
$GL_{2}({\mathbb{Q}})\backslash\operatorname{{\mathop{\rm GL}\nolimits}}_{2}({\mathbb{A}})/UU_{\infty}.$
The analytic structure is characterized by requiring that if $g$ is in $\operatorname{{\mathop{\rm GL}\nolimits}}_{2}({\mathbb{A}}_{\bf f})$, then the map ${\mathfrak{H}}\to X_{U}$ defined by $\gamma(i)\to\operatorname{{\mathop{\rm GL}\nolimits}}_{2}({\mathbb{Q}})g\gamma UU_{\infty},\ \gamma\in\operatorname{{\mathop{\rm GL}\nolimits}}_{2}^{+}({\mathbb{R}}),$ is holomorphic.
A modular form of level $U$ is a function $\phi:\operatorname{{\mathop{\rm GL}\nolimits}}_{2}({\mathbb{A}})\to{\mathbb{C}}$ such that
  • if $\delta\in\operatorname{{\mathop{\rm GL}\nolimits}}_{2}({\mathbb{Q}})$, $x\in\operatorname{{\mathop{\rm GL}\nolimits}}_{2}({\mathbb{A}})$, $u\in U$ and $v={\mbox{\scriptsize$\left(\begin{array}[]{cc}{{a}}&{{b}}\\ {{c}}&{{d}}\end{array}\right)$}}\in U_{\infty}$, then
    $\phi(\delta xuv)={\mathop{\rm det}\nolimits}v(ci+d)^{-k}\phi(x)$
  • if $g\in\operatorname{{\mathop{\rm GL}\nolimits}}_{2}({\mathbb{A}}^{\infty})$, then the function $\mathfrak{H}$$\to{\mathbb{C}}$ defined by
    $\gamma(i)\mapsto({\mathop{\rm det}\nolimits}\gamma)^{-1}(ci+d)^{k}\phi(g\gamma)$
    for $\gamma={\mbox{\scriptsize$\left(\begin{array}[]{cc}{{a}}&{{b}}\\ {{c}}&{{d}}\end{array}\right)$}}\in\operatorname{{\mathop{\rm GL}\nolimits}}_{2}^{+}({\mathbb{R}})$ is a modular form of weight $k$ with respect to $gUg^{-1}\cap\operatorname{{\mathop{\rm GL}\nolimits}}_{2}^{+}({\mathbb{Q}})$.
Note that the first condition ensures that the function in the second condition is well-defined and satisfies the usual transformation property for modular forms. The second condition therefore only amounts to a holomorphy requirement. We let $M_{k}(U)$ denote the space of modular forms of level $U$. The space of cusp forms of level $U$ is defined similarly and denoted $S_{k}(U)$.
Suppose now that $U$ and $U^{\prime}$ are open compact subgroups of $\operatorname{{\mathop{\rm GL}\nolimits}}_{2}({\mathbb{A}}_{\bf f})$, and $g$ is an element of $\operatorname{{\mathop{\rm GL}\nolimits}}_{2}({\mathbb{A}}_{\bf f})$ such that $g^{-1}U^{\prime}g\subset U$. Note that right multiplication by $g$ induces a holomorphic map $X_{U^{\prime}}\to X_{U}$, and inclusions $M_{k}(U)\to M_{k}(U^{\prime})$ and $S_{k}(U)\to S_{k}(U^{\prime})$. We thus obtain an action of $\operatorname{{\mathop{\rm GL}\nolimits}}_{2}({\mathbb{A}}_{\bf f})$ on
${\mathcal{A}}_{k}=\lim_{\stackrel{{\scriptstyle\to}}{{U}}}M_{k}(U)\quad\mbox{and}\quad{\mathcal{A}}_{k}^{0}=\lim_{\stackrel{{\scriptstyle\to}}{{U}}}S_{k}(U).$
[…]
$\Gamma(N)\cdot\gamma(i)\mapsto\operatorname{{\mathop{\rm GL}\nolimits}}_{2}({\mathbb{Q}})\cdot g_{t}\gamma\cdot UU_{\infty}$
for $\gamma\in\operatorname{{\mathop{\rm GL}\nolimits}}_{2}^{+}({\mathbb{R}})$. We identify $M_{k}(U)$ with $\oplus_{t}M_{k}(\Gamma(N))$ via the isomorphism $\beta$ defined by
(12) $\beta(\phi)_{t}(\gamma(i))=({\mathop{\rm det}\nolimits}\gamma)^{-1}(ci+d)^{k}\phi(g_{t}\gamma)$
for $\gamma={\mbox{\scriptsize$\left(\begin{array}[]{cc}{{a}}&{{b}}\\ {{c}}&{{d}}\end{array}\right)$}}\in\operatorname{{\mathop{\rm GL}\nolimits}}_{2}^{+}({\mathbb{R}})$. We have a similar identification for cusp forms. Note that $\eta$ and $\beta$ are independent of the choices of the $g_{t}$. Composing with the isomorphisms defined in §2.3. we obtain isomorphisms
${\mathbb{C}}\otimes\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}{\mathcal{M}}_{{\mathop{\rm dR}\nolimits}}\cong M_{k}(U)\quad\mbox{and}\quad$
${\mathbb{C}}\otimes\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}{\mathcal{M}}_{c,{\mathop{\rm dR}\nolimits}}\cong{\mathbb{C}}\otimes\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}{\mathcal{M}}_{!,{\mathop{\rm dR}\nolimits}}\cong S_{k}(U).$
In particular, we have
${\mathcal{A}}_{k}\cong{\mathbb{C}}\otimes\lim_{\stackrel{{\scriptstyle\to}}{{N}}}\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}{\mathcal{M}}_{N,{\mathop{\rm dR}\nolimits}}{\rm\ and\ }$
${\mathcal{A}}_{k}^{0}\cong{\mathbb{C}}\otimes\lim_{\stackrel{{\scriptstyle\to}}{{N}}}\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}{\mathcal{M}}_{N,!,{\mathop{\rm dR}\nolimits}}\cong{\mathbb{C}}\otimes\lim_{\stackrel{{\scriptstyle\to}}{{N}}}\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}{\mathcal{M}}_{N,c,{\mathop{\rm dR}\nolimits}}.$
Furthermore if $N,N^{\prime}\geq 3$ and $g\in\operatorname{{\mathop{\rm GL}\nolimits}}_{2}({\mathbb{A}}_{\bf f})$ are such that $g^{-1}U_{N^{\prime}}g\subset U_{N}$, then we have a map
${\mathbb{C}}\otimes\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}{\mathcal{M}}_{N,{\mathop{\rm dR}\nolimits}}\to{\mathbb{C}}\otimes\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}{\mathcal{M}}_{N^{\prime},{\mathop{\rm dR}\nolimits}},$
and similarly for ${\mathcal{M}}_{c}$ and $\mathcal{M}$$_{!}$. We shall explain how to recover these from maps of premotivic structures.
3.2. Action on curves
[…]
Lemma 3.1.
  1. (a)
    If $(h,N,N^{\prime})$ is an admissible triple, then $N|N^{\prime}$ and $h^{-1}U_{N^{\prime}}h\subset U_{N}$.
  2. (b)
    If $(h,N,N^{\prime})$ and $(h^{\prime},N^{\prime},N^{\prime\prime})$ are admissible triples, then so is $(h^{\prime}h,N,N^{\prime\prime})$.
  3. (c)
[…]
Suppose now that $(h,N,N^{\prime})$ is an admissible triple. Let $\bar{E}/\bar{X}$ (respectively, $\bar{E}^{\prime}/\bar{X}^{\prime}$) denote the universal generalized elliptic curve with level NN (respectively $N^{\prime}$) structure.
[…]
$({\mathbb{Z}}/N{\mathbb{Z}})^{2}\rightarrow({\mathbb{Z}}/N^{\prime}{\mathbb{Z}})^{2}/(({\mathbb{Z}}/N^{\prime}{\mathbb{Z}})^{2}(N^{\prime}h^{-1})$
and so gives rise to a level NN-structure on $E^{\prime}/G$.
Letting $(\bar{E}^{\prime}/G)_{\rm cont}$ denote the contraction of $\bar{E}^{\prime}/G$ whose cuspidal fibers are NN-gons [D-R, IV.1], we obtain a level NN structure on $(\bar{E}^{\prime}/G)_{\rm cont}\to\bar{X}^{\prime}$. By the universal property of $\bar{E}_{/\bar{X}}$, this defines a map $\chi_{h}=\chi_{h,N,N^{\prime}}:\bar{X}^{\prime}\to\bar{X}$ such that there is an isomorphism $(\bar{E}^{\prime}/G)_{\rm cont}\to\bar{E}\times_{\bar{X}}\bar{X}^{\prime}$ of generalized elliptic curves with level NN-structures. We let $\varepsilon_{h}=\varepsilon_{h,N,N^{\prime}}$ denote the composite of the natural map $\bar{E}^{\prime}\to(\bar{E}^{\prime}/G)_{\rm cont}$ with this isomorphism, so we obtain a commutative diagram
[…]
We now give the concrete description of its analytification in terms of the models in § 2.1.
[…]
$g_{\kappa(t^{\prime})}\in\operatorname{{\mathop{\rm GL}\nolimits}}_{2}({\mathbb{Q}})g^{\prime}_{t^{\prime}}hU\operatorname{{\mathop{\rm GL}\nolimits}}_{2}^{+}({\mathbb{R}})$
[…]
$\begin{array}[]{ccc}X_{N^{\prime},t^{\prime}}&\to&X_{N,\kappa(t^{\prime})}\\ \downarrow&&\downarrow\\ X_{U^{\prime}}&\to&X_{U}\end{array}$
commutes, where the top arrow is defined by $\tau\mapsto\gamma(\tau)$, the bottom one is defined by right multiplication by $h$ and the vertical arrows are $\eta^{\prime}_{t^{\prime}}$ and $\eta_{\kappa(t)}$. Note that $\gamma^{-1}\in M_{2}({\mathbb{Z}})$ and $\gamma\Gamma^{\prime}\gamma^{-1}\subset\Gamma$ so that the map
$\begin{array}[]{rccc}\tilde{\gamma}:&E_{N^{\prime},t^{\prime}}&\to&E_{N,\kappa(t^{\prime})},\\ &(\tau,z)&\mapsto&(\gamma(\tau),(c\tau+d)^{-1}z),\end{array}$
where $\gamma={\mbox{\scriptsize$\left(\begin{array}[]{cc}{{a}}&{{b}}\\ {{c}}&{{d}}\end{array}\right)$}}$, is well-defined. Note also that $\gamma$ is uniquely determined by $(h,N,N^{\prime})$ and $t^{\prime}$ up to left multiplication by an element of $\Gamma$, so $\tilde{\gamma}$ is independent of the choice of $\gamma$.
Lemma 3.2.
Suppose that $(h,N,N^{\prime})$ is an admissible triple and $t^{\prime}\in({\mathbb{Z}}/N^{\prime}{\mathbb{Z}})^{\times}$. The restriction of $\chi_{h}^{\mathop{\rm an}\nolimits}$ to $X_{N^{\prime},t^{\prime}}$ and $\varepsilon_{h}^{\mathop{\rm an}\nolimits}$ to $E_{N^{\prime},t^{\prime}}$ are given by $\gamma:X_{N^{\prime},t^{\prime}}\to X_{N,\kappa(t^{\prime})}$ and $\tilde{\gamma}:E_{N^{\prime},t^{\prime}}\to E_{N,\kappa(t^{\prime})}$.
Proof.
For $\tau^{\prime}\in{\mathfrak{H}}$, let $E_{\tau^{\prime}}={\mathbb{C}}/\Lambda_{\tau^{\prime}}$ denote the fiber of $E^{\prime,{\mathop{\rm an}\nolimits}}$ over $\Gamma^{\prime}\tau^{\prime}\in X_{N^{\prime},t^{\prime}}$, where $\Lambda_{\tau^{\prime}}={\mathbb{Z}}^{2}{\mbox{\scriptsize$\left(\begin{array}[]{c}{{\tau^{\prime}}}\\ {{1}}\end{array}\right)$}}$ and let $G_{\tau^{\prime}}\subset E_{\tau^{\prime}}$ denote the fiber of $G^{\mathop{\rm an}\nolimits}$. Similarly let $E_{\tau}={\mathbb{C}}/\Lambda_{\tau}$ denote the fiber of $E^{\mathop{\rm an}\nolimits}$ over $\Gamma\tau$, where $\tau=\gamma(\tau^{\prime})$ and $\Lambda_{\tau}={\mathbb{Z}}^{2}{\mbox{\scriptsize$\left(\begin{array}[]{c}{{\tau}}\\ {{1}}\end{array}\right)$}}$. It suffices to check the lemma fiberwise, i.e., that
$\tilde{\gamma}_{\tau^{\prime}}:E_{\tau^{\prime}}\to E_{\tau}$
induces an isomorphism $E_{\tau^{\prime}}/G_{\tau^{\prime}}\cong E_{\tau}$ compatible with level NN-structures.
Recall that the level $N^{\prime}$-structure on $E_{\tau^{\prime}}$ is defined by
$(m,n)g^{\prime}_{t^{\prime}}\bmod N^{\prime}\hat{{\mathbb{Z}}}^{2}\ \mapsto\ (m,n)(N^{\prime})^{-1}\left(\begin{array}[]{c}{\tau^{\prime}}\\ {1}\end{array}\right)\bmod\Lambda_{\tau^{\prime}}$
for $(m,n)\in{\mathbb{Z}}^{2}$. Since $h^{-1}=ug_{\kappa(t^{\prime})}^{-1}\gamma_{\bf f}g^{\prime}_{t^{\prime}}$ for some $u\in U$ and multiplication by $ug_{\kappa(t^{\prime})}^{-1}$ is an automorphism of $({\mathbb{Z}}/N^{\prime}{\mathbb{Z}})^{2}$, we see that $G_{\tau^{\prime}}$ is the image of ${\mathbb{Z}}^{2}(N^{\prime}\gamma_{\bf f})g^{\prime}_{t^{\prime}}$. We therefore have
$G_{\tau^{\prime}}={\mathbb{Z}}^{2}\gamma{\mbox{\scriptsize$\left(\begin{array}[]{c}{{\tau^{\prime}}}\\ {{1}}\end{array}\right)$}}/\Lambda_{\tau^{\prime}}.$
Furthermore, since multiplication by $u$ is the identity on $({\mathbb{Z}}/N{\mathbb{Z}})^{2}$, the level NN-structure on $E_{\tau^{\prime}}/G_{\tau^{\prime}}$ is given by
$(m,n)g_{\kappa(t^{\prime})}\bmod N\hat{{\mathbb{Z}}}^{2}\ \mapsto\ (m,n)N^{-1}\gamma{\mbox{\scriptsize$\left(\begin{array}[]{c}{{\tau}}\\ {{1}}\end{array}\right)$}}\bmod{\mathbb{Z}}^{2}\gamma{\mbox{\scriptsize$\left(\begin{array}[]{c}{{\tau^{\prime}}}\\ {{1}}\end{array}\right)$}},$
and the lemma follows. ∎
Lemma $3.$
Suppose that $(h,N,N^{\prime})$ and $(h^{\prime},N^{\prime},N^{\prime\prime})$ are admissible triples. Then $\chi_{h^{\prime}h}=\chi_{h}\circ\chi_{h^{\prime}}$ and $\varepsilon_{h^{\prime}h^{\prime}}=\varepsilon_{h}\circ\varepsilon_{h^{\prime}}$.
Proof.
Since $\bar{E}^{\prime\prime}$ and $\bar{X}^{\prime\prime}$ are flat over ${\mathbb{Z}}[1/N^{\prime\prime}]$, it suffices to check that the desired equalities after extending scalars to $\mathbb{C}$}, so by GAGA it suffices to check the analytifications coincide, so it suffices to check they coincide on $E_{N^{\prime\prime},t^{\prime\prime}}$ and $X_{N^{\prime\prime},t^{\prime\prime}}$ for each $t^{\prime\prime}\in({\mathbb{Z}}/N^{\prime\prime}{\mathbb{Z}})^{\times}$, for their unions are dense subsets.
[…]
and similarly for the $\chi$’s. Furthermore, if we choose $\gamma$ and $\gamma^{\prime}\in\operatorname{{\mathop{\rm GL}\nolimits}}_{2}^{+}({\mathbb{Q}})$ so that $g_{t}\in\gamma_{\bf f}g^{\prime}_{t^{\prime}}hU$ and $g^{\prime}_{t^{\prime}}\in\gamma^{\prime}_{\bf f}g^{\prime\prime}_{t^{\prime\prime}}h^{\prime}U$, then
$g_{t}\in\gamma_{\bf f}\gamma^{\prime}_{\bf f}g^{\prime\prime}_{t^{\prime\prime}}h^{\prime}h(h^{-1}U^{\prime}h)U=\gamma_{\bf f}\gamma^{\prime}_{\bf f}g^{\prime\prime}_{t^{\prime\prime}}h^{\prime}hU.$
So we can take $\gamma^{\prime\prime}=\gamma\gamma^{\prime}$ and apply Lemma 3.2 to conclude the desired equalities. ∎
$3.$ Action on the premotivic structures
Suppose that $(h,N,N^{\prime})$ is an admissible triple. Let $S$ denote the set of primes dividing $Nk!$ and $S^{\prime}$ the set of primes dividing $N^{\prime}k!$. We let $\mathcal{M}$}, ${\mathcal{M}}_{c}$ and $\mathcal{M}$$_{!}$ (respectively, $\mathcal{M}$$^{\prime}$, ${\mathcal{M}}^{\prime}_{c}$ and $\mathcal{M}$$^{\prime}_{!}$) denote the objects of ${\mathcal{P}M}_{\mathbb{Q}}^{S}$ (respectively, ${\mathcal{P}M}_{\mathbb{Q}}^{S^{\prime}}$) structures defined in §2.5 associated to modular forms of weight $k$ and level NN (respectively, level $N^{\prime}$). We shall define compatible morphisms ${\mathcal{M}}_{c}^{S^{\prime}}\to{\mathcal{M}}_{c}^{\prime}$ and ${\mathcal{M}}^{S^{\prime}}\to{\mathcal{M}}^{\prime}$, giving also ${\mathcal{M}}_{!}^{S^{\prime}}\to{\mathcal{M}}_{!}^{\prime}$.
We first describe the action on realizations. The analytification of (13) gives a natural transformation $(\chi_{h}^{\mathop{\rm an}\nolimits})^{*}s^{\mathop{\rm an}\nolimits}_{*}\to(s^{\prime})^{\mathop{\rm an}\nolimits}_{*}(\varepsilon_{h}^{\mathop{\rm an}\nolimits})^{*}$ from which we obtain a map
$(\chi_{h}^{\mathop{\rm an}\nolimits})^{*}{\mathcal{F}}_{\mathop{\rm B}\nolimits}=(\chi_{h}^{\mathop{\rm an}\nolimits})^{*}R^{1}s^{\mathop{\rm an}\nolimits}_{*}{\mathbb{Z}}\to R^{1}(s^{\prime})^{\mathop{\rm an}\nolimits}_{*}{\mathbb{Z}}={\mathcal{F}}_{\mathop{\rm B}\nolimits}^{\prime}.$
On stalks over $x^{\prime}\in X^{\prime,{\mathop{\rm an}\nolimits}}$ this is given by the map $H^{1}(E^{\mathop{\rm an}\nolimits}_{x},{\mathbb{Z}})\to H^{1}(E^{\prime,{\mathop{\rm an}\nolimits}}_{x},{\mathbb{Z}})$ induced by $E^{\prime,{\mathop{\rm an}\nolimits}}_{x^{\prime}}\to E^{\mathop{\rm an}\nolimits}_{x}$ where $x=\chi_{h}(x^{\prime})\in X^{\mathop{\rm an}\nolimits}$. Taking symmetric products gives
$(\chi_{h}^{\mathop{\rm an}\nolimits})^{*}{\mathcal{F}}_{\mathop{\rm B}\nolimits}^{k}\cong\operatorname{{\mathop{\rm Sym}\nolimits}}^{k-2}(\chi_{h}^{\mathop{\rm an}\nolimits})^{*}{\mathcal{F}}_{\mathop{\rm B}\nolimits}\to\operatorname{{\mathop{\rm Sym}\nolimits}}^{k-2}{\mathcal{F}}_{\mathop{\rm B}\nolimits}^{\prime},$
and taking cohomology yields a homomorphism
$h]_{\mathop{\rm B}\nolimits}:{\mathcal{M}}_{\mathop{\rm B}\nolimits}=H^{1}(X^{\mathop{\rm an}\nolimits},{\mathcal{F}}_{\mathop{\rm B}\nolimits}^{k})\to H^{1}(X^{\prime,{\mathop{\rm an}\nolimits}},(\chi_{h}^{\mathop{\rm an}\nolimits})^{*}{\mathcal{F}}_{\mathop{\rm B}\nolimits}^{k})\to H^{1}(X^{\prime,{\mathop{\rm an}\nolimits}},\operatorname{{\mathop{\rm Sym}\nolimits}}^{k-2}{\mathcal{F}}_{\mathop{\rm B}\nolimits}^{\prime})={\mathcal{M}}_{\mathop{\rm B}\nolimits}^{\prime}.$
Since $X^{\prime,{\mathop{\rm an}\nolimits}}\to X^{\mathop{\rm an}\nolimits}$ is proper, the same applies to give $h]_{c,{\mathop{\rm B}\nolimits}}:{\mathcal{M}}_{c,{\mathop{\rm B}\nolimits}}\to{\mathcal{M}}^{\prime}_{c,{\mathop{\rm B}\nolimits}}$. It is straightforward to check that these maps respect complex conjugation.
Similarly we have a morphism of lisse $\ell$-adic sheaves $\chi_{h}^{*}{\mathcal{F}}_{\ell}\to{\mathcal{F}}_{\ell}^{\prime}$, which on stalks $\bar{x}^{\prime}:\operatorname{{\mathop{\rm Spec}\nolimits}}\bar{k}\to X^{\prime}$ is the natural map
$H^{1}(E^{\prime}_{\bar{x}^{\prime}},{\mathbb{Z}}_{\ell})\to H^{1}(E_{\bar{x}},{\mathbb{Z}}_{\ell})$
induced by $E^{\prime}_{\bar{x}^{\prime}}\to E_{\bar{x}}$ where $\bar{x}=\chi_{h}\circ\bar{x}^{\prime}$. Applying $\operatorname{{\mathop{\rm Sym}\nolimits}}^{k-2}$ and taking cohomology yields $G_{\mathbb{Q}}$-linear maps $h]_{\ell}:{\mathcal{M}}_{\ell}\to{\mathcal{M}}_{\ell}^{\prime}$ and (since $\chi_{h}$ is proper) $h]_{c,\ell}:{\mathcal{M}}_{c,\ell}\to{\mathcal{M}}^{\prime}_{c,\ell}$.
From the commutative diagram of log schemes
$\begin{array}[]{ccccc}(\bar{E}^{\prime},{\mathcal{N}}_{E^{\prime}})&\to&(\bar{X}^{\prime},{\mathcal{N}}_{X^{\prime}})&\to&(T^{\prime},{\mathcal{O}}_{T^{\prime}}^{*})\\ \downarrow&&\downarrow&&\downarrow\\ (\bar{E},{\mathcal{N}}_{E})&\to&(\bar{X},{\mathcal{N}}_{X})&\to&(T,{\mathcal{O}}_{T}^{*}),\end{array}$
we obtain a commutative diagram of complexes of sheaves of $(\chi_{h}s^{\prime})^{-1}{\mathcal{O}}_{\bar{X}}$- modules on $\bar{E}^{\prime}$:
$\begin{array}[]{ccccccccc}0&\to&\varepsilon_{h}^{-1}\omega^{\bullet}_{\bar{E}/\bar{X}}\otimes_{\varepsilon_{h}^{-1}{\mathcal{O}}_{\bar{E}}}\varepsilon_{h}^{-1}\bar{s}^{*}\omega^{1}_{\bar{X}/T}&\to&\varepsilon_{h}^{-1}\omega^{\bullet}_{\bar{E}/T}&\to&\varepsilon_{h}^{-1}\omega^{\bullet}_{\bar{E}/\bar{X}}&\to&0\\ &&\downarrow&&\downarrow&&\downarrow&&\\ 0&\to&\omega^{\bullet}_{\bar{E}^{\prime}/\bar{X}^{\prime}}\otimes_{{\mathcal{O}}_{\bar{E}^{\prime}}}\bar{s}^{\prime,*}\omega^{1}_{\bar{X}^{\prime}/T^{\prime}}&\to&\omega^{\bullet}_{\bar{E}^{\prime}/T^{\prime}}&\to&\omega^{\bullet}_{\bar{E}^{\prime}/\bar{X}^{\prime}}&\to&0,\end{array}$
where the top row is obtained by applying $\varepsilon_{h}^{-1}$ to (3), the bottom row is (3) with $N^{\prime}$ instead of NN, and the vertical arrows are given by the canonical $\varepsilon_{h}^{-1}{\mathcal{O}}_{\bar{E}}$-linear maps in each degree. Applying ${\bf R}^{\bullet}\bar{s}^{\prime}_{*}$ and using the natural transformation $\chi_{h}^{-1}\bar{s}_{*}\to\bar{s}^{\prime}_{*}\varepsilon_{h}^{-1}$ gives a commutative diagram
$\begin{array}[]{ccccc}\chi_{h}^{-1}{\bf R}^{1}\bar{s}_{*}\omega^{\bullet}_{\bar{E}/\bar{X}}&\to&{\bf R}^{1}\bar{s}^{\prime}_{*}(\varepsilon_{h}^{-1}\omega^{\bullet}_{\bar{E}/\bar{X}})&\to&{\bf R}^{1}\bar{s}^{\prime}_{*}\omega^{\bullet}_{\bar{E}^{\prime}/\bar{X}^{\prime}})\\ \downarrow&&\downarrow&&\downarrow\\ \hskip-4.26773pt\chi_{h}^{-1}{\bf R}^{2}\bar{s}_{*}(\omega^{\bullet}_{\bar{E}/\bar{X}}\otimes_{{\mathcal{O}}_{\bar{E}}}\bar{s}^{*}\omega^{1}_{\bar{X}/T})\hskip-8.5359pt&\to&\hskip-9.95863pt{\bf R}^{2}s^{\prime}_{*}(\varepsilon_{h}^{-1}(\omega^{\bullet}_{\bar{E}/\bar{X}}\otimes_{{\mathcal{O}}_{\bar{E}}}\bar{s}^{*}\omega^{1}_{\bar{X}/T}))\hskip-8.5359pt&\to&\hskip-9.95863pt{\bf R}^{2}s^{\prime}_{*}(\omega^{\bullet}_{\bar{E}^{\prime}/\bar{X}^{\prime}}\otimes_{{\mathcal{O}}_{\bar{E}^{\prime}}}\bar{s}^{\prime,*}\omega^{1}_{\bar{X}^{\prime}/T^{\prime}}).\end{array}$
The top row gives a $\chi_{h}^{-1}{\mathcal{O}}_{\bar{X}}$-linear morphism $\chi_{h}^{-1}{\mathcal{F}}_{\mathop{\rm dR}\nolimits}\to{\mathcal{F}}^{\prime}_{\mathop{\rm dR}\nolimits}$ compatible with filtrations, becoming the map ${\bf H}^{1}_{\mathop{\rm dR}\nolimits}(E_{x}/k)\to{\bf H}^{1}_{\mathop{\rm dR}\nolimits}(E^{\prime}_{x^{\prime}}/k)$ on a fiber $x^{\prime}:\operatorname{{\mathop{\rm Spec}\nolimits}}k\to X^{\prime}$ with $x=\chi_{h}\circ x^{\prime}$. Moreover the commutativity of the diagram together with the compatibility with projection formulas gives compatibility with Gauss-Manin connections, in the sense that the diagram
$\begin{array}[]{ccccccc}\chi_{h}^{-1}\nabla&:&\chi_{h}^{-1}{\mathcal{F}}_{\mathop{\rm dR}\nolimits}&\to&\chi_{h}^{-1}({\mathcal{F}}_{\mathop{\rm dR}\nolimits}\otimes_{{\mathcal{O}}_{\bar{X}}}\omega^{1}_{\bar{X}/T})&\cong&\chi_{h}^{-1}{\mathcal{F}}_{\mathop{\rm dR}\nolimits}\otimes_{\chi_{h}^{-1}{\mathcal{O}}_{\bar{X}}}\chi_{h}^{-1}\omega^{1}_{\bar{X}/T}\\ &&\downarrow&&&&\downarrow\\ \nabla^{\prime}&:&{\mathcal{F}}^{\prime}_{\mathop{\rm dR}\nolimits}&&\to&&{\mathcal{F}}^{\prime}_{\mathop{\rm dR}\nolimits}\otimes_{{\mathcal{O}}_{\bar{X}^{\prime}}}\omega^{1}_{\bar{X^{\prime}}/T^{\prime}}\end{array}$
commutes. This in turn gives morphisms $\chi_{h}^{-1}{\mathcal{F}}_{\mathop{\rm dR}\nolimits}^{k}\to{\mathcal{F}}_{\mathop{\rm dR}\nolimits}^{\prime,k}$ and $\chi_{h}^{-1}{\mathcal{F}}_{c,{\mathop{\rm dR}\nolimits}}^{k}\to{\mathcal{F}}_{c,{\mathop{\rm dR}\nolimits}}^{\prime,k}$ compatible with filtrations and connections, and taking cohomology yields the maps $h]_{\mathop{\rm dR}\nolimits}:{\mathcal{M}}_{\mathop{\rm dR}\nolimits}\to{\mathcal{M}}^{\prime}_{\mathop{\rm dR}\nolimits}$ and $h]_{c,{\mathop{\rm dR}\nolimits}}:{\mathcal{M}}_{c,{\mathop{\rm dR}\nolimits}}\to{\mathcal{M}}^{\prime}_{c,{\mathop{\rm dR}\nolimits}}$ preserving filtrations.
Suppose now that $\ell$ is a prime not dividing $N^{\prime}k!$. From the discussion of functoriality in [Fa, 4c] and the proof of [Fa, Thm. 6.2], we see that the above construction with $T$ and $T^{\prime}$ replaced by $\mathbb{Z}$$_{\ell}$ yields a morphism $\chi_{h,{\ell\mbox{\scriptsize-crys}}}^{*}{\mathcal{F}}_{\ell\mbox{\scriptsize-crys}}\to{\mathcal{F}}^{\prime}_{\ell\mbox{\scriptsize-crys}}$ in the category ${{\mathcal{M}\mathcal{F}}}^{\nabla}_{[0,1]}(X^{\prime}_{{\mathbb{Z}}_{\ell}})$. For $x^{\prime}:\operatorname{{\mathop{\rm Spec}\nolimits}}{\mathcal{O}}_{F}\to X^{\prime}$ and $x=\chi_{h}\circ x^{\prime}$ with $F$ unramified over $\mathbb{Q}$$_{\ell}$, the pull-back to ${{\mathcal{M}\mathcal{F}}}_{0}({\mathcal{O}}_{F})$ is the natural map $H^{1}_{\mbox{\scriptsize crys}}(E_{x},{\mathcal{O}}_{E_{x},{\mbox{\scriptsize crys}}})\to H^{1}_{\mbox{\scriptsize crys}}(E^{\prime}_{x^{\prime}},{\mathcal{O}}_{E^{\prime}_{x^{\prime}},{\mbox{\scriptsize crys}}})$. Taking symmetric powers and cohomology then yields the maps $h]_{\ell\mbox{\scriptsize-crys}}:{\mathcal{M}}_{\ell\mbox{\scriptsize-crys}}\to{\mathcal{M}}^{\prime}_{\ell\mbox{\scriptsize-crys}}$ and $h]_{c,{\ell\mbox{\scriptsize-crys}}}:{\mathcal{M}}_{\ell\mbox{\scriptsize-crys}}\to{\mathcal{M}}^{\prime}_{\ell\mbox{\scriptsize-crys}}$ in the category ${{\mathcal{M}\mathcal{F}}}_{0}({\mathbb{Z}}_{\ell})$.
Now we sketch the proof that the homomorphisms $h]_{*}$ and $h]_{c,*}$ for $\mathop{\rm B}\nolimits$}, $\mathop{\rm dR}\nolimits$}, $\ell$ and -$\ell\mbox{\scriptsize-crys}$} (for $\ell\not\in S$) are compatible with the comparison isomorphisms.
In the case of $I_{\mathop{\rm B}\nolimits}^{\ell}$, this follows from the functoriality properties of the Betti-étale comparison isomorphisms. In particular, the natural transformations $(f_{*}\cdot)_{\mathop{\rm an}\nolimits}\to f^{\mathop{\rm an}\nolimits}_{*}\cdot_{\mathop{\rm an}\nolimits}$ ([F-K, I.11]) are well-behaved under composition of maps $f:Y\to Z$ of schemes locally of finite type over $\mathbb{C}$}. One deduces that the diagram
$\begin{array}[]{ccccc}(\chi_{h}^{*}{\mathcal{F}}_{\ell}/\ell^{n})_{\mathop{\rm an}\nolimits}&\cong&(\chi^{\mathop{\rm an}\nolimits}_{h})^{*}({\mathcal{F}}_{\ell}/\ell^{n})_{\mathop{\rm an}\nolimits}&\cong&(\chi^{\mathop{\rm an}\nolimits}_{h})^{*}({\mathcal{F}}_{\mathop{\rm B}\nolimits}/\ell^{n})\\ \downarrow&&&&\downarrow\\ ({\mathcal{F}}_{\ell}^{\prime}/\ell^{n})_{\mathop{\rm an}\nolimits}&&\cong&&({\mathcal{F}}_{\mathop{\rm B}\nolimits}^{\prime}/\ell^{n})\end{array}$
commutes. (Though not necessary for the proof, we remark that on stalks at $x^{\prime}\in X^{\prime,{\mathop{\rm an}\nolimits}}$, the diagram becomes
$\begin{array}[]{ccc}H^{1}_{\mathop{\rm et}\nolimits}(E_{x},{\mathbb{Z}}/\ell^{n}{\mathbb{Z}})&\cong&H^{1}(E_{x}^{\mathop{\rm an}\nolimits},{\mathbb{Z}}/\ell^{n}{\mathbb{Z}})\\ \downarrow&&\downarrow\\ H^{1}_{\mathop{\rm et}\nolimits}(E^{\prime}_{x^{\prime}},{\mathbb{Z}}/\ell^{n}{\mathbb{Z}})&\cong&H^{1}(E_{x^{\prime}}^{\prime,{\mathop{\rm an}\nolimits}},{\mathbb{Z}}/\ell^{n}{\mathbb{Z}})\end{array}$
where $x=\chi_{h}^{\mathop{\rm an}\nolimits}(x^{\prime})$ and the vertical maps are induced by $E^{\prime}_{x^{\prime}}\to E_{x}$.) Applying $\operatorname{{\mathop{\rm Sym}\nolimits}}^{k-2}_{{\mathbb{Z}}_{\ell}}$ gives a similar commutative diagram for the sheaves $\mathcal{F}$$_{?}$ replaced by ${\mathcal{F}}_{?}^{k}$, and this gives commutativity of the right-hand square in the diagram
$\begin{array}[]{ccccc}{\mathcal{M}}_{\ell}&\to&H^{1}(X^{\prime}_{\bar{{\mathbb{Q}}}},\chi_{h}^{*}{\mathcal{F}}^{k}_{\ell})&\to&{\mathcal{M}}^{\prime}_{\ell}\\ \downarrow&&\downarrow&&\downarrow\\ {\mathbb{Z}}_{\ell}\otimes{\mathcal{M}}_{\mathop{\rm B}\nolimits}&\to&{\mathbb{Z}}_{\ell}\otimes H^{1}(X^{\prime,{\mathop{\rm an}\nolimits}},\chi_{h}^{{\mathop{\rm an}\nolimits},*}{\mathcal{F}}^{k}_{\mathop{\rm B}\nolimits})&\to&{\mathbb{Z}}_{\ell}\otimes{\mathcal{M}}^{\prime}_{\mathop{\rm B}\nolimits}\end{array}$
where the vertical maps are comparison isomorphisms and the horizontal maps define $h]_{\ell}$ and $h]_{\mathop{\rm B}\nolimits}$. The commutativity of the left-hand square is a similar, but easier, application of the functoriality of the comparison. This gives the compatibility of $h]_{?}$ with $I_{\mathop{\rm B}\nolimits}^{\ell}$, and $h]_{c,?}$ is treated similarly.
For the compatibility of $h]_{\mathop{\rm dR}\nolimits}$ and $h]_{\mathop{\rm B}\nolimits}$ with $I^{\infty}$, one first defines $h]_{\mathop{\rm dR}\nolimits}^{\mathop{\rm an}\nolimits}:{\mathcal{M}}_{\mathop{\rm dR}\nolimits}^{\mathop{\rm an}\nolimits}\to{\mathcal{M}}_{\mathop{\rm dR}\nolimits}^{\prime,{\mathop{\rm an}\nolimits}}$ analogously to $h]_{\mathop{\rm dR}\nolimits}$ and checks that the diagram
$\begin{array}[]{ccc}{\mathbb{C}}\otimes{\mathcal{M}}_{\mathop{\rm dR}\nolimits}&\cong&{\mathcal{M}}_{\mathop{\rm dR}\nolimits}^{\mathop{\rm an}\nolimits}\\ \downarrow&&\downarrow\\ {\mathbb{C}}\otimes{\mathcal{M}}_{\mathop{\rm dR}\nolimits}^{\prime}&\cong&{\mathcal{M}}_{\mathop{\rm dR}\nolimits}^{\prime,{\mathop{\rm an}\nolimits}}\end{array}$
commutes. Furthermore the diagram
$\begin{array}[]{ccc}(\chi_{h}^{\mathop{\rm an}\nolimits})^{-1}({\mathbb{C}}\otimes{\mathcal{F}}_{\mathop{\rm B}\nolimits})&\to&(\chi_{h}^{\mathop{\rm an}\nolimits})^{-1}{\mathcal{G}}\\ \downarrow&&\downarrow\\ {\mathbb{C}}\otimes{\mathcal{F}}_{\mathop{\rm B}\nolimits}^{\prime}&\to&{\mathcal{G}}^{\prime}\end{array}$
commutes, where the horizontal maps are those occuring in the definitions of $h]_{\mathop{\rm B}\nolimits}$ and $h]_{\mathop{\rm dR}\nolimits}^{\mathop{\rm an}\nolimits}$ and the rows arise from (4). (The commutativity follows for example from that of
$\begin{array}[]{ccc}H^{1}(E_{x}^{\mathop{\rm an}\nolimits},{\mathbb{C}})&\cong&{\bf H}^{1}(E_{x}^{\mathop{\rm an}\nolimits},\Omega^{\bullet}_{E_{x}^{\mathop{\rm an}\nolimits}})\\ \downarrow&&\downarrow\\ H^{1}(E_{x^{\prime}}^{\prime,{\mathop{\rm an}\nolimits}},{\mathbb{C}})&\to&{\bf H}^{1}(E_{x^{\prime}}^{\prime,{\mathop{\rm an}\nolimits}},\Omega^{\bullet}_{E_{x^{\prime}}^{\prime,{\mathop{\rm an}\nolimits}}})\end{array}$
at each $x^{\prime}\in X^{\prime,{\mathop{\rm an}\nolimits}}$, $x=\chi_{h}^{\mathop{\rm an}\nolimits}(x^{\prime})$.) One deduces from this the desired compatibility for $h]_{?}$ and $h]_{c,?}$.
The compatibility of $h]_{\mathop{\rm dR}\nolimits}$ and $h]_{\ell\mbox{\scriptsize-crys}}$ with $I^{\ell}_{\mathop{\rm dR}\nolimits}$ for $\ell\not\in S$ follows from the construction of $h]_{\ell\mbox{\scriptsize-crys}}$ based on [Fa], and the same goes for $h]_{c,{\mathop{\rm dR}\nolimits}}$ and $h]_{c,{\ell\mbox{\scriptsize-crys}}}$.
Finally, using the compatibility with pull-back of the natural transformations denoted $\alpha$ in [Fa, Va)], one gets commutativity of the diagram
$\begin{array}[]{ccccc}{\mathbb{V}}(\chi_{h,{\ell\mbox{\scriptsize-crys}}}^{*}{\mathcal{F}}_{\ell\mbox{\scriptsize-crys}})&\cong&\chi_{h}^{*}{\mathbb{V}}({\mathcal{F}}_{\ell\mbox{\scriptsize-crys}})&\cong&\chi_{h}^{*}{\mathcal{F}}_{\ell}\\ \downarrow&&&&\downarrow\\ {\mathbb{V}}({\mathcal{F}}_{\ell\mbox{\scriptsize-crys}}^{\prime})$$\cong&&{\mathcal{F}}^{\prime}_{\ell}\end{array}$
(which on the stalk $\bar{x}^{\prime}:\operatorname{{\mathop{\rm Spec}\nolimits}}\bar{F}\to X^{\prime}$ arising from $x^{\prime}:\operatorname{{\mathop{\rm Spec}\nolimits}}{\mathcal{O}}_{F}\to X^{\prime}$ for $F$ unramified over $\mathbb{Q}$$_{\ell}$ is just
$\begin{array}[]{ccc}{\mathbb{V}}(H^{1}_{\mbox{\scriptsize crys}}(E_{x},{\mathcal{O}}_{E_{x},{\mbox{\scriptsize crys}}}))&\cong&H^{1}_{\mathop{\rm et}\nolimits}(E_{\bar{x}},{\mathbb{Z}}_{\ell})\\ \downarrow&&\downarrow\\ {\mathbb{V}}(H^{1}_{\mbox{\scriptsize crys}}(E^{\prime}_{x^{\prime}},{\mathcal{O}}_{E^{\prime}_{x^{\prime}},{\mbox{\scriptsize crys}}}))&\cong&H^{1}_{\mathop{\rm et}\nolimits}(E^{\prime}_{\bar{x^{\prime}}},{\mathbb{Z}}_{\ell})\end{array}$
where $x=\chi_{h}\circ x^{\prime}$). This in turn gives the commutativity of the right-hand square in
$\begin{array}[]{ccccc}{\mathbb{V}}({\mathcal{M}}_{\ell\mbox{\scriptsize-crys}})&\to&{\mathbb{V}}(H^{1}_{\mbox{\scriptsize crys}}(X^{\prime}_{{\mathbb{Z}}_{\ell}},\chi_{h,{\ell\mbox{\scriptsize-crys}}}^{*}{\mathcal{F}}^{k}_{\ell\mbox{\scriptsize-crys}}))&\to&{\mathbb{V}}({\mathcal{M}}^{\prime}_{\ell\mbox{\scriptsize-crys}})\\ \downarrow&&\downarrow&&\downarrow\\ {\mathcal{M}}_{\ell}&\to&H^{1}_{\mathop{\rm et}\nolimits}(X^{\prime}_{\bar{{\mathbb{Q}}}},\chi_{h}^{*}{\mathcal{F}}^{k}_{\ell})$$\to&{\mathcal{M}}^{\prime}_{\ell}\end{array}$
and the commutativity of the left-hand square is also a consequence of the compatibility of $\alpha$ with pull-back. This gives the compatibility of $h]_{\ell\mbox{\scriptsize-crys}}$ and $h]_{\ell}$ with $I^{\ell}$, and the proof is similar for $h]_{c,?}$.
For each realization ${\mathop{\rm B}\nolimits},{\mathop{\rm dR}\nolimits},\ell$ and -$\ell\mbox{\scriptsize-crys}$}, we have a commutative diagram
[…]
This shows that $h]_{?}$ and $h]_{c,?}$ respect weight filtrations and induce maps $h]_{!,?}:{\mathcal{M}}_{!,?}\to{\mathcal{M}}^{\prime}_{!,?}$. Since $h]_{?}$, $h]_{?,c}$ (and hence $h]_{?,!}$) are compatible with the comparison isomorphisms $I^{\infty}$, $I_{\mathop{\rm B}\nolimits}^{\ell}$, $I_{\mathop{\rm dR}\nolimits}^{\ell}$ and $I^{\ell}$ when $\ell\nmid N^{\prime}k!$, we conclude that $h$ defines morphisms $h]:{\mathcal{M}}^{S^{\prime}}\to{\mathcal{M}}^{\prime}$, $h]_{c}:{\mathcal{M}}_{c}^{S^{\prime}}\to{\mathcal{M}}^{\prime}_{c}$ and $h]_{!}:{\mathcal{M}}_{!}^{S^{\prime}}\to{\mathcal{M}}^{\prime}_{!}$ in ${\mathcal{P}M}^{S^{\prime}}_{\mathbb{Q}}$ where $S^{\prime}=\{\,p\,\mid\,p\nmid N^{\prime}k!\,\}$.
Lemma 3.4.
Suppose that $(h,N,N^{\prime})$ and $(h^{\prime},N^{\prime},N^{\prime\prime})$ are admissible triples. Then
$h^{\prime}]_{N^{\prime},N^{\prime\prime}}\circ[h]_{N,N^{\prime}}=[h^{\prime}h]_{N,N^{\prime\prime}}$
on ${\mathcal{M}}^{S},{\mathcal{M}}^{S}_{c}$ and ${\mathcal{M}}^{S}_{!}$ whenever $S\supseteq S_{N^{\prime\prime}}$.
Proof.
We only need to prove $h^{\prime}]_{B}\circ[h]_{B}=[h^{\prime}h]_{B}$. Using Lemma 3.3, we get the commutative diagram
[…]
which is just
[…]
on the stalk at $x^{\prime\prime}\in X^{\prime\prime,an}$ (where $x^{\prime}=\chi_{h^{\prime}}^{\mathop{\rm an}\nolimits}(x^{\prime\prime})$ and $x=\chi_{h}^{\mathop{\rm an}\nolimits}(x^{\prime})$).
Taking symmetric powers and cohomology, we get the desired commutative diagram. ∎
Note that if $r\in{\mathbb{Z}}$ and $(rI,N,N^{\prime})$ is an admissible triple, then $rN|N^{\prime}$ and $(I,N,N^{\prime})$ is also admissible.
Lemma 3.5.
If $r\in{\mathbb{Z}}$ and $(rI,N,N^{\prime})$ is an admissible triple, then $rI]=r^{k-2}[I]$, $rI]_{c}=r^{k-2}[I]_{c}$ and $rI]_{!}=r^{k-2}[I]_{!}$.
Proof.
It suffices to check the lemma on Betti realizations. Applying Lemma 3.2 to $h=rI$ and $h=I$, we see that $\chi_{rI}^{\mathop{\rm an}\nolimits}=\chi_{I}^{\mathop{\rm an}\nolimits}$ on $X^{\prime,{\mathop{\rm an}\nolimits}}$ and $\varepsilon_{rI}^{\mathop{\rm an}\nolimits}=\varepsilon_{I}^{\mathop{\rm an}\nolimits}r$ on $E^{\prime,{\mathop{\rm an}\nolimits}}$. Since multiplication by $r$ on the fiber $E^{\mathop{\rm an}\nolimits}_{x}$ over $x\in X^{\mathop{\rm an}\nolimits}$ induces multiplication by $r$ on ${\mathcal{F}}_{B,x}=H^{1}(E^{\mathop{\rm an}\nolimits}_{x},{\mathbb{Z}})$, it follows that the map $(\chi^{\mathop{\rm an}\nolimits})^{*}{\mathcal{F}}_{B}\to{\mathcal{F}}_{B}^{\prime}$ arising from $(rI,N,N^{\prime})$ is $r$ times the map arising from $(I,N,N^{\prime})$ (writing simply $\chi^{\mathop{\rm an}\nolimits}$ for $\chi_{rI}^{\mathop{\rm an}\nolimits}=\chi_{I}^{\mathop{\rm an}\nolimits}$). The same then holds for the maps $(\chi^{\mathop{\rm an}\nolimits})^{*}{\mathcal{F}}_{B}^{k}\to{\mathcal{F}}_{B}^{\prime,k}$ with $r$ replaced by $r^{k-2}$, and the lemma follows. ∎
[…]
$r^{k-2}[h]_{N,N^{\prime}}=r^{k-2}[r_{0}r^{-1}I]_{N^{\prime\prime},N^{\prime}}[h_{0}]_{N,N^{\prime\prime}}=r_{0}^{k-2}[I]_{N^{\prime\prime},N^{\prime}}[h_{0}]_{N,N^{\prime\prime}}=r_{0}^{k-2}[h_{0}]_{N,N^{\prime}}$
by Lemmas 3.4 and 3.5. Thus we obtain a morphism $g]=[g]_{N,N^{\prime}}:{\mathcal{M}}^{S^{\prime}}\to{\mathcal{M}}^{\prime}$ by defining
$g]=r^{k-2}[h]:{\mathcal{M}}^{S^{\prime}}\to{\mathcal{M}}^{\prime}$
which is independent of the factorization $g=rh$ with $r\in{\mathbb{Z}}$ and $(h,N,N^{\prime})$ admissible. We also define $g]_{c}=r^{k-2}[h]_{c}$ and $g]_{!}=r^{k-2}[h]_{!}$.
Similarly if $g\in\operatorname{{\mathop{\rm GL}\nolimits}}_{2}({\mathbb{A}}_{\bf f})$ with $g^{-1}U_{N^{\prime}}g\subseteq U_{N}$, we can write $g=rh$ for some $r\in{\mathbb{Q}}$ so that $(h,N,N^{\prime})$ is admissible and obtain morphisms $[g]$, $g]_{c}$ and $g]_{!}$ in $\operatorname{{\bf PM}}_{\mathbb{Q}}^{S^{\prime}}$.
The properties of the action of $g$ on premotivic structures can be summarized in the following theorem.
Theorem 3.6.
Fix positive integers $k\geq 2$, and NN, $N^{\prime}\geq 3$. Let $g$ be an element in $\operatorname{{\mathop{\rm GL}\nolimits}}_{2}({\mathbb{A}}_{\bf f})$ such that $g^{-1}U_{N^{\prime}}g\subseteq U_{N}$.
  1. (a)
    Let $S^{\prime}=S_{N^{\prime}}$. Then $g]=[g]_{N,N^{\prime}}:M^{S^{\prime}}\to M^{\prime}$ is a morphism in $\operatorname{{\bf PM}}^{S^{\prime}}_{\mathbb{Q}}$.
  2. (b)
    If $N^{\prime\prime}\geq 3$ is an integer and $g^{\prime}\in\operatorname{{\mathop{\rm GL}\nolimits}}_{2}({\mathbb{A}}_{\bf f})$ is such that $g^{\prime,-1}U_{N^{\prime\prime}}g^{\prime}\subseteq U_{N}^{\prime}$, then $g^{\prime}]_{N^{\prime\prime},N^{\prime}}\circ[g]_{N^{\prime},N}^{S_{N^{\prime\prime}}}=[g^{\prime}g]_{N^{\prime\prime},N}$.
  3. (c)
[…]
  4. (d)
    Analogous assertions hold for $g]_{c}$ and $g]_{!}$.
3.4. Relation with the action on modular forms
Given $k\geq 2$, we have established in §2.3 an isomorphism
$\alpha:{\mathbb{C}}\otimes\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}{\mathcal{M}}_{{\mathop{\rm dR}\nolimits}}\cong\bigoplus_{t\in({\mathbb{Z}}/N{\mathbb{Z}})^{\times}}M_{k}(\Gamma(N)).$
In (12) we have also described an isomorphism
$\beta:M_{k}(U_{N})\to\bigoplus_{t\in({\mathbb{Z}}/N{\mathbb{Z}})^{\times}}M_{k}(\Gamma(N)).$
Both $\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}M_{N,{\mathop{\rm dR}\nolimits}}\otimes{\mathbb{C}}$ and $M_{k}(U_{N})$ have an action by elements from $M_{2}(\hat{{\mathbb{Z}}})\cap\operatorname{{\mathop{\rm GL}\nolimits}}_{2}({\mathbb{A}}_{\bf f})$.
Proposition 3.7.
The canonical isomorphism
$\beta^{-1}\circ\alpha:\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}{\mathcal{M}}_{N,{\mathop{\rm dR}\nolimits}}\otimes{\mathbb{C}}\to M_{k}(U(N))$
preserves the action of $M_{2}(\hat{{\mathbb{Z}}})\cap\operatorname{{\mathop{\rm GL}\nolimits}}_{2}({\mathbb{A}}_{\bf f})$. The same holds for the isomorphism
$\beta^{-1}\circ\alpha:\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}M_{N,c,{\mathop{\rm dR}\nolimits}}\otimes{\mathbb{C}}\to S_{k}(U(N)).$
Proof.
Fix a $g\in M_{2}(\hat{{\mathbb{Z}}})\cap\operatorname{{\mathop{\rm GL}\nolimits}}_{2}({\mathbb{A}}_{\bf f})$ with $g^{-1}U^{\prime}g\subset U$, we will show that the actions of $g$ on $M_{k}(\Gamma(N))$ induced from the isomorphisms $\alpha$ and $\beta$ are the same. First assume that $(h,N,N^{\prime})$ is an admissible triple.
Let
$\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}{\mathcal{M}}_{{\mathop{\rm dR}\nolimits}}\otimes{\mathbb{C}}\cong H^{0}(\bar{X}^{\mathop{\rm an}\nolimits},{\mathcal{L}}^{\mathop{\rm an}\nolimits})$
$\cong\bigoplus_{t\in({\mathbb{Z}}/N{\mathbb{Z}})^{\times}}H^{0}(\bar{X}_{N,t},i_{t}^{*}{\mathcal{L}}^{\mathop{\rm an}\nolimits})\hookrightarrow\bigoplus_{t\in({\mathbb{Z}}/N{\mathbb{Z}})^{\times}}H^{0}({\mathfrak{H}},\pi_{t}^{*}i_{t}^{*}{\mathcal{L}}^{\mathop{\rm an}\nolimits})$
be the injective map define in § 2.3. Identifying $\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}{\mathcal{M}}_{{\mathop{\rm dR}\nolimits}}\otimes{\mathbb{C}}$ with its image under this map, then the isomorphism $\alpha$ is given by
$f(\tau)(2\pi i)^{k-1}(dz)^{\otimes(k-2)}\otimes d\tau\mapsto f(\tau).$
The restriction of $\alpha$ to $\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}{\mathcal{M}}_{c,{\mathop{\rm dR}\nolimits}}\otimes{\mathbb{C}}$ defines the isomorphism
${\mathbb{C}}\otimes\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}{\mathcal{M}}_{c,{\mathop{\rm dR}\nolimits}}\cong\bigoplus_{t\in({\mathbb{Z}}/N{\mathbb{Z}})^{\times}}S_{k}(\Gamma(N)).$
For $t^{\prime}\in({\mathbb{Z}}/N^{\prime}{\mathbb{Z}})^{\times}$, let $\kappa(t^{\prime})\in({\mathbb{Z}}/N{\mathbb{Z}})^{\times}$ and $\gamma=\gamma_{h,t^{\prime}}={\mbox{\scriptsize$\left(\begin{array}[]{cc}{{a}}&{{b}}\\ {{c}}&{{d}}\end{array}\right)$}}\in\operatorname{{\mathop{\rm GL}\nolimits}}_{2}^{+}({\mathbb{Q}})$ be associated to $(h,N,N^{\prime})$ in Lemma 3.2. Then the action on
$\bigoplus_{t\in({\mathbb{Z}}/N{\mathbb{Z}})^{\times}}H^{0}({\mathfrak{H}},\pi_{t}^{*}i_{t}^{*}{\mathcal{L}}^{\mathop{\rm an}\nolimits})$
induced by $h$ sends $\theta\stackrel{{\scriptstyle\rm def}}{{=}}((2\pi i)^{k-1}f_{t}(\tau)(dz)^{\otimes(k-2)}\otimes d\tau)_{t}$ to
$((2\pi i)^{k-1}f_{\kappa(t^{\prime})}(\gamma_{t^{\prime}}(\tau))d((c\tau+d)^{-1}z)^{\otimes(k-2)}\otimes d\gamma(\tau))_{t^{\prime}}$
$=((2\pi i)^{k-1}(c\tau+d)^{-k}{\mathop{\rm det}\nolimits}\gamma f_{\kappa(t^{\prime})}(\gamma(\tau))dz^{\otimes(k-2)}\otimes d\tau)_{t^{\prime}}.$
This means
(14) $\alpha(g\theta)_{t^{\prime}}(\tau)=(c\tau+d)^{-k}{\mathop{\rm det}\nolimits}\gamma\alpha(\theta)_{\kappa(t^{\prime})}(\gamma(\tau)).$
On the other hand, for $\phi\in M_{k}(U_{N})$, its image $\beta(\phi)\in\oplus_{t\in({\mathbb{Z}}/N{\mathbb{Z}})^{\times}}M_{k}(\Gamma(N))$ is given by
$\beta(\phi)_{t}(x(i))=({\mathop{\rm det}\nolimits}x)^{-1}(c^{\prime}i+d^{\prime})^{k}\phi(g_{t}x),\quad x={\mbox{\scriptsize$\left(\begin{array}[]{cc}{{a^{\prime}}}&{{b^{\prime}}}\\ {{c^{\prime}}}&{{d^{\prime}}}\end{array}\right)$}}\in\operatorname{{\mathop{\rm GL}\nolimits}}_{2}^{+}({\mathbb{R}}).$
Similarly, the image of $h\phi$ is given by
$\beta(h\phi)_{t^{\prime}}(x(i))=({\mathop{\rm det}\nolimits}x)^{-1}(c^{\prime}i+d^{\prime})^{k}(h\phi)(g_{t^{\prime}}x)=({\mathop{\rm det}\nolimits}x)^{-1}(c^{\prime}i+d^{\prime})^{k}\phi(g_{t^{\prime}}xh).$
Using Lemma 3.2 and properties of $\phi$, we verify that
$({\mathop{\rm det}\nolimits}x)^{-1}(c^{\prime}i+d^{\prime})^{k}\phi(g_{t^{\prime}}xh)=\beta(\phi)_{\kappa(t^{\prime})}(x\gamma(i)){\mathop{\rm det}\nolimits}\gamma(cx(i)+d)^{-k}.$
Writing $\tau=x(i)$ for the right action of $x$ on $i$, we have
(15) $\beta(h\phi)_{t^{\prime}}(\tau)=(c\tau+d)^{-k}{\mathop{\rm det}\nolimits}\gamma\beta(\phi)_{\kappa(t^{\prime})}(\gamma(\tau)).$
Combining (14) and (15) we see that if $(\beta^{-1}\circ\alpha)(\theta)=\phi$, then $(\beta^{-1}\circ\alpha)(h\theta)=h\phi$.
Further let $r\in{\mathbb{Z}}$. Then the action of rIrI on both spaces are multiplication by $r^{k-2}$. This proves the proposition. ∎
3.5. Compatibility with the pairings
Proposition 3.8.
Suppose $S$ contains $S_{N}$. Let
$\delta^{L}_{N}:{\mathcal{M}}_{N,c}^{S}\to{\mathop{\rm Hom}\nolimits}({\mathcal{M}}_{N}^{S},{\mathcal{T}}^{\otimes(k-1),S})$
and
$\delta^{L}_{N^{\prime}}:{\mathcal{M}}_{N^{\prime},c}^{S}\to{\mathop{\rm Hom}\nolimits}({\mathcal{M}}_{N^{\prime}}^{S},{\mathcal{T}}^{\otimes(k-1),S})$
be the duality morphisms defined in §2.4. Let $g]^{*}:{\mathop{\rm Hom}\nolimits}({\mathcal{M}}_{N^{\prime}},{\mathcal{T}}^{\otimes(k-1),S})\to{\mathop{\rm Hom}\nolimits}({\mathcal{M}}_{N^{\prime}},{\mathcal{T}}^{\otimes(k-1),S})$ be induced from $[g]$. Then
$\delta^{L}_{N}=||{\mathop{\rm det}\nolimits}g||^{2-k}[U_{N}:U_{N^{\prime}}](g^{*}\circ\delta^{L}_{N^{\prime}}\circ g)$
in ${\mathcal{P}M}^{S}_{\mathbb{Q}}$. A similar relation holds between the duality morphism ${\mathcal{M}}_{N,!}$ and the duality morphism on ${\mathcal{M}}_{N^{\prime},!}$
Proof.
We only need to show the equation on the Betti realization. Let
${(}\ ,\ {)}:{\mathcal{M}}_{c,N}\otimes{\mathcal{M}}_{N}\to{\mathcal{T}}^{\otimes(k-1),S}_{\mathop{\rm B}\nolimits}$
and
${(}\ ,\ {)}^{\prime}:{\mathcal{M}}_{c,N^{\prime}}\otimes{\mathcal{M}}_{N^{\prime}}\to{\mathcal{T}}^{\otimes(k-1),S}_{\mathop{\rm B}\nolimits}$
be the pairings associated to $\delta^{L}_{N}$ and $\delta^{L}_{N^{\prime}}$. We only need to showing that, for $x\in{\mathcal{M}}_{N,c,{\mathop{\rm B}\nolimits}}$ and $y\in{\mathcal{M}}_{N,{\mathop{\rm B}\nolimits}}$, we have
(16) ${(}gx,gy{)}^{\prime}=||{\mathop{\rm det}\nolimits}g||^{2-k}[U_{N}:U_{N^{\prime}}]{(}x,y{)}$
in ${\mathcal{T}}^{\otimes(k-1),S}_{\mathop{\rm B}\nolimits}$.
Write $g=rh$ as before with $r\in{\mathbb{Z}}$ and $(h,N,N^{\prime})$ admissible. Assume that the lemma holds for $h$. Then
${(}gx,gy{)}^{\prime}$ $={(}r^{k-2}hx,r^{k-2}hy{)}^{\prime}=r^{2(k-2)}{(}hx,hy{)}^{\prime}$
$=||{\mathop{\rm det}\nolimits}(rI)||^{2-k}{(}hx,hy{)}^{\prime}=||{\mathop{\rm det}\nolimits}g||^{2-k}[U_{N}:U_{N^{\prime}}]{(}x,y{)}.$
Thus we only need to prove (16) when $(g,N,N^{\prime})$ is admissible.
Using the functoriality of cup products [Br, II.8.2], we have the commutative diagram
[…]
where the first arrow in the top row is the composite
$\chi^{*}{\mathcal{F}}_{B}\otimes\chi^{*}{\mathcal{F}}_{B}\to\chi^{*}({\mathcal{F}}_{B}\otimes{\mathcal{F}}_{B})\stackrel{{\scriptstyle\cup}}{{\longrightarrow}}\chi^{*}R^{2}s_{*}^{{\mathop{\rm an}\nolimits}}{\mathbb{Z}}_{E}$
and the right column is the natural map induced by $h$. Taking the stalks over $x^{\prime}\in X^{\prime,{\mathop{\rm an}\nolimits}}$, the diagram gives
[…]
Next note that the map $\chi:X^{\prime,{\mathop{\rm an}\nolimits}}\to X^{\mathop{\rm an}\nolimits}$, identified with the map
$G_{\mathbb{Q}}\backslash G_{\mathbb{A}}/U^{\prime}U_{\infty}\cong G_{\mathbb{Q}}\backslash G_{\mathbb{A}}/h^{-1}U^{\prime}hU_{\infty}\to G_{\mathbb{Q}}\backslash G_{\mathbb{A}}/UU_{\infty}$
is a covering map of degree $U:h^{-1}U^{\prime}h]=[U:U^{\prime}]$. This gives us the commutative diagram
[…]
Then by functoriality of cup products of cohomology and the Universal Coefficient Theorem [Br, II.15.3] (see also [Ha, Exer. III.8.3]), we get the commutative diagram
[…]
where $H^{2}_{c}(\chi)$ is the composition of the top row in diagram (17). This is what we want. ∎
3.6. Compatibility with changing of levels
Fix $g\in M_{2}({\mathbb{Z}})\cap\operatorname{{\mathop{\rm GL}\nolimits}}_{2}({\mathbb{A}}_{\bf f})$. Let $N,N^{\prime},M$ and $M^{\prime}$ be integers such that $N|M,N^{\prime}|M^{\prime},g^{-1}U_{N^{\prime}}g\subset U_{N}$ and $g^{-1}U_{M^{\prime}}g\subset U_{M}$. Then we have $g]:{\mathcal{M}}_{N}\to{\mathcal{M}}_{N^{\prime}}$ and $g]:{\mathcal{M}}_{M}\to{\mathcal{M}}_{M^{\prime}}$. Let II be the identity matrix in $\operatorname{{\mathop{\rm GL}\nolimits}}_{2}({\mathbb{A}}_{\bf f})$. From $N|M$ we get $I^{-1}U_{M}I\subset U_{N}$ and we have
$I:\bar{E}_{M}{}_{/\bar{X}_{M}}\to\bar{E}_{N}{}_{/\bar{X}_{N}}$
and
$I]:{\mathcal{M}}_{N}\to{\mathcal{M}}_{M}\ {\rm\ for\ }S\supset S_{M}.$
In the same way, we have
$I]:{\mathcal{M}}_{N^{\prime}}\to{\mathcal{M}}_{M^{\prime}}\ {\rm\ for\ }S\supset S_{M^{\prime}}.$
Lemma 3.9.
  1. (a)
    For $S\supseteq S_{M^{\prime}}$, the diagram
[…]
    of morphisms in ${\mathcal{P}M}^{S}$ is commutative.
  2. (b)
    $I]:{\mathcal{M}}_{N,{\mathop{\rm tf}\nolimits}}^{S_{M}}\to{\mathcal{M}}_{M,{\mathop{\rm tf}\nolimits}}^{U_{N}/U_{M}}$ is injective
  3. (c)
    $I]_{\sharp}:{\mathcal{M}}_{N,\sharp}^{S_{M}}\to{\mathcal{M}}_{M,\sharp}^{U_{N}/U_{M}}$ is an isomorphism for $\sharp=c$ or !!.
  4. (d)
    $I]:M_{N}^{S_{M}}\to M_{M}^{U_{N}/U_{M}}$ is an isomorphism.
Proof.
(a) follows directly from Theorem 3.6.
(b). We only need to consider $I]_{B}$. Denote $s_{M}^{\mathop{\rm an}\nolimits}:E_{M}^{\mathop{\rm an}\nolimits}\to X_{M}^{\mathop{\rm an}\nolimits}$ for the universal elliptic curve and denote ${\mathcal{F}}_{M,B}$ for the locally constant sheaf $R^{1}s^{\mathop{\rm an}\nolimits}_{M,*}{\mathbb{Z}}$ on $X_{M}^{\mathop{\rm an}\nolimits}$. Leray spectral sequence gives us
(18) $0\to H^{1}(U_{N}/U_{M},H^{0}(X_{M},{\mathcal{F}}_{M,B}^{k}))\to H^{1}(X_{N},{\mathcal{F}}^{k}_{N,B})\\ \to H^{1}(X_{M},{\mathcal{F}}_{M,B}^{k})^{U_{N}/U_{M}}\to H^{2}(U_{N}/U_{M},H^{0}(X_{M},{\mathcal{F}}_{M,B}^{k})).$
Note that the map $H^{1}(X_{N},{\mathcal{F}}^{k}_{N,B})\to H^{1}(X_{M},{\mathcal{F}}_{M,B}^{k})$ is induced by the projection $p:\bar{E}_{M}{}_{/\bar{X}_{M}}\to\bar{E}_{/\bar{X}}$. Since $U_{N}/U_{M}$ is finite and $H^{0}(X_{M},{\mathcal{F}}_{M,B}^{k})$ is torsion free, we have
$H^{1}(U_{N}/U_{M},H^{0}(X_{M},{\mathcal{F}}_{M,B}^{k}))={\mathop{\rm Hom}\nolimits}(U_{N}/U_{M},H^{0}(X_{M},{\mathcal{F}}_{M,B}^{k}))=0.$
Therefore ${\mathcal{M}}_{N,{\mathop{\rm B}\nolimits}}=H^{1}(X_{N},{\mathcal{F}}^{k}_{N,B})\to H^{1}(X_{M},{\mathcal{F}}_{M,B}^{k})^{U_{N}/U_{M}}=({\mathcal{M}}_{M,{\mathop{\rm B}\nolimits}})^{U_{N}/U_{M}}$ is injective.
(c). The proof is similar to (b). We only need to consider the Betti realization. We know that Leray spectral sequence[Br, IV.9.2] gives an exact sequence
(19) $0\to H^{1}(U_{N}/U_{M},H^{0}_{c}(X_{M},{\mathcal{F}}_{M,B}^{k}))\to H^{1}_{c}(X_{N},{\mathcal{F}}^{k}_{N,B})\\ \to H^{1}_{c}(X_{M},{\mathcal{F}}_{M,B}^{k})^{U_{N}/U_{M}}\to H^{2}(U_{N}/U_{M},H^{0}_{c}(X_{M},{\mathcal{F}}_{M,B}^{k}))$
and $H^{0}_{c}(X_{M},{\mathcal{F}}_{M,B}^{k})=0$ since ${\mathcal{F}}_{M,B}^{k}$ have no non-trivial global sections with compact support. Thus ${\mathcal{M}}_{N,B,c}$ is identified with $({\mathcal{M}}_{M,B,c})^{U_{N}/U_{M}}$.
(d). Since $U_{N}/U_{M}$ is finite, $H^{2}(U_{N}/U_{M},H^{0}(X_{M},{\mathcal{F}}_{M,B}^{k}))$ is torsion. Thus from the proof of (b) we see that $H^{1}(X_{N},{\mathcal{F}}^{k}_{N,B})\otimes{\mathbb{Q}}\to H^{1}(X_{M},{\mathcal{F}}_{M,B}^{k})^{U_{N}/U_{M}}\otimes{\mathbb{Q}}$ is an isomorphism. This prove (d) for the Betti realization which is all we need. ∎
4. The premotivic structure for forms of level NN and character $\psi$
4.1. $\sigma$-constructions
[…]
$S_{\sigma}=\{\lambda\in S_{\bf f}(K)|\ \lambda|k!\mbox{\ or\ }\operatorname{{\mathop{\rm GL}\nolimits}}_{2}({\mathbb{Z}}_{\ell})\not\subset\ker\sigma\ {\rm\ where\ }\lambda|\ell\}.$
Since $U_{N}$ is normal in $U$, by Theorem 3.6, we have a group action of $U$ on ${\mathcal{M}}_{N}$. We regard $\mathcal{V}$} as an object of ${\mathcal{P}M}_{K}^{S_{\sigma}}$ (with trivial additional structures and tautological comparison isomorphisms). Note that we have an action of $U$ on $\mathcal{V}$}, hence on the object ${\mathcal{M}}_{N}\otimes{\mathcal{V}}=({\mathcal{M}}_{N}\otimes{\mathcal{O}}_{K})\otimes_{{\mathcal{O}}_{K}}{\mathcal{V}}$ of ${\mathcal{P}M}_{K}^{S_{\sigma}}$. We can thus define an object
${\mathcal{M}}(\sigma)_{N}=({\mathcal{M}}_{N}\otimes{\mathcal{V}})^{U}$
of ${\mathcal{P}M}_{K}^{S_{\sigma}}$ (see the last paragraph of §1.2). More explicitly, for each $B,{\mathop{\rm dR}\nolimits}$, $\lambda\in S_{\bf f}(K)$ or $\lambda$-crys with $\lambda\in S_{\sigma}$, we have
${\mathcal{M}}(\sigma)_{N,?}=({\mathcal{M}}_{N,?}\otimes{\mathcal{V}})^{U}.$
We also define objects
${\mathcal{M}}(\sigma)_{N,\sharp}=({\mathcal{M}}_{N,\sharp}\otimes{\mathcal{V}})^{U}$
for $\sharp={\mathop{\rm tf}\nolimits}$ or $c$ and define
${\mathcal{M}}(\sigma)_{N,!}=\operatorname{{\mathop{\rm im}\nolimits}}({\mathcal{M}}(\sigma)_{N,c}\to{\mathcal{M}}(\sigma)_{N,{\mathop{\rm tf}\nolimits}})$
where ${\mathcal{M}}(\sigma)_{N,c}\to{\mathcal{M}}(\sigma)_{N}$ is the restriction of the map ${\mathcal{M}}_{N,c}\otimes{\mathcal{V}}\to{\mathcal{M}}_{N}\otimes{\mathcal{V}}$. These premotivic structures will be called the $\sigma$-constructions.
We now show that the $\sigma$-constructions are compatible with the change of levels. More precisely,
Lemma 4.1.
Let $N|N^{\prime}$ with $U_{N}\subset U$.
  1. (a)
    For $S\supseteq S_{N^{\prime}}$, we have ${\mathcal{M}}(\sigma)_{N,c}^{S}\cong{\mathcal{M}}(\sigma)_{N^{\prime},c}^{S}$ in ${\mathcal{P}M}^{S}$. Similarly, ${\mathcal{M}}(\sigma)_{N,!}^{S}\cong{\mathcal{M}}(\sigma)_{N^{\prime},!}^{S}$.
  2. (b)
    For $S\supseteq S_{N^{\prime}}$, we have $M(\sigma)_{N}^{S}\cong M(\sigma)_{N^{\prime}}^{S}$.
Proof.
(a) We only need to prove it for $S=S_{N^{\prime}}$. From $N\mid N^{\prime}$ we have $U_{N^{\prime}}\subset U_{N}$. By Lemma 3.9, we have
${\mathcal{M}}_{N,c}^{S}=({\mathcal{M}}_{N^{\prime},c}^{U_{N}/U_{N^{\prime}}})^{S}.$
Then since $U_{N}$ acts trivially on $\mathcal{V}$}, we have
$({\mathcal{M}}_{N^{\prime},c}^{S}\otimes{\mathcal{V}})^{U_{N}/U_{N^{\prime}}}=({\mathcal{M}}_{N^{\prime},c}^{U_{N}/U_{N^{\prime}}})^{S}\otimes{\mathcal{V}}={\mathcal{M}}_{N,c}^{S}\otimes{\mathcal{V}}.$
Hence
${\mathcal{M}}(\sigma)_{N^{\prime},c}^{S}=({\mathcal{M}}_{N^{\prime},c}^{S}\otimes{\mathcal{V}})^{U}=(({\mathcal{M}}_{N^{\prime},c}^{S}\otimes{\mathcal{V}})^{U_{N}/U_{N^{\prime}}})^{U}=({\mathcal{M}}_{N,c}^{S}\otimes{\mathcal{V}})^{U}={\mathcal{M}}(\sigma)_{N,c}^{S}.$
Since $\mathcal{V}$} is a flat $\mathbb{Z}$}-module, using Lemma 3.9, we similarly find that ${\mathcal{M}}(\sigma)_{N,{\mathop{\rm tf}\nolimits}}^{S}\to{\mathcal{M}}(\sigma)_{N^{\prime},{\mathop{\rm tf}\nolimits}}^{S}$ is injective. Consider the commutative diagram
[…]
By the definition of ${\mathcal{M}}(\sigma)_{N,!}$, the two vertical maps in the upper square are surjective and the two vertical maps in the bottom square are injective. Since the top row is surjective, so is the map in the middle row. Since the bottom row is injective, the map in the middle row is also injective, hence is bijective.
The proof for (b) is the same. ∎
Because of the lemma, we will suppress NN in the notations by using ${\mathcal{M}}(\sigma)_{c}$, ${\mathcal{M}}(\sigma)_{!}$ and $M(\sigma)$ when there is no danger of confusion.
4.2. Action of double cosets
[…]
  • $N\geq 3$,
  • $U_{N}\subset\ker\sigma\cap\ker\sigma^{\prime}$,
  • $U_{N^{\prime}}\subset gU_{N}g^{-1}$ and
  • $S_{N^{\prime}}^{K}\subset S$.
To see that this is possible, first choose NN divisible by 44 so that $S_{N}^{K}\subset S$ and the second condition holds, then choose $N^{\prime}$ so that the third condition holds, and then for each $\ell\not\in S$, replace $N^{\prime}$ by $N^{\prime}/\ell^{r}$ where $\ell^{r}||N^{\prime}$. The condition that $S_{N}^{K}\subset S$ and $g_{\ell}\in\operatorname{{\mathop{\rm GL}\nolimits}}_{2}({\mathbb{Z}}_{\ell})$ ensures that we still have $g^{-1}U_{N^{\prime}}g\subset U_{N}.$ Note that $N|N^{\prime}$ so $N^{\prime}\geq 3$, $U_{N^{\prime}}\subset U_{N}\subset\ker\sigma\cap\ker\sigma^{\prime}$ and $S_{N}^{K}\subset S_{N^{\prime}}^{K}\subset S$. Then from the last section, $g$ gives a map $g]:{\mathcal{M}}_{N}\to{\mathcal{M}}_{N^{\prime}}$. Further $U^{\prime}$ acts on ${\mathcal{M}}_{N^{\prime}}$. Let $U^{\prime}=\coprod_{i}g_{i}U_{1}^{\prime}$ be a coset decomposition. We define a morphism $U^{\prime}gU]_{\tau,N,N^{\prime}}$ by restricting the map
$\begin{array}[]{ccc}({\mathcal{M}}_{N}\otimes{\mathcal{V}})^{S}&\to&({\mathcal{M}}_{N^{\prime}}\otimes{\mathcal{V}}^{\prime})^{S}\\ x\otimes v&\mapsto&\sum_{i}[g_{i}g]x\otimes\sigma^{\prime}(g_{i})\tau(v)\end{array}$
to ${\mathcal{M}}(\sigma)_{N}^{S}$. Similarly define
$U^{\prime}gU]_{\tau,N,N^{\prime},c}:{\mathcal{M}}(\sigma)_{N,c}^{S}\to({\mathcal{M}}_{N^{\prime},c}\otimes{\mathcal{V}}^{\prime})^{S}$
and
$U^{\prime}gU]_{\tau,N,N^{\prime},!}:{\mathcal{M}}(\sigma)_{N,!}\to({\mathcal{M}}_{N^{\prime},!}\otimes{\mathcal{V}}^{\prime})^{S}.$
Lemma 4.2.
$U^{\prime}gU]_{\tau,N,N^{\prime},\sharp}$ defines a morphism ${\mathcal{M}}(\sigma)_{N,\sharp}\to{\mathcal{M}}(\sigma^{\prime})_{N^{\prime},\sharp}$ which is independent of the coset decomposition $U^{\prime}=\coprod_{i}g_{i}U_{1}^{\prime}$, for $\sharp=\emptyset$, $\mathop{\rm tf}\nolimits$}, $c$ and !!.
Proof.
It suffices to prove the lemma on Betti realizations. We assume $\sharp=\emptyset$, the proof in the other cases being the same.
First we prove independence of the decomposition. Let $U^{\prime}=\coprod_{i}g^{\prime}_{i}U_{1}^{\prime}$ be another coset decomposition of $U^{\prime}$. Then $g^{\prime}_{i}=g_{i}h_{i}$ for some $h_{i}\in U_{1}^{\prime}$. Thus we can write $h_{i}=gu_{i}g^{-1}$ for some $u_{i}\in U$. Denote $U^{\prime}gU]^{\prime}_{\tau,N,N^{\prime}}$ for the map arising from $x\otimes v\mapsto\sum_{i}[g^{\prime}_{i}g]x\otimes\sigma^{\prime}(g^{\prime}_{i})\tau(v).$ Then we have
$[U^{\prime}gU]^{\prime}_{\tau,N,N^{\prime}}(x\otimes v)=\sum_{i}[g^{\prime}_{i}g]x\otimes\sigma^{\prime}(g^{\prime}_{i})\tau(v)$
$=$ $\sum_{i}[g_{i}gu_{i}]x\otimes\sigma^{\prime}(g^{\prime}_{i})\tau(v)$
$=$ $\sum_{i}[g_{i}gu_{i}]x\otimes\sigma^{\prime}(g_{i})\sigma^{\prime}(gu_{i}g^{-1})\tau(v)$
$=$ $\sum_{i}[g_{i}gu_{i}]x\otimes\sigma^{\prime}(g_{i})\tau(\sigma(u_{i})v)$
$=$ $\sum_{i}([g_{i}g]\otimes\sigma^{\prime}(g_{i})\tau)(u_{i}x\otimes\sigma(u_{i})v).$
Since $z\in{\mathcal{M}}(\sigma)_{N}$ is invariant under $U$, we have
$[U^{\prime}gU]_{\tau,N,N^{\prime}}(z)$ $=$ $\sum_{i}([g_{i}g]\otimes\sigma^{\prime}(g_{i})\tau)\circ(u_{i}\otimes\sigma(u_{i}))(z)$
$=$ $\sum_{i}([g_{i}g]\otimes\sigma^{\prime}(g_{i})\tau)(z).$
Now we prove that the image is contained in ${\mathcal{M}}(\sigma^{\prime})_{N}$. Let $u^{\prime}\in U^{\prime}$. Then
$u^{\prime}(\sum_{i}[g_{i}g]x\otimes\sigma^{\prime}(g_{i})\tau(v))$ $=$ $\sum_{i}[u^{\prime}g_{i}g]x\otimes\sigma^{\prime}(u^{\prime})\sigma^{\prime}(g_{i})\tau(v)$
$=$ $\sum_{i}[(u^{\prime}g_{i})g]x\otimes\sigma^{\prime}(u^{\prime}g_{i})\tau(v).$
Since $U^{\prime}=u^{\prime}U^{\prime}=\coprod_{i}u^{\prime}g_{i}U_{1}^{\prime}$, $\{u^{\prime}g_{i}\}_{i}$ is also a complete system of coset representatives of $U^{\prime}$ by $U_{1}^{\prime}$, by the independence of coset decomposition,
$\sum_{i}[(u^{\prime}g_{i})g]x\otimes\sigma^{\prime}(u^{\prime}g_{i})\tau(v)=\sum_{i}[g_{i}g]x\otimes\sigma^{\prime}(g_{i})\tau(v).$
4.3. Compositions
[…]
$\left\{\begin{array}[]{l}\ell>k,\\ g_{\ell},g^{\prime}_{\ell}\in\operatorname{{\mathop{\rm GL}\nolimits}}_{2}({\mathbb{Z}}_{\ell}){\rm\ or}\\ \operatorname{{\mathop{\rm GL}\nolimits}}_{2}({\mathbb{Z}}_{\ell})\subset\ker\sigma\cap\ker\sigma^{\prime}\cap\ker\sigma^{\prime\prime}\end{array}\right$.
Fix such an $S$. We choose stable ${\mathcal{O}}_{K}$-lattices ${\mathcal{V}},{\mathcal{V}}^{\prime}$ and $\mathcal{V}$$^{\prime\prime}$ as in the last section. Also choose $N,N^{\prime}$ and $N^{\prime\prime}$ such that
  • $U_{N}\subset\ker\sigma\cap\ker\sigma^{\prime}$,
  • $N\geq 4$,
  • $S_{N^{\prime\prime}}^{K}\subset S$,
  • $g^{-1}U_{N^{\prime}}g\subset U_{N}$ and $g^{\prime}{}^{-1}U_{N^{\prime\prime}}g^{\prime}\subset U_{N^{\prime}}$.
We will give a formula for the composition $U^{\prime\prime}g^{\prime}U^{\prime}]_{\tau^{\prime},N^{\prime},N^{\prime\prime},\sharp}\circ[U^{\prime}gU]_{\tau,N,N^{\prime},\sharp}$ for $\sharp\in\{\phi,c,!\}$. To simplify the notation, we will suppress NN, $N^{\prime}$ and $\sharp$ when there is no danger of confusion.
We first recall the definition of the product formula for double cosets [Sh1, §3.1],[Miy, §2.7]. Let $U^{\prime}_{1}=U^{\prime}\cap gUg^{-1}$ and $U^{\prime\prime}_{1}=U^{\prime\prime}\cap g^{\prime}U^{\prime}{g^{\prime}}^{-1}$. Fix coset decompositions $U^{\prime}=\coprod_{i}g_{i}U^{\prime}_{1}$ and $U^{\prime\prime}=\coprod_{j}g^{\prime}_{j}U^{\prime\prime}_{1}$. We have
$U^{\prime}gU=\coprod_{i}g_{i}gU,\ U^{\prime\prime}g^{\prime}U^{\prime}=\coprod_{j}g^{\prime}_{i}g^{\prime}U^{\prime}.$
Let $U^{\prime\prime}g^{\prime}U^{\prime}gU=\coprod_{n}U^{\prime\prime}g^{\prime}v_{n}gU$ be a double coset decomposition of $U^{\prime\prime}U^{\prime}U$. The product of the double cosets $U^{\prime\prime}g^{\prime}U^{\prime}]$ and $U^{\prime}gU]$ is defined by
$U^{\prime\prime}g^{\prime}U^{\prime}]\cdot[U^{\prime}gU]=\sum_{n}c_{n}[U^{\prime\prime}g^{\prime}v_{n}gU],$
where $c_{n}=\#\{(i,j)\mid g^{\prime}_{j}g^{\prime}g_{i}gU=g^{\prime}v_{n}gU\}$.
Let $U^{\prime}_{2}=U^{\prime}\cap g^{\prime}{}^{-1}U^{\prime\prime}g^{\prime}$ and let $U^{\prime}=\coprod_{k}U^{\prime}_{2}u_{k}U^{\prime}_{1}$ be a double coset decomposition of $U^{\prime}$. For each $k$, let $W_{k}=U^{\prime\prime}\cap(g^{\prime}u_{k}gU(g^{\prime}u_{k}g)^{-1})$ and $W_{k}^{\prime}=W_{k}\cap g^{\prime}U^{\prime}g^{\prime}{}^{-1}$.
Proposition 4.3.
$U^{\prime\prime}g^{\prime}U^{\prime}]\cdot[U^{\prime}gU]=\sum_{k}[W_{k}:W_{k}^{\prime}][U^{\prime\prime}g^{\prime}u_{k}gU].$
Proof.
Let $U^{\prime\prime}=\coprod_{m}w_{k,m}W_{k}$ and $W_{k}=\coprod_{\ell}w^{\prime}_{k,\ell}W^{\prime}_{k}$ be fixed coset decompositions. It is easy to see that we have coset decompositions
(20) $U^{\prime\prime}g^{\prime}U^{\prime}=\coprod_{i,j}g^{\prime}_{j}g^{\prime}g_{i}U^{\prime}_{1}$
and
(21) $U^{\prime\prime}g^{\prime}U^{\prime}=\coprod_{k,m,\ell}w_{k,m}w^{\prime}_{k,\ell}g^{\prime}u_{k}U^{\prime}_{1}.$
Further, for each $u_{k}$ in the double coset decomposition $U^{\prime}=\coprod_{k}U^{\prime}_{2}u_{k}U^{\prime}_{1}$ of $U^{\prime}$, we have
(22) $U^{\prime\prime}g^{\prime}u_{k}gU=\coprod_{m}w_{k,m}g^{\prime}u_{k}gU.$
This is because any $u\in U^{\prime\prime}$ can be written in the form $w_{k,m}w^{\prime}$ for some $m$ and $w^{\prime}\in W_{k}$. Then $ug^{\prime}u_{k}gU=w_{k,m}w^{\prime}g^{\prime}u_{k}gU=w_{k,m}g^{\prime}u_{k}gU.$ Thus we have the union
$U^{\prime\prime}g^{\prime}u_{k}gU=\bigcup_{m}w_{k,m}g^{\prime}u_{k}gU.$
Suppose $w_{k,m_{1}}g^{\prime}u_{k}gU=w_{k,m_{2}}g^{\prime}u_{k}gU$. Then $w_{k,m_{2}}^{-1}w_{k,m_{1}}$ is in $g^{\prime}u_{k}gU(g^{\prime}u_{k}g)^{-1}$. This means that $w_{k,m_{1}}^{-1}w_{k,m_{2}}$ is in $W_{k}$. Therefore $m_{1}=m_{2}$. So the union is disjoint.
(20) and (21) give us two decompositions of $U^{\prime\prime}g^{\prime}U^{\prime}$ into cosets of $U^{\prime}_{1}$. So we have a one-to-one correspondence
(23) $\{g^{\prime}_{j}g^{\prime}g_{i}U^{\prime}_{1}|j,i\}\leftrightarrow\{w_{k,m}w^{\prime}_{k,\ell}g^{\prime}u_{k}U^{\prime}_{1}|k,m,\ell\}.$
Fix a $v_{n}$ in $U^{\prime\prime}g^{\prime}U^{\prime}gU=\coprod_{n}U^{\prime\prime}g^{\prime}v_{n}gU$. Because of (23), we have
$c_{n}=\#\{(k,m,\ell)|w_{k,m}w^{\prime}_{k,\ell}g^{\prime}u_{k}gU=g^{\prime}v_{n}gU\}.$
Given a triple $(k,m,\ell)$ with $w_{k,m}w^{\prime}_{k,\ell}g^{\prime}u_{k}gU=g^{\prime}v_{n}gU$, $u_{k}$ must have the property $U^{\prime\prime}g^{\prime}u_{k}gU=U^{\prime\prime}g^{\prime}v_{n}gU$. Fix such a $u_{k}$. Since $w^{\prime}_{k,\ell}$ is in $W_{k}$, we have $w_{k,m}w^{\prime}_{k,\ell}g^{\prime}u_{k}gU=w_{k,m}g^{\prime}u_{k}gU$. By (22), $g^{\prime}v_{n}gU=w_{k,m_{1}}g^{\prime}u_{k}gU$ for a unique value of $m_{1}$. So $w_{k,m}w^{\prime}_{k,\ell}g^{\prime}u_{k}gU=g^{\prime}v_{n}gU$ gives $w_{k,m}g^{\prime}u_{k}gU=w_{k,m_{1}}g^{\prime}u_{k}gU$, independent of $\ell$. By (22), this holds if and only if $m=m_{1}$ while $\ell$ takes any value. Therefore
$c_{n}$ $=$ $\sum_{k}\#\{(k,m,\ell)|w_{k,m}w^{\prime}_{k,\ell}g^{\prime}u_{k}gU=g^{\prime}v_{n}gU\}$
$=$ $\sum_{k}\#\{(k,m_{1},\ell)|w^{\prime}_{k,\ell}g^{\prime}u_{k}gU=g^{\prime}u_{k}gU\}$
$=$ $\sum_{k}[W_{k}:W^{\prime}_{k}]$
where the sum is over $k$ such that $U^{\prime\prime}g^{\prime}u_{k}gU=U^{\prime\prime}g^{\prime}v_{n}gU$. This is also the coefficient of $U^{\prime\prime}g^{\prime}v_{n}gU]$ given in the proposition. ∎
We now describe the composition of double coset actions. As in the proof of Proposition 4.3, let $U^{\prime\prime}=\coprod_{m,\ell}w_{k,m}w^{\prime}_{k,\ell}W^{\prime}_{k}$. For each pair $(k,\ell)$, define
$\tau_{k,\ell}:{\mathcal{V}}\to{\mathcal{V}}^{\prime\prime},$
$v\mapsto\sigma^{\prime\prime}(w^{\prime}_{k,\ell})\tau^{\prime}\sigma^{\prime}(u_{k})\tau\sigma((g^{\prime}u_{k}g)^{-1}w^{\prime}_{k,\ell}{}^{-1}(g^{\prime}u_{k}g))v,v\in{\mathcal{V}}.$
We will often suppress $\sigma,\sigma^{\prime}$ and $\sigma^{\prime\prime}$ from the notation. Define
$\tau_{k}=\sum_{\ell}\tau_{k,\ell}.$
It is straightforward to check that $\tau_{k,\ell}$ is independent of the choice of the coset representatives $w^{\prime}_{k,\ell}$, and hence so is $\tau_{k}$. Define
$\tilde{\tau}_{k}:W_{k}\to U,u^{\prime\prime}\mapsto(g^{\prime}u_{k}g)^{-1}u^{\prime\prime}(g^{\prime}u_{k}g).$
For $u^{\prime\prime}\in W_{k}$ and $v\in{\mathcal{V}}$, $\{u^{\prime\prime}{}^{-1}w^{\prime}_{k,\ell}\}_{\ell}$ is another set of coset representatives of $W^{\prime}_{k}$ in $W_{k}$. Thus we have
$\tau_{k}(\tilde{\tau}_{k}(u^{\prime\prime})v)$
$=$ $\sum_{\ell}w^{\prime}_{k,\ell}\tau^{\prime}u_{k}\tau(g^{\prime}u_{k}g)^{-1}w^{\prime}_{k,\ell}{}^{-1}(g^{\prime}u_{k}g)(g^{\prime}u_{k}g)^{-1}u^{\prime\prime}(g^{\prime}u_{k}g)v$
$=$ $\sum_{\ell}w^{\prime}_{k,\ell}\tau^{\prime}u_{k}\tau(g^{\prime}u_{k}g)^{-1}w^{\prime}_{k,\ell}{}^{-1}u^{\prime\prime}(g^{\prime}u_{k}g)v$
$=$ $\sum_{\ell}u^{\prime\prime}(u^{\prime\prime}{}^{-1}w^{\prime}_{k,\ell})\tau^{\prime}u_{k}\tau(g^{\prime}u_{k}g)^{-1}(u^{\prime\prime}{}^{-1}w^{\prime}_{k,\ell})^{-1}(g^{\prime}u_{k}g)v$
$=$ $u^{\prime\prime}\tau_{k}(v).$
Thus $U^{\prime\prime}g^{\prime}u_{k}gU]_{\tau_{k}}$ is well-defined.
Proposition $4.$
$U^{\prime\prime}g^{\prime}U^{\prime}]_{\tau^{\prime}}\circ[U^{\prime}gU]_{\tau}=\sum_{k}[U^{\prime\prime}g^{\prime}u_{k}gU]_{\tau_{k}}.$
More precisely,
$U^{\prime\prime}g^{\prime}U^{\prime}]_{\tau^{\prime},N^{\prime},N^{\prime\prime}}\circ[U^{\prime}gU]_{\tau,N,N^{\prime}}=\sum_{k}[U^{\prime\prime}g^{\prime}u_{k}gU]_{\tau_{k},N,N^{\prime\prime}}$
in ${\mathcal{P}M}^{S_{N^{\prime\prime}}}_{K}$.
Proof.
Write $U^{\prime}=\coprod_{i}g_{i}U^{\prime}_{1}$ and $U^{\prime\prime}=\coprod_{j}g^{\prime}_{j}U^{\prime\prime}_{1}$. Then the operators
$g^{\prime}_{j}g^{\prime}g_{i}g]:{\mathcal{M}}_{N}\to{\mathcal{M}}_{N^{\prime}}\to{\mathcal{M}}_{N^{\prime\prime}}$
and
$g^{\prime}_{j}\tau^{\prime}g_{i}\tau:{\mathcal{V}}\to{\mathcal{V}}^{\prime}\to{\mathcal{V}}^{\prime\prime}$
are both well-defined. By definition, we have
$[U^{\prime\prime}g^{\prime}U^{\prime}]_{\tau^{\prime}}\circ[U^{\prime}gU]_{\tau}(x\otimes v)$ $=$ $[U^{\prime\prime}g^{\prime}U^{\prime}]_{\tau^{\prime}}(\sum_{i}[g_{i}g]x\otimes g_{i}\tau(v))$
$=$ $\sum_{i}\sum_{j}[g_{j}^{\prime}g^{\prime}g_{i}g]x\otimes g_{j}^{\prime}\tau^{\prime}(g_{i}\tau(v)).$
We first prove that, for each pair $(i,j)$, if $g^{\prime}_{j}g^{\prime}g_{i}U^{\prime}_{1}=w_{k,m}w^{\prime}_{k,\ell}g^{\prime}u_{k}U^{\prime}_{1}$ under the bijection (23), then
$g^{\prime}_{j}g^{\prime}g_{i}g]x\otimes g^{\prime}_{j}\tau^{\prime}g_{i}\tau v=[w_{k,m}w^{\prime}_{k,\ell}g^{\prime}u_{k}g]x\otimes w_{k,m}w^{\prime}_{k,\ell}\tau_{k}v.$
Note that the 1-1 correspondence $g^{\prime}_{j}g^{\prime}g_{i}gU^{\prime}_{1}\leftrightarrow w_{k,m}w^{\prime}_{k,\ell}g^{\prime}u_{k}gU^{\prime}_{1}$ can be described as follows. Given $g^{\prime}_{j}g^{\prime}g_{i}g$, since $U^{\prime}=\coprod_{k}U^{\prime}_{2}u_{k}U^{\prime}_{1}$, $g_{i}$ is in $U^{\prime}_{2}u_{k}^{\prime}U^{\prime}_{1}$ for a unique $k$. So $g_{i}=w^{\prime}_{2}u_{k}w^{\prime}_{1}$ with $w^{\prime}_{1}\in U^{\prime}_{1}$ and $w^{\prime}_{2}\in U^{\prime}_{2}$. So $w^{\prime}_{1}=gw_{1}g^{-1},w^{\prime}_{2}=g^{\prime}{}^{-1}w^{\prime\prime}_{2}g^{\prime}$ with $w_{1}\in U$ and $w^{\prime\prime}_{2}\in U^{\prime\prime}$. Since $g^{\prime}_{j}w^{\prime\prime}_{2}$ is in $U^{\prime\prime}$, it can be uniquely written as $w_{k,m}w^{\prime}_{k,\ell}w^{\prime\prime}_{3}$ with $w^{\prime\prime}_{3}\in W^{\prime}_{k}$. So $w^{\prime\prime}_{3}=(g^{\prime}u_{k}g)w_{3}(g^{\prime}u_{k}g)^{-1}$ for some $w_{3}\in U$ and $g^{\prime}{}^{-1}w^{\prime\prime}_{3}g^{\prime}=(u_{k}g)w_{3}(u_{k}g)^{-1}$ is in $U^{\prime}$. Based on this description, we have
$g^{\prime}_{j}g^{\prime}g_{i}g$ $=$ $w_{k,m}w^{\prime}_{k,\ell}w^{\prime\prime}_{3}g^{\prime}u_{k}gw_{1}$
$=$ $w_{k,m}w^{\prime}_{k,\ell}g^{\prime}u_{k}gw_{3}w_{1}.$
This gives the correspondence $g^{\prime}_{j}g^{\prime}g_{i}gU^{\prime}_{1}=w_{k,m}w^{\prime}_{k,\ell}g^{\prime}u_{k}gU.$ This description, together with the transformation laws observed by $\tau$ and $\tau^{\prime}$, also gives us
$g^{\prime}_{j}\tau^{\prime}g_{i}\tau$ $=$ $w_{k,m}w^{\prime}_{k,\ell}\tau^{\prime}u_{k}\tau w_{3}w_{1}.$
Note that if $w$ is in $U$, then $(w\otimes w)z=z$ for $z\in{\mathcal{M}}(\sigma)_{N}=({\mathcal{M}}_{N}\otimes{\mathcal{V}})^{U}$ (see the proof of Lemma 4.2). Therefore, for $z\in{\mathcal{M}}(\sigma)_{N}$,
$(\sum_{i,j}[g^{\prime}_{j}g^{\prime}g_{i}g]\otimes g^{\prime}_{j}\tau^{\prime}g_{i}\tau)(z)$
$=$ $(\sum_{k,m,\ell}[w_{k,m}w^{\prime}_{k,\ell}g^{\prime}u_{k}gw_{3}w_{1}]\otimes w_{k,m}w^{\prime}_{k,\ell}\tau^{\prime}u_{k}\tau w_{3}w_{1})(z)$
$=$ $(\sum_{k,m,\ell}[w_{k,m}g^{\prime}u_{k}g]\otimes w_{k,m}\tau_{k,\ell})(z)$
$=$ $(\sum_{k,m}[w_{k,m}g^{\prime}u_{k}g]\otimes w_{k,m}\tau_{k})(z)$
$=$ $(\sum_{k}[U^{\prime\prime}g^{\prime}u_{k}gU]_{\tau_{k}})(z).$
This proves the proposition. ∎
Using Lemma 3.9 and Proposition 4.4, we conclude that $U^{\prime}gU]$ is independent of the choice of NN and $N^{\prime}$. More precisely, if $U^{\prime}gU]_{\tau,M,M^{\prime}}$ is also defined, then $U^{\prime}gU]_{\tau,N,N^{\prime}}^{S}=[U^{\prime}gU]_{\tau,M,M^{\prime}}^{S}$ when $S\supset S_{\operatorname{{\mathop{\rm lcm}\nolimits}}(N^{\prime},M^{\prime})}$.
$4.$ Pairing on $\sigma$-constructions
In this section we will construct perfect pairings between certain $\sigma$-constructions and compute the adjoints of double coset actions with respect to the pairings.
For $N\geq 3$, we let $H=H_{N}$ denote the premotivic structure $H^{0}(X_{N})=H^{0}(X)=H^{0}(\bar{X})$. More precisely, we let ${\mathcal{H}}_{\mathop{\rm B}\nolimits}=H^{0}(X^{\mathop{\rm an}\nolimits},{\mathbb{Z}})$, ${\mathcal{F}}_{\mathop{\rm dR}\nolimits}={\bf H}^{0}(\bar{X},\omega^{\bullet}_{\bar{X}/T})$, ${\mathcal{H}}_{\ell}=H^{0}(X_{\bar{{\mathbb{Q}}}},{\mathbb{Z}}_{\ell})$ and -${\mathcal{H}}_{\ell\mbox{\scriptsize-crys}}=H^{0}_{\mbox{\scriptsize crys}}(X_{{\mathbb{Z}}_{\ell}},{\mathcal{O}}_{X_{{\mathbb{Z}}_{\ell}},{\mbox{\scriptsize crys}}})$ for $\ell\not\in S_{N}$. These come equipped with additional structure and comparison isomorphisms making $\mathcal{H}$} an object of ${\mathcal{P}M}_{\mathbb{Q}}^{S_{N}}$, and we let $H={\mathbb{Q}}\otimes{\mathcal{H}}$.
Let $F={\mathbb{Q}}(\mu_{N})$. The Weil pairing on $({\mathbb{Z}}/N{\mathbb{Z}})^{2}_{X}\cong E[N]$ defines an isomorphism between $({\mathbb{Z}}/N{\mathbb{Z}})_{X}$ and $\mu_{N,X}$, hence a morphism $X\to\operatorname{{\mathop{\rm Spec}\nolimits}}{\mathcal{O}}_{F}$, which one checks induces a morphism ${\mathcal{M}}_{F}^{S_{N}}\to{\mathcal{H}}$ in ${\mathcal{P}M}_{\mathbb{Q}}^{S_{N}}$. One checks also that the map $X^{\mathop{\rm an}\nolimits}\to(\operatorname{{\mathop{\rm Spec}\nolimits}}{\mathcal{O}}_{F})^{\mathop{\rm an}\nolimits}$ sends the component $X_{N,t}$ to $e^{-2\pi it/N}$, where we identify $(\operatorname{{\mathop{\rm Spec}\nolimits}}{\mathcal{O}}_{F})^{\mathop{\rm an}\nolimits}={\mathop{\rm Hom}\nolimits}({\mathcal{O}}_{F},{\mathbb{C}})={\bf I}_{F}$ with the set of primitive $N^{\mathop{\rm th}\nolimits}$ roots in $\mathbb{C}$}. In particular, it follows that ${\mathcal{M}}_{F}^{S_{N}}\cong{\mathcal{H}}$ since the map ${\mathcal{M}}_{F,{\mathop{\rm B}\nolimits}}^{S_{N}}\to{\mathcal{H}}_{\mathop{\rm B}\nolimits}$ is the isomorphism
$H^{0}(X^{\mathop{\rm an}\nolimits},{\mathbb{Z}})=\bigoplus_{t\in({\mathbb{Z}}/N{\mathbb{Z}})^{\times}}H^{0}(X_{N,t},{\mathbb{Z}})=\bigoplus_{t\in({\mathbb{Z}}/N{\mathbb{Z}})^{\times}}{\mathbb{Z}}\cong{\mathbb{Z}}^{{\bf I}_{F}}$
induced by the bijection $t\leftrightarrow e^{-2\pi it/N}$.
Recall that if $(h,N,N^{\prime})$ is an admissible triple in the sense of §3.2, then we obtain a morphism $\chi_{h}:\bar{X}^{\prime}\to\bar{X}$, where $\bar{X}^{\prime}=\bar{X}_{N^{\prime}}$. More generally, if $N,N^{\prime}\geq 3$ and $g\in\operatorname{{\mathop{\rm GL}\nolimits}}_{2}({\mathbb{A}}_{\bf f})$ is such that $g^{-1}U_{N^{\prime}}g\subset U_{N}$, we can define $\chi_{g}=\chi_{h}$ for any $h\in g{\mathbb{Q}}^{\times}$ such that $(h,N,N^{\prime})$ is admissible. Then $\chi_{g}$ induces a morphism $\chi_{g}^{*}:{\mathcal{H}}^{S^{\prime}}\to{\mathcal{H}}^{\prime}$ where $S^{\prime}=S_{N^{\prime}}$ and ${\mathcal{H}}^{\prime}={\mathcal{H}}_{N^{\prime}}$.
Lemma 4.5.
For NN, $N^{\prime}$ and $g$ as above, we have the commutative diagram
[…]
Proof.
It suffices to check the commutativity of the diagram
$\begin{array}[]{ccc}X_{N^{\prime}}&\to&X_{N}\\ \downarrow&&\downarrow\\ \operatorname{{\mathop{\rm Spec}\nolimits}}{\mathcal{O}}_{F^{\prime}}&\to&\operatorname{{\mathop{\rm Spec}\nolimits}}{\mathcal{O}}_{F}\end{array}$
on complex points, and this is immediate from Lemma 3.2. ∎
[…]
We fix a continuous character $\psi:\hat{{\mathbb{Z}}}^{\times}\to K^{\times}$ of finite order. We also view $\psi$ as a character of $\mathbb{A}$$^{\times}$$\mathbb{Q}$$^{\times}$, and write $\psi$ for the corresponding character of $G_{\mathbb{Q}}$. We view $U$ as acting on ${\mathcal{O}}_{\psi}={\mathcal{O}}_{K}$ via $\psi^{-1}\circ{\mathop{\rm det}\nolimits}$.
Lemma 4.6.
Suppose NN is such that $N\geq 3$, $U_{N}\subset U$, the conductor of $\psi$ divides NN. Then the morphism ${\mathcal{M}}_{F}\to{\mathcal{H}}$ induces an isomorphism
${\mathcal{M}}_{\psi}^{S_{N}}\to({\mathcal{O}}_{\psi}\otimes{\mathcal{H}})^{U}$
in ${\mathcal{P}M}_{K}^{S_{N}}$.
Proof.
It follows from Lemma 4.5 that the isomorphism
${\mathcal{O}}_{\psi}\otimes{\mathcal{M}}_{F}^{S_{N}}\cong{\mathcal{O}}_{\psi}\otimes{\mathcal{H}}$
respects the action of $U$, where we view $U$ acting on ${\mathcal{O}}_{\psi}\otimes{\mathcal{M}}_{F}$ via ${\mathop{\rm det}\nolimits}^{-1}:U\to\hat{{\mathbb{Z}}}^{\times}\stackrel{{\scriptstyle\cong}}{{\longrightarrow}}G_{\mathbb{Q}}^{\mathop{\rm ab}\nolimits}$. Taking invariants gives the lemma. ∎
Using the compatibility of the cup product with comparison isomorphisms, one obtains morphisms
(24) $\mathcal{H}$$\otimes{\mathcal{M}}_{\sharp}\to{\mathcal{M}}_{\sharp}$
[…]
commutes, where the morphisms $g]_{\sharp}:{\mathcal{M}}_{\sharp}\to{\mathcal{M}}^{\prime}_{\sharp}$ are those of Theorem 3.6. (Again the commutativity can be easily checked on Betti realizations.) In particular, the morphism of (24) respects the action of $U$.
Next we consider the composite morphism
(26) ${\mathcal{M}}_{\psi}^{S_{N}^{K}}\otimes{\mathcal{M}}_{\sharp}\to{\mathcal{O}}_{\psi}\otimes{\mathcal{H}}\otimes{\mathcal{M}}_{\sharp}\to{\mathcal{O}}_{\psi}\otimes{\mathcal{M}}_{\sharp}$
in ${\mathcal{P}M}_{K}^{S_{N}^{K}}$, where the first map is the one in Lemma 4.6 tensored with $\mathcal{M}$$_{\sharp}$ and the second is (24) tensored with ${\mathcal{O}}_{\psi}$.
Lemma 4.7.
The morphism in (26) is an isomorphism respecting the action of $U$, where the action of of $U$ on is trivial on ${\mathcal{M}}_{\psi}$, defined in Theorem 3.6 on $\mathcal{M}$$_{\sharp}$, and via $\psi^{-1}\circ{\mathop{\rm det}\nolimits}$ on ${\mathcal{O}}_{\psi}$.
Proof.
The $U$-equivariance follows from that of (24) together with Lemma 4.6.
It suffices to check the maps are isomorphisms on Betti realizations. For this, write
${\mathcal{M}}_{\mathop{\rm B}\nolimits}=\bigoplus_{t\in({\mathbb{Z}}/N{\mathbb{Z}})^{\times}}H^{1}(X_{N,t},{\mathcal{F}}_{\mathop{\rm B}\nolimits}^{k}).$
Then for a class $x\in H^{1}(X_{N,t},{\mathcal{F}}_{\mathop{\rm B}\nolimits}^{k})$, we have $b_{\mathop{\rm B}\nolimits}\otimes x\mapsto\psi^{-1}(-t)\otimes x$ under (24). It follows that the map is surjective, hence it is injective since it is between two isomorphic finitely generated ${\mathcal{O}}_{K}$-modules.
The case of ${\mathcal{M}}_{c}$ is similar, and the remaining cases follow from those of $\mathcal{M}$} and ${\mathcal{M}}_{c}$. ∎
We now construct the pairings by defining the duality homomorphisms. Given a representation $\sigma:U\to\operatorname{\operatorname{{\mathop{\rm Aut}\nolimits}}}_{{\mathcal{O}}_{K}}{\mathcal{V}}$, let $\hat{\sigma}\otimes(\psi^{-1}\circ{\mathop{\rm det}\nolimits})$ denote the representation defined by the action of $U$ on ${\mathop{\rm Hom}\nolimits}_{{\mathcal{O}}_{K}}({\mathcal{V}},{\mathcal{O}}_{\psi})$. For $N\geq 3$ such that $U_{N}\subset\ker\sigma$ and the conductor of $\psi$ divides NN, Lemma (4.7) and the duality homomorphism $\delta^{L}_{N}:{\mathcal{M}}_{N,c}\to{\mathop{\rm Hom}\nolimits}({\mathcal{M}}_{N},{\mathcal{T}}^{S_{N},\otimes(k-1)})$ from §2.5 give a morphism
(27) $\begin{array}[]{l}{\mathop{\rm Hom}\nolimits}_{{\mathcal{O}}_{K}}({\mathcal{V}},{\mathcal{O}}_{\psi^{-1}})\otimes{\mathcal{M}}_{N,c}\cong{\mathop{\rm Hom}\nolimits}_{{\mathcal{O}}_{K}}({\mathcal{V}},{\mathcal{M}}_{N,c}\otimes{\mathcal{M}}_{\psi}^{S_{N}^{K}})\\ \to{\mathop{\rm Hom}\nolimits}_{{\mathcal{O}}_{K}}({\mathcal{V}},{\mathop{\rm Hom}\nolimits}_{{\mathcal{O}}_{K}}({\mathcal{M}}_{N},{\mathcal{T}}^{S_{N},\otimes(k-1)}\otimes{\mathcal{M}}_{\psi}^{S_{N}^{K}}))\\ \cong{\mathop{\rm Hom}\nolimits}_{{\mathcal{O}}_{K}}({\mathcal{V}}\otimes{\mathcal{M}}_{N},{\mathcal{M}}_{\psi}(1-k)),\end{array}$
where we write ${\mathcal{M}}_{\psi}(1-k)$ for the object ${\mathcal{T}}^{S_{N},\otimes(k-1)}\otimes{\mathcal{M}}_{\psi}^{S_{N}^{K}}$ of ${\mathcal{P}M}_{K}^{S_{N}^{K}}$. Moreover applying Proposition 3.8 with $g\in U$, we find that the morphism (27) respects the action of $U$ where we view $U$ as acting trivially on ${\mathcal{M}}_{\psi}(1-k)$.
Taking $U$-invariants, we get a morphism
(28) $\delta^{L}_{N}:{\mathcal{M}}_{c}(\hat{\sigma}\otimes(\psi^{-1}\circ{\mathop{\rm det}\nolimits}))_{N}\to{\mathop{\rm Hom}\nolimits}_{{\mathcal{O}}_{K}}({\mathcal{M}}(\sigma)_{N},{\mathcal{M}}_{\psi}(1-k)).$
Tensoring with $\mathbb{Q}$} and normalizing by dividing by $U:U_{N}]$, we get a morphism
$\bar{\delta}^{L}_{N}:M_{c}(\hat{\sigma}\otimes(\psi^{-1}\circ{\mathop{\rm det}\nolimits}))_{N}\to{\mathop{\rm Hom}\nolimits}_{K}(M(\sigma)_{N},M_{\psi}(1-k)).$
Similarly, using the duality morphism ${\mathcal{M}}_{N,!}\to{\mathop{\rm Hom}\nolimits}({\mathcal{M}}_{N,!},{\mathcal{T}}^{S_{N},\otimes(k-1)})$, we obtain morphisms
$\delta^{L}_{N,!}:{\mathcal{M}}_{!}(\hat{\sigma}\otimes(\psi^{-1}\circ{\mathop{\rm det}\nolimits}))_{N}\to{\mathop{\rm Hom}\nolimits}_{{\mathcal{O}}_{K}}({\mathcal{M}}_{!}(\sigma)_{N},{\mathcal{M}}_{\psi}(1-k))$
and
$\bar{\delta}^{L}_{N,!}:M_{!}(\hat{\sigma}\otimes(\psi^{-1}\circ{\mathop{\rm det}\nolimits}))_{N}\to{\mathop{\rm Hom}\nolimits}_{K}(M_{!}(\sigma)_{N},M_{\psi}(1-k)).$
Proposition 4.8.
The morphisms $\bar{\delta}^{L}_{N}$ and $\bar{\delta}^{L}_{N,!}$ are isomorphisms in $\operatorname{{\bf PM}}_{K}^{S_{N}^{K}}$ compatible with the change of levels.
Proof.
It follows from Theorem 2.4 b) that (27) is an isomorphism after tensoring with $\mathbb{Q}$}. The same is true for the natural map
${\mathop{\rm Hom}\nolimits}_{{\mathcal{O}}_{K}}({\mathcal{V}}\otimes{\mathcal{M}}_{N},{\mathcal{M}}_{\psi}(1-k))^{U}\to{\mathop{\rm Hom}\nolimits}_{{\mathcal{O}}_{K}}({\mathcal{M}}(\sigma)_{N},{\mathcal{M}}_{\psi}(1-k)),$
so $\bar{\delta}^{L}_{N}$ is an isomorphism. The compatibility with change of levels follows from applying Proposition 3.8 and (25) with $g=I$. The proof for $\bar{\delta}^{L}_{N,!}$ is the same. ∎
We will therefore omit the subscript NN in the notation for the morphisms $\bar{\delta}^{L}$. (Note however that $\delta^{L}_{N}$ depends on NN, and $\delta^{L}_{N,!}$ depends on NN only in the normalization.)
[…]
Proposition 4.9.
[…]
(29) ${\mathcal{M}}(\hat{\sigma}\otimes(\psi^{-1}\circ{\mathop{\rm det}\nolimits}))_{!}^{S}\to{\mathop{\rm Hom}\nolimits}_{{\mathcal{O}}_{K}}({\mathcal{M}}(\sigma)_{!}^{S},{\mathcal{M}}_{\psi}(1-k)^{S})$
[…]
Proof.
We only need to prove the assertion for Betti realizations, for then $\delta^{L}_{N}=[U:U_{N}]\tilde{\delta}^{L}$ for a unique morphism $\tilde{\delta}^{L}$ with the desired property.
To do this, we give an alternate description of the Betti realizations and pairings which is independent of NN. and define a perfect pairing on the new Betti realizations. Given a sufficiently small $U$ and a $\sigma:U\to\operatorname{{\mathop{\rm Aut}\nolimits}}_{{\mathcal{O}}_{K}}{\mathcal{V}}$, define a locally constant sheaf ${\mathcal{F}}(\sigma)$ on $G_{\mathbb{Q}}\backslash G_{\mathbb{A}}/UU_{\infty}$ by ${\mathcal{F}}(\sigma)=G_{\mathbb{Q}}\backslash(G_{\mathbb{A}}\times{\mathcal{V}})/UU_{\infty}$. Define
${\mathcal{M}}(\sigma)^{\mathop{\rm an}\nolimits}=H^{1}(X^{\mathop{\rm an}\nolimits}_{U},{\mathcal{F}}^{k}_{B}\otimes{\mathcal{F}}(\sigma))$
and
${\mathcal{M}}(\sigma)_{c}^{\mathop{\rm an}\nolimits}=H^{1}_{c}(X^{\mathop{\rm an}\nolimits}_{U},{\mathcal{F}}^{k}_{B}\otimes{\mathcal{F}}(\sigma)).$
Also define ${\mathcal{M}}(\sigma)_{!}^{\mathop{\rm an}\nolimits}$ to be the image of ${\mathcal{M}}(\sigma)_{c}^{\mathop{\rm an}\nolimits}$ in ${\mathcal{M}}(\sigma)^{\mathop{\rm an}\nolimits}_{\mathop{\rm tf}\nolimits}$ where ${\mathcal{M}}(\sigma)^{\mathop{\rm an}\nolimits}_{\mathop{\rm tf}\nolimits}$ is the largest torsion-free quotient of ${\mathcal{M}}(\sigma)^{\mathop{\rm an}\nolimits}$. Then from the pairing ${\mathcal{F}}_{B}^{k}\otimes{\mathcal{F}}_{B}^{k}\to(2\pi i)^{2-k}{\mathbb{Z}}$ we get a pairing
$({\mathcal{F}}^{k}_{B}\otimes{\mathcal{F}}(\hat{\sigma}\otimes(\psi^{-1}\circ{\mathop{\rm det}\nolimits})))\otimes({\mathcal{F}}^{k}_{B}\otimes{\mathcal{F}}(\sigma))\to(2\pi i)^{2-k}{\mathbb{Z}}\otimes{\mathcal{F}}(\psi^{-1}\circ{\mathop{\rm det}\nolimits}).$
Here ${\mathcal{F}}(\psi^{-1}\circ{\mathop{\rm det}\nolimits})G_{\mathbb{Q}}\backslash(G_{\mathbb{A}}\times{\mathcal{O}}_{\psi})/UU_{\infty}$ is the sheaf on $X^{\mathop{\rm an}\nolimits}=G_{\mathbb{Q}}\backslash G_{\mathbb{A}}/UU_{\infty}$ defined by ${\mathcal{O}}_{\psi}$.
We have an isomorphism of locally constant sheaves
[…]
where $\Gamma_{U}$ acts on ${\mathcal{O}}_{\psi^{-1}}$ through its canonical embedding into $U$ and then take the inverse. Since ${\mathop{\rm det}\nolimits}(\Gamma_{U})=1$, the action of $\Gamma_{U}$ on ${\mathcal{O}}_{\psi}$ is trivial. Thus the sheaf on the right is constant. So the same is true for the sheaf on the left.
Thus by Poincaré duality, we have a pairing
$[\ ,\ ]^{\mathop{\rm an}\nolimits}_{c}:\$${\mathcal{M}}(\hat{\sigma}\otimes(\psi^{-1}\circ{\mathop{\rm det}\nolimits}))_{c}^{\mathop{\rm an}\nolimits}\otimes{\mathcal{M}}(\sigma)^{\mathop{\rm an}\nolimits}$
$\to H^{2}_{c}(X^{\mathop{\rm an}\nolimits},({\mathcal{F}}^{k}_{B}\otimes{\mathcal{F}}(\hat{\sigma}\otimes(\psi^{-1}\circ{\mathop{\rm det}\nolimits})))\otimes({\mathcal{F}}^{k}_{B}\otimes{\mathcal{F}}(\sigma)))$
$\to H^{2}_{c}(X^{\mathop{\rm an}\nolimits},(2\pi i)^{2-k}{\mathbb{Z}}\otimes{\mathcal{O}}_{\psi^{-1}})$
$\cong H^{2}_{c}(X,{\mathbb{Z}})\otimes(2\pi i)^{2-k}{\mathcal{O}}_{\psi^{-1}}$
$\cong H^{2}_{c}(X,{\mathbb{Z}})\otimes(2\pi i)^{2-k}{\mathcal{M}}_{\psi,B}$
$\cong{\mathcal{M}}_{\psi}(1-k)_{B}$
which is perfect on torsion-free quotients. We then obtain a pairing
$^{\mathop{\rm an}\nolimits}_{!}:{\mathcal{M}}(\hat{\sigma}\otimes(\psi^{-1}\circ{\mathop{\rm det}\nolimits}))_{!}^{\mathop{\rm an}\nolimits}\otimes{\mathcal{M}}(\sigma)_{!}^{\mathop{\rm an}\nolimits}\to{\mathcal{M}}_{\psi}(1-k)_{B}$
by using the commutative diagram
[…]
where all maps are the natural ones except that the bottom map is the pairing
${\mathcal{M}}(\sigma)_{c}^{\mathop{\rm an}\nolimits}\otimes_{{\mathcal{O}}_{K}}{\mathcal{M}}(\hat{\sigma}\otimes(\psi^{-1}\circ{\mathop{\rm det}\nolimits}))^{\mathop{\rm an}\nolimits}\to{\mathcal{M}}_{\psi}(1-k)_{B}$
proceeded by the order-reversing isomorphism and followed by multiplication by $(-1)^{k-1}$.
[…]
$\tilde{\delta}^{\mathop{\rm an}\nolimits}:{\mathcal{M}}(\hat{\sigma}\otimes(\psi^{-1}\circ{\mathop{\rm det}\nolimits}))_{!}^{\mathop{\rm an}\nolimits}\to{\mathop{\rm Hom}\nolimits}_{{\mathcal{O}}_{K}}({\mathcal{M}}(\sigma)_{!}^{\mathop{\rm an}\nolimits},{\mathcal{M}}_{\psi}(1-k)_{B})$
[…]
Lemma 4.10.
There is an isomorphism ${\mathcal{M}}(\sigma)_{!}^{\mathop{\rm an}\nolimits}\cong{\mathcal{M}}(\sigma)_{N,!,B}$ identifying $\tilde{\delta}^{\mathop{\rm an}\nolimits}\otimes{\mathbb{Q}}$ with the pairing $\bar{\delta}^{L}_{!}$.
Proof.
Define ${\mathcal{F}}^{\prime}(\sigma)_{N}=G_{\mathbb{Q}}\backslash(G_{\mathbb{A}}\times{\mathcal{V}})/U_{N}U_{\infty}$. The same argument as in Lemma 3.9 shows that the natural projection $X_{N}^{\mathop{\rm an}\nolimits}\to X^{\mathop{\rm an}\nolimits}=X_{U}^{\mathop{\rm an}\nolimits}$ induces a map
${\mathop{\rm res}\nolimits}:H^{1}(X^{\mathop{\rm an}\nolimits},{\mathcal{F}}_{B}^{k}\otimes{\mathcal{F}}(\sigma))\to H^{1}(X_{N}^{\mathop{\rm an}\nolimits},{\mathcal{F}}_{N,B}^{k}\otimes{\mathcal{F}}^{\prime}(\sigma)_{N})^{U/U_{N}}$
with finite kernel and cokernel, and an isomorphism
${\mathop{\rm res}\nolimits}:H^{1}_{c}(X^{\mathop{\rm an}\nolimits},{\mathcal{F}}_{B}^{k}\otimes{\mathcal{F}}(\sigma))\to H^{1}_{c}(X_{N}^{\mathop{\rm an}\nolimits},{\mathcal{F}}_{N,B}^{k}\otimes{\mathcal{F}}^{\prime}(\sigma)_{N})^{U/U_{N}}.$
The sheaf ${\mathcal{F}}^{\prime}(\sigma)_{N}$ on $X_{N}$ is constant and the right hand sides become ${\mathcal{M}}(\sigma)_{N,B}$ and ${\mathcal{M}}(\sigma)_{N,c,B}$. This proves ${\mathcal{M}}(\sigma)_{\mathop{\rm tf}\nolimits}^{\mathop{\rm an}\nolimits}\hookrightarrow{\mathcal{M}}(\sigma)_{N,{\mathop{\rm tf}\nolimits},B}$ and ${\mathcal{M}}(\sigma)_{c}^{\mathop{\rm an}\nolimits}\cong{\mathcal{M}}(\sigma)_{N,c,B}$. We then get the desired isomorphism from the commutative diagram
[…]
in which the two top vertical maps are surjective and two bottom vertical maps are injective.
The compatibility of the pairings follows from the commutativity of the diagram
[…]
which is proved by the same argument as in Proposition 3.8 (with $g=I$). ∎
We now compute the adjoints of double coset operators under the pairings. Suppose now that $\sigma$, $\sigma^{\prime}$, $g$, $\tau$ and $S$ are as in §4.2. Recall that we defined morphisms
$U^{\prime}gU]_{\tau,\sharp}:M(\sigma)_{\sharp}^{S}\to M(\sigma^{\prime})_{\sharp}^{S}$
for $\sharp=\emptyset$, $c$ and !!. (We omit the subscripts $N,N^{\prime}$ of which they are independent.) It is easy to check that we then also have morphisms
[…]
Denote this operator by $U^{\prime}gU]_{\tau,\sharp}^{T}$.
Proposition 4.11.
The morphism $U^{\prime}gU]_{\tau,c}^{T}$ (respectively, $U^{\prime}gU]_{\tau,!}^{T}$) is the adjoint of $U^{\prime}gU]_{\tau}$ (respectively, $U^{\prime}gU]_{\tau,!}$) with respect to the pairing $\bar{\delta}^{L}$, (respectively, $\bar{\delta}^{L}_{!}$), i.e., we have the commutative diagram
[…]
and a similar diagram for $\bar{\delta}^{L}_{!}$.
Proof.
We first make some simplifications. Let $\sigma:U\to\operatorname{{\mathop{\rm Aut}\nolimits}}_{K}V$, $\sigma^{\prime}:U^{\prime}\to\operatorname{{\mathop{\rm Aut}\nolimits}}_{K}V^{\prime}$ and $\tau:V\to V^{\prime}$ be as in the definition of $U^{\prime}gU]_{\tau}$. Consider the diagram
[…]
Here we have used the notations
$U^{iv}=gUg^{-1},V^{iv}=V,$
$\sigma^{iv}:U^{iv}\to\operatorname{{\mathop{\rm Aut}\nolimits}}_{K}V^{iv},\sigma^{iv}(u^{iv})v^{iv}=\sigma(g^{-1}u^{iv}g)v^{iv},$
$g^{iv}=g,\tau^{iv}={\mathop{\rm id}\nolimits}:V\to V^{iv},$
$U^{\prime\prime\prime}=U^{\prime}\cap gUg^{-1},V^{\prime\prime\prime}=V,\sigma^{\prime\prime\prime}=\sigma_{|U^{\prime\prime\prime}}:U^{\prime\prime\prime}\to\operatorname{{\mathop{\rm Aut}\nolimits}}_{K}V,$
$g^{\prime\prime\prime}=1,\tau^{\prime\prime\prime}={\mathop{\rm id}\nolimits},$
$U^{\prime\prime}=U^{\prime}\cap gUg^{-1},V^{\prime\prime}=V^{\prime},\sigma^{\prime\prime}=\sigma^{\prime}_{|U^{\prime\prime}}:U^{\prime\prime}\to\operatorname{{\mathop{\rm Aut}\nolimits}}_{K}V^{\prime\prime},$
$g^{\prime\prime}=1,\tau^{\prime\prime}=\tau\quad{\rm\ and\ }$
$g^{\prime}=1,\tau^{\prime}={\mathop{\rm id}\nolimits}:V^{\prime\prime}=V^{\prime}\to V^{\prime}.$
Then we have morphisms
$U^{iv}g^{iv}U]_{\tau^{iv}}:M(\sigma)\to M(\sigma^{iv}),$
$U^{\prime\prime\prime}g^{\prime\prime\prime}U^{iv}]_{\tau^{\prime\prime\prime}}:M(\sigma^{iv})\to M(\sigma^{\prime\prime\prime}),$
$U^{\prime\prime}g^{\prime\prime}U^{\prime\prime\prime}]_{\tau^{\prime\prime}}:M(\sigma^{\prime\prime\prime})\to M(\sigma^{\prime\prime})\quad{\rm\ and\ }$
$U^{\prime}g^{\prime}U^{\prime\prime}]_{\tau^{\prime}}:M(\sigma^{\prime\prime})\to M(\sigma^{\prime}).$
Also
$\tau=\tau^{\prime}\circ\tau^{\prime\prime}\circ\tau^{\prime\prime\prime}\circ\tau^{iv}.$
Applying Proposition 4.4, we have
$[U^{\prime}g^{\prime}U^{\prime\prime}]_{\tau^{\prime}}\circ[U^{\prime\prime}g^{\prime\prime}U^{\prime\prime\prime}]_{\tau^{\prime\prime}}\circ[U^{\prime\prime\prime}g^{\prime\prime\prime}U^{iv}]_{\tau^{\prime\prime\prime}}\circ[U^{iv}g^{iv}U]_{\tau^{iv}}$ $=$ $[U^{\prime}gU]_{\tau}.$
We similarly verify that
[…]
$\circ[U^{iv}(||{\mathop{\rm det}\nolimits}g{{}^{\prime\prime\prime}}||g{{}^{\prime\prime\prime}})^{-1}U{{}^{\prime\prime\prime}}]_{(\tau{{}^{\prime\prime\prime}})^{t}\otimes\psi({\mathop{\rm det}\nolimits}(g{{}^{\prime\prime\prime}})),c}$
$\circ[U{{}^{\prime\prime\prime}}(||{\mathop{\rm det}\nolimits}g{{}^{\prime\prime}}||g{{}^{\prime\prime}})^{-1}U{{}^{\prime\prime}}]_{(\tau{{}^{\prime\prime}})^{t}\otimes\psi({\mathop{\rm det}\nolimits}(g{{}^{\prime\prime}})),c}$
$\circ[U{{}^{\prime\prime}}(||{\mathop{\rm det}\nolimits}g{{}^{\prime}}||g{{}^{\prime}})^{-1}U{{}^{\prime}}]_{(\tau{{}^{\prime}})^{t}\otimes\psi({\mathop{\rm det}\nolimits}(g{{}^{\prime}})),c}$
[…]
$\circ[(U^{\prime}\cap(gUg^{-1}))1(U^{\prime}\cap(gUg^{-1}))]_{\tau^{t}\otimes\psi(1),c}\circ[(U^{\prime}\cap((gUg^{-1}))1U{{}^{\prime}}]_{{\mathop{\rm id}\nolimits}\otimes\psi(1),c}$
[…]
Thus to prove the proposition for $\bar{\delta}^{L}$, we only need to consider the four composition factors, i.e., to prove it in each of the following four special cases.
  1. Case 1. $U^{\prime}=gUg^{-1}$ and $\tau={\mathop{\rm id}\nolimits}$,
  2. Case 2. $U^{\prime}\subset U$, $g=1$ and $\tau={\mathop{\rm id}\nolimits}$,
  3. Case 3. $U^{\prime}=gUg^{-1}$ and $g=1$, and
  4. Case 4. $U^{\prime}\supset U$, $g=1$ and $\tau={\mathop{\rm id}\nolimits}$.
Furthermore, we only need to consider the Betti realization. In this case the desired commutativity is equivalent to that of the diagram
[…]
where [,][\ ,\ ] and [,][\ ,\ $^{\prime}$ are the pairings induced by $\bar{\delta}^{L}$ for $\sigma$ and $\sigma^{\prime}$.
Case 1: When $U^{\prime}=gUg^{-1}$ and $\tau={\mathop{\rm id}\nolimits}$, we have $U^{\prime}gU=gU$. Let $U_{N}$ and $U_{N^{\prime}}$ be as chosen in the definition of $U^{\prime}gU]_{\mathop{\rm id}\nolimits}=[U^{\prime}gU]_{{\mathop{\rm id}\nolimits},N,N^{\prime}}$. So the action of $U^{\prime}gU]_{{\mathop{\rm id}\nolimits}}$ on $M_{N,B}\otimes V$ is
$U^{\prime}gU]_{{\mathop{\rm id}\nolimits}}(x\otimes v)=[g]x\otimes v,x\in M_{N,B},v\in V.$
Using Proposition 3.8, the definition of [,][\ ,\ ] and the fact that $\tau^{t}={\mathop{\rm id}\nolimits}$, we see that [,][\ ,\ ] (resp. [,][\ ,\ $^{\prime}$) is from the composite of the maps in the top row (resp. bottom row) of the following commutative diagram
[…]
Restricted to the suitable $U$- or $U^{\prime}$-invariants, it gives
[…]
We also note that
(31) $Ug^{-1}U^{\prime}]_{{\mathop{\rm id}\nolimits}\otimes(\psi\circ{\mathop{\rm det}\nolimits}(g)),c}=([U^{\prime}gU]_{{\mathop{\rm id}\nolimits}\otimes(\psi^{-1}\circ{\mathop{\rm det}\nolimits}(g)),c})^{-1}$
as maps $M(\hat{\sigma}^{\prime}\otimes(\psi^{-1}\circ{\mathop{\rm det}\nolimits}))_{c,B}\to M(\hat{\sigma}\otimes(\psi^{-1}\circ{\mathop{\rm det}\nolimits}))_{c,B}$. Combining (30) and (31) proves the formula in this case.
Case 4. Assume $U^{\prime}\supset U$ with $g=1$ and $\tau={\mathop{\rm id}\nolimits}$. So $\sigma=\sigma^{\prime}_{|U}$. Let $U_{N}$ be a sufficiently small subgroup of $U^{\prime}\cap(g^{-1}Ug)=U$ and let $U^{\prime}=\coprod_{i}g_{i}U$ be a coset decomposition. Then the action of $U^{\prime}gU]_{\tau}$ on $M(\sigma)_{B}$ is the restriction of the map
$U^{\prime}gU]_{\tau}(x\otimes v)=\sum_{i}[g_{i}]x\otimes g_{i}v,\ x\in M_{N,B},\ v\in V.$
For each fixed $i$, consider the setting of the proposition with $U=U^{\prime}=U_{N}$, $g=g_{i}$, $V=V^{\prime}$ being the given $V$ (restricted to $U_{N}$) and $\tau=g_{i}$. Then for any $u\in U$ and $v\in V$, $\tau((g_{i}^{-1}ug_{i})v)=g_{i}(g_{i}^{-1}ug_{i})v=ug_{i}v=u\tau(v).$ Thus $U_{N}g_{i}U_{N}]_{\tau,\sharp}:M(\sigma|_{U_{N}})_{\sharp,B}\to M(\sigma|_{U_{N}})_{\sharp,B}$ is defined. In fact, we are in Case 1 of the proposition. Let $_{N}^{N}$ be the pairing
$M((\hat{\sigma}\otimes(\psi^{-1}\circ{\mathop{\rm det}\nolimits}))|_{U_{N}})_{c,B}\otimes M(\sigma|_{U_{N}})_{B}\to M_{\psi}(1-k)_{B}.$
Then by Case 1 of the proposition, we have
$U_{N}g_{i}^{-1}U_{N}]_{g_{i}^{t}\otimes(\psi^{-1}\circ{\mathop{\rm det}\nolimits}(g_{i}))}\alpha,\beta]_{N}^{N}=[\alpha,[U_{N}g_{i}U_{N}]_{g_{i}}\beta]_{N}^{N}$
for $\alpha\in M(\hat{\sigma}\otimes(\psi^{-1}\circ{\mathop{\rm det}\nolimits})|_{U_{N}})_{N,c,B}$ and $\beta\in M(\sigma|_{U_{N}})_{N,B}.$ In other words,
$(g_{i}^{-1}\otimes(g_{i}^{t}\otimes(\psi^{-1}\circ{\mathop{\rm det}\nolimits}(g_{i})))\alpha,\beta]_{N}=[\alpha,(g_{i}\otimes g_{i})\beta]_{N}.$
Take $\alpha\in M(\hat{\sigma}\otimes(\psi^{-1}\circ{\mathop{\rm det}\nolimits}))_{N,c,B}$ and $\beta\in M(\sigma)_{N,B}$. Then we have
$\alpha,\beta]_{N}^{N}=[\alpha,(g_{i}\otimes g_{i})\beta]_{N}^{N}.$
Going back to our original $U^{\prime}\supseteq U$, $g=1$ and $\tau={\mathop{\rm id}\nolimits}$, we have
$\alpha,[U^{\prime}1U]_{\mathop{\rm id}\nolimits}\beta]_{N}^{N}=[\alpha,\sum_{i}(g_{i}\otimes g_{i})\beta]_{N}^{N}=[U^{\prime}:U][\alpha,\beta]_{N}^{N}.$
Dividing it by $U^{\prime}:U_{N}]$, we have
$[\alpha,[U^{\prime}1U]_{\mathop{\rm id}\nolimits}\beta]^{\prime}=\frac{1}{[U^{\prime}:U_{N}]}[\alpha,[U^{\prime}1U]_{\mathop{\rm id}\nolimits}\beta]_{N}^{N}=\frac{[U^{\prime}:U]}{[U^{\prime}:U_{N}]}[\alpha,\beta]_{N}^{N}$
$=$ $\frac{1}{[U:U_{N}]}[\alpha,\beta]_{N}^{N}=[\alpha,\beta]=[[U1U^{\prime}]_{{\mathop{\rm id}\nolimits}\otimes(\psi^{-1}\circ{\mathop{\rm det}\nolimits}(1))}\alpha,\beta].$
This proves the proposition in this case.
Case 2: This case follows by exchanging $U$ and $U^{\prime}$ in Case 4.
Case 3. In this case $\tau:V\to V^{\prime}$ is a morphism of $U$-modules and the formula follows directly from the definitions. ∎
4.5. Premotivic structure of level NN and character $\psi$
[…]
Recall that the isomorphism
${\mathbb{C}}\otimes\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}{\mathcal{M}}_{M,{\mathop{\rm dR}\nolimits}}\cong\oplus_{t^{\prime}\in({\mathbb{Z}}/M{\mathbb{Z}})^{\times}}M_{k}(\Gamma(M))\cong M_{k}(U_{M})$
[…]
$f\left(x{\mbox{\scriptsize$\left(\begin{array}[]{cc}{{a}}&{{b}}\\ {{c}}&{{d}}\end{array}\right)$}}\right)=\psi(a_{N})f(x)$
for all ${\mbox{\scriptsize$\left(\begin{array}[]{cc}{{a}}&{{b}}\\ {{c}}&{{d}}\end{array}\right)$}}\in U_{0}(N)$. It is straightforward to check that this is precisely the space $M_{k}(N,\psi)$ of classical modular forms of weight $k$, level NN and character $\psi$, i.e., the set of forms $f\in M_{k}(\Gamma(M))$ such that
$f\left(\frac{a\tau+b}{c\tau+d}\right)=\psi(d)(c\tau+d)^{k}f(\tau)$
for all $\tau\in{\mathfrak{H}}$, $\gamma={\mbox{\scriptsize$\left(\begin{array}[]{cc}{{a}}&{{b}}\\ {{c}}&{{d}}\end{array}\right)$}}\in\Gamma_{0}(N)=U_{0}(N)\cap\operatorname{{\mathop{\rm GL}\nolimits}}_{2}^{+}({\mathbb{Q}})$. We thus obtain an isomorphism
${\mathbb{C}}\otimes_{K}\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}{\mathcal{M}}(N,\psi)_{M,{\mathop{\rm dR}\nolimits}}\cong M_{k}(N,\psi).$
The same holds for ${\mathcal{M}}(N,\psi)_{M,{\mathop{\rm tf}\nolimits}}$, and replacing this by ${\mathcal{M}}(N,\psi)_{M,c}$ or ${\mathcal{M}}(N,\psi)_{M,!}$ gives the space of cusp forms $S_{k}(N,\psi)$. Taking the product over all embeddings $K\to{\mathbb{C}}$, we obtain isomorphisms
(32) $\begin{array}[]{lllll}{\mathbb{C}}\otimes\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}{\mathcal{M}}(N,\psi)_{M,{\mathop{\rm dR}\nolimits}}&\cong&{\mathbb{C}}\otimes\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}{\mathcal{M}}(N,\psi)_{M,{\mathop{\rm tf}\nolimits},{\mathop{\rm dR}\nolimits}}&\cong&M_{k}(N,\psi)^{{\bf I}_{K}}\\ {\mathbb{C}}\otimes\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}{\mathcal{M}}(N,\psi)_{M,c,{\mathop{\rm dR}\nolimits}}&\cong&{\mathbb{C}}\otimes\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}{\mathcal{M}}(N,\psi)_{M,!,{\mathop{\rm dR}\nolimits}}&\cong&S_{k}(N,\psi)^{{\bf I}_{K}}\end{array}$
of ${\mathbb{C}}^{{\bf I}_{K}}\cong{\mathbb{C}}\otimes K$-modules.
Tensoring the $q$-expansion maps of (6) with $K$, we regard the $q$-expansion of a form in $M_{k}(N,\psi)^{{\bf I}_{K}}$ as taking values in $(R^{\prime}\otimes{\mathbb{C}}\otimes K)[[q^{1/M}]]$ where $R^{\prime}={\mathbb{Z}}[1/M,\mu_{M}]$.
Lemma 4.12.
For any $S\supset S_{N}^{K}$, the isomorphisms of (32) identify
$\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}{\mathcal{M}}(N,\psi)_{M,{\mathop{\rm dR}\nolimits}}^{S}=\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}{\mathcal{M}}(N,\psi)_{M,{\mathop{\rm tf}\nolimits},{\mathop{\rm dR}\nolimits}}^{S}$
(respectively,
$\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}{\mathcal{M}}(N,\psi)_{M,c,{\mathop{\rm dR}\nolimits}}^{S}\cong\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}{\mathcal{M}}(N,\psi)_{M,!,{\mathop{\rm dR}\nolimits}}^{S})$
with the set of forms in $M_{k}(N,\psi)^{{\bf I}_{K}}$ (respectively, $S_{k}(N,\psi)^{{\bf I}_{K}}$) having $q$-expansions in ${\mathcal{O}}_{S}[[q]]$.
Proof.
We may assume $S=S_{N}^{K}$. Tensoring (6) with ${\mathcal{O}}_{S}$ and taking $U_{0}(N)$-invariants gives a pull-back diagram
[…]
In particular, if $f\in M_{k}(N,\psi)^{{\bf I}_{K}}$ has $q$-expansion in ${\mathcal{O}}_{S}[[q]]\subset(R^{\prime}\otimes{\mathcal{O}}_{S})[[q^{1/M}]]$, then $f$ is in the image of $\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}{\mathcal{M}}(N,\psi)_{M,{\mathop{\rm dR}\nolimits}}$. Conversely, if $f$ is in the image of $\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}{\mathcal{M}}(N,\psi)_{M,{\mathop{\rm dR}\nolimits}}$, then its $q$-expansion is in $(R^{\prime}\otimes{\mathcal{O}}_{S})[[q^{1/M}]]$. Viewing $f\in M_{k}(U_{M})\otimes K$, we see that $f$ is invariant under the action of ${\mbox{\scriptsize$\left(\begin{array}[]{cc}{{1}}&{{1}}\\ {{0}}&{{1}}\end{array}\right)$}}\in\operatorname{{\mathop{\rm GL}\nolimits}}_{2}({\mathbb{A}}_{\bf f})$ and under that of $g_{t^{\prime}}$ for each $t^{\prime}\in({\mathbb{Z}}/M{\mathbb{Z}})^{\times}$. Unravelling the effect of $\left(\begin{array}[]{cc}{{1}}&{{1}}\\ {{0}}&{{1}}\end{array}\right)$ on $q$-expansions shows that $f$ has $q$-expansion in $(R^{\prime}\otimes{\mathbb{C}}\otimes K)[[q]]$, hence in $(R^{\prime}\otimes{\mathcal{O}}_{S})[[q]]$. The commutativity of the diagram (6) together with the invariance under the $g_{t^{\prime}}$ shows that the image of $f$ in $({\mathbb{C}}\otimes K)[[q]]$ is independent of $t^{\prime}$, i.e., independent of the map $R^{\prime}\to{\mathbb{C}}$. We conclude that the $q$-expansion of $f$ is invariant under $\operatorname{\operatorname{{\mathop{\rm Aut}\nolimits}}}R^{\prime}$, hence has coefficients in ${\mathcal{O}}_{S}$. ∎
5. Premotivic structure of a newform
5.1. Hecke actions
Using the notations in §4.5 and letting $S$ be a superset of $S_{N}^{K}$, we now consider the action of Hecke operators on ${\mathcal{M}}(N,\psi)_{M,!}$ as an object in ${\mathcal{P}M}^{S}_{K}$.
Fix a rational prime pp. By Lemma 4.2, $U{\mbox{\scriptsize$\left(\begin{array}[]{cc}{{p}}&{{0}}\\ {{0}}&{{1}}\end{array}\right)$}}_{p}U]_{\psi(p_{p})^{-1}}$ defines an endomorphism of ${\mathcal{M}}(N,\psi)_{M,!}$ regarded as an object in ${\mathcal{P}M}^{S_{pN}}$. This induces an endomorphism
$U{\mbox{\scriptsize$\left(\begin{array}[]{cc}{{p}}&{{0}}\\ {{0}}&{{1}}\end{array}\right)$}}_{p}U]_{\psi(p_{p})^{-1}}:M(N,\psi)_{M,!}\to M(N,\psi)_{M,!}$
in $\operatorname{{\bf PM}}^{S_{pN}}$.
Lemma 5.1.
  1. (a)
    The action of the usual Hecke operator $T_{p}$ on $M_{k}(N,\psi)$ corresponds to that of the double coset operator $U{\mbox{\scriptsize$\left(\begin{array}[]{cc}{{p}}&{{0}}\\ {{0}}&{{1}}\end{array}\right)$}}_{p}U]_{\psi(p_{p})^{-1}}$ on $M(N,\psi)_{M}$, regarded as an object in $\operatorname{{\bf PM}}^{S_{pN}}_{K}$.
  2. (b)
    If $p\nmid N$, then the operator $U{\mbox{\scriptsize$\left(\begin{array}[]{cc}{{p}}&{{0}}\\ {{0}}&{{p}}\end{array}\right)$}}_{p}U]_{\psi(p_{p})^{-2}}$ acts on $M(N,\psi)_{M,c}$ via $\psi(p_{p})^{-1}p^{k-2}$. This corresponds to the usual Hecke operator $S_{p}$ on $M_{k}(N,\psi)$.
Proof.
(a) Denote $g={\mbox{\scriptsize$\left(\begin{array}[]{cc}{{p}}&{{0}}\\ {{0}}&{{1}}\end{array}\right)$}}_{p}$ and let $U=\coprod_{i}g_{i}(gUg^{-1})\cap U$ be a coset representation. Then we have double coset decomposition $UgU=\coprod h_{i}U$ with $h_{i}=g_{i}g$. The usual Hecke operator $T_{p}$ on $M_{k}(\Gamma_{1}(M))$ corresponds to that of the double coset operator $T^{\mathop{\rm ad}\nolimits}_{p}=[U{\mbox{\scriptsize$\left(\begin{array}[]{cc}{{p}}&{{0}}\\ {{0}}&{{1}}\end{array}\right)$}}_{p}U]$ on the adelic modular forms $M_{k}(U_{1}(M))$ [D-I, §11.1]. By definition,
$T_{p}^{\mathop{\rm ad}\nolimits}=[U{\mbox{\scriptsize$\left(\begin{array}[]{cc}{{p}}&{{0}}\\ {{0}}&{{1}}\end{array}\right)$}}_{p}U]=\sum_{i}h_{i}.$
By Proposition 3.7, this corresponds to the operator $T_{p}^{\rm mot}\stackrel{{\scriptstyle\rm def}}{{=}}\sum_{i}[h_{i}]$ on ${\mathbb{C}}\otimes\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}{\mathcal{M}}_{M,!,{\mathop{\rm dR}\nolimits}}$. Therefore the operator $\sum_{i}h_{i}$ on $(V\otimes M_{k}(U_{1}(M)))^{U}$ corresponds to the operator $\sum_{i}h_{i}$ on $(V\otimes{\mathbb{C}}\otimes\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}{\mathcal{M}}_{M,!,{\mathop{\rm dR}\nolimits}})^{U}={\mathbb{C}}\otimes{\mathcal{M}}(N,\psi)_{M,!,{\mathop{\rm dR}\nolimits}}$. The second operator is defined by
$\sum_{i}h_{i}(\sum_{j}x_{j}\otimes v_{j})=\sum_{i,j}[g_{i}g]x_{j}\otimes g_{i}gv_{j}=\sum_{i,j}[g_{i}g]x_{j}\otimes g_{i}\psi(p_{p})^{-1}v_{j}=[UgU]_{\psi(p_{p})^{-1}}.$
On the other hand, if $f$ is in $M_{k}(U,\psi)$, then $1\otimes f$ is in $(V\otimes M_{k}(U_{1}(M)))^{U}$ and $f\mapsto 1\otimes f$ defines an isomorphism $S_{k}(U,\psi)\cong(V\otimes M_{k}(U_{1}(M)))^{U}$. Then we have
$\sum_{i}h_{i}(1\otimes f)=\sum_{i}(\psi(h_{i})^{-1}\otimes h_{i}f)=1\otimes(\sum_{i}\psi(h_{i})^{-1}h_{i}f).$
[…]
$\sum_{i}\left[{\mbox{\scriptsize$\left(\begin{array}[]{cc}{{p^{\prime}}}&{{0}}\\ {{0}}&{{p^{\prime}}}\end{array}\right)$}}\right]x_{i}\otimes\psi^{-1}(p^{\prime})v_{i}=\sum_{i}x_{i}\otimes v_{i}.$
Thus
$\sum_{i}\left[{\mbox{\scriptsize$\left(\begin{array}[]{cc}{{p^{\prime}}}&{{0}}\\ {{0}}&{{p^{\prime}}}\end{array}\right)$}}\right]x_{i}\otimes v_{i}=\sum_{i}x_{i}\otimes\psi(p^{\prime})v_{i}=\psi(p^{\prime})\sum_{i}x_{i}\otimes v_{i}.$
Then we have
[…]
Because of the lemma, we can write $T_{p}$ (resp. $S_{p}$) for the endomorphisms of $\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}M(N,\psi)_{M,{\mathop{\rm dR}\nolimits}}$, $\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}M(N,\psi)_{M,c,{\mathop{\rm dR}\nolimits}}$ and $\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}M(N,\psi)_{M,!,{\mathop{\rm dR}\nolimits}}$ defined by the coset operator $U{\mbox{\scriptsize$\left(\begin{array}[]{cc}{{p}}&{{0}}\\ {{0}}&{{1}}\end{array}\right)$}}_{p}U]_{\psi(p_{p})^{-1}}$ (resp. $U{\mbox{\scriptsize$\left(\begin{array}[]{cc}{{a}}&{{0}}\\ {{0}}&{{a}}\end{array}\right)$}}U]_{\psi(p_{p})^{-2}}$).
5.2. Compatibility of the Hecke action with the pairings
Let $w$ denote ${\mbox{\scriptsize$\left(\begin{array}[]{cc}{{0}}&{{-1}}\\ {{N}}&{{0}}\end{array}\right)$}}_{N}$ in $\prod_{p|N}\operatorname{{\mathop{\rm GL}\nolimits}}_{2}({\mathbb{Z}}_{p})$. Let $V^{\prime}$ be the one dimensional representation
${\mathop{\rm Hom}\nolimits}(V,K)\otimes K_{\psi^{-1}\circ{\mathop{\rm det}\nolimits}}$
of $U$. We abbreviate $\sigma^{\prime}$ for $\hat{\sigma}_{\psi}\otimes(\psi^{-1}\circ{\mathop{\rm det}\nolimits})$. Fix a basis $v_{0}$ of $\mathcal{V}$}. Define $\omega:{\mathcal{V}}\to{\mathcal{V}}^{\prime}$ by sending $v_{0}$ to $\hat{v}_{0}\otimes 1$, where $\hat{v}_{0}\in{\mathop{\rm Hom}\nolimits}({\mathcal{V}},{\mathcal{O}})$ is such that $\hat{v}_{0}(v_{0})=1$.
Lemma 5.2.
  1. (a)
    The operator $UwU]_{\omega}$ defines an isomorphism
    ${\mathcal{M}}(N,\psi)_{M}\to{\mathcal{M}}(\sigma^{\prime})_{M}$
    in ${\mathcal{P}M}^{S_{N}^{K}}_{K}$ and $UwU]_{\omega}^{-1}=[Uw^{-1}U]_{\omega^{t}\otimes\psi({\mathop{\rm det}\nolimits}w)}$. Similarly for ${\mathcal{M}}(N,\psi)_{M,c}$ and ${\mathcal{M}}(N,\psi)_{M,!}$.
  2. (b)
    The operator $UwU]_{\omega}:{\mathcal{M}}(\psi)_{M,!}\to{\mathcal{M}}(N,\sigma^{\prime})_{M,!}$ is adjoint to
    $N^{k-2}[Uw^{-1}U]_{\omega^{t}\otimes\psi({\mathop{\rm det}\nolimits}w))}:{\mathcal{M}}(\sigma^{\prime})_{M,!}\to{\mathcal{M}}(N,\psi)_{M,!}$
    and coincides with $\psi(-1_{\infty})N^{k-2}[Uw^{-1}U]_{\omega}$.
Proof.
[…]
For $u={\mbox{\scriptsize$\left(\begin{array}[]{cc}{{a}}&{{b}}\\ {{c}}&{{d}}\end{array}\right)$}}\in U$ we have
$\sigma(w^{-1}uw)v_{0}=\sigma{\mbox{\scriptsize$\left(\begin{array}[]{cc}{{d}}&{{-\frac{c}{N_{N}}}}\\ {{-N_{N}b}}&{{a}}\end{array}\right)$}}v_{0}=\psi^{-1}(d_{N})v_{0}$
and
$\sigma^{\prime}(u)(\hat{v}_{0}\otimes 1)=\psi^{-1}((({\mathop{\rm det}\nolimits}u)^{-1}d)_{N})\hat{v}_{0}\otimes\psi^{-1}({\mathop{\rm det}\nolimits}u)\cdot 1=\psi^{-1}(d_{N})\hat{v}_{0}\otimes 1.$
since $\psi({\mathop{\rm det}\nolimits}u/({\mathop{\rm det}\nolimits}u)_{N})=1$. Therefore we have
$\omega(\sigma(w^{-1}uw)v_{0})=\sigma^{\prime}(u)\omega(v_{0}).$
Also $w^{-1}U_{M^{2}}w\subseteq U_{M}$. So the operator
$UwU]_{\omega,M,M^{2}}:{\mathcal{M}}(N,\psi)_{M}\to{\mathcal{M}}(\sigma^{\prime})_{M^{2}}$
is well-defined and defines an morphism in ${\mathcal{P}M}^{S_{N}}_{K}$. Since ${\mathcal{M}}(\sigma^{\prime})_{M^{2}}\cong{\mathcal{M}}(\sigma^{\prime})_{M}$ by Lemma 4.1 (a), we have
$UwU]_{\omega}:{\mathcal{M}}(N,\psi)_{M}\to{\mathcal{M}}(\sigma^{\prime})_{M}.$
To prove that the map is an isomorphism, consider $w={\mbox{\scriptsize$\left(\begin{array}[]{cc}{{0}}&{{1}}\\ {{-N}}&{{0}}\end{array}\right)$}}_{N}$. The same argument as above shows that we have a homomorphism
$U(-w)U]_{\omega^{\prime}}:{\mathcal{M}}(\sigma^{\prime})_{M}\to{\mathcal{M}}(\sigma^{\prime\prime})_{M}$
in ${\mathcal{P}M}^{S_{N}}_{K}$ where $\sigma^{\prime\prime}$ is the representation of $U$ on ${\mathcal{V}}^{\prime\prime}={\mathop{\rm Hom}\nolimits}(\hat{{\mathcal{V}}}^{\prime},{\mathcal{O}})\otimes{\mathcal{O}}_{\psi^{-1}\circ{\mathop{\rm det}\nolimits}}$ which is canonically isomorphic to $\sigma_{\psi}$ and where $\omega^{\prime}:{\mathcal{V}}^{\prime}\to{\mathcal{V}}^{\prime\prime}$ sends $\hat{v}_{0}$ to $\hat{\hat{v}}_{0}\otimes 1$ which is identified with $v_{0}$ under the above isomorphism. Thus we have
$U(-w)U]_{\omega^{\prime}}\circ[UwU]_{\omega}:{\mathcal{M}}(N,\psi)_{M}\to{\mathcal{M}}(N,\psi)_{M}.$
Since $U(-w)U][UwU]=[U(-w)wU]=[U(N_{N}I)U],$ the composition formula in Proposition 4.4 gives us
$U(-w)U]_{\omega^{\prime}}\circ[UwU]_{\omega}=[U(NI)U]_{1}=N^{k-2}.$
Since NN is invertible in $\mathcal{O}$}, $UwU]_{\omega}$ is an isomorphism in ${\mathcal{P}M}^{S_{N}^{K}}_{K}$ and
$UwU]_{\omega}^{-1}=N^{2-k}[U(-w)U]_{\omega^{\prime}}=[Uw^{-1}U]_{\omega^{t}\otimes\psi({\mathop{\rm det}\nolimits}w)}.$
The statements for ${\mathcal{M}}_{c}$ and $\mathcal{M}$$_{!}$ follow since $UwU]_{\omega}$ commutes with the weight filtration maps.
(b) By Lemma 4.11, the operator $UwU]_{\omega}:{\mathcal{M}}(N,\psi)_{M,!}\to{\mathcal{M}}(\sigma^{\prime})_{M,!}$ is adjoint to $N^{k-2}[Uw^{-1}U]_{\omega^{\prime}}:{\mathcal{M}}(\sigma^{\prime})_{M,!}\to{\mathcal{M}}(N,\psi)_{M,!}$.
On the other hand,
$w={\mbox{\scriptsize$\left(\begin{array}[]{cc}{{0}}&{{-N}}\\ {{1}}&{{0}}\end{array}\right)$}}_{N}={\mbox{\scriptsize$\left(\begin{array}[]{cc}{{-1}}&{{0}}\\ {{0}}&{{-1}}\end{array}\right)$}}_{N}{\mbox{\scriptsize$\left(\begin{array}[]{cc}{{N}}&{{0}}\\ {{0}}&{{N}}\end{array}\right)$}}_{N}w^{-1}.$
Thus by Proposition 4.4,
$[UwU]_{\omega}$ $=$ $[Uw^{-1}U]_{\omega}\circ[U{\mbox{\scriptsize$\left(\begin{array}[]{cc}{{N}}&{{0}}\\ {{0}}&{{N}}\end{array}\right)$}}_{N}U]_{1}\circ[U{\mbox{\scriptsize$\left(\begin{array}[]{cc}{{-1}}&{{0}}\\ {{0}}&{{-1}}\end{array}\right)$}}_{N}U]_{1}$
$=$ $\psi^{-1}(-1_{\infty})N^{k-2}[Uw^{-1}U]_{\omega}$
since $\sigma({\mbox{\scriptsize$\left(\begin{array}[]{cc}{{-1}}&{{0}}\\ {{0}}&{{-1}}\end{array}\right)$}}_{N})=\psi^{-1}(-1_{N})=\psi^{-1}(-1_{f})=\psi(-1_{\infty})$. ∎
Composing the operator $UwU]_{\omega}$ with the duality morphism $\bar{\delta}^{L}_{N}$ defined in (29), we obtain a duality morphism
(34) $\hat{\delta}^{L}_{N}:{\mathcal{M}}(N,\psi)_{M,!}\to{\mathop{\rm Hom}\nolimits}_{\mathcal{O}}({\mathcal{M}}(N,\psi)_{M,!},{\mathcal{M}}_{\psi}(1-k))$
that becomes an isomorphism after tensoring with $\mathbb{Q}$}.
Let $U^{\prime}=U_{0}(N^{\prime})$ and let $\psi^{\prime}$ be a character of $\mathbb{A}$$^{\times}$ with conductor dividing $N^{\prime}$. Then we can construct the premotivic structure ${\mathcal{M}}(N^{\prime},\psi^{\prime})_{M^{\prime},!}$ associated to $\psi^{\prime}$ in the same way as we construct ${\mathcal{M}}(N,\psi)_{M,!}$. Given $U^{\prime}gU]_{\tau,M,M^{\prime}}:{\mathcal{M}}(N,\psi)_{M,!}\to{\mathcal{M}}(N^{\prime},\psi^{\prime})_{M^{\prime},!}$ where $M^{\prime}$ is chosen with the additional condition $g^{-1}U_{M^{\prime}}g\subseteq U_{M}$, let $U^{\prime}gU]_{\tau}^{t}$ denote the adjoint of $U^{\prime}gU]_{\tau}$ with respect to the pairing morphism $\hat{L}$.
Lemma 5.3.
  1. (a)
[…]
    in ${\mathcal{P}M}^{S_{M^{\prime\prime}}}$ for any $M^{\prime\prime}$ with $g^{-1}U_{M^{\prime\prime}}g\subseteq U_{M^{\prime}}$.
  2. (b)
    For any prime $p$, $T_{p}$ is self-adjoint under the duality morphism $\hat{\delta}^{L}$ in ${\mathcal{P}M}^{S_{pN}}$.
Proof.
(a) We only need to verify for the Betti realization. Let [,][\ ,\ ], [,][\ ,\ $^{\prime}$, $\langle\ ,\ \rangle$ and $\langle\ ,\ \rangle^{\prime}$ be the pairings on Betti realizations induced by $\bar{\delta}^{L}_{N},\bar{\delta}^{L}_{N^{\prime}},\hat{\delta}^{L}_{N}$ and $\hat{\delta}^{L}_{N^{\prime}}$. By definition we have the commutative diagram
[…]
Then from Proposition 4.11 and Lemma 5.2
$\langle[U^{\prime}gU]_{\tau}$ $a,b\rangle^{\prime}=[[U^{\prime}gU]_{\tau}a,[U^{\prime}w^{\prime}U^{\prime}]_{\omega^{\prime}}b]^{\prime}$
$=[a,[U^{\prime}gU]_{\tau}^{T}\circ[U^{\prime}w^{\prime}U^{\prime}]_{\omega^{\prime}}b]$
$=\langle a,[UwU]_{\omega}^{-1}\circ[U^{\prime}gU]^{T}\circ[U^{\prime}w^{\prime}U^{\prime}]_{\omega^{\prime}}b\rangle$
$=\langle a,[Uw^{-1}U]_{\omega^{t}\otimes\psi({\mathop{\rm det}\nolimits}w)}\circ[U^{\prime}gU]^{T}\circ[U^{\prime}w^{\prime}U^{\prime}]_{\omega^{\prime}}b\rangle$
[…]
This proves (a).
(b) Note that $T_{p}^{T}=[U{\mbox{\scriptsize$\left(\begin{array}[]{cc}{{1}}&{{0}}\\ {{0}}&{{p_{p}}}\end{array}\right)$}}U]_{\psi(p_{p})^{-1}}$ by Proposition 4.11. Then using Proposition 4.4 we find that $UwU][U{\mbox{\scriptsize$\left(\begin{array}[]{cc}{{p_{p}}}&{{0}}\\ {{0}}&{{1}}\end{array}\right)$}}U]$ and $U{\mbox{\scriptsize$\left(\begin{array}[]{cc}{{1}}&{{0}}\\ {{0}}&{{p_{p}}}\end{array}\right)$}}U][UwU]$ are both $U{\mbox{\scriptsize$\left(\begin{array}[]{cc}{{0}}&{{-1}}\\ {{p_{p}N_{N}}}&{{0}}\end{array}\right)$}}U]$. It follows that
$UwU]_{\omega}T_{p}=T_{p}^{T}[UwU]_{\omega}.$
Then from (a) we have
$T_{p}^{t}=[UwU]_{\omega}^{-1}T_{p}^{T}[UwU]_{\omega}=T_{p}.$
Lemma 5.4.
The duality morphism
$\hat{\delta}^{L}_{N}:{\mathcal{M}}(N,\psi)_{!,M}\to{\mathop{\rm Hom}\nolimits}_{{\mathcal{O}}_{K}}({\mathcal{M}}_{!}(N,\psi)_{M},{\mathcal{M}}_{\psi}(1-k))$
has sign 1-1.
Proof.
Recall that for ${\mathcal{M}}(N,\psi)_{!,M}$ to be non-zero, we must have $\psi(-1_{\infty})=(-1)^{k-2}$. Fix such a $\psi$. By Lemma 5.2 and the fact that [,][\ ,\ ] has sign $(-1)^{k-1}$, we have, for $x,\ y\in{\mathcal{M}}(N,\psi)_{M,!,B}$,
$\langle x,y\rangle$ $=$ $[[UwU]_{\omega}x,y]$
$=$ $\psi(-1_{\infty})[N^{k-2}[Uw^{-1}U]_{\omega}x,y]$
$=$ $-[y,N^{k-2}[Uw^{-1}U]_{\omega}x]$
$=$ $-[[UwU]_{\omega}y,x]$
$=$ $-\langle y,x\rangle.\hskip 227.62204pt\qed$
5.3. The action of Hecke rings
Let $\tilde{{\mathbb{T}}}$ denote the polynomial algebra over ${\mathcal{O}}={\mathcal{O}}_{S}$ generated by the variables $t_{p}$, where $p$ runs over rational primes. Then $\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}M(\psi)_{N,!,{\mathop{\rm dR}\nolimits}}$ becomes a $\tilde{{\mathbb{T}}}-$module with $t_{p}$ acting by $U{\mbox{\scriptsize$\left(\begin{array}[]{cc}{{p}}&{{0}}\\ {{0}}&{{1}}\end{array}\right)$}}_{p}U]_{1}$. Since the modular forms with Fourier coefficients in $\mathcal{O}$} is stable under the action of the Hecke operators $T_{p}$, by Lemma 4.12 (c), $\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}{\mathcal{M}}(\psi)_{N,!,{\mathop{\rm dR}\nolimits}}$ is also a $\tilde{{\mathbb{T}}}-$module. Let $\mathfrak{a}$} denote the annihilator in $\tilde{{\mathbb{T}}}$ of $\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}{\mathcal{M}}(N,\psi)_{M,!,{\mathop{\rm dR}\nolimits}}$ and let ${\mathbb{T}}=\tilde{{\mathbb{T}}}/{\mathfrak{a}}$. Similarly let ${\mathfrak{a}}^{\prime}\subset{\mathfrak{a}}$ be the annihilator of $\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}M(N,\psi)_{M,{\mathop{\rm dR}\nolimits}}$ and let ${\mathbb{T}}^{\prime}=\tilde{{\mathbb{T}}}/{\mathfrak{a}}^{\prime}$. Then ${\mathbb{T}}\subset\operatorname{{\mathop{\rm End}\nolimits}}_{K}(\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}{\mathcal{M}}(N,\psi)_{M,!,{\mathop{\rm dR}\nolimits}})$ and ${\mathbb{T}}^{\prime}\subset\operatorname{{\mathop{\rm End}\nolimits}}_{K}(\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}M(N,\psi)_{M,{\mathop{\rm dR}\nolimits}})$ .
Lemma $5.$
  1. (a)
    The $\mathbb{T}$}-linear homomorphism
    $\varphi:\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}{\mathcal{M}}(N,\psi)_{M,!,{\mathop{\rm dR}\nolimits}}\to{\mathop{\rm Hom}\nolimits}({\mathbb{T}},{\mathcal{O}}),$
    $\varphi(f)(T)=a_{1}(T(f)),\ f\in\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}{\mathcal{M}}(N,\psi)_{M,!,{\mathop{\rm dR}\nolimits}},T\in{\mathbb{T}},$
    with $a_{1}(T(f))$ the first Fourier coefficient of $T(f)$, is an isomorphism.
  2. (b)
    For any $p\nmid N$, ${\mathbb{Q}}\otimes{\mathbb{T}}$ is generated by the images of $T_{\ell}$ for $\ell\neq p$.
  3. (c)
[…]
    $\varphi:\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}M(N,\psi)_{M,{\mathop{\rm dR}\nolimits}}\to{\mathop{\rm Hom}\nolimits}({\mathbb{Q}}\otimes{\mathbb{T}}^{\prime},K),$
    $\varphi(f)(T)=a_{1}(T(f)),\ f\in\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}M(N,\psi)_{M,{\mathop{\rm dR}\nolimits}},T\in{\mathbb{T}}^{\prime}$
    is an isomorphism.
  4. (d)
    For any $p\nmid N$, ${\mathbb{Q}}\otimes{\mathbb{T}}^{\prime}$ is generated by the images of $T_{\ell}$ for $\ell\neq p$.
Proof.
(a) The proof is the same as the proof of Theorem 2.2 in [Ri].
(b) Fix a $p\nmid N$. Let ${\mathbb{T}}_{1}$ be the subalgebra of ${\mathbb{Q}}\otimes{\mathbb{T}}$ generated by the images of $t_{\ell}$ for $\ell\neq p$. Consider the map
$\phi:\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}M(N,\psi)_{M,!,{\mathop{\rm dR}\nolimits}}\cong{\mathop{\rm Hom}\nolimits}({\mathbb{Q}}\otimes{\mathbb{T}},K)\to{\mathop{\rm Hom}\nolimits}({\mathbb{T}}_{1},K)$
where the second map is the restriction, and thus is surjective. We only need to prove that this map is injective.
[…]
The proof of (c) (resp. (d)) is the same as the proof of (a) (resp. (b)). ∎
Proposition 5.6.
  1. (a)
    $\mathbb{T}$} acts on ${\mathcal{M}}(N,\psi)_{M,!}$ as an object in ${\mathcal{P}M}^{S_{N}^{K}}_{K}$.
  2. (b)
    $\mathbb{T}$$^{\prime}$ acts on $M(N,\psi)_{M}$ as an object in $\operatorname{{\bf PM}}^{S_{N}^{K}}_{K}$.
  3. (c)
    $\mathfrak{a}$$^{\prime}$ is also the annihilator of ${\mathcal{M}}(N,\psi)_{M,c,{\mathop{\rm dR}\nolimits}}$ and $\mathbb{T}$$^{\prime}$ acts on ${\mathcal{M}}(N,\psi)_{M,c}$ as an object in ${\mathcal{P}M}^{S_{N}^{K}}_{K}$.
  4. (d)
[…]
Proof.
(a) We first prove that $\mathfrak{a}$} is annihilates each of the realizations of $M(N,\psi)_{M,!}$. Let $c:{\mathbb{C}}\to{\mathbb{C}},\ v\mapsto\bar{v}$ be the complex conjugation. By definition,
$\overline{\operatorname{{\mathop{\rm Fil}\nolimits}}}^{k-1}M(N,\psi)_{M,!,{\mathop{\rm dR}\nolimits}}=I^{\infty}(({\mathop{\rm id}\nolimits}\otimes c)((I^{\infty})^{-1}(\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}M(N,\psi)_{M,!,{\mathop{\rm dR}\nolimits}}))).$
Here
${\mathop{\rm id}\nolimits}\otimes c:M(N,\psi)_{M,!,B}\otimes{\mathbb{C}}\to M(N,\psi)_{M,!,B}\otimes{\mathbb{C}},\ x\otimes v\mapsto x\otimes\bar{v}$
is clearly $\tilde{{\mathbb{T}}}-$linear, as is $I^{\infty}$ by Lemma 4.2. Thus $\mathfrak{a}$} annihilates $\overline{\operatorname{{\mathop{\rm Fil}\nolimits}}}^{k-1}M_{!}(N,\psi)_{M,{\mathop{\rm dR}\nolimits}}$. So $\mathfrak{a}$} is the annihilator of
$M(N,\psi)_{M,!,{\mathop{\rm dR}\nolimits}}\otimes{\mathbb{C}}=(\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}M(N,\psi)_{M,!,{\mathop{\rm dR}\nolimits}}\otimes{\mathbb{C}})\oplus(\overline{\operatorname{{\mathop{\rm Fil}\nolimits}}}^{k-1}M(N,\psi)_{M,!,{\mathop{\rm dR}\nolimits}}\otimes{\mathbb{C}}).$
Since $I^{\infty}$ is $\tilde{{\mathbb{T}}}-$linear, $\mathfrak{a}$} is the annihilator of $M(N,\psi)_{M,!,B}\otimes{\mathbb{C}}$. Since the action of $\tilde{{\mathbb{T}}}$ on $\mathbb{C}$} is trivial, $\mathfrak{a}$} is the annihilator of $M(N,\psi)_{M,!,{\mathop{\rm dR}\nolimits}}$ and $M(N,\psi)_{M,!,B}$. Then it is also the annihilator of $M(N,\psi)_{M,!,\lambda}=M(N,\psi)_{M,!,B}\otimes_{K}K_{\lambda}$ and $M(N,\psi)_{M,!,{\mathop{\rm dR}\nolimits}}\otimes{\mathbb{Q}}_{p}$.
We next show that $\mathbb{T}$} acts on ${\mathcal{M}}(N,\psi)_{M,!}$ as an object in ${\mathcal{P}M}^{S_{N}^{K}}_{K}$.
Fix a finite prime pp. Recall that ${\mathcal{M}}(N,\psi)_{M,!}$ is in ${\mathcal{P}M}^{S_{N}^{K}}_{K}$ and, by Lemma 4.2, $T_{p}:{\mathcal{M}}(N,\psi)_{M,!}\to{\mathcal{M}}(N,\psi)_{pM,!}$ is defined in ${\mathcal{P}M}^{S_{pM}^{K}}_{K}={\mathcal{P}M}^{S_{pN}^{K}}_{K}$. If $p|Nk!$, then $S_{pN}=S_{N}$. So $T_{p}$ is a morphism in ${\mathcal{P}M}^{S_{N}^{K}}_{K}$.
If $p\nmid Nk!$, then for each realization $\mathcal{M}$$_{?}$ of ${\mathcal{M}}(N,\psi)_{M,!}\in{\mathcal{P}M}^{S_{N}}$, the corresponding realization of ${\mathcal{M}}(N,\psi)_{pM,!}\in{\mathcal{P}M}^{S_{pN}}$ is given by
$\left\{\begin{array}[]{ll}{\mathcal{M}}_{?},&?=B,\lambda{\rm\ or\ }\lambda{\rm-crys}{\rm\ with\ }\lambda\nmid pNk!,\\ {\mathcal{M}}_{?}[1/p],&?={\mathop{\rm dR}}\nolimits\end{array}\right$.
Thus by Lemma 4.2, $T_{p}$ acts on these realizations and observes comparison isomorphisms among them. So to prove the proposition, we only need to give actions of $T_{p}$ on ${\mathcal{M}}_{\lambda\mbox{\scriptsize-crys}}$ with $\lambda\nmid p$ and ${\mathcal{M}}_{\mathop{\rm dR}\nolimits}$, and show that these actions observe the comparison isomorphisms $I_{\mathop{\rm dR}\nolimits}^{\lambda}$ and $I^{\lambda\mbox{\scriptsize-crys}}$. Fix a $\lambda\mid p$. By Lemma 5.5, $T_{p}$ is a polynomial of $T_{q},\ q\neq p,$ with coefficients in $K$. Thus there is $r\in{\mathbb{Z}}_{>0}$ such that $rT_{p}$ is a polynomial of $T_{q},\ q\neq p,$ with coefficients in ${\mathcal{O}}_{K}$. Since each $T_{q},\ q\neq p$, acts on ${\mathcal{M}}_{\lambda\mbox{\scriptsize-crys}}$ by Lemma 4.2, we have an action of $rT_{p}$ on ${\mathcal{M}}_{\lambda\mbox{\scriptsize-crys}}$ which we provisionally denote by $F$. Further, $T_{p}:{\mathcal{M}}_{\lambda}\to{\mathcal{M}}_{\lambda}$ gives $rT_{p}{\mathcal{M}}_{\lambda}\subseteq r{\mathcal{M}}_{\lambda}$. Since each $T_{q},\ q\neq p,$ observes $I^{\lambda\mbox{\scriptsize-crys}}$, we have
${\mathbb{V}}(F({\mathcal{M}}_{\lambda\mbox{\scriptsize-crys}}))=F({\mathbb{V}}({\mathcal{M}}_{\lambda\mbox{\scriptsize-crys}}))\cong F({\mathcal{M}}_{p})\subseteq r{\mathcal{M}}_{p}\cong r{\mathbb{V}}({\mathcal{M}}_{\lambda\mbox{\scriptsize-crys}})={\mathbb{V}}(r{\mathcal{M}}_{\lambda\mbox{\scriptsize-crys}}).$
Thus $F({\mathcal{M}}_{\lambda\mbox{\scriptsize-crys}})\subseteq r{\mathcal{M}}_{\lambda\mbox{\scriptsize-crys}}.$ Since ${\mathcal{M}}_{\lambda\mbox{\scriptsize-crys}}$ for $\lambda\nmid Nk!$ is torsion free, we obtain a map $r^{-1}F:{\mathcal{M}}_{\lambda\mbox{\scriptsize-crys}}\to{\mathcal{M}}_{\lambda\mbox{\scriptsize-crys}}$ that observes $I^{\lambda\mbox{\scriptsize-crys}}:{\mathbb{V}}({\mathcal{M}}_{\lambda\mbox{\scriptsize-crys}})\to{\mathcal{M}}_{\lambda}.$ Then via $I_{\mathop{\rm dR}\nolimits}^{\lambda}:{\mathcal{M}}_{\mathop{\rm dR}\nolimits}\otimes{\mathcal{O}}_{\lambda}\to{\mathcal{M}}_{\lambda\mbox{\scriptsize-crys}}$, we obtain a map $r^{-1}F:{\mathcal{M}}_{\mathop{\rm dR}\nolimits}\otimes{\mathcal{O}}_{\lambda}\to{\mathcal{M}}_{\mathop{\rm dR}\nolimits}\otimes{\mathcal{O}}_{\lambda}$. Note that $T_{\ell}$ acts on $M_{\mathop{\rm dR}\nolimits}$ for all $\ell$ and, by the first part of the proof, the action of $T_{p}$ on $M_{\mathop{\rm dR}\nolimits}$ agrees with $r^{-1}F$. Since ${\mathcal{M}}_{\mathop{\rm dR}\nolimits}\otimes{\mathbb{Z}}_{p}=\bigoplus_{\lambda|p}{\mathcal{M}}_{\mathop{\rm dR}\nolimits}\otimes{\mathcal{O}}_{\lambda}$, we have the following commutative diagram
[…]
This shows that the image of ${\mathcal{M}}_{\mathop{\rm dR}\nolimits}$ under $T_{p}$ is contained in ${\mathcal{M}}_{\mathop{\rm dR}\nolimits}[1/p]\cap({\mathcal{M}}_{\mathop{\rm dR}\nolimits}\otimes{\mathbb{Z}}_{p})={\mathcal{M}}_{\mathop{\rm dR}\nolimits}$. Thus if we define $T_{p}:{\mathcal{M}}_{{\lambda\mbox{\scriptsize-crys}}}\to{\mathcal{M}}_{{\lambda\mbox{\scriptsize-crys}}}$ by $r^{-1}F$, then it is compatible with $I^{\lambda\mbox{\scriptsize-crys}}$ and with $I_{\mathop{\rm dR}\nolimits}^{\lambda\mbox{\scriptsize-crys}}$. So we have seen that $T_{p}:{\mathcal{M}}_{!}(N,\psi)_{M}\to{\mathcal{M}}_{!}(N,\psi)_{pM}$ in ${\mathcal{P}M}^{S_{pN}}$ extends to a morphism $T_{p}:{\mathcal{M}}_{!}(N,\psi)_{M}\to{\mathcal{M}}_{!}(N,\psi)_{M}$ in ${\mathcal{P}M}^{S_{N}^{K}}_{K}$.
Thus we have shown that $T_{p}$ is a morphism in ${\mathcal{P}M}^{S_{N}^{K}}_{K}$ for every $p$ and proven (a).
The argument for (b) is the same as for (a).
To prove (c), note that by (b) and the fact that $T_{p}$ is self-adjoint under the pairing between $M(N,\psi)_{M}$ and $M(N,\psi)_{M,c}$ defined in Lemma 5.3, $\mathfrak{a}$$^{\prime}$ is also the annihilator of $M(N,\psi)_{M,c}$. The rest of the proof is the same as for part (a).
For (d), we only need to prove that each $T_{p}$ commutes with these maps. This follows from Lemma 4.2. ∎
5.4. Premotivic structure for a newform
Now suppose that $g$ is a (normalized) eigenform in $\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}{\mathcal{M}}(N,\psi)_{N,!,{\mathop{\rm dR}\nolimits}}$ for the action of $\mathbb{T}$}. So for $T\in{\mathbb{T}}$ we have $T(g)=a_{1}(T(g))g$. Let $K=K_{g}$ denote the field generated by $a_{1}(T(g)),\ T\in{\mathbb{T}}$ and let $I_{g}$ denote the kernel of the map ${\mathbb{T}}\to K,\ T\mapsto a_{1}(T(g))$. Let ${\mathcal{M}}_{g}\subset{\mathcal{M}}(N,\psi)_{M,!}$ denote the intersection of the kernels of elements of $I_{g}$. By Proposition 5.6, ${\mathcal{M}}_{g}$ is in ${\mathcal{P}}{\mathcal{M}}^{S}_{K}$ for any $S\supset S_{N}^{K}$. Let $M_{g}={\mathcal{M}}_{g}\otimes K\in\operatorname{{\bf PM}}^{S}_{K}$.
Lemma 5.7.
$M_{g}$ is a premotivic structure of rank 2 over $K$ and $\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}M_{g,{\mathop{\rm dR}\nolimits}}=Kg$. If $g$ is a new form, then the pairing on $M(N,\psi)_{M,!}$ restricts to a perfect alternating pairing on $M_{g}$, i.e., an isomorphism
$\wedge^{2}_{K}M_{g}\to M_{\psi}(1-k).$
Proof.
[…]
The pairing on $M_{g}$ is restricted from an alternating pairing. So it is still alternating. To show that the pairing on $M_{g}$ is perfect, we only need to show that the pairing on $M_{g,{\mathop{\rm dR}\nolimits}}\otimes{\mathbb{C}}$ is perfect.
By (34), $\langle M(N,\psi)_{M,!,{\mathop{\rm dR}\nolimits}},M(N,\psi)_{M,!,{\mathop{\rm dR}\nolimits}}\rangle\neq 0$. For $\alpha\in I_{g}$ and $x\in M(N,\psi)_{M,!,{\mathop{\rm dR}\nolimits}}$, we have $\langle M_{g,{\mathop{\rm dR}\nolimits}},\alpha x\rangle=\langle\alpha M_{g,{\mathop{\rm dR}\nolimits}},x\rangle=0$. Thus $\langle M_{g,{\mathop{\rm dR}\nolimits}},I_{g}M(N,\psi)_{M,!,{\mathop{\rm dR}\nolimits}}\rangle=0$. By strong multiplicity one theorem,
$\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}M(N,\psi)_{M,!,{\mathop{\rm dR}\nolimits}}=Kg+I_{g}\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}M(N,\psi)_{M,!,{\mathop{\rm dR}\nolimits}}.$
Then
${\mathbb{C}}\otimes M(N,\psi)_{M,!,{\mathop{\rm dR}\nolimits}}$
$\cong$ $({\mathbb{C}}\otimes\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}M_{!}(N,\psi)_{M,{\mathop{\rm dR}\nolimits}})+(I^{\infty}({\mathop{\rm id}\nolimits}\otimes c)(I^{\infty})^{-1})({\mathbb{C}}\otimes\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}M_{!}(N,\psi)_{M,{\mathop{\rm dR}\nolimits}})$
$\cong$ ${\mathbb{C}}g+{\mathbb{C}}\otimes I_{g}\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}M_{!}(N,\psi)_{M,{\mathop{\rm dR}\nolimits}}+(I^{\infty}({\mathop{\rm id}\nolimits}\otimes c)(I^{\infty})^{-1})(g\otimes{\mathbb{C}})$
$+(I^{\infty}({\mathop{\rm id}\nolimits}\otimes c)(I^{\infty})^{-1})({\mathbb{C}}\otimes I_{g}\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}M_{!}(N,\psi)_{M,{\mathop{\rm dR}\nolimits}}).$
Thus
$0\neq$ $\langle{\mathbb{C}}g,{\mathbb{C}}\otimes M_{!}(N,\psi)_{M,{\mathop{\rm dR}\nolimits}}\rangle=\langle{\mathbb{C}}g,{\mathbb{C}}g+(I^{\infty}({\mathop{\rm id}\nolimits}\otimes c)(I^{\infty})^{-1})(g\otimes{\mathbb{C}})\rangle$
$=$ $\langle{\mathbb{C}}g,{\mathbb{C}}\otimes M_{g,{\mathop{\rm dR}\nolimits}}\rangle.$
Since $M_{g,{\mathop{\rm dR}\nolimits}}$ is two dimensional and the pairing in alternating, this implies that the pairing on ${\mathbb{C}}\otimes M_{g,{\mathop{\rm dR}\nolimits}}$ is nondegenerate, as is needed. ∎
$5.$ The $L$-function
Suppose that $f(\tau)=\sum a_{n}e^{2\pi in\tau}$ is a newform of weight $k$, conductor $N_{f}$ and character $\psi_{f}$. Associated to $f$ is the $L$-function with Euler product factorization:
$L(f,s)=\sum_{n\geq 1}a_{n}n^{-s}=\prod_{p\nmid N_{f}}(1-a_{p}p^{-s}+\psi_{f}(p)p^{k-1-2s})^{-1}\prod_{p|N_{f}}(1-a_{p}p^{-s})^{-1}.$
Recall that ${\mathcal{A}}_{k}^{0}$ denote the representation of $\operatorname{{\mathop{\rm GL}\nolimits}}_{2}({\mathbb{A}}_{\bf f})$ defined by
$\lim_{\stackrel{{\scriptstyle\rightarrow}}{{N}}}\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}M_{N,!,{\mathop{\rm dR}\nolimits}}\otimes{\mathbb{C}}.$
Then there is an irreducible subrepresentation $\pi(f)$ of ${\mathcal{A}}_{k}^{0}$ of central character $\psi_{f}||\ ||^{2-k}$ such that $f$ spans the image of $\pi(f)^{U_{1}(N_{f})}$ under the isomorphism
$({\mathcal{A}}_{k}^{0})^{U_{1}(N_{f})}\cong S_{k}(\Gamma_{1}(N_{f}))$
(Here we view $\psi_{f}$ as a character on ${\mathbb{A}}_{f}^{\times}\subset{\mathbb{A}}^{\times}$). Moreover, we have the decomposition
${\mathcal{A}}_{k}^{0}=\bigoplus_{f}\pi(f)$
where $f$ runs over newforms of weight $k$ of any conductor and character. For each $f$ we have a factorization
$\pi(f)\cong\otimes_{p}\pi_{p}(f)$
where $\pi_{p}(f)$ is an irreducible admissible representation of $\operatorname{{\mathop{\rm GL}\nolimits}}_{2}({\mathbb{Q}}_{p})$. For each $p$, we let $c_{p}=v_{p}(N_{f})$, so $p^{c_{p}}$ is the conductor of $\pi_{p}(f)$. We let
$V_{p}=\left\{\left.{\mbox{\scriptsize$\left(\begin{array}[]{cc}{{a}}&{{b}}\\ {{c}}&{{d}}\end{array}\right)$}}\in\operatorname{{\mathop{\rm GL}\nolimits}}_{2}({\mathbb{Z}}_{p})\right|a-1\equiv c\equiv 0\bmod p^{c_{p}}\right\}.$
Then $\pi_{p}(f)^{V_{p}}$ is one-dimensional and we consider the double coset operators on it defined by
$t_{p}=V_{p}{\mbox{\scriptsize$\left(\begin{array}[]{cc}{{p}}&{{0}}\\ {{0}}&{{1}}\end{array}\right)$}}V_{p},\quad t_{p}^{\prime}=V_{p}{\mbox{\scriptsize$\left(\begin{array}[]{cc}{{1}}&{{0}}\\ {{0}}&{{p}}\end{array}\right)$}}V_{p},\quad s_{p}=V_{p}{\mbox{\scriptsize$\left(\begin{array}[]{cc}{{p}}&{{0}}\\ {{0}}&{{p}}\end{array}\right)$}}V_{p}.$
Then we have $t_{p}=\psi_{f}(p_{p})a_{p}$ and $s_{p}=\psi_{f}(p_{p})p^{k-2}$ on $\pi_{p}(f)^{V_{p}}$. We recall the following general facts about irreducible admissible representations of $\operatorname{{\mathop{\rm GL}\nolimits}}_{2}({\mathbb{Q}}_{p})$:
Lemma 5.8.
Suppose that $c_{p}>0$ and $\pi_{p}(f)$ has minimal conductor among its twists.
  1. (a)
    $\pi_{p}(f)$ is special if and only if $t_{p}t_{p}^{\prime}=s_{p}$;
  2. (b)
    $\pi_{p}(f)$ is principal series if and only if $t_{p}t_{p}^{\prime}=ps_{p}$;
  3. (c)
    $\pi_{p}(f)$ is supercuspidal if and only if $t_{p}=0$.
Suppose now that $g$ is a newform of weight $k$, level $N_{g}$ and character $\psi_{g}$, with coefficients in a number field $K$. For each embedding $\tau:K\to{\mathbb{C}}$, we have a newform $\tau(g)$ as above, and so a representation $\pi_{\tau(g)}$ of $\operatorname{{\mathop{\rm GL}\nolimits}}_{2}({\mathbb{A}}_{\bf f})$. The $L$-functions attached to $M_{g}$ are related to those attached to the $\tau(g)$ by the formula
$L(M_{g}\otimes M_{\psi_{g}^{-1}},\tau,s)=L(\tau(g),s).$
In fact, a stronger statement is true: For any primes $p$ of $\mathbb{Q}$} and $\lambda$ of $K$, the representation $D_{{\mathop{\rm pst}\nolimits}}(M_{\lambda}|G_{p})^{{\mathop{\rm ss}\nolimits}}$ of $^{\prime}W_{p}$ is $K$-rational and corresponds via local Langlands to $\pi_{p}(\tau(g))$ (where we extend scalars to $\mathbb{C}$} via $\tau$ and normalize the local Langlands correspondence as in [Ca0]). This is due to Eichler, Shimura and Igusa for $\lambda$ not dividing $pN_{g}$, by Langlands, Deligne and Carayol [Ca0] for $\lambda\nmid p$, and by Scholl [Scho2] and Saito [Sai] in general. In particular $M_{g}$ is $L$-admissible everywhere; moreover $M_{g}=M^{S_{N}}$ for an object $M$ of $\operatorname{{\bf PM}}_{K}$ which is $L$-admissible everywhere. Using the compatibility with the local Langlands correspondence, it also follows from the functional equations for the $L(\tau(g),s)$ that Conjecture 1.2 holds for $M_{g}$.
Finally we recall some basic properties of the local Langlands correspondence. In the following, we identify characters of ${\mathbb{Q}}_{p}^{\times}$ with those of $^{\prime}W_{p}$ via class field theory.
Lemma 5.9.
Suppose that $\pi$ is an irreducible admissible representation of $\operatorname{{\mathop{\rm GL}\nolimits}}_{2}({\mathbb{Q}}_{p})$ and $\rho:{}^{\prime}W_{p}\to\operatorname{\operatorname{{\mathop{\rm Aut}\nolimits}}}_{\mathbb{C}}V$ corresponds to ${\mathop{\rm det}\nolimits}||^{1/2}\pi$ via local Langlands. Then the following hold:
  1. (a)
    The conductors of $\pi$ and $\rho$ coincide.
  2. (b)
    The central character of $\pi$ is the determinant of $\rho$.
  3. (c)
    For any character $\theta$ of ${\mathbb{Q}}_{p}^{\times}$, ${\mathop{\rm det}\nolimits}||^{1/2}(\theta\circ{\mathop{\rm det}\nolimits})\pi$ corresponds to $\theta\rho$.
  4. (d)
    The trace of $\rho(\operatorname{{\mathop{\rm Frob}\nolimits}}_{p})$ on $(V/NV)^{I_{p}}$ is the eigenvalue of $t_{p}$ on $\pi^{V_{p}}$.
  5. (e)
    $\pi$ is supercuspidal if and only if $\rho$ is irreducible.
These properties already determine the correspondence explicitly if $\pi$ is principal series or special. In particular, if $\pi=I(\mu_{1},\mu_{2})$, then $\rho|_{W_{p}}$ is equivalent to $\mu_{1}\oplus\mu_{2}$ and $N\neq 0$ if and only if $\pi$ is special. We shall also need the explicit description of the correspondence due to Gérardin [] in certain cases where $\pi$ is supercuspidal. For this we refer to sections 3.2 and 4.2 of [C-D-T].
6. The adjoint premotivic structure
6.1. Realizations of the adjoint premotivic structure
$A_{f}={\mathop{\rm ad}\nolimits}^{0}M_{f}$ is defined to be the kernel of the trace map
${\mathop{\rm Hom}\nolimits}_{K}(M_{f},M_{f})\to K.$
It is a premotivic structure in $\operatorname{{\bf PM}}^{S}_{K}$ for $S\supseteq S_{N}$.
For $B,{\mathop{\rm dR}\nolimits}$ or $\lambda$, $A_{f,?}$ has an integral structure given by
${\mathcal{A}}_{f,?}=\{a\in\operatorname{{\mathop{\rm End}\nolimits}}({\mathcal{M}}_{f,?})|\operatorname{{\mathop{\rm tr}\nolimits}}(a)=0\}.$
The extra structures on the realizations of ${\mathcal{A}}_{f}$ are obtained by restrictions from those of $\operatorname{{\mathop{\rm End}\nolimits}}({\mathcal{M}}_{f})$. For example the filtration on ${\mathcal{A}}_{f,{\mathop{\rm dR}\nolimits}}$ is given by
$\operatorname{{\mathop{\rm Fil}\nolimits}}^{n}{\mathcal{A}}_{f,{\mathop{\rm dR}\nolimits}}=\{a\in{\mathcal{A}}_{\mathop{\rm dR}\nolimits}\subseteq\operatorname{{\mathop{\rm End}\nolimits}}({\mathcal{M}}_{f,{\mathop{\rm dR}\nolimits}})|a(\operatorname{{\mathop{\rm Fil}\nolimits}}^{i}{\mathcal{M}}_{f,{\mathop{\rm dR}\nolimits}})\subseteq\operatorname{{\mathop{\rm Fil}\nolimits}}^{n+i}{\mathcal{M}}_{f,{\mathop{\rm dR}\nolimits}},\forall j\}$
$=$ $\hskip-8.5359pt\left\{\hskip-5.69046pt\begin{array}[]{ll}{\mathcal{A}}_{f,{\mathop{\rm dR}\nolimits}},&n\leq-(k-1),\\ \{a\in{\mathcal{A}}_{f,{\mathop{\rm dR}\nolimits}}|a(\operatorname{{\mathop{\rm Fil}\nolimits}}^{0}{\mathcal{M}}_{f,{\mathop{\rm dR}\nolimits}})\subseteq\operatorname{{\mathop{\rm Fil}\nolimits}}^{0}{\mathcal{M}}_{f,{\mathop{\rm dR}\nolimits}}\},&-(k-1)<n\leq 0,\\ \{a\in{\mathcal{A}}_{f,{\mathop{\rm dR}\nolimits}}|a({\mathcal{M}}_{f,{\mathop{\rm dR}\nolimits}})\subseteq\operatorname{{\mathop{\rm Fil}\nolimits}}^{0}{\mathcal{M}}_{f,{\mathop{\rm dR}\nolimits}},\ a(\operatorname{{\mathop{\rm Fil}\nolimits}}^{0}{\mathcal{M}}_{f,{\mathop{\rm dR}\nolimits}})=0\},\hskip-5.69046pt&\ 0<n\leq k-1,\\ 0,&n>k-1.\end{array}\right$.
The filtration on $\operatorname{{\mathop{\rm End}\nolimits}}({\mathcal{M}}_{f,{\lambda\mbox{\scriptsize-crys}}})$ is the same. Since the the weight is from 1k1-k to k1k-1, we need p1>2(k1)p-1>2(k-1) instead of $p>k$ (plus a twist of the crystalline realization) in order to get an object in ${{\mathcal{M}\mathcal{F}}}^{0}$ and to make ${\mathcal{A}}_{f,{\lambda\mbox{\scriptsize-crys}}}$ well-defined.
Lemma 6.1.
There is an isomorphism ${\mathop{\rm det}\nolimits}_{K}A_{f}\cong K$ in $\operatorname{{\bf PM}}^{S}_{K}$ which restricts to isomorphisms ${\mathop{\rm det}\nolimits}_{{\mathcal{O}}_{K}}{\mathcal{A}}_{f,?}[1/Nk!]\cong{\mathcal{O}}_{K}[1/Nk!]$ for $\in\{{\mathop{\rm B}\nolimits},{\mathop{\rm dR}\nolimits},\lambda\}$.
Proof.
Setting $M_{f}^{*}={\mathop{\rm Hom}\nolimits}_{K}(M_{f},K)$ and noting that $M_{f}$ has $K$-rank 2 we find
${\mathop{\rm det}\nolimits}_{K}A_{f}\otimes_{K}K\cong$ ${\mathop{\rm det}\nolimits}_{K}(A_{f}\oplus K)\cong{\mathop{\rm det}\nolimits}_{K}{\mathop{\rm Hom}\nolimits}_{K}(M_{f},M_{f})\cong{\mathop{\rm det}\nolimits}_{K}(M_{f}^{*}\otimes_{K}M_{f})$
$\cong$ ${\mathop{\rm det}\nolimits}_{K}(M_{f}^{*})^{\otimes 2}\otimes_{K}{\mathop{\rm det}\nolimits}_{K}(M_{f})^{\otimes 2}$
$\cong$ $\left({\mathop{\rm det}\nolimits}_{K}(M_{f})^{*}\otimes_{K}{\mathop{\rm det}\nolimits}_{K}(M_{f})\right)^{\otimes 2}$
$\cong$ $K^{\otimes 2}\cong K.$
The integral isomorphisms are proved in the same way. ∎
We note that $A_{f}$ and ${\mathop{\rm Hom}\nolimits}_{K}(A_{f},K)$ are isomorphic. There is in fact a canonical isomorphism defined by the pairing
(36) $\alpha\otimes\beta\mapsto\operatorname{{\mathop{\rm tr}\nolimits}}(\alpha\circ\beta)$
on each realization of $A_{f}$.
Note that if we replace $K$ by $K^{\prime}\supset K$ and $S$ by a subset $S^{\prime}$ of the primes over those in $S$, then $A_{f}$ is replaced by $(A_{f}\otimes_{K}K^{\prime})^{S^{\prime}}$. If $\psi^{\prime}$ is a character ${\mathbb{A}}^{\times}\to K^{\times}$ of conductor $D^{\prime}$ and $f\otimes\psi^{\prime}$ denotes the eigenform (of weight $k$, conductor dividing $N_{f}D^{\prime}{}^{2}$ and character $\psi(\psi^{\prime})^{2}$) associated to the eigenform $\sum_{(n,D^{\prime})=1}\psi^{\prime}(n_{n})a_{n}e^{2\pi inz}$. We will relate ${\mathcal{M}}_{f}$ to the premotivic structure associated to $f\otimes\psi^{\prime}$.
Lemma 6.2.
Let $S\supseteq S_{f}$ be such that $D^{\prime}$ is invertible in ${\mathcal{O}}_{S}$. We have ${\mathcal{M}}_{f\otimes\psi^{\prime}}\cong{\mathcal{M}}_{f}\otimes_{{\mathcal{O}}}{\mathcal{M}}_{\psi^{\prime}}$ in ${\mathcal{P}M}^{S}_{K}$. We also have $A_{f\otimes\psi^{\prime}}\cong A_{f}$ in $\operatorname{{\bf PM}}^{S}_{K}$.
It is sometimes convenient to assume that $f$ has minimal conductor among its twists.
Proof.
[…]
With this setup, and $U=U_{0}(N),U^{\prime}=U_{0}(N^{\prime}),U^{\prime\prime}=U_{0}(N^{\prime\prime})$, we define
$\theta:{\mathcal{M}}(N,\psi)_{D^{\prime}M}\otimes{\mathcal{M}}_{\psi^{\prime}}\to{\mathcal{M}}(N^{\prime\prime},\psi\psi^{\prime}{}^{2})_{(D^{\prime})^{3}M}$
to be the composite from the left column of the following diagram.
[…]
We will show that it restricts to an isomorphism ${\mathcal{M}}_{f}\otimes{\mathcal{M}}_{\psi^{\prime}}\to{\mathcal{M}}_{f\otimes\psi^{\prime}}$ stated in the lemma.
We first show that $\theta$ is well-defined. In the top half of the diagram, under the isomorphism
${\mathcal{M}}_{D^{\prime}M}\otimes{\mathcal{V}}_{\psi^{-1}}\otimes{\mathcal{M}}_{F}\otimes{\mathcal{O}}_{\psi^{\prime}{}^{-1}}\cong{\mathcal{M}}_{D^{\prime}M}\otimes{\mathcal{V}}_{\psi^{-1}}\otimes{\mathcal{H}}^{0}(X_{D^{\prime}M})\otimes{\mathcal{O}}_{\psi^{\prime}}\to{\mathcal{M}}_{D^{\prime}M}\otimes{\mathcal{V}}_{\psi^{-1}}\otimes{\mathcal{O}}_{\psi^{\prime}}$
where the first map is from the proof of Lemma 4.6 and the second map is from cup product, we have
${\mathcal{M}}(N,\psi)_{D^{\prime}M}\otimes{\mathcal{M}}_{\psi^{\prime}{}^{-1}}$ $\cong({\mathcal{M}}_{D^{\prime}M}\otimes{\mathcal{V}}_{\psi^{-1}})^{U}\otimes({\mathcal{H}}^{0}(X_{D^{\prime}M})\otimes{\mathcal{O}}_{\psi^{\prime}})^{U^{\prime}}$
$\subseteq({\mathcal{M}}_{D^{\prime}M}\otimes{\mathcal{V}}_{\psi^{-1}}\otimes{\mathcal{H}}^{0}(X_{D^{\prime}M})\otimes{\mathcal{O}}_{\psi^{\prime}})^{U^{\prime}}$
$\to({\mathcal{M}}_{D^{\prime}M}\otimes{\mathcal{V}}_{\psi^{-1}}\otimes{\mathcal{O}}_{\psi^{\prime}})^{U^{\prime}}.$
Here the first map is by Lemma 4.6, the inclusion map is because $U^{\prime}\subseteq U$ and the third map is because the cup product is $U^{\prime}$-equivariant.
[…]
$\left(\begin{array}[]{cc}{{a+D^{\prime}N^{\prime}c}}&{{b+D^{\prime}{}^{-1}(d-a)-D^{\prime}{}^{-2}N^{\prime}c}}\\ {{N^{\prime}c}}&{{-D^{\prime}{}^{-1}N^{\prime}c+d}}\end{array}\right)$
for some ${\mbox{\scriptsize$\left(\begin{array}[]{cc}{{a}}&{{b}}\\ {{N^{\prime}c}}&{{d}}\end{array}\right)$}}\in U^{\prime}$. Since $u$ is in $U^{\prime\prime}$, we have $N^{\prime\prime}|N^{\prime}c$, so $\frac{N^{\prime\prime}}{N^{\prime}}|c$. Then $\frac{N}{{\rm gcd}(D^{\prime},N)}\frac{N^{\prime\prime}}{N^{\prime}}|\frac{N}{{\rm gcd}(D^{\prime},N)}c$. Since $\frac{N}{{\rm gcd}(D^{\prime},N)}\frac{N^{\prime\prime}}{N^{\prime}}=\frac{N^{\prime\prime}}{D^{\prime}}$, we have $D^{\prime}|\frac{N}{{\rm gcd}(D^{\prime},N)}c$. From this and the $(1,2)$-entry of $u$ we get $d-a\equiv 0\mod D^{\prime}$. Thus
$\tau(\sigma^{\prime}(g^{\prime}{}^{-1}ug^{\prime})v)$ $=\tau(\psi^{-1}(a_{N})\psi^{\prime}{}^{-1}\circ{\mathop{\rm det}\nolimits}(g^{\prime}{}^{-1}ug^{\prime})v)$
$=\tau(\psi^{-1}(a_{N})\psi^{\prime}(ad)^{-1}v)$
$=\tau(\psi^{-1}(a_{N})\psi^{\prime}(a)^{-2}v)$
$=\psi^{-1}(a_{N})\psi^{\prime}(a)^{-2}v$
$=\sigma^{\prime\prime}(\tau(v)).$
This shows that
$U^{\prime\prime}{\mbox{\scriptsize$\left(\begin{array}[]{cc}{{1}}&{{1/D^{\prime}}}\\ {{0}}&{{1}}\end{array}\right)$}}U^{\prime}]_{1}:({\mathcal{M}}_{D^{\prime}M}\otimes{\mathcal{V}}_{\psi^{-1}}\otimes{\mathcal{O}}_{\psi^{\prime}})^{U^{\prime}}\to{\mathcal{M}}(N,\psi^{-1}\psi^{\prime}{}^{-2})_{(D^{\prime})^{3}M}$
is well-defined. We also note that, letting $\Lambda\subset\hat{{\mathbb{Z}}}^{\times}$ be a complete set of coset representatives of $({\mathbb{Z}}/D^{\prime}{\mathbb{Z}})^{\times},$ then $g_{d}={\mbox{\scriptsize$\left(\begin{array}[]{cc}{{1}}&{{0}}\\ {{0}}&{{d}}\end{array}\right)$}},d\in\Lambda$ gives a coset decomposition $)U^{\prime\prime}=\coprod_{d}g_{d}(U^{\prime\prime}\cap(g^{\prime}U^{\prime}g^{\prime}{}^{-1}$)).
Thus $\theta$ is well-defined and is the restriction of the composite of the maps in the bottom row in the above diagram. To prove the first statement of the lemma, we only need to show
(37) $\theta_{\mathop{\rm dR}\nolimits}(f\otimes b_{\mathop{\rm dR}\nolimits})=D^{\prime}\psi^{\prime}(-1)f_{\psi^{\prime}}.$
Tracing through the maps in the above diagram, we see that
$f\otimes b_{\mathop{\rm dR}\nolimits}\in\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}{\mathcal{M}}(N,\psi)_{D^{\prime}M,{\mathop{\rm dR}\nolimits}}\otimes{\mathcal{M}}_{\psi^{\prime},{\mathop{\rm dR}\nolimits}}$
is sent to
$\sum_{a\mod D^{\prime}}\sum_{d\mod D^{\prime}}[g_{d}g]\zeta_{D^{\prime}}^{-a}f\otimes g_{d}~(1\otimes\psi^{\prime}(a))\in{\mathcal{M}}_{D^{\prime}{}^{3}M}\otimes{\mathcal{V}}_{\psi^{-1}\psi^{\prime}{}^{-2}}.$
Under the isomorphism
${\mathbb{C}}\otimes\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}{\mathcal{M}}_{D^{\prime}M,{\mathop{\rm dR}\nolimits}}\cong\oplus_{t\in({\mathbb{Z}}/D^{\prime}M{\mathbb{Z}})^{\times}}M(\Gamma(D^{\prime}M)),$
[…]
$\gamma=\gamma_{d,t^{\prime}}=\left({\mbox{\scriptsize$\left(\begin{array}[]{cc}{{1}}&{{0}}\\ {{0}}&{{t^{\prime}{}^{-1}}}\end{array}\right)$}}{\mbox{\scriptsize$\left(\begin{array}[]{cc}{{1}}&{{D^{\prime}{}^{-1}}}\\ {{0}}&{{d}}\end{array}\right)$}}{\mbox{\scriptsize$\left(\begin{array}[]{cc}{{1}}&{{0}}\\ {{0}}&{{d^{-1}t^{\prime}}}\end{array}\right)$}}\right)^{-1}={\mbox{\scriptsize$\left(\begin{array}[]{cc}{{1}}&{{-d^{-1}t^{\prime}D^{\prime}{}^{-1}}}\\ {{0}}&{{1}}\end{array}\right)$}}.$
Then $g_{d}g]\otimes g_{d}$ sends $(\zeta_{D^{\prime}}^{-t^{-1}a}f_{t})_{t}\otimes(1\otimes\psi^{\prime}(a))$ to
$\left(\zeta_{D^{\prime}}^{-d(t^{\prime})^{-1}a}f_{d^{-1}t^{\prime}}|_{{\mbox{\scriptsize$\left(\begin{array}[]{cc}{{1}}&{{-d^{-1}t^{\prime}D^{\prime}{}^{-1}}}\\ {{0}}&{{1}}\end{array}\right)$}}}\right)_{t^{\prime}}\otimes\psi^{\prime}(a)$
[…]
$\sum_{a\mod D^{\prime}}\zeta_{D^{\prime}}^{-a}\sum_{d\mod D^{\prime}}[g_{d}g]f\otimes g_{d}\tau(1\otimes\psi^{\prime}(a))$
in the above diagram is
$\sum_{a\mod D^{\prime}}\sum_{d\mod D^{\prime}}\zeta_{D^{\prime}}^{-d^{-1}t^{\prime}a}f_{d^{-1}t^{\prime}}|_{{\mbox{\scriptsize$\left(\begin{array}[]{cc}{{1}}&{{-d^{-1}t^{\prime}D^{\prime}{}^{-1}}}\\ {{0}}&{{1}}\end{array}\right)$}}}\otimes\psi^{\prime}(a).$
Since $f$ is from $\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}{\mathcal{M}}(N,\psi)_{D^{\prime}M,{\mathop{\rm dR}\nolimits}}$, by Lemma 4.12 and its proof, the components of $(f_{t})_{t}$ are the same. Then we have
$\sum_{a\mod D^{\prime}}$ $\sum_{d\mod D^{\prime}}\zeta_{D^{\prime}}^{-d^{-1}t^{\prime}a}f_{d^{-1}t^{\prime}}|_{{\mbox{\scriptsize$\left(\begin{array}[]{cc}{{1}}&{{-d^{-1}t^{\prime}D^{\prime}{}^{-1}}}\\ {{0}}&{{1}}\end{array}\right)$}}}\otimes\psi^{\prime}(a)$
$=\sum_{d\mod D^{\prime}}(\sum_{a\mod D^{\prime}}\zeta_{D^{\prime}}^{-d^{-1}t^{\prime}a}\otimes\psi^{\prime}(a))f_{d^{-1}t^{\prime}}|_{{\mbox{\scriptsize$\left(\begin{array}[]{cc}{{1}}&{{-d^{-1}t^{\prime}D^{\prime}{}^{-1}}}\\ {{0}}&{{1}}\end{array}\right)$}}}$
$=\sum_{d\mod D^{\prime}}G_{\psi^{\prime}}\psi^{\prime}(-d^{-1}t^{\prime})^{-1}f_{\kappa(t^{\prime})}|_{{\mbox{\scriptsize$\left(\begin{array}[]{cc}{{1}}&{{-d^{-1}t^{\prime}D^{\prime}{}^{-1}}}\\ {{0}}&{{1}}\end{array}\right)$}}}$
$=D^{\prime}G_{\psi^{\prime}{}^{-1}}^{-1}\psi^{\prime}(-1)\sum_{d\mod D^{\prime}}\psi^{\prime}(-d^{-1}t^{\prime})^{-1}f_{\kappa(t^{\prime})}|_{{\mbox{\scriptsize$\left(\begin{array}[]{cc}{{1}}&{{-d^{-1}t^{\prime}D^{\prime}{}^{-1}}}\\ {{0}}&{{1}}\end{array}\right)$}}}$
[…]
Since $\psi^{\prime}$ is a character, we have
${\mathop{\rm Hom}\nolimits}(M_{f}\otimes M_{\psi^{\prime}},M_{f}\otimes M_{\psi^{\prime}})={\mathop{\rm Hom}\nolimits}(M_{f},M_{f}\otimes M_{\psi^{\prime}}\otimes M_{\psi^{\prime}}^{\otimes(-1)})={\mathop{\rm Hom}\nolimits}(M_{f},M_{f}).$
It then follows that $A_{f\otimes\psi^{\prime}}\cong A_{f}$. ∎
6.2. Euler factors and functional equation
For each prime $p$, we let $\delta_{p}$ denote the dimension of $M_{f,\lambda}^{I_{p}}$ for any $\lambda$ not dividing $p$, so
$\delta_{p}=\left\{\begin{array}[]{ll}2,\mbox{if $p\nmid N_{f}$,}\\ 1,\mbox{if $p|N_{f}$ and $a_{p}\neq 0$,}\\ 0,\mbox{if $p|N_{f}$ and $a_{p}=0.$}\end{array}\right$.
We set $L_{p}^{\mathop{\rm nv}\nolimits}(A_{f},s)=L_{p}(A_{f},s)$ if $\delta_{p}>0$, and $L_{p}^{\mathop{\rm nv}\nolimits}(A_{f},s)=1$ if $\delta_{p}=0$. We let $\Sigma_{e}=\Sigma_{e}(f)$ denote the set of primes $p$ such that $\delta_{p}=0$ and $L_{p}(A_{f},s)\neq 1$. , and set
$L^{\mathop{\rm nv}\nolimits}(A_{f},s)=\prod_{p}L_{p}^{\mathop{\rm nv}\nolimits}(A_{f},s)=\prod_{p\not\in\Sigma_{e}(f)}L_{p}(A_{f},s).$
We call the primes in $\Sigma_{e}$ exceptional for $f$.
Recall that if $\delta_{p}=2$, then writing
$L_{p}(f,s)=(1-a_{p}p^{-s}+\psi(p)p^{k-1-2s})^{-1}=(1-\alpha_{p}p^{-s})^{-1}(1-\beta_{p}p^{-s})^{-1},$
we have
$L_{p}(A_{f},s)=(1-\alpha_{p}\beta_{p}^{-1}p^{-s})^{-1}(1-p^{-s})^{-1}(1-\alpha_{p}^{-1}\beta_{p}p^{-s})^{-1}.$
If $\delta_{p}=1$, then
(38) $L_{p}(A_{f},s)=\left\{\begin{array}[]{ll}(1-p^{-1-s})^{-1}&\mbox{if $\pi_{p}(f)$ is special;}\\ (1-p^{-s})^{-1}&\mbox{if $\pi_{p}(f)$ is principal series.}\end{array}\right$.
To see this, note that if $\pi_{p}(f)$ is special, then $\rho_{f,\lambda}|_{G_{p}}$ has matrix $\chi\otimes{\mbox{\scriptsize$\left(\begin{array}[]{cc}{{\chi_{\ell}}}&{{*}}\\ {{0}}&{{1}}\end{array}\right)$}}$ and is indecomposible. Here $\chi$ is unramified and $\chi_{\ell}$ is the cyclotomic character. So ${\mathop{\rm ad}\nolimits}^{0}\rho_{f,\lambda}$ has matrix $\left(\begin{array}[]{ccc}\chi_{\ell}&*&*\\ &1&*\\ &&\chi_{\ell}^{-1}\end{array}\right)$. So ${\mathop{\rm ad}\nolimits}^{0}\rho_{f,\lambda}^{I_{p}}$ is one dimensional acted by $G_{p}$ through the character $\chi_{\ell}$. Since $\chi_{\ell}(\operatorname{{\mathop{\rm Frob}\nolimits}}_{p})=p^{-1}$, we have $L_{p}(A_{f},s)=(1-p^{-1-s})^{-1}$.
If $\pi_{p}(f)$ is principal series, then $\rho_{f,\lambda}|_{G_{p}}$ has matrix $\left(\begin{array}[]{cc}{{\chi_{1}}}&{{0}}\\ {{0}}&{{\chi_{2}}}\end{array}\right)$ with $\chi_{1}$ ramified and $\chi_{2}$ unramified. So ${\mathop{\rm ad}\nolimits}^{0}\rho_{f,\lambda}$ has matrix $\left(\begin{array}[]{ccc}\chi_{1}\chi_{2}^{-1}&&\\ &1&\\ &&\chi_{1}^{-1}\chi_{2}\end{array}\right)$. So ${\mathop{\rm ad}\nolimits}^{0}\rho_{f,\lambda}^{I_{p}}$ is one dimensional with trivial $G_{p}$-action. So $L_{p}(A_{f},s)=(1-p^{-s})^{-1}$.
Shimura [Sh2] proved that $L(A_{f},s)$ extends to an entire function on the complex plane. Recall that we regard $L(M,s)$ as taking values in $K\otimes{\mathbb{C}}$. Each embedding $\tau:K\to{\mathbb{C}}$ gives a map $K\otimes{\mathbb{C}}\to{\mathbb{C}}$ and we write $L(M,\tau,s)$ for the composite with $L(M,s)$. Moreover, the work of Gelbart-Jacquet [G-J] and others (see [Schm]) shows that
$\begin{array}[]{rcl}\Lambda(A_{f},s)&=&L(A_{f},s)\Gamma_{\mathbb{R}}(s+1)\Gamma_{\mathbb{C}}(s+k-1)\\ &=&2^{2-k-s}\pi^{(1-2k-3s)/2}L(A_{f},s)\Gamma(\frac{s+1}{2})\Gamma(s+k-1)\end{array}$
satisfies the functional equation
$\Lambda(A_{f},s)=\epsilon(A_{f},s)\Lambda(A_{f},1-s),$
where $\epsilon(A_{f},s)$ is as defined by Deligne [De2]. Here we have used that $A_{f}$ and ${\mathop{\rm Hom}\nolimits}_{K}(A_{f},K)$ are isomorphic (using (36)).
6.3. $\Sigma$-variations for non-exceptional primes
Let $\Sigma$ denote a finite set of rational primes. We assume that no prime in $S$ divides a prime in $\Sigma$ and $\Sigma_{e}\subseteq\Sigma$. The case when $\Sigma_{e}\not\subseteq\Sigma$ will be considered in §6.5. We define a premotivic structure $M_{f}{{}^{\Sigma}}$ exactly as above, but we replace $N_{f}$ by $N{{}^{\Sigma}}=N_{f}\displaystyle\prod_{p\in\Sigma}p^{\delta_{p}}$ and $f(z)$ by the eigenform $f{{}^{\Sigma}}(z)=\displaystyle\sum a_{n}^{\prime}e^{2\pi inz}$, where $a_{n}^{\prime}=0$ if $n$ is divisible by a prime in $\Sigma$, and $a_{n}^{\prime}=a_{n}$ otherwise. We will next related the premotivic structure of $f{{}^{\Sigma}}$ to the premotivic structure of $f$.
For positive integers $m$ dividing $N^{\Sigma}/N_{f}=\prod_{p\in\Sigma}p^{\delta_{p}}$, we let
$\epsilon_{m}=m^{-1}\left[U^{\Sigma}{\mbox{\scriptsize$\left(\begin{array}[]{cc}{{m^{-1}}}&{{0}}\\ {{0}}&{{1}}\end{array}\right)$}}U\right]_{1}=m^{1-k}\left[U{{}^{\Sigma}}{\mbox{\scriptsize$\left(\begin{array}[]{cc}{{1}}&{{0}}\\ {{0}}&{{m}}\end{array}\right)$}}U\right]_{1}:M(N,\psi)\to M(N{{}^{\Sigma}},\psi)$
and define the endomorphism $\phi_{m}$ of $M(N,\psi)$ by
  • $\phi_{1}=1$, $\phi_{p}=-T_{p}$, $\phi_{p^{2}}=pS_{p}$;
  • $\phi_{m_{1}m_{2}}=\phi_{m_{1}}\phi_{m_{2}}$ if $(m_{1},m_{2})=1$.
Proposition 6.3.
The morphism
$\gamma=\sum_{m}\epsilon_{m}\phi_{m}:M(N,\psi)\to M(N{{}^{\Sigma}},\psi)$
induces an isomorphism $M_{f}\to M_{f}^{\Sigma}$ in $\operatorname{{\bf PM}}^{S}$. Furthermore, $\gamma_{\mathop{\rm dR}\nolimits}(f)=f{{}^{\Sigma}}$.
Proof.
We first reduce the proof to the case when $\Sigma$ contains a single prime.
Let $\Sigma^{\prime}$ be a subset of $\Sigma$. Define $U^{\Sigma^{\prime}}$ and $M(N^{\Sigma^{\prime}},\psi)$ as above. Let $\Sigma_{0}\subsetneq\Sigma_{1}$ be subsets of $\Sigma$ and let $m$ be a positive integer dividing $N^{\Sigma_{1}}/N^{\Sigma_{0}}$. Define
$\epsilon_{m}^{\Sigma_{0},\Sigma_{1}}:m^{-1}[U^{\Sigma_{1}}{\mbox{\scriptsize$\left(\begin{array}[]{cc}{{m^{-1}}}&{{0}}\\ {{0}}&{{1}}\end{array}\right)$}}U^{\Sigma_{0}}]_{1}=m^{1-k}[U^{\Sigma_{1}}{\mbox{\scriptsize$\left(\begin{array}[]{cc}{{1}}&{{0}}\\ {{0}}&{{m}}\end{array}\right)$}}U^{\Sigma_{0}}]_{1}.$
Also define $\phi_{m}^{\Sigma_{0}}$ to be the $\phi_{m}$ defined above acting on $M(N^{\Sigma_{0}},\psi)$ and define $\gamma_{m}^{\Sigma_{0},\Sigma_{1}}=\epsilon_{m}^{\Sigma_{0},\Sigma_{1}}\circ\phi_{m}^{\Sigma_{0}}.$
Lemma 6.4.
Let $\Sigma_{0}\subsetneq\Sigma_{1}\subsetneq\Sigma_{2}$ be subsets of $\Sigma$. Let $m_{1}$ (resp. $m_{2}$) be positive integers dividing $N^{\Sigma_{1}}/N^{\Sigma_{0}}$ (resp. $N^{\Sigma_{2}}/N^{\Sigma_{1}}$). The following equations hold.
  1. (a)
    $\epsilon_{m_{1}m_{2}}^{\Sigma_{0},\Sigma_{2}}=\epsilon_{m_{2}}^{\Sigma_{1},\Sigma_{2}}\circ\epsilon_{m_{1}}^{\Sigma_{0},\Sigma_{1}}.$
  2. (b)
    $\phi_{m_{2}}^{\Sigma_{1}}\circ\epsilon_{m_{1}}^{\Sigma_{0},\Sigma_{1}}=\epsilon_{m_{1}}^{\Sigma_{0},\Sigma_{1}}\circ\phi_{m_{2}}^{\Sigma_{0}}.$
  3. (c)
    $\gamma^{\Sigma_{0},\Sigma_{2}}_{m_{1}m_{2}}=\gamma^{\Sigma_{1},\Sigma_{2}}_{m_{2}}\circ\gamma^{\Sigma_{0},\Sigma_{1}}_{m_{1}}.$
Proof.
(a). We only need to prove
$U^{\Sigma_{2}}{\mbox{\scriptsize$\left(\begin{array}[]{cc}{{1}}&{{0}}\\ {{0}}&{{m_{2}}}\end{array}\right)$}}U^{\Sigma_{1}}]_{1}\circ[U^{\Sigma_{1}}{\mbox{\scriptsize$\left(\begin{array}[]{cc}{{1}}&{{0}}\\ {{0}}&{{m_{1}}}\end{array}\right)$}}U^{\Sigma_{0}}]_{1}=[U^{\Sigma_{2}}{\mbox{\scriptsize$\left(\begin{array}[]{cc}{{1}}&{{0}}\\ {{0}}&{{m_{1}m_{2}}}\end{array}\right)$}}U^{\Sigma_{0}}]_{1}.$
In Proposition 4.4 take $U=U^{\Sigma_{0}},U^{\prime}=U^{\Sigma_{1}},U^{\prime\prime}=U^{\Sigma_{2}},g={\mbox{\scriptsize$\left(\begin{array}[]{cc}{{1}}&{{0}}\\ {{0}}&{{m_{1}}}\end{array}\right)$}},g^{\prime}={\mbox{\scriptsize$\left(\begin{array}[]{cc}{{1}}&{{0}}\\ {{0}}&{{m_{2}}}\end{array}\right)$}}$ and $\tau=\tau^{\prime}={\mathop{\rm id}\nolimits}$. Then we have $U^{\prime}_{1}=U^{\prime}$. So in the double coset decomposition $U^{\prime}=\coprod_{k}U^{\prime}_{2}u_{k}U^{\prime}_{1}$ there is only one $u_{k}$ which we can take II. Further $W_{1}=W_{1}^{\prime}=U^{\prime\prime}$ and $\tau_{1}=1$. This gives us the desired equation.
(b). Because of (a) and similar multiplicity property of $\phi_{m}$, we only need to prove (b) in the special case $\Sigma_{1}\backslash\Sigma_{0}=\{p\}$ and $\Sigma_{2}\backslash\Sigma_{1}=\{q\}$ where $p$ and $q$ are primes.
So we just need to prove the commutativity of the diagram
[…]
for $0\leq i,j\leq 2$.
For this we first display the following elementary relations. Fix a prime number pp. For each integer $i\geq 0$, let $V_{i}=\{{\mbox{\scriptsize$\left(\begin{array}[]{cc}{{a}}&{{b}}\\ {{c}}&{{d}}\end{array}\right)$}}\in\operatorname{{\mathop{\rm GL}\nolimits}}_{2}({\mathbb{Z}}_{p})|c\equiv 0\mod p^{i}\}.$ Let $A_{t}={\mbox{\scriptsize$\left(\begin{array}[]{cc}{{1}}&{{0}}\\ {{0}}&{{p^{t}}}\end{array}\right)$}},t=0,1,2$. Let $B_{0}=I,B_{1}={\mbox{\scriptsize$\left(\begin{array}[]{cc}{{p}}&{{0}}\\ {{0}}&{{1}}\end{array}\right)$}},B_{2}={\mbox{\scriptsize$\left(\begin{array}[]{cc}{{p}}&{{0}}\\ {{0}}&{{p}}\end{array}\right)$}}$. Then
(39) $V_{2}IV_{0}][V_{0}BV_{0}]=[V_{2}BV_{0}]+[V_{2}AV_{0}]$
and
(40) $V_{2}BV_{2}][V_{2}A_{t}V_{0}]=\left\{\begin{array}[]{ll}[V_{2}BV_{0}],&t=0,\\ p[V_{2}BA_{t}V_{0}]=p^{k-1}[V_{2}A_{t-1}V_{0}],&0<t\leq 2.\end{array}\right$.
If $i>0$, then
(41) $V_{i+s}B_{t}V_{i+s}][V_{i+s}IV_{i}]=[V_{i+s}B_{t}V_{i}]=[V_{i+s}IV_{i}][V_{i}B_{t}V_{i}],0\leq s,t\leq 2$
and
(42) $V_{i+1}BV_{i+1}][V_{i+1}AV_{i}]=p[V_{i+1}pIV_{i}]=p^{k-1}[V_{i+1}IV_{i}].$
These relations can be easily verified applying Proposition 4.3 to double cosets in $\operatorname{{\mathop{\rm GL}\nolimits}}_{2}({\mathbb{Q}}_{p})$ instead of to $\operatorname{{\mathop{\rm GL}\nolimits}}_{2}({\mathbb{A}}_{\bf f})$.
We now continue with the proof of Lemma 6.4 (b) and verify
(43) $\epsilon_{p^{j}}^{\Sigma_{0},\Sigma_{1}}\circ\phi_{q^{i}}^{\Sigma_{0}}=\phi_{q^{i}}^{\Sigma_{1}}\circ\epsilon_{p^{j}}^{\Sigma_{0},\Sigma_{1}}.$
Since a double coset product in $\operatorname{{\mathop{\rm GL}\nolimits}}_{2}({\mathbb{A}}_{\bf f})$ is determined by its local factors at each finite prime, we only need to verify the equation at each prime. The equation at primes different from $p$ and $q$ follows from $V_{i+s}IV_{i}][V_{i}IV_{i}]=[V_{i+s}IV_{i+s}][V_{i+s}IV_{i}]$ which is obvious. At the prime $p$, the equation follows from
$V_{i+s}A_{t}V_{i}][V_{i}IV_{i}]=[V_{i+s}A_{t}V_{i}]=[V_{i+s}IV_{i+s}][V_{i+s}A_{t}V_{i}],0\leq s,t\leq 2,$
which is again obvious. At the prime $q$, the equation follows from (41) by choosing $(s,t)=(0,i)$. This proves (43) and hence (b) of Lemma 6.4.
(c) follows from (a) and (b):
(44) $\gamma^{\Sigma_{0},\Sigma_{2}}_{m_{1}m_{2}}=\alpha^{\Sigma_{0},\Sigma_{2}}_{m_{1}m_{2}}\circ\beta^{\Sigma_{0},\Sigma_{2}}_{m_{1}m_{2}}=\alpha^{\Sigma_{1},\Sigma_{2}}_{m_{2}}\circ\alpha^{\Sigma_{0},\Sigma_{1}}_{m_{1}}\circ\beta^{\Sigma_{1},\Sigma_{2}}_{m_{2}}\circ\beta^{\Sigma_{0},\Sigma_{1}}_{m_{1}}\\ =\alpha^{\Sigma_{1},\Sigma_{2}}_{m_{2}}\circ\beta^{\Sigma_{1},\Sigma_{2}}_{m_{2}}\circ\alpha^{\Sigma_{0},\Sigma_{1}}_{m_{1}}\circ\beta^{\Sigma_{0},\Sigma_{1}}_{m_{1}}=\gamma^{\Sigma_{1},\Sigma_{2}}_{m_{2}}\circ\gamma^{\Sigma_{0},\Sigma_{1}}_{m_{1}}.$
This completes the proof of Lemma 6.4. ∎
We continue with the proof of Proposition 6.3. For the given $\Sigma$ and $m$, let $p_{1},\ldots,p_{r}$ be the primes in $\Sigma$. Let $\Sigma_{0}=\phi$ and $\Sigma_{i}=\{p_{1},\ldots,p_{i}\},\ 1\leq i\leq r$. With this convention, we have $\epsilon_{m}=\epsilon_{m}^{\Sigma_{0},\Sigma_{r}}$ and $\phi_{m}=\phi_{m}^{\Sigma_{p}}$. Any positive divisor of $\prod_{i=1}^{r}p_{i}^{\delta_{p_{i}}}$ is of the form $\prod_{i=1}^{r}p^{e_{i}},0\leq e_{i}\leq\delta_{p_{i}}$. Recall that before the lemma, we define
$\gamma_{\prod_{i}p_{i}^{e_{i}}}^{\Sigma_{0},\Sigma_{r}}=\epsilon_{\prod_{i}p_{i}^{e_{i}}}^{\Sigma_{0},\Sigma_{r}}\circ\phi_{\prod_{i}p_{i}^{e_{i}}}^{\Sigma_{0},\Sigma_{r}}.$
Using Lemma 6.4, we obtain
$\gamma_{\prod_{i}p_{i}^{e_{i}}}^{\Sigma_{0},\Sigma_{r}}=\circ_{i=1}^{r}\gamma_{p_{i}^{e_{i}}}^{\Sigma_{i-1},\Sigma_{i}}.$
We therefore have
$\gamma=\sum_{i=1}^{r}\sum_{e_{i}=0}^{\delta_{p_{i}}}\gamma_{\prod_{i}p_{i}^{e_{i}}}^{\Sigma_{0},\Sigma_{r}}=\sum_{i=1}^{r}\sum_{e_{i}=0}^{\delta_{p_{i}}}\circ_{i=1}^{r}\gamma_{p_{i}^{e_{i}}}^{\Sigma_{i-1},\Sigma_{i}}=\circ_{i=0}^{r}(\sum_{e_{i}=0}^{\delta_{p_{i}}}\gamma_{p_{i}^{e_{i}}}^{\Sigma_{i-1},\Sigma_{i}}).$
Thus to prove Proposition 6.3, we only need to prove it in the case when $\Sigma$ contains one prime pp. Let $\Sigma_{0}=\phi$ and $\Sigma_{1}=\Sigma$. We consider three cases depending on $\delta_{p}=0,1$ or 22.
Case 1: When $\delta_{p}=0$, then $N^{\Sigma}=N$ and $\gamma$ is just the identity map. We have also seen that $a_{p}=0$. Thus $a_{pn}=0$. So $f=f^{\Sigma}$ and the proposition is proved in this case.
Case 2: When $\delta_{p}=1$,
$\gamma=\epsilon_{1}\circ\phi_{1}+\epsilon_{p}\circ\phi_{p}.$
By Lemma 6.4, for $q\neq p$,
$T_{q}\circ\gamma$ $=-\phi_{q}^{\Sigma_{1}}\circ\epsilon_{1}^{\Sigma_{1},\Sigma_{0}}\circ\phi_{1}^{\Sigma_{0}}-\phi_{q}^{\Sigma_{1}}\circ\epsilon_{p}^{\Sigma_{1},\Sigma_{0}}\circ\phi_{0}^{\Sigma_{0}}$
$=-\epsilon_{1}^{\Sigma_{1},\Sigma_{0}}\circ\phi_{q}^{\Sigma_{0}}\circ\phi_{1}^{\Sigma_{0}}-\epsilon_{p}^{\Sigma_{1},\Sigma_{0}}\circ\phi_{q}^{\Sigma_{1}}\circ\phi_{0}^{\Sigma_{0}}$
$=-\epsilon_{1}^{\Sigma_{1},\Sigma_{0}}\circ\phi_{1}^{\Sigma_{0}}\circ\phi_{q}^{\Sigma_{0}}-\epsilon_{p}^{\Sigma_{1},\Sigma_{0}}\circ\phi_{0}^{\Sigma_{0}}\circ\phi_{q}^{\Sigma_{1}}$
$=\gamma\circ T_{q}.$
When $q=p$, by (41) and (42), the double coset product of the local factor at primes other than $p$ involves II and is obviously zero. The local factor at $p$ is
$[V_{i+1}BV_{i+1}]([V_{i+1}A_{0}V_{i}][V_{i}B_{0}V_{i}]-p^{1-k}[V_{i+1}A_{1}V_{i}][V_{i+1}BV_{i}])$
$=[V_{i+1}BV_{i}][V_{i}IV_{i}]-p^{1-k}p[V_{i+1}pIV_{i}][V_{i}BV_{i}]$
$=[V_{i+1}IV_{i}][V_{i}BV_{i}]-p^{2-k}p^{k-2}[V_{i+1}IV_{i}][V_{i}BV_{i}]=0.$
So we get
$T_{p}\circ\gamma$ $=-\phi_{p}^{\Sigma_{1}}\circ\epsilon_{1}^{\Sigma_{0},\Sigma_{1}}\circ\phi_{1}^{\Sigma_{0}}-\phi_{p}^{\Sigma_{1}}\circ\epsilon_{p}^{\Sigma_{0},\Sigma_{1}}\circ\phi_{p}^{\Sigma_{0}}$
$=-\epsilon_{1}^{\Sigma_{0},\Sigma_{1}}\circ\phi_{p}^{\Sigma_{0}}\circ\phi_{1}^{\Sigma_{0}}-p^{1-k}p^{k-2}\epsilon_{1}^{\Sigma_{0},\Sigma_{1}}\circ\phi_{p}^{\Sigma_{0}}$
$=-p\epsilon_{1}^{\Sigma_{0},\Sigma_{1}}\circ\phi_{p}^{\Sigma_{0}}+p^{1-k}p^{k-2}\epsilon_{1}^{\Sigma_{0},\Sigma_{1}}\circ\phi_{p}^{\Sigma_{0}}=0.$
This shows that $\gamma_{\mathop{\rm dR}\nolimits}f$ is a multiple of $f{{}^{\Sigma}}$. So to prove the proposition, we only need to prove the last statement there.
We consider the action of $\gamma=\epsilon_{1}\circ\phi_{1}+\epsilon_{p}\circ\phi_{p}$ on $f$ as a classical modular form. First $\epsilon_{1}\circ\phi_{1}=[U{{}^{\Sigma}}IU]_{1}$ acts on ${\mathcal{M}}(N,\psi)_{M}\subseteq{\mathcal{M}}_{M}\otimes{\mathcal{V}}$ by $I\otimes 1$, where $M\geq 3$ is chosen such that $N|M$ and $S_{M}=S_{N}$. The action corresponding to II on the classical forms is II. Since $f$ has rational coefficients, its image in $\oplus_{t\in({\mathbb{Z}}/M{\mathbb{Z}})^{\times}}M(M,\psi)$ is of the form $(f_{0}(z))_{t}$ with $f_{0}$ independent of $t$. Then, in the notations of Lemma 3.2 and Proposition 3.7, for each $t^{\prime}\in({\mathbb{Z}}/pM{\mathbb{Z}})^{\times}$ we have $\kappa(t^{\prime})=t^{\prime}\mod M$ and $\tilde{\gamma}:E_{pM,t^{\prime}}\to E_{M,t}$ defined by $\gamma=\gamma_{t^{\prime}}=I.$ So $(f_{0}(z))_{t}$ is sent to $(f_{0}(z))_{t^{\prime}}$, $t^{\prime}\in({\mathbb{Z}}/pM{\mathbb{Z}})^{\times}$ by $\epsilon_{1}\circ\phi_{1}$. Next $U{{}^{\Sigma}}=U{{}^{\Sigma}}\cap({\mbox{\scriptsize$\left(\begin{array}[]{cc}{{1}}&{{0}}\\ {{0}}&{{p}}\end{array}\right)$}}U{\mbox{\scriptsize$\left(\begin{array}[]{cc}{{1}}&{{0}}\\ {{0}}&{{p}}\end{array}\right)$}}^{-1}$. So $\epsilon_{p}=p^{1-k}[U{{}^{\Sigma}}{\mbox{\scriptsize$\left(\begin{array}[]{cc}{{1}}&{{0}}\\ {{0}}&{{p}}\end{array}\right)$}}U]_{1}$ acts on ${\mathcal{M}}_{pM}\otimes{\mathcal{V}}$ by ${\mbox{\scriptsize$\left(\begin{array}[]{cc}{{1}}&{{0}}\\ {{0}}&{{p}}\end{array}\right)$}}\otimes 1$. The corresponding action on $\oplus_{t}M_{k}(\Gamma_{0}(pN))$ is given by taking, for each $t^{\prime}\in({\mathbb{Z}}/p^{2}M{\mathbb{Z}})^{\times}$, $\kappa(t^{\prime})=t^{\prime}\mod pM$ and $\tilde{\gamma}_{t^{\prime}}:E_{p^{2}M,t^{\prime}}\to E_{pM,t}$ given by $\gamma_{t^{\prime}}={\mbox{\scriptsize$\left(\begin{array}[]{cc}{{1}}&{{0}}\\ {{0}}&{{p^{-1}}}\end{array}\right)$}}.$ So $\epsilon_{p}$ sends $(f_{0})_{t}\in\oplus_{t}M_{k}(\Gamma_{0}(pM))$ to $(f_{0}(pz))_{t^{\prime}}$. Since $\phi_{p}=-T_{p}$, $\epsilon_{p}\circ\phi_{p}$ sends $(f_{0}(z))_{t}$ to $(-f_{0}{}_{|T_{p}})_{t^{\prime}}$. So by [Miy, Lemma 4.6.5], $\gamma$ sends $(f_{0}(z))_{t}$ to $(f_{0}{{}^{\Sigma}}(z))_{t^{\prime}}$, as is needed.
Case 3: When $\delta_{p}=2$, we have $p\nmid N$ and
$\gamma=\epsilon_{1}^{\Sigma_{0},\Sigma_{1}}\circ\phi_{1}^{\Sigma_{0}}+\epsilon_{p}^{\Sigma_{0},\Sigma_{1}}\circ\phi_{p}^{\Sigma_{0}}+\epsilon_{p^{2}}^{\Sigma_{0},\Sigma_{1}}\circ\phi_{p^{2}}^{\Sigma_{0}}.$
When $q\neq p$, the relation $T_{q}\gamma=\gamma T_{q}$ is proved in the same way as in Case 2. When $q=p$, we only need to consider the local factor of $T_{p}\gamma$ at the prime pp. The double coset product of the local factor is
$-[V_{2}BV_{2}][V_{2}IV_{0}][V_{0}IV_{0}]+p^{1-k}[V_{2}BV_{2}][V_{2}A_{1}U_{0}][V_{0}BV_{0}]$
$-p^{k-2}[V_{2}BV_{2}][V_{2}A_{2}V_{0}][V_{0}B_{2}V_{0}].$
By (40) with $t=0$, the first term is $V_{2}BV_{0}]$. By (40) with $t=1$ and (39), the second term is $V_{2}BV_{0}]-[V_{2}AV_{0}]$. By (40) with $t=2$, the third term is $V_{2}AV_{0}].$ Thus the equation is zero, as needed. It follows that $T_{p}\circ\gamma=0$.
We now show that $\gamma$ sends $f$ to $f{{}^{\Sigma}}$. This will complete the proof of Proposition 6.3. As in the case when $\delta_{p}=1$, we see that the action of $\epsilon_{1}\circ\phi_{1}+\epsilon_{p}\circ\phi_{p}$ on the classical modular forms sends $(f_{0}(z))_{t}\in\oplus_{t\in({\mathbb{Z}}/M{\mathbb{Z}})^{\times}}M_{k}(\Gamma_{0}(pM),\psi)$ to $(f_{0}(z)-(f_{0}{}_{|T_{p}}(pz))_{t^{\prime}}\in\oplus_{t^{\prime}\in({\mathbb{Z}}/p^{2}M{\mathbb{Z}})^{\times}}M_{k}(\Gamma_{0}(p^{2}M),\psi)$. Further, the action of $\epsilon_{p^{2}}=p^{2(1-k)}[U{{}^{\Sigma}}{\mbox{\scriptsize$\left(\begin{array}[]{cc}{{1}}&{{0}}\\ {{0}}&{{p^{2}}}\end{array}\right)$}}U]_{1}$ sends $(f_{0}(z))_{t}\in\oplus_{t\in({\mathbb{Z}}/M{\mathbb{Z}})^{\times}}M_{k}(M,\psi)$ to $(f_{0}(p^{2}z))_{t^{\prime}}\in\oplus_{t^{\prime}\in({\mathbb{Z}}/p^{2}M{\mathbb{Z}})^{\times}}M_{k}(p^{2}M,\psi)$. Thus $\epsilon_{p^{2}}\circ\phi_{p^{2}}=\psi(p_{p})^{-1}p^{k-1}\epsilon_{p^{2}}$ sends $(f_{0}(z))_{t}\in\oplus_{t\in({\mathbb{Z}}/M{\mathbb{Z}})^{\times}}M_{k}(M,\psi)$ to
$(\psi(p_{p})^{-1}p^{k-1}f_{0}(p^{2}z))_{t^{\prime}}\in\oplus_{t^{\prime}\in({\mathbb{Z}}/p^{2}M{\mathbb{Z}})^{\times}}M_{k}(p^{2}M),\psi).$
Therefore $\gamma$ send $(f_{0}(z))_{t}$ to
$(f_{0}(z)-f_{0}{}_{|T_{p}}(pz)+\psi(p_{p})^{-1}p^{k-1}f_{0}(p^{2}z))_{t^{\prime}}=(f_{0}{{}^{\Sigma}}(z))_{t^{\prime}}.$
This is what we want. ∎
Proposition 6.5.
Let $\gamma^{t}$ denote the adjoint of $\gamma$. Then
$\gamma^{t}\circ\gamma=\phi_{N{{}^{\Sigma}}/N_{f}}\prod_{p\in\Sigma}L_{p}^{\mathop{\rm nv}\nolimits}(A_{f},1)^{-1}$
on $M_{f}$.
Proof.
As in the proof of Proposition 6.3, we can again specialize to the case when $\Sigma$ has a single prime pp. Then we only need to consider the place at pp.
When $\delta_{p}=0$, $\gamma$ is the identity map and the formula is clear.
When $\delta_{p}=1$, we have
$\gamma=\epsilon_{1}\circ\phi_{1}+\epsilon_{p}\circ\phi_{p}=[\epsilon_{1},\epsilon_{p}]\left[\begin{array}[]{l}\phi_{1}\\ \phi_{p}\end{array}\right].$
By Lemma 5.3, we find that the adjoints are given by
$\begin{array}[]{l}\epsilon_{1}^{t}=[U{{}^{\Sigma}}IU]_{1}^{t}=[U{\mbox{\scriptsize$\left(\begin{array}[]{cc}{{p}}&{{0}}\\ {{0}}&{{1}}\end{array}\right)$}}U{{}^{\Sigma}}]_{1},\\ \phi_{1}^{t}=[UIU]_{1}^{t}=[UIU]_{1},\\ \epsilon_{p}^{t}=p^{1-k}[Uw^{-1}U]_{\omega^{t}\otimes\psi({\mathop{\rm det}\nolimits}w)}\circ[U{\mbox{\scriptsize$\left(\begin{array}[]{cc}{{1}}&{{0}}\\ {{0}}&{{p}}\end{array}\right)$}}U{{}^{\Sigma}}]_{1}\circ[U{{}^{\Sigma}}w^{\prime}U{{}^{\Sigma}}]_{\omega^{\prime}}=p^{-1}[UIU{{}^{\Sigma}}]_{1}{\rm\ and}\\ \phi_{p}^{t}=-[U{\mbox{\scriptsize$\left(\begin{array}[]{cc}{{p_{p}}}&{{0}}\\ {{0}}&{{1}}\end{array}\right)$}}U]_{\psi(p_{p})^{-1}}^{t}=-[U{\mbox{\scriptsize$\left(\begin{array}[]{cc}{{p_{p}}}&{{0}}\\ {{0}}&{{1}}\end{array}\right)$}}U]_{\psi(p_{p})^{-1}}.\end{array}$
Then by Proposition 4.4 and Lemma 5.8,
$\gamma^{t}\gamma$ $=[\phi_{1}^{t},\phi_{p}^{t}]\left[\begin{array}[]{ll}\epsilon_{1}^{t}\epsilon_{1}&\epsilon_{1}^{t}\epsilon_{p}\\ \epsilon_{p}^{t}\epsilon_{1}&\epsilon_{p}^{t}\epsilon_{p}\\ \end{array}\right]\left[\begin{array}[]{l}\phi_{1}\\ \phi_{p}\end{array}\right]$
$=[1,-T_{p}]\left[\begin{array}[]{cc}T_{p}&1\\ 1&p^{-k}[U{\mbox{\scriptsize$\left(\begin{array}[]{cc}{{1}}&{{0}}\\ {{0}}&{{p}}\end{array}\right)$}}U]_{1}\end{array}\right]\left[\begin{array}[]{l}1\\ -T_{p}\end{array}\right]$
$=\left[0,1-\left\{\begin{array}[]{ll}p^{-k}s_{p},&\mbox{if $\pi_{p}$ is special}\\ p^{-k}ps_{p},&\mbox{if $\pi_{p}$ is principal series}\end{array}\right\}\right]\left[\begin{array}[]{l}1\\ -T_{p}\end{array}\right]$
$=\left\{\begin{array}[]{ll}(1-p^{-2})\phi_{p},&\mbox{if $\pi_{p}$ is special,}\\ (1-p^{-1})\phi_{p},&\mbox{if $\pi_{p}$ is principal series.}\end{array}\right$.
By (38), this is what we need.
When $\delta_{p}=2$, we have $p\nmid N$ and $N{{}^{\Sigma}}=p^{2}N$. Also
$\gamma=\epsilon_{1}\circ\phi_{1}+\epsilon_{p}\circ\phi_{p}+\epsilon_{p^{2}}\circ\phi_{p^{2}}=[\epsilon_{1},\epsilon_{p},\epsilon_{p^{2}}]\left[\begin{array}[]{l}\phi_{1}\\ \phi_{p}\\ \phi_{p^{2}}\end{array}\right].$
By Lemma 5.3,
$\epsilon_{1}^{t}$ $=[U{{}^{\Sigma}}IU]_{1}^{t}=[Uw^{-1}U]_{\omega^{t}\otimes\psi({\mathop{\rm det}\nolimits}w)}\circ[UIU{{}^{\Sigma}}]_{1}\circ[U{{}^{\Sigma}}w^{\prime}U{{}^{\Sigma}}]_{\omega^{\prime}}$
$=\psi(p)^{2}[U{\mbox{\scriptsize$\left(\begin{array}[]{cc}{{p^{2}}}&{{0}}\\ {{0}}&{{1}}\end{array}\right)$}}U{{}^{\Sigma}}]_{1},$
$\phi_{1}^{t}$ $=[UIU]_{1}^{t}=[UIU]_{1},$
$\epsilon_{p}^{t}$ $=p^{1-k}[U{{}^{\Sigma}}{\mbox{\scriptsize$\left(\begin{array}[]{cc}{{1}}&{{0}}\\ {{0}}&{{p}}\end{array}\right)$}}U]_{1}^{t}=p^{-1}\psi(p)[U{\mbox{\scriptsize$\left(\begin{array}[]{cc}{{p}}&{{0}}\\ {{0}}&{{1}}\end{array}\right)$}}U{{}^{\Sigma}}]_{1},$
$\phi_{p}^{t}$ $=-[U{\mbox{\scriptsize$\left(\begin{array}[]{cc}{{p_{p}}}&{{0}}\\ {{0}}&{{1}}\end{array}\right)$}}U]_{\psi(p_{p})^{-1}}^{t}=-[U{\mbox{\scriptsize$\left(\begin{array}[]{cc}{{p_{p}}}&{{0}}\\ {{0}}&{{1}}\end{array}\right)$}}U]_{\psi(p_{p})^{-1}},$
$\epsilon_{p^{2}}^{t}$ $=p^{2(1-k)}[U{{}^{\Sigma}}{\mbox{\scriptsize$\left(\begin{array}[]{cc}{{1}}&{{0}}\\ {{0}}&{{p^{2}}}\end{array}\right)$}}U]_{1}^{t}=p^{2(1-k)}[U{\mbox{\scriptsize$\left(\begin{array}[]{cc}{{p^{2}}}&{{0}}\\ {{0}}&{{p^{2}}}\end{array}\right)$}}U{{}^{\Sigma}}]_{1}=p^{-2}[UIU{{}^{\Sigma}}]_{1}{\rm\ and\ }$
$\phi_{p^{2}}^{t}$ $=p[U{\mbox{\scriptsize$\left(\begin{array}[]{cc}{{p_{p}}}&{{0}}\\ {{0}}&{{p_{p}}}\end{array}\right)$}}U]_{\psi(p_{p})^{-2}}^{t}=p^{k-1}\psi(p_{p})^{-1}[UIU]_{1}^{t}=\psi(p)p^{k-1}[UIU]_{1}.$
So by Proposition 4.4,
$\gamma^{t}\gamma=[\phi_{1}^{t},\phi_{p}^{t},\phi_{p^{2}}^{t}]\left[\begin{array}[]{lll}\epsilon_{1}^{t}\epsilon_{1}&\epsilon_{1}^{t}\epsilon_{p}&\epsilon_{1}^{t}\epsilon_{p^{2}}\\ \epsilon_{p}^{t}\epsilon_{1}&\epsilon_{p}^{t}\epsilon_{p}&\epsilon_{p}^{t}\epsilon_{p^{2}}\\ \epsilon_{p^{2}}^{t}\epsilon_{1}&\epsilon_{p^{2}}^{t}\epsilon_{p}&\epsilon_{p^{2}}^{t}\epsilon_{p^{2}}\end{array}\right]\left[\begin{array}[]{l}\phi_{1}\\ \phi_{p}\\ \phi_{p^{2}}\end{array}\right]$
$=[1,-T_{p},\psi(p)p^{k-1}]\hskip-2.84544pt\left[\hskip-4.26773pt\begin{array}[]{ccc}T_{p}^{2}-(p+1)\psi(p)p^{k-2}\hskip-22.76228pt&T_{p}&p^{-1}(p+1)\\ T_{p}&p^{-1}(p+1)&\psi(p)^{-1}p^{-k}T_{p}\\ p^{-1}(p+1)&(\psi(p)p^{k})^{-1}T_{p}&\hskip-5.69046pt(p^{2k}\psi(p)^{2})^{-1}(T_{p}^{2}-(p+1)\psi(p)p^{k-2})\end{array}\hskip-4.26773pt\right]\hskip-4.26773pt\left[\hskip-4.26773pt\begin{array}[]{l}1\\ -T_{p}\\ \psi(p)p^{k-1}\end{array}\hskip-5.69046pt\right]$
$=[0,0,(1-p^{-1})((1+p^{-1})^{2}-\psi(p)^{-1}p^{-k}T_{p}^{2})]\left[\begin{array}[]{l}1\\ -T_{p}\\ \psi(p)p^{k-1}\end{array}\right]$
$=(1-p^{-1})((1+p^{-1})^{2}-\psi(p)^{-1}p^{-k}T_{p}^{2})\psi(p)p^{k-1}.$
So the action of $\gamma^{t}\gamma$ on ${\mathcal{M}}(N,\psi)$ is
$(1-p^{-1})$ $((1+p^{-1})^{2}-\psi(p)^{-1}p^{-k}a_{p}^{2})\psi(p)p^{k-1}$
$=(1-p^{-1})((1+p^{-1})^{2}-\psi(p)^{-1}p^{-k}(\alpha_{p}+\beta_{p})^{2})\psi(p)p^{k-1}$
$=(1-p^{-1})((1+p^{-1})^{2}-\psi(p)^{-1}p^{-k}(\alpha_{p}^{2}+2\psi(p)p^{k-1}+\beta_{p}^{2}))\psi(p)p^{k-1}$
$=(1-p^{-1})(1+p^{-2}-\psi(p)^{-1}p^{-k}(\alpha_{p}^{2}+\beta_{p}^{2}))\psi(p)p^{k-1}.$
On the other hand,
$\phi_{p^{2}}p[U{\mbox{\scriptsize$\left(\begin{array}[]{cc}{{p_{p}}}&{{0}}\\ {{0}}&{{p_{p}}}\end{array}\right)$}}U]$ $(1-\alpha_{p}\beta_{p}^{-1}p^{-1})(1-p^{-1})(1-\alpha_{p}^{-1}\beta_{p}p^{-1})$
$=p^{k-1}\psi(p)(1-p^{-1})(1+p^{-2}-(\alpha_{p}^{2}+\beta_{p}^{2})/(\alpha_{p}\beta_{p}p))$
$=p^{k-1}\psi(p)(1-p^{-1})(1+p^{-2}-(\alpha_{p}^{2}+\beta_{p}^{2})\psi(p)^{-1}p^{-k}).$
This proves the proposition when $\delta_{p}=2$. ∎
6.4. Integral structure for a $\Sigma$-variation
For a newform $f$ and a finite set of primes $\Sigma$ invertible in $\mathcal{O}$}, we set
${\mathcal{M}}_{f}^{\Sigma}=M_{f}^{\Sigma}\cap{\mathcal{M}}(N{{}^{\Sigma}},\psi^{\Sigma})_{!}.$
Note that $\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}{\mathcal{M}}_{f,{\mathop{\rm dR}\nolimits}}^{\Sigma}={\mathcal{O}}\cdot f^{\Sigma}$ by Proposition 6.3. The image of $\wedge^{2}_{\mathcal{O}}{\mathcal{M}}_{f,B}^{\Sigma}$ defines an integral structure for $M_{\psi}(1-k)$ of the form $\eta_{f}^{\Sigma}\otimes_{\mathcal{O}}{\mathcal{M}}_{\psi}(1-k)$ for some fractional $\mathcal{O}$}-ideal $\eta_{f}^{\Sigma}\subset K$. We call $\eta_{f}^{\Sigma}$ the (naive, $\Sigma$-finite) congruence $\mathcal{O}$}-ideal of $f$. Note that $\eta_{f}^{\Sigma}$ is well-behaved under extension of scalars ${\mathcal{O}}^{\prime}\otimes_{\mathcal{O}}\cdot$ if $K\subset K^{\prime}$ and $S^{\prime}$ is contained in the set of primes dividing those in $S$.
Proposition $6.$
If ${\mathcal{M}}_{f,\lambda}/\lambda$ is an irreducible $({\mathcal{O}}/\lambda)[G_{\mathbb{Q}}]$-module for every $\lambda$ in $S$, then the map $M_{f}\to M_{f}^{\Sigma}$ induces an isomorphism ${\mathcal{M}}_{f}\to{\mathcal{M}}_{f}^{\Sigma}$ in ${\mathcal{P}M}^{S}$.
Proof.
Fix a $\lambda\in S$. We will identify realizations of ${\mathcal{M}}_{f}$ with their images in ${\mathcal{M}}_{f}^{\Sigma}$ through the isomorphism $M_{f}\to M_{f}^{\Sigma}$. Since ${\mathcal{M}}_{f,\lambda}$ is compact and $\lambda^{n}{\mathcal{M}}_{f,\lambda}^{\Sigma},n\in{\mathbb{Z}}$ is an open covering of $M_{f,\lambda}^{\Sigma}$, there is $n=n_{\lambda}\in{\mathbb{Z}}$ such that ${\mathcal{M}}_{f,\lambda}\subset\lambda^{n}{\mathcal{M}}_{f,\lambda}^{\Sigma}$ and ${\mathcal{M}}_{f,\lambda}\not\subset\lambda^{n+1}{\mathcal{M}}_{f,\lambda}^{\Sigma}$. So
$\lambda^{n+1}{\mathcal{M}}_{f,\lambda}^{\Sigma}\subset\lambda^{n+1}{\mathcal{M}}_{f,\lambda}^{\Sigma}+{\mathcal{M}}_{f,\lambda}\subset\lambda^{n}{\mathcal{M}}_{f,\lambda}^{\Sigma}.$
Since ${\mathcal{M}}_{f,\lambda}^{\Sigma}/\lambda$ and hence $\lambda^{n}{\mathcal{M}}_{f,\lambda}^{\Sigma}/\lambda^{n+1}{\mathcal{M}}_{f,\lambda}$ are irreducible, one of the inclusions must be an equality. The left inclusion being equal means ${\mathcal{M}}_{f,\lambda}\subset\lambda^{n+1}{\mathcal{M}}_{f,\lambda}^{\Sigma}$, a contradiction. So we must have $\lambda^{n+1}{\mathcal{M}}_{f,\lambda}^{\Sigma}+{\mathcal{M}}_{f,\lambda}\subset\lambda^{n}{\mathcal{M}}_{f,\lambda}^{\Sigma}.$ Then by Nakayama Lemma, we have ${\mathcal{M}}_{f,\lambda}=\lambda^{n}{\mathcal{M}}_{f,\lambda}^{\Sigma}.$
Define ${\mathfrak{b}}=\prod_{\lambda\in S}\lambda^{n_{\lambda}}$. Then $\mathfrak{b}$} is a fractional ideal of $\mathcal{O}$}. Then ${\mathcal{O}}_{\lambda}\otimes({\mathfrak{b}}\otimes_{\mathcal{O}}{\mathcal{M}}_{f,B}^{\Sigma})\cong{\mathcal{M}}_{f,\lambda}^{\Sigma}.$ Then by the comparison isomorphisms between ${\mathcal{M}}_{f,B}^{\Sigma}$ and ${\mathcal{M}}_{f,\lambda}^{\Sigma}$ for $\lambda\in S$, we have
${\mathcal{O}}_{\lambda}\otimes({\mathfrak{b}}\otimes_{\mathcal{O}}{\mathcal{M}}_{f,B}^{\Sigma})\cong{\mathcal{O}}_{\lambda}\otimes{\mathcal{M}}_{B}^{\prime},\lambda\in S.$
Therefore ${\mathcal{M}}_{f,B}\cong{\mathfrak{b}}\otimes{\mathcal{M}}_{f,B}^{\Sigma}.$ Similarly, using comparison isomorphisms between ${\mathcal{M}}{{}^{\Sigma}}_{f,\lambda}$ and ${\mathcal{M}}{{}^{\Sigma}}_{f,{\lambda\mbox{\scriptsize-crys}}}$, and between ${\mathcal{M}}_{f,{\mathop{\rm dR}\nolimits}}^{\Sigma}$ and ${\mathcal{M}}_{f,{\ell\mbox{\scriptsize-crys}}}^{\Sigma}$, we see that ${\mathcal{M}}_{f,{\mathop{\rm dR}\nolimits}}\cong{\mathfrak{b}}\otimes{\mathcal{M}}_{f,{\mathop{\rm dR}\nolimits}}^{\Sigma}$ in ${\mathcal{P}M}^{S,S}$. In particular, $\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}{\mathcal{M}}_{f,{\mathop{\rm dR}\nolimits}}\cong{\mathfrak{b}}\otimes\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}{\mathcal{M}}_{f,{\mathop{\rm dR}\nolimits}}^{\Sigma}.$ Since we have seen above that the image of $\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}{\mathcal{M}}_{f,{\mathop{\rm dR}\nolimits}}$ under the isomorphism $M_{f}\to M_{f}^{\Sigma}$ is ${\mathcal{O}}\cdot f^{\Sigma}$ which is also $\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}{\mathcal{M}}_{f,{\mathop{\rm dR}\nolimits}}^{\Sigma}$, we conclude that ${\mathfrak{b}}={\mathcal{O}}$. ∎
Combining this with Proposition 6.3, we obtain:
Corollary 6.7.
For $f$ and $S$ as in proposition 6.6, we have
$\eta_{f}^{\Sigma}=\eta_{f}^{\emptyset}\prod_{p\in\Sigma}L_{p}^{\mathop{\rm nv}\nolimits}(A_{f},1)^{-1}.$
Proof.
We only need to consider the case when $\Sigma=\{p\}$. Consider the following commutative diagram
[…]
Since $\gamma:{\mathcal{M}}_{f,B}\to{\mathcal{M}}{{}^{\Sigma}}_{f,B}$ is an isomorphism, we have
$\langle\gamma{\mathcal{M}}_{f,B},\gamma{\mathcal{M}}_{f,B}\rangle{{}^{\Sigma}}=\langle{\mathcal{M}}{{}^{\Sigma}}_{f,B},{\mathcal{M}}{{}^{\Sigma}}_{f,B}\rangle{{}^{\Sigma}}=\eta_{f}{{}^{\Sigma}}\otimes_{\mathcal{O}}{\mathcal{M}}_{\psi}(1-k)_{B}.$
On the other hand, by Proposition 6.5,
$\langle\gamma{\mathcal{M}}_{f,B},\gamma{\mathcal{M}}_{f,B}\rangle{{}^{\Sigma}}=\langle\gamma^{t}\circ\gamma{\mathcal{M}}_{f,B},{\mathcal{M}}_{f,B}\rangle=\langle\phi_{p^{\delta_{p}}}{\mathcal{M}}_{f,B},{\mathcal{M}}_{f,B}\rangle L_{p}^{\mathop{\rm nv}\nolimits}(A_{f},1)^{-1}.$
When $\delta_{p}=0$, $\phi_{p^{\delta_{p}}}=1$ and we have
$\eta_{f}{{}^{\Sigma}}\otimes_{\mathcal{O}}{\mathcal{M}}_{\psi}(1-k)_{B}$ $=\langle{\mathcal{M}}_{f,B},{\mathcal{M}}_{f,B}\rangle L_{p}^{\mathop{\rm nv}\nolimits}(A_{f},1)^{-1}$
$=\eta_{f}^{\phi}L_{p}^{\mathop{\rm nv}\nolimits}(A_{f},1)^{-1}\otimes_{\mathcal{O}}{\mathcal{M}}_{\psi}(1-k)_{B}.$
When $\delta_{p}=1$, $\phi_{p^{\delta_{p}}}=-T_{p}$ and we have
$\eta_{f}{{}^{\Sigma}}\otimes_{\mathcal{O}}{\mathcal{M}}_{\psi}(1-k)_{B}$ $=\psi(p)^{-1}a_{p}\langle{\mathcal{M}}_{f,B},{\mathcal{M}}_{f,B}\rangle L_{p}^{\mathop{\rm nv}\nolimits}(A_{f},1)^{-1}$
$=\eta_{f}^{\phi}L_{p}^{\mathop{\rm nv}\nolimits}(A_{f},1)^{-1}\otimes_{\mathcal{O}}{\mathcal{M}}_{\psi}(1-k)_{B}.$
When $\delta_{p}=2$, $\phi_{p^{\delta_{p}}}=pS_{p}$ and we have
$\eta_{f}{{}^{\Sigma}}\otimes_{\mathcal{O}}{\mathcal{M}}_{\psi}(1-k)_{B}$ $=\psi(p)p^{k-1}\langle{\mathcal{M}}_{f,B},{\mathcal{M}}_{f,B}\rangle L_{p}^{\mathop{\rm nv}\nolimits}(A_{f},1)^{-1}$
$=\psi(p)p^{k-1}\eta_{f}^{\phi}L_{p}^{\mathop{\rm nv}\nolimits}(A_{f},1)^{-1}\otimes_{\mathcal{O}}{\mathcal{M}}_{\psi}(1-k)_{B}.$
This proves the corollary since $p$ is invertible in $\mathcal{O}$}. ∎
6.5. Exceptional primes
[…]
Let $F$ denote the quadratic unramified extension of ${\mathbb{Q}}_{p}$ in $\bar{{\mathbb{Q}}}_{p}$, and let $H_{p}$ denote the subgroup of $G_{p}$ fixing $F$.
Lemma 6.8.
Let $f$ have minimal conductor among its twists. If $\pi_{p}(f)$ is supercuspidal, then the following are equivalent:
  1. (a)
    $p$ is in $\Sigma_{e}$;
  2. (b)
    $L_{p}(A_{f},s)=(1+p^{-s})^{-1}$;
  3. (c)
    $\bar{K}_{\lambda}\otimes_{K_{\lambda}}M_{f,\lambda}|_{I_{p}}$ is reducible for some (hence all) $\lambda\in S$;
  4. (d)
    $\bar{K}_{\lambda}\otimes_{K_{\lambda}}M_{f,\lambda}|_{G_{p}}$ is induced from a character of $H_{p}$ for some (hence all) $\lambda\in S$;
Proof.
(a) $\Rightarrow$ (c): If $p$ is in $\Sigma_{e}$ then $L_{p}(A_{f},s)\neq 1$. This means that $A_{f,\lambda}^{I_{p}}$ is not zero which means that ${\mathop{\rm Hom}\nolimits}(M_{f,\lambda},M_{f,\lambda})^{I_{p}}\supsetneq K_{\lambda}$. By Schur’s lemma, this shows that, as a representation of $I_{p}$, $\bar{K}_{\lambda}\otimes_{K_{\lambda}}M_{f,\lambda}$ is reducible.
(c) $\Rightarrow$ (d): Assume that $\bar{K}_{\lambda}\otimes_{K_{\lambda}}M_{f,\lambda}|_{I_{p}}$ is reducible. Since the image of $I_{p}$ is finite, $\bar{K}_{\lambda}\otimes_{K_{\lambda}}M_{f,\lambda}|_{I_{p}}$ decomposes into $V_{1}\oplus V_{2}$ of representations of $I_{p}$. Let $\operatorname{{\mathop{\rm Frob}\nolimits}}_{p}$ be a lift of the Frobenius in $G_{p}/I_{p}$. Since $M_{\lambda}|_{G_{p}}$ is irreducible, the two terms $V_{1}$ and $V_{2}$ are exchanged by $\operatorname{{\mathop{\rm Frob}\nolimits}}_{p}$ and preserved by $\operatorname{{\mathop{\rm Frob}\nolimits}}_{p}^{2}$. So $V_{1}$ and $V_{2}$ are preserved by the subgroup of $G_{p}$ generated by $I_{p}$ and $\operatorname{{\mathop{\rm Frob}\nolimits}}_{p}^{2}$, which is $H_{p}$. Further $M_{f,\lambda}$ is the representation of $G_{p}$ induced by the representation $V_{1}$ of $H_{p}$.
(d) $\Rightarrow$ (b): If $\bar{K}_{\lambda}\otimes_{K_{\lambda}}M_{f,\lambda}|_{G_{p}}$ is induced from a character of $H_{p}$ for some $\lambda\in S$. Then the action of $D_{p}$ on $\bar{K}_{\lambda}\otimes_{K_{\lambda}}A_{\lambda}$ has the matrix $\left(\begin{array}[]{ccc}\varepsilon&&\\ &\chi&\\ &&\varepsilon^{-1}\end{array}\right)$ where $\varepsilon$ is a ramified character and $\chi$ is an unramified quadratic character. Thus $A_{\lambda}^{I_{p}}=K_{\lambda}(\chi)$ and $\chi(\operatorname{{\mathop{\rm Frob}\nolimits}}_{p})=-1$. This shows that $L_{p}(A,s)=(1+p^{-s})^{-1}$.
(b) $\Rightarrow$ (a): Clear. ∎
Suppose now that $p$ is in $\Sigma_{e}$ and $\tau$ is an embedding of $K$ in $\mathbb{C}$}. Note that $c_{p}$ is even by the lemma (see VI.2 of [Se2]). Moreover the explicit description of $\pi_{p}(\tau(f))$ in this case is given in Section 3 of []. In particular, we have that
$\pi_{p}(\tau(f))^{V_{p}^{\prime}}|_{\operatorname{{\mathop{\rm GL}\nolimits}}_{2}({\mathbb{Z}}_{p})}\cong\Theta(\varepsilon)$
for some character
$\varepsilon:{\mathcal{O}}_{F}^{\times}\to({\mathcal{O}}_{F}/p^{c_{p}/2})^{\times}\to{\mathbb{C}}^{\times},$
where $V_{p}^{\prime}$ denotes the set of matrices in $\operatorname{{\mathop{\rm GL}\nolimits}}_{2}({\mathbb{Z}}_{p})$ congruent to II mod $p^{c_{p}/2}$, and $\Theta(\varepsilon)$ is defined in §3.2 of [C-D-T]. Our hypotheses ensure that $\varepsilon/\varepsilon\circ\operatorname{{\mathop{\rm Frob}\nolimits}}_{p}$ has conductor $(p{\mathcal{O}}_{F})^{c_{p}/2}$. Let $g_{p}={\mbox{\scriptsize$\left(\begin{array}[]{cc}{{1}}&{{0}}\\ {{0}}&{{p^{c_{p}/2}}}\end{array}\right)$}}$. We recall the following properties of $\Theta(\varepsilon)$:
Lemma 6.9.
Let $B$ denote the set of matrices $\left(\begin{array}[]{cc}{{a}}&{{b}}\\ {{c}}&{{d}}\end{array}\right)$ in $\operatorname{{\mathop{\rm GL}\nolimits}}_{2}({\mathbb{Z}}_{p})$ with $c\equiv 0\bmod p$. Let $D$ denote the set of matrices in $\operatorname{{\mathop{\rm GL}\nolimits}}_{2}({\mathbb{Z}}_{p})$ which are diagonal $\bmod p^{c_{p}/2}$, and ${\mathbb{C}}_{\psi}={\mathbb{C}}$ with the action of $D$ defined by ${\mbox{\scriptsize$\left(\begin{array}[]{cc}{{a}}&{{b}}\\ {{c}}&{{d}}\end{array}\right)$}}\mapsto\psi(a)$.
  1. (a)
    $\Theta(\varepsilon)|_{B}$ is irreducible.
  2. (b)
    $\Theta(\varepsilon^{-1})\cong\hat{\Theta}(\varepsilon)\cong\Theta(\varepsilon)\otimes\varepsilon^{-1}\circ{\mathop{\rm det}\nolimits}$.
  3. (c)
    ${\mathop{\rm Hom}\nolimits}_{{\mathbb{C}}[D]}({\mathbb{C}}_{\psi},\Theta(\varepsilon))$ is one-dimensional.
  4. (d)
    If $\mathcal{V}$} is a model over $\mathcal{O}$} for $\Theta(\varepsilon)$ and $\lambda$ is a prime of $\mathcal{O}$} not dividing p1p-1, then ${\mathcal{V}}/\lambda{\mathcal{V}}$ is absolutely irreducible.
Proof.
(a), (b) and (d) are part of Lemma 3.2.1 of [C-D-T]. (For (d), note that $p$ is already assumed to be invertible in $\mathcal{O}$}.) To prove (c), we use that $\pi_{p}(\tau(f))^{V_{p}}$ is one-dimensional with $\left(\begin{array}[]{cc}{{a}}&{{0}}\\ {{0}}&{{1}}\end{array}\right)$ acting via $\psi(a)$ for $a\in{\mathbb{Z}}_{p}^{\times}$. Conjugating by $g_{p}$ gives the result. ∎
Suppose now that $\Sigma_{0}$ is a subset of $\Sigma_{e}$. For any prime $p$, we let ${\mathcal{V}}_{p}^{1}={\mathcal{O}}$ with ${\mbox{\scriptsize$\left(\begin{array}[]{cc}{{a}}&{{b}}\\ {{c}}&{{d}}\end{array}\right)$}}\in V_{p}$ acting via $\psi^{-1}(a)$.
Lemma 6.10.
There exists a field $K_{0}$ such that if $K_{0}\subset K$, then the following holds for all primes $p$ in $\Sigma_{0}$: The representation $\Theta(\varepsilon^{-1})$ has a locally free model ${\mathcal{V}}_{p}^{0}$ over $\mathcal{O}$} such that
  1. (a)
    there is an isomorphism of ${\mathcal{O}}[\operatorname{{\mathop{\rm GL}\nolimits}}_{2}({\mathbb{Z}}_{p})]$-modules
    $\omega_{p}:{\mathcal{V}}_{p}^{0}\cong{\mathop{\rm Hom}\nolimits}_{\mathcal{O}}({\mathcal{V}}_{p}^{0},{\mathcal{O}}(\psi^{-1}\circ{\mathop{\rm det}\nolimits})).$
  2. (b)
    there is a surjection of ${\mathcal{O}}[V_{p}]$-modules
    $\tau_{p}:{\mathop{\rm res}\nolimits}_{g_{p}\operatorname{{\mathop{\rm GL}\nolimits}}_{2}({\mathbb{Z}}_{p})g_{p}^{-1}}^{V_{p}}g_{p}{\mathcal{V}}_{p}^{0}\to{\mathcal{V}}_{p}^{1}.$
Proof.
It suffices to prove the lemma for $\Sigma_{0}=\{p\}$. We first assume $K$ is sufficiently large that $\Theta(\varepsilon^{-1})$ is defined over $K$. Then it has a model $\mathcal{V}$$^{\prime}$ over $\mathcal{O}$} and by part (b) of Lemma 6.9, there are non-zero homomorphisms
$\omega:{\mathcal{V}}\to{\mathop{\rm Hom}\nolimits}_{\mathcal{O}}({\mathcal{V}},{\mathcal{O}}(\psi^{-1}\circ{\mathop{\rm det}\nolimits}))\quad\mbox{and}\quad\tau:{\mathop{\rm res}\nolimits}_{g_{p}\operatorname{{\mathop{\rm GL}\nolimits}}_{2}({\mathbb{Z}}_{p})g_{p}^{-1}}^{V_{p}}g_{p}{\mathcal{V}}\to{\mathcal{V}}_{p}^{1}.$
These are isomorphisms after tensoring with ${\mathcal{O}}_{\lambda}$ for all but finitely many $\lambda$ in $S$. Using Lemma 3.3.1 of [C-D-T], we may enlarge $K$ so that for the remaining $\lambda$, there are $\operatorname{{\mathop{\rm GL}\nolimits}}_{2}({\mathbb{Z}}_{p}]$-stable lattices ${\mathcal{V}}^{\prime}_{\lambda}$ in ${\mathcal{V}}\otimes_{\mathcal{O}}{\mathcal{O}}_{\lambda}$ and isomorphisms
$\omega^{\prime}_{\lambda}:{\mathcal{V}}^{\prime}_{\lambda}\to{\mathop{\rm Hom}\nolimits}_{\mathcal{O}}({\mathcal{V}}^{\prime}_{\lambda},{\mathcal{O}}(\psi^{-1}\circ{\mathop{\rm det}\nolimits})).$
Now replace $\mathcal{V}$} with $\mathcal{V}$$^{\prime}$ so that ${\mathcal{V}}^{\prime}\otimes_{\mathcal{O}}{\mathcal{O}}_{\lambda}={\mathcal{V}}^{\prime}\lambda$ for these $\lambda$ and ${\mathcal{V}}\otimes_{\mathcal{O}}{\mathcal{O}}_{\lambda}$ otherwise. The image of $\omega$ and $\tau$ are now of the form
${\mathfrak{a}}{\mathop{\rm Hom}\nolimits}_{\mathcal{O}}({\mathcal{V}},{\mathcal{O}}(\psi^{-1}\circ{\mathop{\rm det}\nolimits}))\quad\mbox{and\ }{\mathfrak{b}}{\mathcal{V}}_{p}^{1}$
for some non-zero ideals $\mathfrak{a}$} and $\mathfrak{b}$} of $\mathcal{O}$}. To complete the proof, we further enlarge $K$ so that these become principal. ∎
Now let $\Sigma$ be any finite set of primes. Define
$N_{\Sigma_{0}}^{\Sigma}=\prod_{p\not\in\Sigma_{0}\cup\Sigma}p^{c_{p}}\prod_{p\in\Sigma}p^{c_{p}+\delta_{p}},$
and let $U^{\Sigma}_{\Sigma_{0}}=U_{0}(N^{\Sigma}_{\Sigma_{0}})$. Write $U^{\Sigma}_{\Sigma_{0}}=\prod_{p}U_{p}$ and consider the representation $\sigma^{\Sigma}_{\Sigma_{0}}$ defined by
${\mathcal{V}}^{\Sigma}_{\Sigma_{0}}=\otimes_{\mathcal{O}}{\mathcal{V}}_{p,\Sigma_{0}}^{\Sigma},$
where ${\mathcal{V}}^{\Sigma}_{p,\Sigma_{0}}={\mathcal{V}}_{p}^{0}$ if $p\in\Sigma_{0}-\Sigma$ and ${\mathcal{V}}_{p}^{1}$ (restricted to $U_{p}$) otherwise. Here the tensor product is over all primes $p$, but the action is trivial for all but finitely many pp. Note that if $\Sigma_{0}\subset\Sigma$, then $U^{\Sigma}_{\Sigma_{0}}=U^{\Sigma}$ and $\sigma^{\Sigma}_{\Sigma_{0}}=\sigma(N,\psi)$.
Suppose now that $\Sigma\subset\Sigma^{\prime}$. For positive integers $m$ dividing $N_{\Sigma^{\prime}}/N_{\Sigma}$, define
$\epsilon_{\Sigma_{0},m}^{\Sigma,\Sigma^{\prime}}=m^{1-k}[U^{\Sigma^{\prime}}{\mbox{\scriptsize$\left(\begin{array}[]{cc}{{1}}&{{0}}\\ {{0}}&{{m^{\prime}}}\end{array}\right)$}}U^{\Sigma}]_{\tau}:{\mathcal{M}}(\sigma^{\Sigma}_{\Sigma_{0}})_{!}\to{\mathcal{M}}(\sigma^{\Sigma^{\prime}}_{\Sigma_{0}})_{!},$
where $m^{\prime}=\prod_{p}m^{\prime}_{p}$ with $m^{\prime}_{p}=p^{c_{p}/2}$ if $p\in\Sigma_{0}\cap\Sigma^{\prime}-\Sigma$ and $m^{\prime}_{p}=p^{v_{p}(m)}$ otherwise, and $\tau=\otimes\tau_{p}$ where the tensor product is over primes $p\in\Sigma_{0}\cap\Sigma^{\prime}-\Sigma$. We now define
$\gamma_{\Sigma_{0}}^{\Sigma,\Sigma^{\prime}}=\sum_{m}\epsilon_{\Sigma_{0},m}^{\Sigma,\Sigma^{\prime}}\phi_{m}.$
Note that if $\Sigma_{0}\subset\Sigma$, then $\epsilon_{\Sigma_{0},m}^{\Sigma,\Sigma^{\prime}}$ is $\epsilon_{m}$ and $\gamma_{\Sigma_{0}}^{\Sigma,\Sigma^{\prime}}$ is the same $\gamma$ as in Proposition 6.3.
For each prime $p$, we define the Hecke operator $T_{p}$ on $M_{!}(\sigma^{\Sigma}_{\Sigma_{0}})$ (in what category) by the double coset operator
$\left[U_{\Sigma_{0}}^{\Sigma}{\mbox{\scriptsize$\left(\begin{array}[]{cc}{{p}}&{{0}}\\ {{0}}&{{1}}\end{array}\right)$}}_{p}U_{\Sigma_{0}}^{\Sigma}\right]_{\psi^{-1}(p_{p})}$
if $p\not\in\Sigma_{0}$ and we set $T_{p}=0$ if $p\in\Sigma_{0}$. For primes $p$ not dividing $N^{\Sigma}$, we define $S_{p}$ by
$\left[U_{\Sigma_{0}}^{\Sigma}{\mbox{\scriptsize$\left(\begin{array}[]{cc}{{p}}&{{0}}\\ {{0}}&{{p}}\end{array}\right)$}}_{p}U_{\Sigma_{0}}^{\Sigma}\right]_{\psi^{-2}(p_{p})}.$
Note that if $\Sigma\subset\Sigma$, then this coincides with the previous definition.
Lemma 6.11.
For any $p$ in $\Sigma_{0}-\Sigma$, the morphism $\epsilon=\epsilon_{\Sigma_{0},p}^{\Sigma,\Sigma\cup\{p\}}$ is injective and $T_{p}=0$ on its image. Furthermore, its cokernel is torsion-free if p1p-1 is invertible in $\mathcal{O}$}.
Proof.
Let ${\mathcal{M}}={\mathcal{M}}_{Np^{c_{p}/2}}$, ${\mathcal{M}}^{\prime}={\mathcal{M}}_{Np^{c_{p}}}$, $U_{\Sigma_{0}}^{\Sigma}, U=U$$^{\prime}=U_{\Sigma_{0}}^{\Sigma\cup\{p\}}$, ${\mathcal{V}}={\mathcal{V}}_{\Sigma_{0}}^{\Sigma}$, ${\mathcal{V}}^{\prime}={\mathcal{V}}_{\Sigma_{0}}^{\Sigma\cup\{p\}}$.
We first prove the assertions concerning the kernel and cokernel. Note that $\epsilon$ is the map
$g_{p}\otimes\tau_{p}:({\mathcal{M}}_{!}\otimes_{\mathcal{O}}{\mathcal{V}})^{U}\to({\mathcal{M}}^{\prime}_{!}\otimes_{\mathcal{O}}{\mathcal{V}}^{\prime})^{U^{\prime}},$
and the latter object can be identified with
$\operatorname{Ind}_{U^{\prime}}^{U}({\mathcal{M}}^{\prime}_{!}\otimes_{\mathcal{O}}{\mathcal{V}}^{\prime})^{U}$
where the inclusion $U^{\prime}\to U$ is defined by $u^{\prime}\mapsto g_{p}^{-1}u^{\prime}g_{p}.$ It therefore suffices to prove the map
$\begin{array}[]{ccc}{\mathcal{M}}_{!}\otimes_{\mathcal{O}}{\mathcal{V}}&\to&\operatorname{Ind}_{U^{\prime}}^{U}({\mathcal{M}}^{\prime}_{!}\otimes_{\mathcal{O}}{\mathcal{V}}^{\prime})\\ x\otimes y&\mapsto&(u\mapsto g_{p}ux\otimes\tau_{p}uy)\end{array}$
is injective, with torsion-free cokernel if $p-1\in{\mathcal{O}}^{\times}$. But this map can be written as a composite
${\mathcal{M}}_{!}\otimes_{\mathcal{O}}{\mathcal{V}}\to{\mathcal{M}}_{!}\otimes_{\mathcal{O}}\operatorname{Ind}_{U^{\prime}}^{U}{\mathcal{V}}^{\prime}\cong\operatorname{Ind}_{U^{\prime}}^{U}({\mathop{\rm res}\nolimits}_{U}^{U^{\prime}}({\mathcal{M}}_{!}\otimes_{\mathcal{O}}{\mathcal{V}}^{\prime}))\to\operatorname{Ind}_{U^{\prime}}^{U}({\mathcal{M}}^{\prime}_{!}\otimes_{\mathcal{O}}{\mathcal{V}}^{\prime}),$
where the first map is defined by sending $y$ to $(u\mapsto\tau_{p}uy)$ for $y\in{\mathcal{V}}$, and the last map is the induction of $g_{p}\otimes 1$. Since $\mathcal{M}$$_{!}$, $\mathcal{M}$$^{\prime}_{!}$, $\mathcal{V}$} and $\mathcal{V}$$^{\prime}$ are torsion-free, it suffices to prove that
${\mathcal{V}}\to\operatorname{Ind}_{U^{\prime}}^{U}{\mathcal{V}}^{\prime}\quad\mbox{and}\quad{\mathcal{M}}\to{\mathcal{M}}^{\prime}$
are injective, with torsion-free cokernel if $p-1\in{\mathcal{O}}^{\times}$.
For the first map, the injectivity follows from the fact that the map is non-zero and ${\mathcal{V}}\otimes_{\mathcal{O}}K$ is irreducible. The torsion-freeness then amounts to the injectivity after tensoring with ${\mathcal{O}}/\lambda$ for all $\lambda$ in $S$. The map is non-zero since $\tau_{p}\otimes_{\mathcal{O}}{\mathcal{O}}/\lambda$ is surjective and factors through it, and so the injectivity follows from Lemma 6.9 (d). For the second map, the injectivity and torsion-freeness follows for example by considering the composite
$g_{p}^{-1}\circ g_{p}:{\mathcal{M}}\to{\mathcal{M}}^{\prime}\to{\mathcal{M}}_{Np^{3c_{p}/2},!},$
which we have already seen is injective with torsion-free cokernel, by Lemma 3.9 (c).
To complete the proof of the lemma, it suffices to prove $T_{p}=0$ on the image of
$((\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}{\mathcal{M}}_{\mathop{\rm dR}\nolimits}\otimes{\mathbb{C}})\otimes{\mathcal{V}})^{U}\to((\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}{\mathcal{M}}^{\prime}_{\mathop{\rm dR}\nolimits}\otimes{\mathbb{C}})\otimes{\mathcal{V}}^{\prime})^{U^{\prime}},$
or equivalently, on the image of
$g_{p}\otimes\tau_{p}:(\pi(g)\otimes{\mathcal{V}})^{U}\to(\pi(g)\otimes{\mathcal{V}}^{\prime})^{U^{\prime}}$
for each newform $g$ of weight $k$. But by the definition of $\Theta(\varepsilon)$, $(\pi(g)\otimes{\mathcal{V}})^{U}=0$ unless $\pi_{p}(g)$ is supercuspidal of conductor exactly divisible by $p^{c_{p}}$, so the result follows from Lemma 5.8. ∎
It follows from the lemma and the definitions that $\gamma_{\Sigma_{0}}^{\Sigma,\Sigma\cup\Sigma_{0}}$ is injective and commutes with the action of the Hecke operators $T_{p}$ for all primes pp. In particular, for $p\in S$, the operator $T_{p}$ can be written as a polynomial in $T_{q}$ for $q\neq p$, and the same arguments as in Proposition 5.6 show that $\mathbb{T}$} acts on ${\mathcal{M}}(\sigma_{\Sigma_{0}}^{\Sigma})_{!}$ as an object of ${\mathcal{P}M}^{S}_{K}$. Now define ${\mathcal{M}}_{f,\Sigma_{0}}^{\Sigma}$ and $M_{f,\Sigma_{0}}^{\Sigma}$ exactly as before, i.e., as the intersection of the kernels of elements of $I_{f}$. Note that if $\Sigma_{0}\subset\Sigma$, then $M_{f,\Sigma_{0}}^{\Sigma}=M_{f}^{\Sigma}$.
Lemma 6.12.
The operator $\gamma=\gamma_{\Sigma_{0}}^{\emptyset,\Sigma}$ restricts to an isomorphism
$M_{f,\Sigma_{0}}^{\emptyset}\to M_{f,\Sigma_{0}}^{\Sigma}.$
Furthermore if $S$ is as in Proposition 6.6 and $p-1\in{\mathcal{O}}^{\times}$ for all $p\in\Sigma_{0}$, then it induces an isomorphism
${\mathcal{M}}_{f,\Sigma_{0}}^{\emptyset}\to{\mathcal{M}}_{f,\Sigma_{0}}^{\Sigma}.$
Proof.
By Proposition 6.6, Lemma 6.11 and the commutativity of the diagram
$\begin{array}[]{ccc}{\mathcal{M}}_{f,\Sigma_{0}}^{\emptyset}&\to&{\mathcal{M}}_{f,\Sigma_{0}}^{\Sigma}\\ \downarrow&&\downarrow\\ {\mathcal{M}}_{f,\Sigma_{0}}^{\Sigma_{0}}&\to&{\mathcal{M}}_{f,\Sigma_{0}}^{\Sigma_{0}\cup\Sigma},\end{array}$
it suffices to prove that the lemma in the case $\Sigma=\Sigma_{0}$. Furthermore, by lemma 6.11, it suffices to prove that
$\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}M_{f,\Sigma_{0},{\mathop{\rm dR}\nolimits}}^{\emptyset}\neq 0,$
and this follows from the definition of $\sigma_{\Sigma_{0}}^{\Sigma}$ in terms of $\pi_{p}(f)$. ∎
Define $w=w^{\Sigma}_{\Sigma_{0}}={\mbox{\scriptsize$\left(\begin{array}[]{cc}{{0}}&{{-1}}\\ {{N}}&{{0}}\end{array}\right)$}}_{N}\in\operatorname{{\mathop{\rm GL}\nolimits}}_{2}({\mathbb{A}}_{\bf f})$ where $N=N^{\Sigma}_{\Sigma_{0}}$. Let $\sigma=\sigma_{\Sigma_{0}}^{\Sigma}$ and ${\mathcal{V}}={\mathcal{V}}_{\Sigma_{0}}^{\Sigma}=\otimes{\mathcal{V}}_{p,\Sigma_{0}}^{\Sigma}$, and denote by $\sigma^{\prime}$ the representation of $U=U_{\Sigma_{0}}^{\Sigma}$ defined by
${\mathcal{V}}^{\prime}={\mathop{\rm Hom}\nolimits}_{\mathcal{O}}({\mathcal{V}},{\mathcal{O}}(\psi^{-1}\circ{\mathop{\rm det}\nolimits}))\cong\otimes{\mathop{\rm Hom}\nolimits}_{\mathcal{O}}({\mathcal{V}}_{p,\Sigma_{0}}^{\Sigma},{\mathcal{O}}(\psi^{-1}\circ{\mathop{\rm det}\nolimits})).$
Define $\omega:{\mathcal{V}}\to{\mathcal{V}}^{\prime}$ as the tensor product of the maps $\omega_{p}$, where $\omega_{p}$ is defined in Lemma 6.10 if $p\in\Sigma-\Sigma$, and by sending a generator $v_{0}$ to the map $v_{0}\mapsto 1$ otherwise. We then have that
$\omega(\sigma(w^{-1}uw)v)=\sigma^{\prime}(u)\omega(v)$
for all $u\in U$ and $v\in{\mathcal{V}}$, so the operator $UwU]_{\omega}$ is well-defined and induces an isomorphism
${\mathcal{M}}(\sigma)_{!}\to{\mathcal{M}}(\sigma^{\prime})_{!}.$
Composing with the isomorphism
$M(\sigma^{\prime})_{!}\to{\mathop{\rm Hom}\nolimits}_{K}(M(\sigma)_{!},M_{\psi}(1-k)),$
we obtain a perfect pairing
$M(\sigma)_{!}\otimes_{K}M(\sigma)_{!}\to M_{\psi}(1-k).$
Lemma 6.13.
Suppose that $p\in\Sigma_{0}-\Sigma$. Let $\epsilon=\epsilon_{\Sigma_{0},p}^{\Sigma,\Sigma\cup\{p\}}$ and $\epsilon^{t}$ its transpose with respect to the above pairings. Then $\epsilon^{t}\epsilon$ is a scalar divisible by $p+1$.
Proof.
[…]
$M(\sigma^{\prime})_{!}\otimes_{K}M(\sigma^{\prime})_{!}\to M_{\psi}(1-k)$
exactly as we did for $M(\sigma)_{!}$, using the same $w$ and $\omega$. Letting
$\epsilon_{1}=\left[U^{\prime}1U\right]_{1};\quad\epsilon_{2}=\left[U^{\prime\prime}g_{p}U^{\prime}\right]_{\tau_{p}},$
we find that $\epsilon=\epsilon_{2}\circ\epsilon_{1}$, $\epsilon_{1}^{t}\epsilon_{1}=p+1$ and $\epsilon_{2}^{t}\epsilon_{2}=[U^{\prime}1U^{\prime}]_{\delta}$ for some $U^{\prime}$-linear $\delta:{\mathcal{V}}\to{\mathcal{V}}$. By Lemma 6.9 (a), $\sigma^{\prime}$ is absolutely irreducible, so $\delta$ is a scalar in $\mathcal{O}$} and $\epsilon^{t}\epsilon=(p+1)\delta$. ∎
The same proof as that of Proposition 6.5 now gives:
Proposition 6.14.
Suppose that $\Sigma_{0}\subset\Sigma_{e}$ and $p+1\in{\mathcal{O}}^{\times}$ for all $p\in\Sigma_{e}-\Sigma_{0}$. Suppose that $\Sigma$ is any finite set of primes invertible in $\mathcal{O}$}, and let $\gamma=\gamma_{\Sigma_{0}}^{\emptyset,\Sigma}$. Then
$\gamma^{t}\circ\gamma=\beta_{f}^{\Sigma}\prod_{p\in\Sigma}L_{p}(A_{f},1)^{-1}$
on $M_{f,\Sigma_{0}}^{\emptyset}$ for some non-zero $\beta_{f}^{\Sigma}$ in $\mathcal{O}$}.
Note also that together with the injectivity of $\gamma_{\Sigma_{0}}^{\Sigma,\Sigma\cup\Sigma_{0}}$, Lemma 6.13 implies that the pairing on $M_{!}(\sigma_{\Sigma_{0}}^{\Sigma})$ is alternating and induces an isomorphism
$\wedge^{2}_{K}M_{f,\Sigma_{0}}^{\Sigma}\to M_{\psi}(1-k).$
We let $\eta_{f,\Sigma_{0}}^{\Sigma}$ be the fractional ideal in $K$ such that $\wedge^{2}_{K}{\mathcal{M}}_{f,\Sigma_{0}}^{\Sigma}$ maps isomorphically to $\eta_{f,\Sigma_{0}}^{\Sigma}{\mathcal{M}}_{\psi}(1-k)$. Note that if $\Sigma_{0}\subset\Sigma$, then $\eta_{f,\Sigma_{0}}^{\Sigma}=\eta_{f}^{\Sigma}$. The same proof as that of Corollary 6.7 now gives:
Corollary 6.15.
Suppose that $f$ and $S$ are as in Proposition 6.6 and $\Sigma_{0}\subset\Sigma_{e}$ is such that $p-1\in{\mathcal{O}}^{\times}$ for all $p\in\Sigma_{0}$, and $p+1\in{\mathcal{O}}^{\times}$ for all $p\in\Sigma_{e}-\Sigma_{0}$. Then
$\eta_{f,\Sigma_{0}}^{\Sigma}\subset\eta_{f}^{\emptyset}\prod_{p\in\Sigma}L_{p}(A_{f},1)^{-1}.$
7. The Taylor-Wiles construction
Our method for computing the Selmer group of $A_{f}(1)_{\lambda}$ is based on that of Wiles [Wi] and his work with Taylor [T-W]. We begin by recalling some definitions and results from Galois cohomology. Then we give an axiomatic formulation of the method of [Wi] and [T-W], made possible by the simplifications due to Faltings ([T-W], appendix), Lenstra [Le] and one of the authors [Di5]. Finally we verify these axioms are satisfied in the context of modular forms of higher weight.
7.1. Galois cohomology
In this section, we fix a prime $\ell$ and suppose that $K$ is a finite extension of $\mathbb{Q}$$_{\ell}$ with ring of integers $\mathcal{O}$}, uniformizer $\lambda$ and residue field $\kappa$. We let $\Sigma_{0}$ denote the set of primes different from $\ell$. We let $\chi_{\ell}:G_{{\mathbb{Q}}}\to{\mathbb{Z}}_{\ell}^{\times}$ denote the $\ell$-adic cyclotomic character.
Suppose that for $i=1,2$,
$\rho_{i}:G_{\mathbb{Q}}\to\operatorname{\operatorname{{\mathop{\rm Aut}\nolimits}}}_{K}V_{i}$
is a continuous geometric representation whose restriction to $G_{\ell}$ is crystalline and short, where short means that $\operatorname{{\mathop{\rm Fil}\nolimits}}^{\ell}D_{\mathop{\rm pst}\nolimits}(V_{i})=0$ and $\operatorname{{\mathop{\rm Fil}\nolimits}}^{0}D_{\mathop{\rm pst}\nolimits}(V_{i})=D_{\mathop{\rm pst}\nolimits}(V_{i})$ and that $V_{i}$ has no nonzero quotient $V^{\prime}$ so that $V^{\prime}(\ell-1)$ is unramified. We let $V={\mathop{\rm Hom}\nolimits}_{K}(V_{1},V_{2})$. Suppose that $L_{i}$ is a $G_{\mathbb{Q}}$-stable $\mathcal{O}$}-lattice in $V_{i}$, and let
$A=(K/{\mathcal{O}})\otimes_{\mathcal{O}}{\mathop{\rm Hom}\nolimits}_{\mathcal{O}}(L_{1},L_{2}).$
Below we shall define an $\mathcal{O}$}-submodule $H^{1}_{w}(G_{p},A)\subset H^{1}(G_{p},A)$ for each prime pp. For each finite set of primes $\Sigma\subset\Sigma_{0}$, we define
$H^{1}_{\Sigma}(G_{\mathbb{Q}},A)\subset H^{1}(G_{\mathbb{Q}},A)$
as the $\mathcal{O}$}-submodule of elements with restrictions in $H^{1}_{w}(G_{p},A)$ for every $p\in\Sigma$.
First we define $H^{1}_{w}(G_{p},A_{n})$ for each prime $p$ and integer $n\geq 0$, where $A_{n}=A[\lambda^{n+1}]$. If $p\neq\ell$, then we simply set
$H^{1}_{w}(G_{p},A_{n})=H^{1}(G_{{\mathbb{F}}_{p}},A_{n}^{I_{p}})=\ker(H^{1}(G_{p},A_{n})\to H^{1}(I_{p},A_{n})).$
To treat the case $p=\ell$, we let $\mathcal{F}\mathcal{R}$ denote the full subcategory of finite ${\mathbb{Z}}_{\ell}G_{\ell}$-modules whose objects are quotients of $G_{\ell}$-stable lattices in short crystalline representations. Thus $\mathcal{F}\mathcal{R}$ is stable under taking direct sums, subobjects and quotients, and the $L_{i}/\lambda^{n+1}L_{i}$ are objects of it. Recall that there is a natural $\mathcal{O}$}-linear isomorphism between $H^{1}(G_{\ell},A_{n})$ and the $\mathcal{O}$}-module of extensions of $({\mathcal{O}}/\lambda^{n})[G_{\ell}]$-modules
$0\to\lambda^{-n-1}L_{2}/L_{2}\to E\to L_{1}/\lambda^{n+1}L_{1}\to 0.$
We define $H^{1}_{w}(G_{\ell},A_{n})$ to be the set of elements of $H^{1}(G_{\ell},A_{n})$ corresponding to extensions in $\mathcal{F}\mathcal{R}$. Using the stability of $\mathcal{F}\mathcal{R}$ under direct sums, subobjects and quotients, one checks that $H^{1}_{w}(G_{\ell},A_{n})$ is an $\mathcal{O}$}-submodule, and furthermore that it is the preimage of $H^{1}_{w}(G_{\ell},A_{n+1})$ under the natural map. Now define $H^{1}_{w}(G_{p},A)=\displaystyle{\lim_{\longrightarrow}}\,H^{1}_{w}(G_{p},A_{n})$ for every prime pp.
Recall that Bloch and Kato define a divisible $\mathcal{O}$}-submodule $H^{1}_{\bf f}(G_{p},A)\subset H^{1}(G_{p},A)$ for each $p$ (see [B-K] or [F-P]). Their Selmer group
$H^{1}_{\bf f}(G_{\mathbb{Q}},A)\subset H^{1}(G_{\mathbb{Q}},A)$
is then the $\mathcal{O}$}-submodule of elements with restrictions in $H^{1}_{\bf f}(G_{p},A)$ for all primes pp. If $p\neq\ell$, it is clear that
$H^{1}_{\bf f}(G_{p},A)\subset H^{1}_{w}(G_{p},A)$
and equality holds if $A^{I_{p}}$ is divisible.
Proposition 7.1.
$H^{1}_{w}(G_{\ell},A)=H^{1}_{\bf f}(G_{\ell},A)$ is divisible of $\mathcal{O}$}-corank
$d=\dim_{K}H^{0}(G_{\ell},V)+\dim_{K}V-\dim_{K}\operatorname{{\mathop{\rm Fil}\nolimits}}^{0}D_{\ell\mbox{\scriptsize-crys}}(V).$
Proof.
The divisibility of $H^{1}_{\bf f}(G_{\ell},A)$ follows from its definition, and the corank is computed in [B-K] or [F-P]. One shows as in [Wi] that
$H^{1}_{\bf f}(G_{\ell},A)\subset H^{1}_{w}(G_{\ell},A),$
so it suffices to prove that
$\dim_{\kappa}H^{1}_{w}(G_{\ell},A)[\lambda]\leq d.$
Since the kernel of $H^{1}(G_{\ell},A_{0})\to H^{1}(G_{\ell},A)$ is contained in $H^{1}_{w}(G_{\ell},A_{0})$, we have
$\begin{array}[]{rl}\dim_{\kappa}H^{1}_{w}(G_{\ell},A)[\lambda]&=\dim_{\kappa}H^{1}_{w}(G_{\ell},A_{0})-\dim_{\kappa}(\kappa\otimes_{\mathcal{O}}A^{G_{\ell}})\\ &=\dim_{\kappa}H^{1}_{w}(G_{\ell},A_{0})-\dim_{\kappa}A_{0}^{G_{\ell}}+\dim_{K}H^{0}(G_{\ell},V).\end{array}$
Therefore we need to prove that
$\dim_{\kappa}H^{1}_{w}(G_{\ell},A_{0})-\dim_{\kappa}H^{0}(G_{\ell},A_{0})\leq\dim_{K}V-\dim_{K}\operatorname{{\mathop{\rm Fil}\nolimits}}^{0}D_{\ell\mbox{\scriptsize-crys}}(V).$
We now sketch the proof that equality holds by the theory of Fontaine and Laffaille [F-L]. Recall from §1.1 that the category ${{\mathcal{M}\mathcal{F}}}_{{\mathop{\rm tor}\nolimits}}^{0}$ is equivalent, via the functor
${\mathcal{D}}\mapsto{\mathbb{V}}({\mathcal{D}}):={\mathop{\rm Hom}\nolimits}(\underline{U}_{S}({\mathcal{D}}),{\mathbb{Q}}_{\ell}/{\mathbb{Z}}_{\ell}),$
to $\mathcal{F}\mathcal{R}$. Let $\kappa$-$\mathcal{M}\mathcal{F}$}} denote the category of $\kappa$-modules in $\mathcal{M}\mathcal{F}$}} so that the functor $\mathbb{V}$} defines an equivalence between $\text{$\kappa$-${{\mathcal{M}\mathcal{F}}}$}^{0}$ and the category of $\kappa$-modules in $\mathcal{F}\mathcal{R}$. Choosing objects ${\mathcal{D}}_{1}$ and ${\mathcal{D}}_{2}$ in $\kappa$-$\mathcal{M}\mathcal{F}$}} so that ${\mathbb{V}}({\mathcal{D}}_{1})\cong L_{1}/\lambda L_{1}$ and ${\mathbb{V}}({\mathcal{D}}_{2})\cong\lambda^{-1}L_{2}/\lambda L_{2}$, we have
$\dim_{\kappa}\operatorname{{\mathop{\rm Fil}\nolimits}}^{j}{\mathcal{D}}_{i}=\dim_{K}\operatorname{{\mathop{\rm Fil}\nolimits}}^{j}D_{\mbox{\scriptsize crys}}(V_{i})$
for $i=1,2$ and $j=0,\ldots,\ell-1$. The dimension formula then follows from the exact sequence
$\begin{array}[]{rcl}0\to{\mathop{\rm Hom}\nolimits}_{\text{$\kappa$-${{\mathcal{M}\mathcal{F}}}$}}({\mathcal{D}}_{1},{\mathcal{D}}_{2})\to&{\mathop{\rm Hom}\nolimits}_{\kappa,\operatorname{{\mathop{\rm Fil}\nolimits}}}({\mathcal{D}}_{1},{\mathcal{D}}_{2})&\\ \to&{\mathop{\rm Hom}\nolimits}_{\kappa}({\mathcal{D}}_{1},{\mathcal{D}}_{2})&\to\operatorname{{\mathop{\rm Ext}\nolimits}}^{1}_{\text{$\kappa$-${{\mathcal{M}\mathcal{F}}}$}}({\mathcal{D}}_{1},{\mathcal{D}}_{2})\to 0\end{array}$
whose construction we leave as an exercise (see also diagram (68) for a similar computation). ∎
Now we specialize to the case where $\rho_{1}=\rho_{2}$, each having dimension 22 over $K$ with $\dim_{K}\operatorname{{\mathop{\rm Fil}\nolimits}}^{1}D_{\mbox{\scriptsize crys}}(V_{i})$ =1=1 (so also $\ell>2$). We assume also that $L_{1}=L_{2}$, and let
$A^{0}=(K/{\mathcal{O}})\otimes_{\mathcal{O}}{\mathop{\rm ad}\nolimits}^{0}_{\mathcal{O}}L_{1}$
where ${\mathop{\rm ad}\nolimits}^{0}_{\mathcal{O}}L_{1}$ denotes the kernel of the trace map $\operatorname{{\mathop{\rm End}\nolimits}}_{\mathcal{O}}L\to{\mathcal{O}}$. We let $A^{0}_{n}=A^{0}[\lambda^{n+1}]$, identify $A^{0}$ with a submodule of AA, and define $H^{1}_{w}(G_{p},A^{0}_{n})$, $H^{1}_{w}(G_{p},A^{0})$ and $H^{1}_{w}(G_{\mathbb{Q}},A^{0})$ as the preimage of the corresponding group for AA. From the proposition above, we get
Corollary 7.2.
Under the above hypotheses, $H^{1}_{f}(G_{\ell},A^{0})=H^{1}_{w}(G_{\ell},A^{0})$ and
$\dim_{\kappa}H^{1}_{w}(G_{\ell},A_{0}^{0})=\dim_{\kappa}H^{0}(G_{\ell},A_{0}^{0})+1.$
Moreover if $H^{0}(I_{p},A^{0})$ is divisible for all $p\neq\ell$, then
$H^{1}_{\emptyset}(G_{\mathbb{Q}},A^{0})=H^{1}_{\bf f}(G_{\mathbb{Q}},A^{0}).$
7.2. An axiomatic formulation
Fix a continuous, odd, irreducible representation
$\rho_{0}:G_{\mathbb{Q}}\to\operatorname{\operatorname{{\mathop{\rm Aut}\nolimits}}}_{\kappa}(V_{0})$
where $V_{0}$ is two-dimensional over $\kappa$. We impose the following two technical hypotheses on the representations $\rho_{0}$ we consider:
  • The restriction of $\rho_{0}$ to $G_{F}$ is absolutely irreducible where $F$ is the quadratic subfield of ${\mathbb{Q}}(\mu_{\ell})$.
  • The Serre weight $k$ of $\rho_{0}$ is at most $\ell-1$.
We let $\psi$ denote the $\mathcal{O}$}-valued character of $G_{\mathbb{Q}}$ having finite order not divisible by $\ell$ so that $\psi\chi_{\ell}^{1-k}$ has reduction ${\mathop{\rm det}\nolimits}\rho_{0}$. Let $K_{\psi}$ denote $K$ with an action of $G_{\mathbb{Q}}$ via $\psi$, and define ${\mathcal{O}}_{\psi}$ similarly.
We consider continuous pseudo-geometric $\ell$-adic representations
$\rho:G_{\mathbb{Q}}\to\operatorname{\operatorname{{\mathop{\rm Aut}\nolimits}}}_{K_{\rho}}(V_{\rho})$
where $V_{\rho}$ is two-dimensional over a finite extension $K_{\rho}$ of $K$ contained in $\bar{K}$, $\rho$ has determinant $\psi\chi_{\ell}^{1-k}$ and reduction isomorphic to $\rho_{0}$ over $\bar{\kappa}$. We let ${\mathcal{O}}_{\rho}$ denote the ring of integers of $K_{\rho}$. We say such a representation $\rho$ is an allowable lift of $\rho_{0}$ if its restriction to $G_{\ell}$ is short and crystalline.
Suppose we are given a set $\mathcal{N}$} of allowable lifts. We assume the $\bar{K}$-isomorphism classes of the elements of $\mathcal{N}$} are distinct. For each $\rho$, we let $\Sigma_{\rho}$ denote the set of primes at which $\rho$ is not minimally ramified (see [Di4]). For each set of primes $\Sigma$ contained in $\Sigma_{0}$, we let ${\mathcal{N}}^{\Sigma}$ denote the set of $\rho$ in $\mathcal{N}$} such that $\Sigma_{\rho}\subset\Sigma$ and write $V^{\Sigma}$ for the direct sum over ${\mathcal{N}}^{\Sigma}$ of the $\bar{V}_{\rho}=\bar{K}\otimes_{K_{\rho}}V_{\rho}$. We assume that ${\mathcal{N}}^{\Sigma}$ is finite if $\Sigma$ is finite.
A trellis for $\mathcal{N}$} is an ${\mathcal{O}}G_{\mathbb{Q}}$-submodule $L$ of $V^{\Sigma_{0}}$ such that for each finite set $\Sigma\subset\Sigma_{0}$, the $\mathcal{O}$}-module $L^{\Sigma}=L\cap V^{\Sigma}$ is finitely generated and the map $\bar{K}\otimes_{{\mathcal{O}}}L^{\Sigma}\to V^{\Sigma}$ is an isomorphism. One checks that if $\rho$ in $\mathcal{N}$} is such that $K_{\rho}=K$, then $L_{\rho}=L\cap V_{\rho}$ is a lattice in $V_{\rho}$. (Write $V_{\rho}$ as an intersection of kernels of endomorphisms of $V^{\Sigma}$ defined by elements of ${\mathcal{O}}G_{\mathbb{Q}}$.)
A system of perfect pairings $\varphi$ for $L$ is an ${\mathcal{O}}[G_{\mathbb{Q}}]$-isomorphism
$\varphi^{\Sigma}:L^{\Sigma}\to{\mathop{\rm Hom}\nolimits}_{{\mathcal{O}}}(L^{\Sigma},{\mathcal{O}}_{\psi}(1-k))$
for each finite $\Sigma\subset\Sigma_{0}$. Note that for each $\rho$ in ${\mathcal{N}}^{\Sigma}$, $\varphi^{\Sigma}$ induces a $K_{\rho}G_{\mathbb{Q}}$-isomorphism
$\wedge^{2}_{K_{\rho}}V_{\rho}\to K_{\rho}\otimes_{K}K_{\psi}(1-k)$
which we denote by $\varphi_{\rho}^{\Sigma}$.
We say that a prime $q$ is horizontal if the following hold
  • $q\equiv 1\bmod\ell$;
  • $\rho_{0}$ is unramified at $q$;
  • $\rho_{0}(\operatorname{{\mathop{\rm Fr}\nolimits}}_{q})$ has distinct eigenvalues.
If $Q$ is a finite set of horizontal primes, we let $\Delta_{Q}$ denote the maximal quotient of $\prod\nolimits_{q\in Q}({\mathbb{Z}}/q{\mathbb{Z}})^{\times}$ of $\ell$-power order.
Theorem 7.3.
Let $\mathcal{N}$} be a set of allowable lifts of $\rho_{0}$ (with distinct $\bar{K}$-isomorphism classes and finite ${\mathcal{N}}^{\Sigma}$ for each finite $\Sigma\subset\Sigma_{0}$), $L$ a trellis for $\mathcal{N}$} and $\varphi$ a system of perfect pairings for $L$. Suppose that
  • $\mathcal{N}$$^{\emptyset}\neq\emptyset$;
  • if $\Sigma\subset\Sigma^{0}$ is a finite set of primes and $\rho\in{\mathcal{N}}^{\emptyset}$, then
    $\varphi_{\rho}^{\Sigma}=\varphi_{\rho}^{\emptyset}\beta_{\rho}^{\Sigma}\prod_{p\in\Sigma}L_{p}({\mathop{\rm ad}\nolimits}^{0}_{K}V_{\rho},1)^{-1}$
    for some $\beta_{\rho}^{\Sigma}$ in ${\mathcal{O}}_{\rho}$;
  • if $Q$ is a finite set of horizontal primes, then
    ${\mathcal{N}}^{Q}\leq\#{\mathcal{N}}^{\emptyset}\cdot\#\Delta_{Q}$
    and $\beta_{\rho}^{Q}\in{\mathcal{O}}$ is independent of $\rho\in{\mathcal{N}}^{\emptyset}$.
Then every allowable lift of $\rho_{0}$ is isomorphic over $\bar{K}$ to some $\rho$ in $\mathcal{N}$}. Furthermore if $K_{\rho}=K$, then the lengths of
$H^{1}_{\Sigma}(G_{\mathbb{Q}},A_{\rho}^{0})\quad\mbox{and}\quad{\mathcal{O}}_{\psi}(1-k)/\mu_{\rho}^{\Sigma}(\wedge_{{\mathcal{O}}}^{2}L_{\rho})$
coincide for any finite subset $\Sigma$ of $\Sigma_{0}$ containing $\Sigma_{\rho}$.
Proof.
One checks that to prove the theorem, we can enlarge $K$ and so assume that $\kappa$ contains the eigenvalues of the elements of the image of $\rho_{0}$. Note also that the hypotheses ensure the existence of an element $\rho_{\mathop{\rm min}\nolimits}$ of $\mathcal{N}$$^{\emptyset}$. We may assume also that $K=K_{\rho_{\mathop{\rm min}\nolimits}}$ and we write simply $V_{\mathop{\rm min}\nolimits}$ and $L_{\mathop{\rm min}\nolimits}$ for $V_{\rho_{\mathop{\rm min}\nolimits}}$ and $L_{\rho_{\mathop{\rm min}\nolimits}}$.
We first recall the results we need from the deformation theory of Galois representations. See [dS-L], [Ma] and the appendix of [C-D-T] for more details. We let $\mathcal{C}$} denote the category of complete local Noetherian $\mathcal{O}$}-algebras.
Recall that if AA is an object of $\mathcal{C}$} with maximal ideal $\mathfrak{m}$}, then an AA-deformation of $V_{0}$ is an isomorphism class of free AA-modules $M$ endowed with continuous $AG_{\mathbb{Q}}$-action such that $M/{\mathfrak{m}}M$ is $(A/{\mathfrak{m}})G_{\mathbb{Q}}$-isomorphic to $(A/{\mathfrak{m}})\otimes_{\kappa}V_{0}$.
Suppose that $\Sigma$ is a finite subset of $\Sigma_{0}$. We say that $M$ is of type $\Sigma$ if the following hold:
  • the $AG_{\mathbb{Q}}$-module $M$ is minimally ramified outside $\Sigma$;
  • for every $n>0$, the ${\mathbb{Z}}_{\ell}G_{\ell}$-module $M/{\mathfrak{m}}^{n}M$ is an object of the category $\mathcal{F}\mathcal{R}$;
  • the $AG_{\mathbb{Q}}$-module $\wedge^{2}_{A}M$ is isomorphic to $A\otimes_{{\mathcal{O}}}{\mathcal{O}}_{\psi}(1-k)$.
Consider the functor on $\mathcal{C}$} which associates to AA the set of AA-deformations of $\rho_{0}$ of type $\Sigma$. By the results of Mazur and Ramakrishna, this functor is representable by an object of $\mathcal{C}$}. We denote this object $R^{\Sigma}$ and let $M^{\Sigma}$ denote the universal deformation. We recall also that $R^{\Sigma}$ is topologically generated by the elements $t_{g}^{\Sigma}$ for $g$ in $G_{\mathbb{Q}}$, where $t_{g}^{\Sigma}$ denotes the trace of the endomorphism $g$ of the free $R^{\Sigma}$-module $M^{\Sigma}$. In particular, $R^{\Sigma}$ has residue field $\kappa$.
If $\Sigma_{1}\subset\Sigma_{2}$, then $M^{\Sigma_{1}}$ is an $R^{\Sigma_{1}}$-deformation of $V_{0}$ of type $\Sigma_{2}$ and hence gives rise to a natural surjection $R^{\Sigma_{2}}\to R^{\Sigma_{1}}$.
Suppose now that $\rho$ is in $\mathcal{N}$} and $\Sigma_{\rho}\subset\Sigma$. Then ${\mathcal{O}}_{\rho}$ is an object of $\mathcal{C}$} and there is an ${\mathcal{O}}_{\rho}$-deformation $M$ of $\rho_{0}$ of type $\Sigma$ so that $V_{\rho}$ is $K_{\rho}G_{\mathbb{Q}}$-isomorphic to $K_{\rho}\otimes_{{\mathcal{O}}_{\rho}}M$. We thus obtain a continuous $\mathcal{O}$}-algebra homomorphism
$\theta_{\rho}^{\Sigma}:R^{\Sigma}\to K_{\rho}$
so that $K_{\rho}\otimes_{R^{\Sigma}}M^{\Sigma}$ is isomorphic to $V_{\rho}$. The maps $\theta_{\rho}^{\Sigma}$ for varying $\Sigma\supset\Sigma_{\rho}$ are compatible with the natural surjections $R^{\Sigma_{2}}\to R^{\Sigma_{1}}$ defined above. Note also that if $K_{\rho}=K$, then $A={\mathcal{O}}$ and $\theta_{\rho}^{\Sigma}$ defines a surjection $R^{\Sigma}\to{\mathcal{O}}$. In that case we have a natural isomorphism
(45) ${\mathop{\rm Hom}\nolimits}_{{\mathcal{O}}}({\mathfrak{p}}_{\rho}^{\Sigma}/({\mathfrak{p}}_{\rho}^{\Sigma})^{2},K/{\mathcal{O}})\cong H^{1}_{\Sigma}(G_{\mathbb{Q}},A_{\rho}^{0})$
of $\mathcal{O}$}-modules where ${\mathfrak{p}}_{\rho}^{\Sigma}$ is the kernel of $\theta_{\rho}^{\Sigma}$. In particular this is the case for $\rho=\rho_{\mathop{\rm min}\nolimits}$ and any finite $\Sigma\subset\Sigma_{0}$.
We regard $V^{\Sigma}$ as a module for $R^{\Sigma}$ via
(46) $R^{\Sigma}\to\prod_{\rho\in{\mathcal{N}}^{\Sigma}}K_{\rho}$
defined by the maps $\theta_{\rho}^{\Sigma}$. Note that if $g$ is in $G_{\mathbb{Q}}$, then $t_{g}^{\Sigma}$ acts on $V^{\Sigma}$ via the endomorphism
$\operatorname{{\mathop{\rm tr}\nolimits}}(\rho(g))=g+\psi(g)\chi_{\ell}^{1-k}(g)g^{-1}$
which is given by an element of ${\mathcal{O}}G_{\mathbb{Q}}$. It follows that $L^{\Sigma}$ is stable under the action of $R^{\Sigma}$ and that $\phi^{\Sigma}$ is $R^{\Sigma}$-linear.
We define the finite flat $\mathcal{O}$}-algebra $T^{\Sigma}$ to be the image of $R^{\Sigma}$ in $\operatorname{{\mathop{\rm End}\nolimits}}_{{\mathcal{O}}}L^{\Sigma}$. The maps $\theta_{\rho}^{\Sigma}$ induce an isomorphism of finite $\bar{K}$-algebras
$\bar{K}\otimes_{{\mathcal{O}}}T^{\Sigma}\to\prod_{\rho\in{\mathcal{N}}^{\Sigma}}\bar{K}$
such that $t_{g}^{\Sigma}\mapsto(\operatorname{{\mathop{\rm tr}\nolimits}}\rho(g))_{\rho\in{\mathcal{N}}^{\Sigma}}$ for $g$ in $G_{\mathbb{Q}}$. In particular $T^{\Sigma}$ is reduced and
$\operatorname{{\mathop{\rm rank}\nolimits}}_{{\mathcal{O}}}L^{\Sigma}=2\cdot\#{\mathcal{N}}^{\Sigma}=2\cdot\operatorname{{\mathop{\rm rank}\nolimits}}_{{\mathcal{O}}}T^{\Sigma}.$
Suppose that $\rho$ is an element of ${\mathcal{N}}^{\Sigma}$ such that $K_{\rho}=K$. Write ${\mathcal{P}}_{\rho}^{\Sigma}$ for the image of ${\mathfrak{p}}_{\rho}^{\Sigma}$ in $T^{\Sigma}$ and $I_{\rho}^{\Sigma}$ for the annihilator of ${\mathcal{P}}_{\rho}^{\Sigma}$ in $T^{\Sigma}$. Note that ${\mathcal{P}}_{\rho}^{\Sigma}$ (respectively, $I_{\rho}^{\Sigma}$) is the set of elements in $T^{\Sigma}$ whose image in $\prod\bar{K}$ has trivial component at $\rho$ (respectively, at each $\rho^{\prime}\neq\rho$).
Now consider the $\mathcal{O}$}-module
$\Omega_{\rho}^{\Sigma}=L^{\Sigma}/(L^{\Sigma}[{\mathcal{P}}_{\rho}^{\Sigma}]+L^{\Sigma}[I_{\rho}^{\Sigma}]).$
We define $\eta_{\rho}^{\Sigma}$ as the annihilator of the finite torsion $\mathcal{O}$}-module
${\mathcal{O}}_{\psi}(1-k)/\varphi_{\rho}^{\Sigma}(\wedge_{{\mathcal{O}}}^{2}L_{\rho}).$
Lemma 7.4.
The $\mathcal{O}$}-module $\Omega_{\rho}^{\Sigma}$ is isomorphic to $({\mathcal{O}}/\eta_{\rho}^{\Sigma})^{2}$.
Proof.
Note that the kernel of the projection $L^{\Sigma}\to V_{\rho}$ coincides with that of the composite
$L^{\Sigma}\to{\mathop{\rm Hom}\nolimits}_{{\mathcal{O}}}(L^{\Sigma},{\mathcal{O}}_{\psi}(1-k))\to{\mathop{\rm Hom}\nolimits}_{{\mathcal{O}}}(L_{\rho},{\mathcal{O}}_{\psi}(1-k))$
where the first map is $\varphi^{\Sigma}$ and the second is the natural surjection. Denoting this kernel by $L_{\rho}^{\perp}$, we have $L_{\rho}\subset L^{\Sigma}[{\mathcal{P}}_{\rho}^{\Sigma}]$ and $L_{\rho}^{\perp}\subset L^{\Sigma}[I_{\rho}^{\Sigma}]$. Furthermore both inclusions are equalities since $L^{\Sigma}[{\mathcal{P}}_{\rho}^{\Sigma}]$ and $L^{\Sigma}[I_{\rho}^{\Sigma}]$ have trivial intersection. Therefore the $\mathcal{O}$}-module $\Omega_{\rho}^{\Sigma}$ is isomorphic to the cokernel of the map
$L_{\rho}\to{\mathop{\rm Hom}\nolimits}_{{\mathcal{O}}}(L_{\rho},{\mathcal{O}}_{\psi}(1-k))$
induced by $\varphi^{\Sigma}$, and this is isomorphic to
${\mathop{\rm Hom}\nolimits}_{{\mathcal{O}}}(L_{\rho},{\mathcal{O}}_{\psi}(1-k))\otimes_{{\mathcal{O}}}{\mathcal{O}}/\eta_{\rho}^{\Sigma}$
(in fact, canonically so as an ${\mathcal{O}}G_{\mathbb{Q}}$-module). ∎
Suppose that $q$ is a horizontal element of $\Sigma$ and choose an ordering $\alpha_{1},\alpha_{2}$ of the eigenvalues of $\rho_{0}(\operatorname{{\mathop{\rm Fr}\nolimits}}_{q})$. We then have a unique decomposition
$M^{\Sigma}=M^{\Sigma}_{1}\oplus M^{\Sigma}_{2}$
as an $R^{\Sigma}G_{q}$-module such that for $i=1,2$, $\kappa\otimes_{R^{\Sigma}}M^{\sigma}_{1}$ is unramified with $\operatorname{{\mathop{\rm Fr}\nolimits}}_{q}$ acting via $\alpha_{i}$. We let $\xi_{q}^{\Sigma}$ denote the character $G_{q}\to(R^{\Sigma})^{\times}$ defined by the action of $G_{q}$ on $M^{\Sigma}_{1}$. The restriction of $\xi_{q}^{\Sigma}$ to the inertia group $I_{q}$ factors through
$I_{q}\to{\mathbb{Z}}_{q}^{\times}\to\Delta_{\{q\}}$
where the first map is gotten from local class field theory and the second is the natural projection. We thus obtain a homomorphism $\Delta_{\{q\}}\to(R^{\Sigma})^{\times}$. If $Q$ is a finite set of horizontal primes, then we regard $R_{Q}$ as an ${\mathcal{O}}\Delta_{Q}$-algebra via these homomorphisms and so regard $L^{Q}$ as an ${\mathcal{O}}\Delta_{Q}$-module.
Now let ${\mathfrak{p}}^{Q}$ denote the augmentation ideal of ${\mathcal{O}}[\Delta_{Q}]$, i.e., the kernel of the map ${\mathcal{O}}[\Delta_{Q}]\to{\mathcal{O}}$ defined by $\delta\mapsto 1$ for $\delta$ in $\Delta_{Q}$. Let $S^{Q}$ denote the image of ${\mathcal{O}}[\Delta_{Q}]$ in the ring of endomorphisms of $L^{Q}$ and write ${\mathcal{P}}^{Q}$ for the image of ${\mathfrak{p}}^{Q}$ and $I^{Q}$ for the annihilator of ${\mathcal{P}}^{Q}$ in $S^{Q}$. Now consider the $\mathcal{O}$}-module
$\Omega^{Q}=L^{Q}/(L^{Q}[{\mathcal{P}}^{Q}]+L^{Q}[I^{Q}]).$
Lemma 7.5.
The $\mathcal{O}$}-rank of $L^{Q}[{\mathcal{P}}^{Q}]$ is $\mathcal{N}$$^{\emptyset}$ and the $\mathcal{O}$}-length of $\Omega^{Q}$ is at least $\Delta_{Q}\cdot\#{\mathcal{N}}^{\emptyset}$.
Proof.
First note that if $\rho$ is in ${\mathcal{N}}^{Q}$, then $\rho$ is in $\mathcal{N}$$^{\emptyset}$ if and only if $\Delta_{Q}$ acts trivially on $V_{\rho}$. It follows that $L^{\emptyset}\subset L^{Q}[{\mathcal{P}}^{Q}]$ and $(L^{\emptyset})^{\perp}\subset L^{Q}[I^{Q}]$ where $(L^{\emptyset})^{\perp}$ is the kernel of the projection $L^{Q}\to V^{\emptyset}$. This kernel is the same as that of the composite
$L^{Q}\to{\mathop{\rm Hom}\nolimits}_{{\mathcal{O}}}(L^{Q},{\mathcal{O}}_{\psi}(1-k))\to{\mathop{\rm Hom}\nolimits}_{{\mathcal{O}}}(L^{\emptyset},{\mathcal{O}}_{\psi}(1-k))$
where the first map is $\varphi^{Q}$ and the second is the natural surjection. Since $L^{Q}[{\mathcal{P}}^{Q}]$ and $L^{Q}[I^{Q}]$ have trivial intersection and
$L^{\emptyset}\oplus(L^{\emptyset})^{\perp}$
has finite index in $L^{Q}$, it follows that $L^{\emptyset}=L^{Q}[{\mathcal{P}}^{Q}]$ and $(L^{\emptyset})^{\perp}=L^{Q}[I^{Q}]$. Therefore the $\mathcal{O}$}-module $L^{Q}[{\mathcal{P}}^{Q}]$ has rank $\mathcal{N}$$^{\emptyset}$. Moreover $\Omega^{Q}$ is isomorphic to the cokernel of the endomorphism $\nu_{Q}$ of $L^{\emptyset}$ obtained by composing the inclusion $L^{\emptyset}\to L^{Q}$ with its adjoint with respect to $\varphi^{Q}$ and $\varphi^{\emptyset}$. Our hypotheses on the pairings ensure that that $\nu_{Q}=\#\Delta_{Q}\nu_{Q}^{\prime}$ for some $\nu_{Q}^{\prime}$ in ${\mathcal{O}}G_{\mathbb{Q}}$. ∎
The following is proved exactly as in [T-W] except that we use Corollary 7.2.
Lemma 7.6.
There exists an integer $r\geq 0$ and sets of horizontal primes $Q_{n}$ for each $n\geq 1$ such that the following hold:
  • $Q_{n}=r$;
  • $q\equiv 1\bmod\ell^{n}$ for each $q\in Q_{n}$;
  • $R^{Q_{n}}$ is generated by $r$ elements as an $\mathcal{O}$}-algebra.
We are now ready to prove that $R^{\emptyset}$ is a complete intersection over which $L^{\emptyset}$ is free of rank two. We let $r$ and $Q=Q_{n}$ for $n\geq 1$ be as in lemma 7.6. Lemma 7.5 allows us to apply the implication (a) $\Rightarrow$ (c) of Theorem 2.4 of [Di5] to conclude that $L^{Q}$ is a free ${\mathcal{O}}[\Delta_{Q}]$-module of rank $\mathcal{N}$$^{\emptyset}$. Since the adjoint of the natural inclusion $L^{\emptyset}\to L^{Q}$ is surjective, it induces an isomorphism $L^{Q}/{\mathfrak{p}}^{Q}L^{Q}\to L^{\emptyset}$ of $R^{Q}$-modules.
Setting $A=\kappa[[S_{1},\ldots,S_{r}]]$, $B=\kappa[[X_{1},\ldots,X_{r}]]$, $R=\kappa\otimes_{{\mathcal{O}}}R^{\emptyset}$ and $H=\kappa\otimes_{{\mathcal{O}}}L^{\emptyset}$, we shall define $B$-modules $H_{n}$ and maps $\phi_{n}:A\to B$, $\psi_{n}:B\to R$ and $\pi_{n}:H_{n}\to H$ satisfying the hypotheses of Theorem 1.3 of [Di5]. We first choose surjective $\kappa$-algebra homomorphisms $A\to\kappa[\Delta_{Q_{n}}]$ and $B\to R_{n}$ where $R_{n}=\kappa\otimes_{{\mathcal{O}}}R^{Q_{n}}$. Define $\psi_{n}$ as the composite $B\to R_{n}\to R$ and define $\phi_{n}:A\to B$ so the diagram
$\begin{array}[]{ccc}A&\to&B\\ \downarrow&&\downarrow\\ \kappa[\Delta_{Q_{n}}]&\to&R_{n}\end{array}$
commutes. We consider $L_{n}=\kappa\otimes_{{\mathcal{O}}}L^{Q_{n}}$ as a $B$-module via $B\to R_{n}$, and define $H_{n}$ as $L_{n}/{\mathfrak{m}}_{A}^{n}L_{n}$ and $\pi_{n}$ as the map induced by $L^{Q_{n}}\to L^{\emptyset}$. We can then apply Theorem 1.3 of [Di5] to conclude that $R$ is a complete intersection over which $H$ is a free module. It then follows that $R^{\emptyset}$ is a complete intersection over which $L^{\emptyset}$ is a free module of rank 2.
We now apply the implication (c) $\Rightarrow$ (b) of Theorem 2.4 of [Di5] to the $R^{\emptyset}$-module $L^{\emptyset}$ and prime ideal ${\mathfrak{p}}_{\rho_{\mathop{\rm min}\nolimits}}^{\emptyset}$. We thus obtain the formula
$\begin{array}[]{rl}2\cdot\operatorname{{\mathop{\rm length}\nolimits}}_{{\mathcal{O}}}H^{1}_{\emptyset}(G_{\mathbb{Q}},A_{\mathop{\rm min}\nolimits}^{0})&=2\cdot\operatorname{{\mathop{\rm length}\nolimits}}_{{\mathcal{O}}}{\mathfrak{p}}_{\rho_{\mathop{\rm min}\nolimits}}^{\emptyset}/({\mathfrak{p}}_{\rho_{\mathop{\rm min}\nolimits}}^{\emptyset})^{2}\\ =\operatorname{{\mathop{\rm length}\nolimits}}_{{\mathcal{O}}}\Omega_{\rho_{\mathop{\rm min}\nolimits}}^{\emptyset}&=2\cdot v(\eta_{\rho_{\mathop{\rm min}\nolimits}}^{\emptyset})\end{array}$
where the first equality follows from (45) and the last from Lemma 7.4.
Suppose now that $\Sigma$ is a finite subset of $\Sigma_{0}$. Applying (45) and Lemma 7.4 again, together with the inequality
$\operatorname{{\mathop{\rm length}\nolimits}}_{{\mathcal{O}}}H^{1}_{\Sigma}(G_{\mathbb{Q}},A_{\mathop{\rm min}\nolimits}^{0})\leq\operatorname{{\mathop{\rm length}\nolimits}}_{{\mathcal{O}}}H^{1}_{\emptyset}(G_{\mathbb{Q}},A_{\mathop{\rm min}\nolimits}^{0})-\sum_{p\in\Sigma}v(L_{p}({\mathop{\rm ad}\nolimits}^{0}_{K}V_{\rho},1))$
obtained from a Galois cohomology argument, we find that
$\begin{array}[]{rl}2\cdot\operatorname{{\mathop{\rm length}\nolimits}}_{{\mathcal{O}}}{\mathfrak{p}}_{\rho_{\mathop{\rm min}\nolimits}}^{\Sigma}/({\mathfrak{p}}_{\rho_{\mathop{\rm min}\nolimits}}\Sigma)^{2}&=2\cdot\operatorname{{\mathop{\rm length}\nolimits}}_{{\mathcal{O}}}H^{1}_{\Sigma}(G_{\mathbb{Q}},A_{\mathop{\rm min}\nolimits}^{0})\\ \leq 2\cdot v(\eta_{\rho_{\mathop{\rm min}\nolimits}}^{\Sigma})&=\operatorname{{\mathop{\rm length}\nolimits}}_{{\mathcal{O}}}\Omega_{\rho_{\mathop{\rm min}\nolimits}}^{\Sigma}.\end{array}$
We can then apply the implication (a) $\Rightarrow$ (c) of Theorem 2.4 of [Di5] to conclude $R^{\Sigma}$ is a complete intersection over which $L^{\Sigma}$ is a free module of rank 2.
The second assertion of the theorem follows from another application of Theorem 2.4 of [Di5], (45) and lemma 7.4. To deduce the first assertion of the theorem, note that the map
$\bar{K}\otimes_{{\mathcal{O}}}R^{\Sigma}\to\prod_{\rho\in{\mathcal{N}}^{\Sigma}}\bar{K}$
induced by (46) is an isomorphism. Every allowable lift of $\rho_{0}$ arises, up to $\bar{K}$-isomorphism, from a $\bar{K}$-linear map $\bar{K}\otimes_{{\mathcal{O}}}R^{\Sigma}$ for some $\Sigma$. It therefore arises from $\theta_{\rho}^{\Sigma}$ for some $\rho$ in $\mathcal{N}$}. ∎
7.3. The construction
Fix a number field $K_{0}\subset{\mathbb{C}}$. Let $\lambda_{0}$ denote the prime over $\ell$ determined by our fixed embeddings $\bar{{\mathbb{Q}}}\to{\mathbb{C}}$ and $\bar{{\mathbb{Q}}}\to\bar{{\mathbb{Q}}}_{\ell}$, and let $S=\{\lambda_{0}\}$. Let ${\mathcal{O}}_{0}$ denote the valuation ring of $\lambda_{0}$ in $_{0}, KK$ the completion at $\lambda_{0}$ of $K_{0}$, $\mathcal{O}$} its ring of integers, $\lambda$ a uniformizer and $\kappa$ its residue field.
We consider a representation
$\rho_{0}:G_{\mathbb{Q}}\to\operatorname{\operatorname{{\mathop{\rm Aut}\nolimits}}}_{\kappa}(V_{0})$
as in the preceding section. We assume also that $\rho_{0}$ has minimal conductor among its twists by characters $G_{\mathbb{Q}}\to\bar{\kappa}^{\times}$. Let $k$ denote the Serre weight and $\psi$ the character associated to $\rho_{0}$, as in the preceding section.
If $f$ is a newform, we let $K_{0,f}$ denote the extension of $K_{0}$ generated by the Fourier coefficients of $f$, and $\lambda_{f}$ the prime over $\lambda_{0}$ determined by our embeddings. Define $K_{f}$ as the completion, ${\mathcal{O}}_{f}$ as its ring of integers and $\kappa_{f}$ as its residue field.
We assume that $\rho_{0}$ is modular, in the sense that
$\kappa_{f}\otimes_{\kappa}V_{0}\cong\kappa_{f}\otimes_{{\mathcal{O}}_{f}}{\mathcal{M}}_{f,\lambda_{f}}$
for some newform $f$ (which can have any weight level and character). We let $\mathcal{N}$} denote the set of representations
$\rho_{f}:G_{\mathbb{Q}}\to\operatorname{\operatorname{{\mathop{\rm Aut}\nolimits}}}_{K_{f}}M_{f,\lambda_{f}}$
such that where $f$ is a newform of weight $k$ character $\psi$ and level $N_{f}$ not divisible by $\ell$. Thus $\mathcal{N}$} is in bijection with the set of modular forms $f$ such that $\rho_{f}$ is an allowable lift of $\rho_{0}$. From the work of Ribet and others, one knows that $\mathcal{N}$$_{\emptyset}$ is non-empty. (Combine the main results of [D-T2] and [Di2].)
Let $\Sigma_{1}\subset\Sigma_{0}$ denote the set of primes at which $\rho_{0}$ is ramified. We let $\Sigma_{2}\subset\Sigma_{1}$ denote the following set of exceptional primes for $\rho_{0}$: those $p\equiv-1\bmod\ell$ such that $\rho_{0}|G_{p}$ is absolutely irreducible, but $\rho_{0}|I_{p}$ is not For each such $p$, $\bar{\kappa}\otimes_{\kappa}\rho_{0}$ is induced from a character of $H_{p}$, and we let $\xi_{p}$, $\xi_{p}^{\prime}$ denote the pair of characters ${\mathcal{O}}_{F}^{\times}\to\bar{\kappa}^{\times}$ defined as in §6.5. We then write
$\sigma_{p}^{\mathop{\rm min}\nolimits}:\operatorname{{\mathop{\rm GL}\nolimits}}_{2}({\mathbb{Z}}_{p})\to\operatorname{\operatorname{{\mathop{\rm Aut}\nolimits}}}_{K_{0}}V_{p}^{\mathop{\rm min}\nolimits}$
for the representation associated to $\tilde{\xi}_{p}$, $e_{p}^{\mathop{\rm min}\nolimits}$ for a distinguished vector and $\mu_{p}^{\mathop{\rm min}\nolimits}$ for a normalized pairing. Here we assume $K_{0}$ is sufficiently large so that
  • $K_{0}$ contains the values of $\tilde{\psi}$;
  • for each $p$ in $\Sigma_{2}$, $\sigma_{p}^{\mathop{\rm min}\nolimits}$, $e_{p}^{\mathop{\rm min}\nolimits}$ and $\mu_{p}^{\mathop{\rm min}\nolimits}$ are defined over $K_{0}$.
Write the Artin conductor $N(\rho_{0})$ as $\prod_{p}p^{c_{p}}$. Note that $c_{p}$ is even if $p$ is in $\Sigma_{2}$. We let $c_{p}^{\mathop{\rm min}\nolimits}=0$ if $p$ is in $\Sigma_{2}$ and $c_{p}^{\mathop{\rm min}\nolimits}=c_{p}$ otherwise. For any prime $p\neq\ell$, we let $c_{p}^{{\mathop{\rm max}\nolimits}}=c_{p}+\delta_{p}$ where $\delta_{p}=\dim_{\kappa}V_{0}^{I_{p}}$. For each finite $\Sigma\subset\Sigma_{0}$, we let $c_{p}^{\Sigma}=c_{p}^{{\mathop{\rm max}\nolimits}}$ or $c_{p}^{\mathop{\rm min}\nolimits}$ according to whether or not $p$ is in $\Sigma$. We let $U_{p}^{*}$ denote the set of matrices ${\mbox{\scriptsize$\left(\begin{array}[]{cc}{{a}}&{{b}}\\ {{c}}&{{d}}\end{array}\right)$}}\in\operatorname{{\mathop{\rm GL}\nolimits}}_{2}({\mathbb{Z}}_{p})$ with $c\in p^{c_{p}^{*}}$, where * can be $\mathop{\rm min}\nolimits$} or $\mathop{\rm max}\nolimits$}}. We let $U_{p}^{\Sigma}=U_{p}^{{\mathop{\rm max}\nolimits}}$ or $U_{p}^{\mathop{\rm min}\nolimits}$ according to whether or not $p$ is in $\Sigma$. We then set
$U^{\Sigma}=\prod_{p}U_{p}^{\Sigma}=\prod_{p\in\Sigma}U_{p}^{{\mathop{\rm max}\nolimits}}\prod_{p\not\in\Sigma}U_{p}^{\mathop{\rm min}\nolimits}=U_{0}(N^{\Sigma}),$
where $N^{\Sigma}=\prod_{p}c_{p}^{\Sigma}$.
We have already defined a representation
$\sigma_{p}^{\mathop{\rm min}\nolimits}:U_{p}^{\mathop{\rm min}\nolimits}\to\operatorname{\operatorname{{\mathop{\rm Aut}\nolimits}}}_{K_{0}}V_{p}^{\mathop{\rm min}\nolimits}$
for $p\in\Sigma_{2}$. If $c_{p}^{\mathop{\rm min}\nolimits}>0$, we define $\sigma_{p}^{\mathop{\rm min}\nolimits}$ by $V_{p}^{\mathop{\rm min}\nolimits}=K_{0}$ with $\left(\begin{array}[]{cc}{{a}}&{{b}}\\ {{c}}&{{d}}\end{array}\right)$ acting via $\psi_{p}(d)$, where $\psi_{p}=\tilde{\psi}|_{{\mathbb{Z}}_{p}^{\times}}$. If $p$ does not divide $N(\rho_{0})$ then $\sigma_{p}^{\mathop{\rm min}\nolimits}$ is trivial. For any $p$, we define $\sigma_{p}^{{\mathop{\rm max}\nolimits}}:U_{p}^{{\mathop{\rm max}\nolimits}}\to K_{0}^{\times}$ by $\psi_{p}(d)$. We let $\sigma_{p}^{\Sigma}=\sigma_{p}^{{\mathop{\rm max}\nolimits}}$ or $\sigma_{p}^{\mathop{\rm min}\nolimits}$ according to whether or not $p\in\Sigma$. We then define the representation $V_{0}^{\Sigma}$ of $U^{\Sigma}$ as the tensor product over all $p$ of $V_{p}^{\Sigma}$. Similarly we define the ${\mathcal{O}}_{0}$-lattice ${\mathcal{V}}^{\Sigma}$ in $V_{0}^{\Sigma}$ as the tensor product over $\mathcal{O}$} of ${\mathcal{V}}_{p}^{\Sigma}$.
If $\rho\in{\mathcal{N}}^{\Sigma}$, then $U_{f}^{\Sigma}=U^{\Sigma}$ and $\sigma_{f}^{\Sigma}=K_{0,f}\otimes_{K_{0}}\sigma^{\Sigma}$. We therefore have a natural inclusion $V_{\rho}\to K_{\rho}\otimes_{K}M(\sigma^{\Sigma})_{!,\lambda_{0}}$ induced by the isomorphism $M_{f}\to M_{f}^{\Sigma}$. These give rise to an inclusion
$V^{\Sigma}\to\bar{K}\otimes_{K}M(\sigma^{\Sigma})_{!,\lambda_{0}}.$
We define $L^{\Sigma}$ as the preimage of ${\mathcal{M}}(\sigma^{\Sigma})_{!,\lambda_{0}}$ in $V^{\Sigma}$, where the integral structure is defined as in §6.4 using ${\mathcal{V}}^{\Sigma}$.
Let $\tilde{{\mathbb{T}}}^{\Sigma}$ denote the polynomial algebra over $\mathcal{O}$} generated by the variables $t_{p}$, where $p$ runs over rational primes $p$ such that $p\in\Sigma$ or $p\not\in\Sigma_{1}$. We extend the natural action of $\mathcal{O}$} on $V^{\Sigma}$ to one of $\tilde{{\mathbb{T}}}^{\Sigma}$ by defining the action of $t_{p}$ on $V_{\rho}$ to be 0 or $\operatorname{{\mathop{\rm tr}\nolimits}}(\rho(\operatorname{{\mathop{\rm Frob}\nolimits}}_{p}))$ according to whether or not $p\in\Sigma$. Arguing as in the proof of Theorem 7.3, we see that $L^{\Sigma}$ is stable under the action of $\tilde{{\mathbb{T}}}^{\Sigma}$.
We also regard ${\mathcal{M}}(\sigma^{\Sigma})_{!,\lambda_{0}}$ as a module for $\tilde{{\mathbb{T}}}^{\Sigma}$ where $t_{p}$ acts as the usual double coset operator. We let ${\mathfrak{m}}^{\Sigma}$ denote the maximal ideal of $\tilde{{\mathbb{T}}}^{\Sigma}$ which is the kernel of the homomorphism defined by $t_{p}\mapsto 0$ or $\operatorname{{\mathop{\rm tr}\nolimits}}(\rho_{0}(\operatorname{{\mathop{\rm Frob}\nolimits}}_{p}))$ according to whether or not $p$ is in $\Sigma$. We let ${\mathcal{M}}_{!}^{\Sigma}$ denote the localization of ${\mathcal{M}}(\sigma^{\Sigma})_{!,\lambda_{0}}$ at ${\mathfrak{m}}^{\Sigma}$. Thus ${\mathcal{M}}_{!}^{\Sigma}$ can be canonically identified with a quotient (also a direct summand) of ${\mathcal{M}}(\sigma^{\Sigma})_{!,\lambda_{0}}$.
Lemma $7.$
The composites of the natural maps
$\begin{array}[]{ccccc}L^{\Sigma}&\to&{\mathcal{M}}(\sigma^{\Sigma})_{!,\lambda_{0}}&\to&{\mathcal{M}}_{!}^{\Sigma}\\ V^{\Sigma}&\to&\bar{K}\otimes_{\mathcal{O}}{\mathcal{M}}(\sigma^{\Sigma})_{!,\lambda_{0}}&\to&\bar{K}\otimes_{\mathcal{O}}{\mathcal{M}}_{!}^{\Sigma}\end{array}$
are isomorphisms of $\tilde{{\mathbb{T}}}^{\Sigma}$-modules.
Proof.
Let $\bar{{\mathbb{T}}}^{\Sigma}$ denote the image of ${\mathbb{T}}^{\Sigma}$ in $\operatorname{{\mathop{\rm End}\nolimits}}_{\mathcal{O}}{\mathcal{M}}(\sigma^{\Sigma})_{!,\lambda_{0}}$. Let $I^{\Sigma}$ denote the kernel of $\bar{{\mathbb{T}}}^{\Sigma}\to\bar{{\mathbb{T}}}^{\Sigma}_{{\mathfrak{m}}^{\Sigma}}$, or equivalently, the annihilator of ${\mathcal{M}}_{!}^{\Sigma}$.
As in [D-D-T, §4.2] (see also [C-D-T, Lemma 5.1.1]), one gets an explicit description of $\bar{K}\otimes_{\mathcal{O}}\bar{{\mathbb{T}}}^{\Sigma}_{{\mathfrak{m}}^{\Sigma}}$: There is an isomorphism
$\bar{K}\otimes_{\mathcal{O}}\bar{{\mathbb{T}}}^{\Sigma}_{{\mathfrak{m}}^{\Sigma}}\to\prod_{\rho\in{\mathcal{N}}^{\Sigma}}\bar{K}$
defined by $(1\otimes t_{p})\mapsto 0$ or $\operatorname{{\mathop{\rm tr}\nolimits}}(\rho(\operatorname{{\mathop{\rm Frob}\nolimits}}_{p}))_{\rho}$ according to whether or not $p$ is in $\Sigma$. We can then identify $V^{\Sigma}$ with the $\bar{{\mathbb{T}}}^{\Sigma}$-module
$\bar{K}\otimes_{\mathcal{O}}{\mathcal{M}}(\sigma^{\Sigma})_{!,\lambda_{0}}[\bar{K}\otimes_{\mathcal{O}}I^{\Sigma}].$
It follows that the second composite in the statement of the lemma is an isomorphism. The first is therefore injective and
$L^{\Sigma}={\mathcal{M}}(\sigma^{\Sigma})_{!,\lambda_{0}}[I^{\Sigma}],$
so it is also surjective. ∎
Suppose now that $\Sigma\subset\Sigma^{\prime}\subset\Sigma_{0}$. For each exceptional prime $p$ in $\Sigma^{\prime}-\Sigma$, we have a map $\tau_{p}:V_{p}^{{\mathop{\rm max}\nolimits}}\to V_{p}^{\mathop{\rm min}\nolimits}$ restricting to ${\mathcal{V}}_{p}^{{\mathop{\rm max}\nolimits}}\to{\mathcal{V}}_{p}^{\mathop{\rm min}\nolimits}$ (see §6.5). We define $\tau:{\mathcal{V}}^{\Sigma}\to{\mathcal{V}}^{\Sigma^{\prime}}$ as the tensor product of the $\tau_{p}$. Let
$N^{\prime}=N^{\Sigma^{\prime}}/\prod_{p}p^{c_{p}/2}\quad\mbox{and}\quad N=N^{\Sigma}\prod_{p}p^{c_{p}/2}$
[…]
$\gamma=\sum_{d}\frac{1}{d^{\prime}}\left[U^{\Sigma^{\prime}}\alpha_{d}U^{\Sigma}\right]_{\tau}\beta_{d}.$
Then for each $\rho\in{\mathcal{N}}^{\Sigma}$, $\gamma$ induces the identity map on $V_{\rho}$. Furthermore $\gamma$ restricts to a map ${\mathcal{M}}(\sigma^{\Sigma})_{!}\to{\mathcal{M}}(\sigma^{\Sigma^{\prime}})_{!}$. It follows that $L^{\Sigma}\subset L^{\Sigma^{\prime}}$. To prove that $L={\displaystyle\bigcup_{\Sigma}}L^{\Sigma}$ is a trellis for $\mathcal{N}$}, we just need the following:
Proposition 7.8.
If $\Sigma\subset\Sigma^{\prime}\subset\Sigma^{0}$, then $L^{\Sigma^{\prime}}/L^{\Sigma}$ is torsion-free.
Proof.
It suffices to consider the case $\Sigma^{\prime}=\Sigma\cup\{p\}$ for some $p$ not in $\Sigma\cup\{\ell\}$. We let $c=c_{p}$, $\gamma={{\mathop{\rm max}\nolimits}}\{1,c\}$ and $\delta=\delta_{p}$.
Choose an integer $N\geq 3$ such that $(p\ell,N)=1$ and
$V_{N}^{p}=\ker(\operatorname{{\mathop{\rm GL}\nolimits}}_{2}(\hat{{\mathbb{Z}}})\to\operatorname{{\mathop{\rm GL}\nolimits}}_{2}({\mathbb{Z}}/N{\mathbb{Z}})\times\operatorname{{\mathop{\rm GL}\nolimits}}_{2}({\mathbb{Z}}_{p}))$
is contained in $\ker\sigma^{\Sigma}$. We now define subgroups $V_{p}$ and $V_{p}^{\prime}$ of $\operatorname{{\mathop{\rm GL}\nolimits}}_{2}({\mathbb{Z}}_{p})$ such that $V_{p}\subset\ker\sigma_{p}^{\mathop{\rm min}\nolimits}$ is normal in $U_{p}^{\mathop{\rm min}\nolimits}$, $V_{p}^{\prime}\subset\ker\sigma_{p}^{{\mathop{\rm max}\nolimits}}$ is normal in $U_{p}^{{\mathop{\rm max}\nolimits}}$ and
$V_{p}^{\prime}\subset\bigcap_{i=0}^{\delta}\alpha_{p^{i}}V_{p}\alpha_{p^{i}}^{-1}.$
(Recall that $\alpha_{p^{i}}={\mbox{\scriptsize$\left(\begin{array}[]{cc}{{p^{-i}}}&{{0}}\\ {{0}}&{{1}}\end{array}\right)$}}$ unless $p\in\Sigma_{2}$, in which case $\alpha_{1}={\mbox{\scriptsize$\left(\begin{array}[]{cc}{{p^{-c/2}}}&{{0}}\\ {{0}}&{{1}}\end{array}\right)$}}_{p}$.) If $p$ is in $\Sigma_{2}$, we define $V_{p}$ as the kernel of $\operatorname{{\mathop{\rm GL}\nolimits}}_{2}({\mathbb{Z}}_{p})\to\operatorname{{\mathop{\rm GL}\nolimits}}_{2}({\mathbb{Z}}/p^{c/2})$ and $V_{p}^{\prime}=\alpha_{1}V_{p}\alpha_{1}^{-1}$. Otherwise, we define $V_{p}$ as the set of matrices in $\operatorname{{\mathop{\rm GL}\nolimits}}_{2}({\mathbb{Z}}_{p})$ congruent to $\left(\begin{array}[]{cc}{1}&{*}\\ {0}&{1}\end{array}\right)$ mod $p^{c}$, and
$V_{p}^{\prime}=\left\{\left.\,\left(\begin{array}[]{cc}{w}&{x}\\ {y}&{z}\end{array}\right)\,\right|\,w\equiv z\equiv 1\bmod p^{\gamma},\,y\equiv 0\bmod p^{c+\delta}\,\right\}.$
We then let $V_{N}^{p}V_{p}, V=V$$^{\prime}=V_{N}^{p}V_{p}^{\prime}$. Let $\tilde{{\mathbb{T}}}$ denote the common subring of $\tilde{{\mathbb{T}}}^{\Sigma}$ and $\tilde{{\mathbb{T}}}^{\Sigma^{\prime}}$ generated by the variables $t_{p}$ for $p$ not dividing $Np\ell$, and let ${\mathfrak{m}}={\mathfrak{m}}^{\Sigma}\cap\tilde{{\mathbb{T}}}={\mathfrak{m}}^{\Sigma^{\prime}}\cap\tilde{{\mathbb{T}}}$.
We wish to prove that the map
$\kappa\otimes_{{\mathcal{O}}}L^{\Sigma}\to\kappa\otimes_{{\mathcal{O}}}L^{\Sigma^{\prime}}$
is injective. The natural map $L^{\Sigma}\to{\mathcal{M}}(\sigma^{\Sigma})_{!,{\mathfrak{m}}}$ is injective, so it suffices to prove the map
$\kappa\otimes_{{\mathcal{O}}}{\mathcal{M}}(\sigma^{\Sigma})_{!,{\mathfrak{m}}}\to\kappa\otimes_{\mathcal{O}}{\mathcal{M}}(\sigma^{\Sigma^{\prime}})_{!,{\mathfrak{m}}}$
defined in the same way as $\gamma$ is injective. Therefore it suffices to prove that
$\bigoplus_{i=0}^{\delta}\kappa\otimes_{\mathcal{O}}{\mathcal{M}}(\sigma^{\Sigma})_{!,{\mathfrak{m}}}\to\kappa\otimes_{\mathcal{O}}{\mathcal{M}}(\sigma^{\Sigma^{\prime}})_{!,{\mathfrak{m}}}$
is injective, where the map is defined by
$\bigoplus_{i=0}^{\delta}p^{-i}\left[U^{\Sigma^{\prime}}\alpha_{p^{i}}U^{\Sigma}\right]_{\tau}.$
Since $\rho_{0}$ is irreducible, $\mathfrak{m}$} is not in the support of the kernel of ${\mathcal{M}}(V)_{c}\to{\mathcal{M}}(V)_{!}$ or ${\mathcal{M}}(V^{\prime})_{c}\to{\mathcal{M}}(V^{\prime})_{!}$. We can therefore identify the above map with the map
$\bigoplus_{i=0}^{\delta}\kappa\otimes_{\mathcal{O}}({\mathcal{V}}^{\Sigma}\otimes_{{\mathcal{O}}_{0}}{\mathcal{M}}(V))_{!,{\mathfrak{m}}}^{U^{\Sigma}}\to\kappa\otimes_{\mathcal{O}}({\mathcal{V}}^{\Sigma^{\prime}}\otimes_{{\mathcal{O}}_{0}}{\mathcal{M}}(V^{\prime}))^{U^{\Sigma^{\prime}}}_{!,{\mathfrak{m}}}$
induced by
$\bigoplus_{i=0}^{\delta}p^{-i}\left[V^{\prime}\alpha_{p^{i}}V\right]_{\tau}.$
We shall prove that in fact if $\delta<2$, then
$\gamma^{\prime}:\bigoplus_{i=0}^{\delta}({\mathcal{V}}_{0}^{\Sigma}\otimes_{{\mathcal{O}}_{0}}{\mathcal{M}}(V))_{!}^{U^{\Sigma}}\to({\mathcal{V}}_{0}^{\Sigma^{\prime}}\otimes_{{\mathcal{O}}_{0}}{\mathcal{M}}(V^{\prime}))_{!}^{U^{\Sigma^{\prime}}}$
is injective, where ${\mathcal{V}}_{0}^{\Sigma}=\kappa\otimes_{{\mathcal{O}}_{0}}{\mathcal{V}}^{\Sigma}$ and ${\mathcal{V}}_{0}^{\Sigma^{\prime}}=\kappa\otimes_{{\mathcal{O}}_{0}}{\mathcal{V}}^{\Sigma^{\prime}}$, and that the same holds after localization at $\mathfrak{m}$} if $\delta=2$.
First suppose $\delta=0$. If $p$ is not in $\Sigma_{2}$, then $\gamma^{\prime}$ is the identity and there is nothing to prove. So assume $p$ is in $\Sigma_{2}$. By Shapiro’s Lemma, we can replace the target of $\gamma^{\prime}$ with
$\left(\operatorname{Ind}_{\alpha U^{\Sigma^{\prime}}\alpha^{-1}}^{U^{\Sigma}}\alpha^{-1}{\mathcal{V}}_{0}^{\Sigma^{\prime}}\otimes_{{\mathcal{O}}_{0}}{\mathcal{M}}(V)_{!}\right)^{U^{\Sigma}},$
where $\alpha=\alpha_{1}$ and the map is defined so the composite
${\mathcal{V}}_{0}^{\Sigma}\to\operatorname{Ind}_{\alpha U^{\Sigma^{\prime}}\alpha^{-1}}^{U^{\Sigma}}\alpha^{-1}{\mathcal{V}}_{0}^{\Sigma^{\prime}}\to\alpha^{-1}{\mathcal{V}}_{0}^{\Sigma^{\prime}}$
is $\alpha^{-1}\tau$. Therefore it suffices to prove that
$\kappa\otimes_{{\mathcal{O}}_{0}}{\mathcal{V}}_{p}^{\mathop{\rm min}\nolimits}\to\kappa\otimes_{{\mathcal{O}}_{0}}\operatorname{Ind}_{\alpha U^{\Sigma^{\prime}}\alpha^{-1}}^{U^{\Sigma}}\alpha^{-1}{\mathcal{V}}_{p}^{{{\mathop{\rm max}\nolimits}}}$
is injective. This holds because $\kappa\otimes_{{\mathcal{O}}_{0}}{\mathcal{V}}_{p}^{\mathop{\rm min}\nolimits}$ is irreducible and $\kappa\otimes_{{\mathcal{O}}_{0}}\tau_{p}$ is non-zero (since it is surjective by its construction).
Now suppose $\delta=1$. Let ${\mathcal{V}}_{0}={\mathcal{V}}_{0}^{\Sigma}={\mathcal{V}}_{0}^{\Sigma^{\prime}}$, $V^{\prime\prime}=\alpha_{p}^{-1}V^{\prime}\alpha_{p}$ and $A={\mathcal{V}}_{0}\otimes_{{\mathcal{O}}_{0}}{\mathcal{M}}_{!}(V^{\prime\prime})$. Then $V^{\prime\prime}$ is normal in both $U^{\Sigma}$ and $\alpha_{p}^{-1}U^{\Sigma}\alpha_{p}$, and both groups act on AA, where the action of $\alpha_{p}^{-1}U^{\Sigma}\alpha_{p}$ on ${\mathcal{V}}_{0}$ is defined by first conjugating by $\alpha_{p}$. Note that the actions agree on
$\alpha_{p}^{-1}U^{\Sigma^{\prime}}\alpha_{p}=U^{\Sigma}\cap\alpha_{p}^{-1}U^{\Sigma}\alpha_{p}.$
We claim that the map
$A^{\alpha_{p}^{-1}U^{\Sigma}\alpha_{p}}\oplus A^{U^{\Sigma}}\to A^{\alpha_{p}^{-1}U^{\Sigma^{\prime}}\alpha_{p}}$
is injective. To prove this, suppose there were a nonzero vector in AA fixed by both $\alpha_{p}^{-1}U^{\Sigma}\alpha_{p}$ and $A^{U^{\Sigma}}$. Then there would be a nonzero vector in ${\mathcal{M}}_{!}(V^{\prime\prime})$ on which $\left(\begin{array}[]{cc}{{a}}&{{b}}\\ {{c}}&{{d}}\end{array}\right)$ acts via $\psi_{p}^{-1}(d)$ for all matrices in $U_{p}^{\mathop{\rm min}\nolimits}$ or $\alpha_{p}^{-1}U_{p}^{\mathop{\rm min}\nolimits}\alpha_{p}$. This contradicts the fact that $\psi_{p}$ has conductor $p^{c}$. It follows that the map
$B^{U^{\Sigma}}\oplus A^{U^{\Sigma}}\to B^{U^{\Sigma^{\prime}}}$
where $B={\mathcal{V}}_{0}\otimes_{{\mathcal{O}}_{0}}{\mathcal{M}}_{!}(V^{\prime})$ and the map $A\to B$ is defined by $p^{-1}\left[V^{\prime}\alpha_{p}V^{\prime\prime}\right].$ Since $V/V^{\prime}$ and $V/V^{\prime\prime}$ have order $p$, we can identify ${\mathcal{M}}(V^{\prime})_{!}^{V}$ and ${\mathcal{M}}(V^{\prime\prime})_{!}^{V}$ with ${\mathcal{M}}(V)_{!}$ and conclude $\gamma^{\prime}$ is injective.
Finally, suppose that $\delta=2$. Again we let ${\mathcal{V}}_{0}={\mathcal{V}}_{0}^{\Sigma}={\mathcal{V}}_{0}^{\Sigma^{\prime}}$, $V^{\prime\prime}=\alpha_{p}^{-1}V^{\prime}\alpha_{p}$ and $A={\mathcal{V}}_{0}\otimes_{{\mathcal{O}}_{0}}{\mathcal{M}}_{!}(V^{\prime\prime})$. We also define $\tilde{U}$ to be the subgroup of $U^{\Sigma}$ generated by $U^{\Sigma^{\prime}}$ and $\alpha_{p}^{-1}U^{\Sigma}\alpha_{p}$. Then we have an exact sequence
$0\to A^{U^{\Sigma}}\to A^{\alpha_{p}^{-1}\tilde{U}\alpha_{p}}\oplus A^{\tilde{U}}\to A^{\alpha_{p}^{-1}U^{\Sigma^{\prime}}\alpha_{p}}$
with maps defined by $a\mapsto(a,-a)$ and $(a_{1},a_{2})\mapsto a_{1}+a_{2}$. This gives an exact sequence
(47) $0\to A^{U^{\Sigma}}\to B^{\tilde{U}}\oplus A^{\tilde{U}}\to B^{U^{\Sigma^{\prime}}}$
[…]
$(a,b)\mapsto\left[\tilde{V}\cdot 1\cdot V\right]_{1}(a)+\left[\tilde{V}\alpha_{p}V\right]_{1}(b).$
So we also get the exactness of
(48) $0\to C^{U^{\Sigma}}\times C^{U^{\Sigma}}\to D^{\tilde{U}}$.
Now let $A_{c}$, $B_{c}$, $C_{c}$ and $D_{c}$ denote the $\tilde{{\mathbb{T}}}-$modules defined in the same way as AA, $B$, $C$ and $D$, but using ${\mathcal{M}}_{c}$ instead of $\mathcal{M}$$_{!}$. Then the natural maps $C_{c}\to A_{c}^{V}$, $D_{c}\to B_{c}^{\tilde{V}}$ and $D_{c}\to A_{c}^{\tilde{V}}$ are isomorphisms. Since $\mathfrak{m}$} is not Eisenstein, it follows that so are the maps $C_{\mathfrak{m}}^{U^{\Sigma}}\to A_{\mathfrak{m}}^{U}$, $D_{\mathfrak{m}}^{\tilde{U}}\to B_{\mathfrak{m}}^{\tilde{U}}$ and $D_{\mathfrak{m}}^{\tilde{U}}\to A_{\mathfrak{m}}^{\tilde{U}}$. We now obtain the desired injectivity on combining the localization at $\mathfrak{m}$} of the exact sequences (47) and (48) (cf. the discussion after Lemma 2.5 of [Wi]). ∎
Recall that in §6.5, we defined a perfect pairing on $M(\sigma^{\Sigma})_{!}$, and we now let $\phi^{\Sigma}$ denote the paring
$M(\sigma^{\Sigma})_{!,\lambda_{0}}\otimes_{K}M(\sigma^{\Sigma})_{!,\lambda_{0}}\to K.$
Lemma 7.9.
The restriction of $\phi^{\Sigma}$ defines a perfect pairing $L^{\Sigma}\otimes_{\mathcal{O}}L^{\Sigma}\to{\mathcal{O}}$.
Proof.
First we apply Lemma 3 of [D-T2] to choose a prime $r\geq 5$ at which no lift of $\rho_{0}$ is ramified. It follows that $L^{\Sigma}$ can be identified with $L^{\Sigma\cup\{r\}}$ compatibly with the map
${\mathcal{M}}(\sigma^{\Sigma})_{!,\lambda_{0}}\to{\mathcal{M}}(\sigma^{\Sigma\cup\{r\}})_{!,\lambda_{0}}.$
According to Proposition 6.14, the restriction to $L^{\Sigma}$ of the transpose of this map with respect to our pairings is given by an element of $\tilde{{\mathbb{T}}}^{\Sigma}-{\mathfrak{m}}^{\Sigma}$, and so by Lemma 7.7, is an automorphism. So to prove the lemma, we can assume $r$ is in $\Sigma$.
Let $U^{\prime}$ denote the set of $\left(\begin{array}[]{cc}{a}&{b}\\ {c}&{d}\end{array}\right)$ in $U^{\Sigma}$ such that $d_{r}\equiv 1\bmod r$, and similarly define $U^{\prime\prime}$ requiring $a_{r}\equiv 1\bmod r$. We let $\sigma^{\prime}$ and $\sigma^{\prime\prime}$ denote the restrictions of $\sigma^{\Sigma}$ to $U^{\prime}$ and $U^{\prime\prime}$ respectively. If we define an action of $\tilde{{\mathbb{T}}}^{\Sigma}$ on ${\mathcal{M}}(\sigma^{\prime})_{!}$ and ${\mathcal{M}}(\sigma^{\prime\prime})_{!}$ in the usual way and the natural inclusions
${\mathcal{M}}(\sigma)_{!}\to{\mathcal{M}}(\sigma^{\prime})_{!}\quad\mbox{and}\quad{\mathcal{M}}(\sigma)_{!}\to{\mathcal{M}}(\sigma^{\prime\prime})_{!}$
are $\tilde{{\mathbb{T}}}^{\Sigma}$-linear. We record the following as a lemma for future reference:
Lemma 7.10.
The natural maps induce isomorphisms
${\mathcal{M}}(\sigma)_{!,{\mathfrak{m}}^{\Sigma}}\to{\mathcal{M}}(\sigma^{\prime})_{!,{\mathfrak{m}}^{\Sigma}}\quad\mbox{and}\quad{\mathcal{M}}(\sigma)_{!,{\mathfrak{m}}^{\Sigma}}\to{\mathcal{M}}(\sigma^{\prime\prime})_{!,{\mathfrak{m}}^{\Sigma}}.$
To prove this, note that we can replace !! by $c$ since ${\mathfrak{m}}^{\Sigma}$ is not Eisenstein. So we must prove the inclusion
${\mathcal{M}}(\sigma^{\prime})_{c}^{U^{\Sigma}}\to{\mathcal{M}}(\sigma^{\prime})_{c}$
localized at ${\mathfrak{m}}^{\Sigma}$ is an isomorphism, and similarly for $\sigma^{\prime\prime}$. Let $u$ be a generator for $U^{\Sigma}/U^{\prime}$, then $u$ acts on ${\mathcal{M}}(\sigma^{\prime})_{c}$ as $p^{2-k}S_{p}$ for some prime $p$ at which $\rho_{0}$ is unramified and ${\mathop{\rm det}\nolimits}\rho_{0}(\operatorname{{\mathop{\rm Frob}\nolimits}}_{p})=p^{k-1}$. It follows that $u$ acts as an element of $1+{\mathfrak{m}}^{\Sigma}$. Since $u$ has order r1r-1, which is not divisible by $\ell$, it follows that $u=1$ on ${\mathcal{M}}(\sigma^{\prime})_{c,{\mathfrak{m}}^{\Sigma}}$. The same argument works for $\sigma^{\prime\prime}$ and this completes the proof of Lemma 7.10.
Since $U^{\prime}$ is sufficiently small, we have a perfect pairing between ${\mathcal{M}}(\sigma^{\prime})_{!}$ and ${\mathcal{M}}(\hat{\sigma}^{\prime}\otimes\psi^{-1}\circ{\mathop{\rm det}\nolimits})_{!}$ given by Proposition 4.8. We also have an isomorphism
${\mathcal{M}}(\sigma^{\prime\prime})_{!}\to{\mathcal{M}}(\hat{\sigma}^{\prime}\otimes\psi^{-1}\circ{\mathop{\rm det}\nolimits})_{!}$
defined as in §6.5, and so we obtain a perfect pairing between ${\mathcal{M}}(\sigma^{\prime})_{!}$ and ${\mathcal{M}}(\sigma^{\prime\prime})_{!}$. It is straightforward to check that the resulting isomorphism
${\mathcal{M}}(\sigma^{\prime\prime})_{!,\lambda_{0}}\to{\mathop{\rm Hom}\nolimits}_{{\mathcal{O}}}({\mathcal{M}}(\sigma^{\prime})_{!,\lambda_{0}},{\mathcal{O}})$
is $\tilde{{\mathbb{T}}}^{\Sigma}$-linear, and that the composite
${\mathcal{M}}(\sigma)_{!,\lambda_{0}}\otimes_{\mathcal{O}}{\mathcal{M}}(\sigma)_{!,\lambda_{0}}\to{\mathcal{M}}(\sigma^{\prime})_{!,\lambda_{0}}\otimes_{\mathcal{O}}{\mathcal{M}}(\sigma^{\prime\prime})_{!,\lambda_{0}}\to{\mathcal{O}}$
is $(r-1)\phi^{\Sigma}$. It follows from Lemmas 7.7 and 7.10 that $\phi^{\Sigma}$ restricts to a perfect pairing on $L^{\Sigma}$. ∎
In view of Proposition 6.14, we only need to prove
${\mathcal{N}}^{Q}\leq\#\Delta_{Q}\#{\mathcal{N}}^{\emptyset}$
to enable us to apply Theorem 7.3. We shall in fact prove:
Lemma 7.11.
For any finite set of primes $\Sigma\subset\Sigma_{0}$ and horizontal prime $q\not\in\Sigma$, we have
${\mathcal{N}}^{\Sigma\cup\{q\}}=\#\Delta_{\{q\}}\#{\mathcal{N}}^{\Sigma}.$
Proof.
First note that we may assume $\Sigma$ contains an auxiliary prime $r$ as in the proof of Lemma 7.9. Let $\tilde{{\mathbb{T}}}=\tilde{{\mathbb{T}}}^{\Sigma}=\tilde{{\mathbb{T}}}^{\Sigma\cup\{q\}}$. Choose an eigenvalue $\alpha_{q}$ of $\rho_{0}(\operatorname{{\mathop{\rm Frob}\nolimits}}_{q})$ and let $\mathfrak{m}$} denote the maximal ideal which is the kernel of the map defined by $t_{q}\mapsto\alpha_{q}$, $t_{p}\mapsto 0$ for $p\in\Sigma$, and $t_{p}\mapsto\operatorname{{\mathop{\rm tr}\nolimits}}\rho_{0}(\operatorname{{\mathop{\rm Frob}\nolimits}}_{p})$ otherwise.
We let $U_{0}$ denote the set of $\left(\begin{array}[]{cc}{{a}}&{{b}}\\ {{c}}&{{d}}\end{array}\right)$ in $U^{\Sigma}$ such that $c_{q}\equiv 0\bmod q$ and $d_{r}\equiv 1\bmod r$, and let $U_{1}$ denote the set of $\left(\begin{array}[]{cc}{{a}}&{{b}}\\ {{c}}&{{d}}\end{array}\right)$ in $U_{0}$ such that $c_{q}$ has trivial image in $\Delta_{\{q\}}$. We identify $\Delta_{\{q\}}$ with $U_{0}/U_{1}$. Let $\sigma_{i}$ denote the restriction of $\sigma^{\Sigma}$ to $U_{i}$ for $i=0,1$. Let $\bar{{\mathbb{T}}}_{i}$ denote the image of $\tilde{{\mathbb{T}}}$ in $\operatorname{{\mathop{\rm End}\nolimits}}_{K}M(\sigma_{i})_{!,\lambda_{0}}$ for $i=0,1$.
As in [D-D-T, §4.2] (see also §6.4 of [C-D-T]), one can give an explicit description of $\bar{K}\otimes_{\mathcal{O}}\bar{{\mathbb{T}}}_{0,{\mathfrak{m}}}$ which shows that its dimension over $\bar{K}$ is ${\mathcal{N}}^{\Sigma}$. Similarly one finds the dimension of $\bar{K}\otimes_{\mathcal{O}}\bar{{\mathbb{T}}}_{1,{\mathfrak{m}}}$ is ${\mathcal{N}}^{\Sigma\cup\{q\}}$. Since $M(\sigma_{i})_{!,\lambda_{0}}$ is free of rank two over $K\otimes_{\mathcal{O}}\bar{{\mathbb{T}}}_{i}$ (see §6.5), we need to prove that
$\operatorname{{\mathop{\rm rank}\nolimits}}_{\mathcal{O}}{\mathcal{M}}(\sigma_{1})_{!,{\mathfrak{m}}}=\#\Delta_{\{q\}}\cdot\operatorname{{\mathop{\rm rank}\nolimits}}_{\mathcal{O}}{\mathcal{M}}(\sigma_{0})_{!,{\mathfrak{m}}}.$
We shall show that ${\mathcal{M}}(\sigma_{1})_{!,{\mathfrak{m}}}$ is free over ${\mathcal{O}}[\Delta_{\{q\}}]$ and its $\Delta_{\{q\}}$-invariants can be identified with ${\mathcal{M}}(\sigma_{0})_{!,{\mathfrak{m}}}$. The case $k=2$, trivial $\psi$ and $\Sigma_{2}=\emptyset$ is done exactly as in §2 of [T-W], so assume $k>2$, $\psi$ is not trivial or $\Sigma_{2}$ is not empty. It follows from the Hochschild-Serre spectral sequence that
$H^{i}(\Delta_{Q},{\mathcal{M}}(\sigma_{1})_{c})=\left\{\begin{array}[]{ll}{\mathcal{M}}(\sigma_{0})_{c},&\mbox{if $i=0$;}\\ 0,&\mbox{if $i>0$.}\end{array}\right$.
This holds as $\tilde{{\mathbb{T}}}-$modules, so it holds after completing at $\mathfrak{m}$}. We can then replace !! by $c$ and deduce the freeness of ${\mathcal{M}}(\sigma_{1})_{!,{\mathfrak{m}}}$ from cohomological vanishing. ∎
7.4. Consequences
Theorem 7.12.
Suppose $\rho:G_{\mathbb{Q}}\to\operatorname{\operatorname{{\mathop{\rm Aut}\nolimits}}}_{K}V$ is a continuous geometric representation whose restriction to $G_{\ell}$ is ramified, crystalline and short. If $\rho_{0}$ is modular and its restriction to $G_{F}$ is absolutely irreducible, where $F$ is the quadratic subfield of ${\mathbb{Q}}(\mu_{\ell})$, then $\rho$ is modular.
This follows from applying Theorem 7.3 to the set $\mathcal{N}$} constructed in the preceding section for the twist $\rho_{0}\otimes_{k}\psi^{\prime}$ of minimal conductor, where $\psi^{\prime}$ is unramified at $\ell$. Writing $\tilde{\ }$ for Teichmuller liftings, we conclude that $\rho\otimes_{K}\tilde{\psi}^{\prime}\psi$ is modular, where $\psi$ is a character of $\ell$-power order such that $\chi_{\ell}^{1-k}\psi^{2}{\mathop{\rm det}\nolimits}\rho$ has order not divisible by $\ell$.
Suppose that $K_{0}$ is a number field with ring of integers ${\mathcal{O}}_{0}$, and $f=\sum a_{n}e^{2\pi inz}$ is a newform of weight $k\geq 2$ and level NN with coefficients $a_{n}$ in $K_{0}$. For primes $\lambda$ of $K_{0}$, we write $\bar{\rho}_{f,\lambda}$ for the representation of $G_{\mathbb{Q}}$ on
$\bar{\kappa}\otimes_{{\mathcal{O}}}{\mathcal{M}}_{f,\lambda},$
where ${\mathcal{O}}={\mathcal{O}}_{0,\lambda}$ and $\bar{\kappa}$ is an algebraic closure of ${\mathcal{O}}_{0}/\lambda$. We use $\ell$ to denote the rational prime divisible by $\lambda$. Let $S_{f}$ denote the set of primes $\lambda$ of $K_{0}$ such that the following hold:
  • $\lambda$ does not divide $Nk!$;
  • $\bar{\rho}_{f,\lambda}$ has irreducible restriction to $G_{F}$, where $F$ is the quadratic subfield of ${\mathbb{Q}}(\mu_{\ell})$.
The latter condition excludes only finitely many primes by the following two lemmas:
Lemma 7.13.
For all but finitely many primes $\lambda$, $\bar{\rho}_{f,\lambda}$ is irreducible.
Proof.
Suppose that $\lambda$ does not divide $Nk!$ and $\bar{\rho}_{f,\lambda}$ is reducible. Its semisimplification is of the form $\psi_{1}\oplus\psi_{2}$ where $\psi_{1}$ and $\psi_{2}$ are characters of $\operatorname{{\mathop{\rm Gal}\nolimits}}({\mathbb{Q}}(\mu_{N\ell})/{\mathbb{Q}})$. The representation is necessarily ordinary at $\ell$ (see [E]), so one of the characters is unramified at $\ell$ and the other has restriction $\chi_{\ell}^{k-1}$ on $I_{\ell}$. It follows that
$a_{p}\equiv p^{k-1}+1\bmod\lambda$
for all $p\equiv 1\bmod N$. If this holds for infinitely many $\lambda$, then we get $a_{p}=p^{k-1}+1$ for all such $p$, violating the Ramanujan conjecture (a theorem of Deligne [De1]). ∎
Lemma 7.14.
Suppose that $\lambda$ does not divide N(2k1)(2k3)k!N(2k-1)(2k-3)k!. If $\bar{\rho}_{f,\lambda}$ is irreducible, then so is its restriction to $G_{F}$.
Proof.
Consider the restriction of $\bar{\rho}_{f,\lambda}$ to $I_{\ell}$. By results of Deligne and Fontaine (see [E]), this restriction has semisimplification of the form $\chi_{\ell}^{k-1}\oplus 1$ or $\psi_{\ell}^{k-1}\oplus\psi_{\ell}^{\ell(k-1)}$, where $\psi_{\ell}$ is a fundamental character of level 2, according to whether or not $a_{\ell}$ is a unit mod $\lambda$.
Suppose that $\bar{\rho}_{f,\lambda}$ is irreducible but its restriction to $G_{F}$ is not. Then $\bar{\rho}_{f,\lambda}$ is induced from a character of $G_{F}$, and $\bar{\rho}_{f,\lambda}|_{I_{\ell}}$ is induced from a character of its subgroup of index 2. It follows that the ratio of the characters into which $\bar{\rho}_{f,\lambda}|_{I_{\ell}}$ decomposes is quadratic. Since $\psi_{\ell}$ has order $\ell^{2}-1$, this forces either $(\ell-1)|2(k-1)$ or $(\ell+1)|2(k-1)$ and we arrive at a contradiction. ∎
Theorem 7.15.
Let $f$ be a newform of weight $k\geq 2$ and level NN with coefficients in $K_{0}$. Suppose that $\lambda$ is a prime in $S_{f}$ (as defined above), and let ${\mathcal{O}}={\mathcal{O}}_{0,\lambda}$ and $K=K_{0,\lambda}$. Suppose that $\Sigma$ is a finite set of primes not containing $\ell$ such that $M_{f,\lambda}$ is minimally ramified outside $\Sigma$. Then the $\mathcal{O}$}-module
$H^{1}_{\Sigma}(G_{\mathbb{Q}},K/{\mathcal{O}}\otimes_{{\mathcal{O}}}{\mathop{\rm ad}\nolimits}^{0}_{\mathcal{O}}{\mathcal{M}}_{f,\lambda})$
has length $v_{\lambda}(\eta_{f}^{\Sigma})$.
Applying Theorem 7.3 to the set $\mathcal{N}$} constructed in the preceding section for the twist $\rho_{0}\otimes_{k}\psi^{\prime}$ of minimal conductor, we conclude that the theorem holds for a twist of $f$, hence for $f$ itself. See the discussion before Proposition 6.6 for the definition of $\eta_{f}^{\Sigma}$.
8. The Bloch-Kato conjecture for $A_{f}$ and $A_{f}(1)$
In this section we explain how to deduce the $\lambda$-part of the Bloch-Kato conjecture [B-K] for $A_{f}$ and $B_{f}=A_{f}(1)$, where $f$ is a newform of weight $k\geq 2$, conductor $N\geq 1$, with coefficients in the number field $K$, and $\lambda$ is a prime in $S_{f}$ (the set defined in §7.4). By Lemma 6.2 we can assume that $f$ has minimal conductor among its twists and we shall do so in this section. Our formulation of the conjecture follows Fontaine and Perrin-Riou [F-P], generalized to motives with coefficients in $K$. For a more systematic discussion of the Bloch-Kato conjecture for motives with coefficients we refer to [B-F]. If $A_{f}$ is the scalar extension of a premotivic structure with coefficients in a subfield $K^{\prime}\subseteq K$ then Theorem 4.1 and Lemma 9 of [B-F] shows that the conjecture over $K$ implies the one over $K^{\prime}$ (in the context of Deligne’s conjecture this was already noted in [De3, Rem 2.10]). So we need not be concerned with finding the smallest coefficient field for $A_{f}$.
8.1. Order of vanishing
Suppose that $M$ is an $L$-admissible object of $\operatorname{{\bf PM}}_{K}$ and let $M^{D}={\mathop{\rm Hom}\nolimits}_{K}(M,K(1))$. We recall the conjectured order of vanishing of $L(M,s)$ at $s=0$ [F-P, III. 4.2.2].
Conjecture 8.1.
Let $\tau:K\to{\mathbb{C}}$ be an embedding and $\lambda$ any finite prime of $K$. Then
$\operatorname{{\mathop{\rm ord}\nolimits}}_{s=0}L(M,\tau,s)=\dim_{K_{\lambda}}H^{1}_{\bf f}({\mathbb{Q}},M^{D}_{\lambda})-\dim_{K_{\lambda}}H^{0}({\mathbb{Q}},M^{D}_{\lambda}).$
Theorem 8.2.
Conjecture 8.1 holds for both $M=A_{f}$ and $M=B_{f}$ if $\lambda$ is in $S_{f}$. More precisely, we have $\operatorname{{\mathop{\rm ord}\nolimits}}_{s=0}L(A_{f},\tau,s)=\operatorname{{\mathop{\rm ord}\nolimits}}_{s=0}L(B_{f},\tau,s)=0$ and
$H^{0}({\mathbb{Q}},A_{f,\lambda})\cong H^{1}_{\bf f}({\mathbb{Q}},A_{f,\lambda})\cong H^{1}_{\bf f}({\mathbb{Q}},B_{f,\lambda})\cong H^{0}({\mathbb{Q}},B_{f,\lambda})\cong 0.$
if $\lambda\in S_{f}$.
Proof.
Lemma 8.7 below shows that
$L(A_{f},\tau,1)=L^{\mathop{\rm nv}\nolimits}(A_{f},\tau,1)\prod_{p\in\Sigma_{e}(f)}L_{p}(A_{f},\tau,1)$
is a nonzero multiple of the Petersson inner product of $f$ with itself and hence it follows that $L(B_{f},\tau,0)=L(A_{f},\tau,1)\neq 0$ for each $\tau$. It follows from the functional equation (6.2) that
(49) $L(A_{f},\tau,0)=\frac{(k-1)\epsilon(A_{f})}{2\pi^{2}}L(A_{f},\tau,1)\neq 0$
for each $\tau$ as well. The absolute irreducibility of $M_{f,\lambda}$ for each $\lambda$ implies that $\operatorname{{\mathop{\rm End}\nolimits}}_{K_{\lambda}[G_{\mathbb{Q}}]}(M_{f,\lambda})=K_{\lambda}$, so $H^{0}({\mathbb{Q}},A_{f,\lambda})=0$, and since $M_{f,\lambda}$ is not isomorphic to $M_{f,\lambda}(1)$, we also have $H^{0}({\mathbb{Q}},B_{f,\lambda})=0$. It follows then from [F-P, II.2.2.2] (see also [Fl1, Cor. 1.5]) that
$\dim_{K_{\lambda}}H^{1}_{\bf f}({\mathbb{Q}},A_{f,\lambda})=\dim_{K_{\lambda}}H^{1}_{\bf f}({\mathbb{Q}},B_{f,\lambda})$
for all $\lambda$ and hence that Theorem 8.2 is implied by the vanishing of $H^{1}_{\bf f}({\mathbb{Q}},A_{f,\lambda})$. Theorem 7.15 shows that
$H^{1}_{\bf f}({\mathbb{Q}},A_{f,\lambda}/{\mathop{\rm ad}\nolimits}^{0}_{{\mathcal{O}}_{K,\lambda}}{\mathcal{M}}_{f,\lambda})\subset H^{1}_{\Sigma}({\mathbb{Q}},A_{f,\lambda}/{\mathop{\rm ad}\nolimits}^{0}_{{\mathcal{O}}_{K,\lambda}}{\mathcal{M}}_{f,\lambda})$
is finite for $\lambda$ in $S_{f}$. Since the kernel of
$H^{1}_{\bf f}({\mathbb{Q}},A_{f,\lambda})\to H^{1}_{\bf f}({\mathbb{Q}},A_{f,\lambda}/{\mathop{\rm ad}\nolimits}^{0}_{{\mathcal{O}}_{K,\lambda}}{\mathcal{M}}_{f,\lambda})$
is finitely generated over ${\mathcal{O}}_{\lambda}$ we deduce $H^{1}_{\bf f}({\mathbb{Q}},A_{f,\lambda})=0$ and Theorem 8.2 follows. ∎
8.2. Deligne’s period
We now recall the formulation in [F-P] of Deligne’s conjecture [De3] for the “transcendental part” of $L(M,0)$ for $M=A_{f}$ or $B_{f}$. The authors there actually discuss the more general conjecture of Beilinson concerning the leading coefficient $L^{*}(M,0)$ for premotivic structures arising from motives, but their formulation relies on the conjectural existence of a category of mixed motives with certain properties. We restrict our attention to those $M$, such as $A_{f}$ and $B_{f}$, for which $L(M,0)\neq 0$ and which are critical in the sense of Deligne. In that case Beilinson’s conjecture reduces (conjecturally) to Deligne’s, which can be stated without reference to the category of mixed motives.
Under these hypotheses, the fundamental line for $M$ is the $K$-line defined by
$\Delta_{\bf f}(M)={\mathop{\rm Hom}\nolimits}_{K}({\mathop{\rm det}\nolimits}_{K}M_{B}^{+},{\mathop{\rm det}\nolimits}_{K}t_{M})$
where +{\ $^{+}$ indicates the subspace fixed by $F_{\infty}$ and $t_{M}=M_{\mathop{\rm dR}\nolimits}/\operatorname{{\mathop{\rm Fil}\nolimits}}^{0}M_{\mathop{\rm dR}\nolimits}$. Furthermore the composite
${\mathbb{R}}\otimes M_{B}^{+}\to({\mathbb{C}}\otimes M_{B})^{+}\stackrel{{\scriptstyle(I^{\infty})^{-1}}}{{\longrightarrow}}{\mathbb{R}}\otimes M_{{\mathop{\rm dR}\nolimits}}\to{\mathbb{R}}\otimes t_{M}$
is an ${\mathbb{R}}\otimes K$-linear isomorphism. Its determinant over ${\mathbb{R}}\otimes K$ defines a basis for ${\mathbb{R}}\otimes\Delta_{\bf f}(M)$ called the Deligne period, denoted $c^{+}(M)$.
Conjecture 8.3.
There exists a basis $b(M)$ for $\Delta_{\bf f}(M)$ such that
$L(M,0)(1\otimes b(M))=c^{+}(M).$
For example, the conjecture holds for $M={\mathbb{Q}}(2)$ with the basis $b({\mathbb{Q}}(2))$ of
$\Delta_{\bf f}({\mathbb{Q}}(2))={\mathop{\rm Hom}\nolimits}_{\mathbb{Q}}({\mathbb{Q}}(2\pi i)^{2},{\mathbb{Q}}[2])$
defined by $(2\pi i)^{2}\mapsto-24\iota^{-2}$. There are various rationality results for $L(A_{f},0)$ and $L(B_{f},0)$ in the literature (see for example [Schm, Th. 2.3]) although the precise relationship with Conjecture 8.3 for $M=A_{f}$ or $B_{f}$ is not always clear. In this section we recall the proof of Conjecture 8.3 for $M=A_{f}$ and $B_{f}$ and give convenient natural descriptions for $b(A_{f})$ and $b(B_{f})$.
We begin by observing that $A_{f,B}^{+}$ and $t_{A_{f}}$ are one-dimensional over $K$. Furthermore, complex conjugation
$F_{\infty}:M_{f,B}\to M_{f,B}$
has trace zero and commutes with $F_{\infty}$, so it is a basis for $A_{f,B}^{+}$. Note also that the natural map
$A_{f,{\mathop{\rm dR}\nolimits}}\to{\mathop{\rm Hom}\nolimits}_{K}(\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}M_{f,{\mathop{\rm dR}\nolimits}},M_{f,{\mathop{\rm dR}\nolimits}}/\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}M_{f,{\mathop{\rm dR}\nolimits}})$
factors through an isomorphism
$t_{A_{f}}\to{\mathop{\rm Hom}\nolimits}_{K}(\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}M_{f,{\mathop{\rm dR}\nolimits}},M_{f,{\mathop{\rm dR}\nolimits}}/\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}M_{f,{\mathop{\rm dR}\nolimits}}).$
The fundamental line $\Delta_{\bf f}(A_{f})$ can therefore be identified with
${\mathop{\rm Hom}\nolimits}_{K}(\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}M_{f,{\mathop{\rm dR}\nolimits}}\otimes{\mathbb{Q}}\cdot F_{\infty},M_{f,{\mathop{\rm dR}\nolimits}}/\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}M_{f,{\mathop{\rm dR}\nolimits}}).$
We shall describe $b(A_{f})$ by specifying the image of the canonical basis $f\otimes F_{\infty}$ for $\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}M_{f,{\mathop{\rm dR}\nolimits}}\otimes{\mathbb{Q}}\cdot F_{\infty}$ where we view $f$ as an element of $M_{f,{\mathop{\rm dR}\nolimits}}$ by Lemma 5.7. Recall that we defined in (34) a perfect alternating pairing
$\langle\cdot,\cdot\rangle\ :\ M_{f}\otimes_{K}M_{f}\to M_{\psi}(1-k),$
and this induces an isomorphism
$M_{f,{\mathop{\rm dR}\nolimits}}/\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}M_{f,{\mathop{\rm dR}\nolimits}}\to{\mathop{\rm Hom}\nolimits}_{K}(\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}M_{f,{\mathop{\rm dR}\nolimits}},M_{\psi}(1-k)_{\mathop{\rm dR}\nolimits}).$
We shall eventually define $b(A_{f})$ by specifying the element $\langle f,b(A_{f})(f\otimes F_{\infty})\rangle$ of $M_{\psi}(1-k)_{\mathop{\rm dR}\nolimits}$.
We can make a similar analysis of the fundamental line $\Delta_{\bf f}(B_{f})$. One finds that $B_{f,B}^{+}$ and $t_{B_{f}}$ are two-dimensional over $K$. Note that $B_{f,B}^{+}$ can be identified with $A_{f,B}^{-}\otimes{\mathbb{Q}}(1)_{B}$ and that the natural map
$A_{f,B}^{-}\to{\mathop{\rm Hom}\nolimits}_{K}(M_{f,B}^{+},M_{f,B}^{-})\oplus{\mathop{\rm Hom}\nolimits}_{K}(M_{f,B}^{-},M_{f,B}^{+})$
defined by restrictions is an isomorphism. We therefore have an isomorphism
${\mathop{\rm det}\nolimits}_{K}B_{f,B}^{+}\to K(2)_{B}$
which is canonical up to sign. To fix the choice of sign, we use $\alpha\wedge\alpha^{-1}$ as a basis for ${\mathop{\rm det}\nolimits}_{K}A_{f,B}^{-}$ where $\alpha:M_{f,B}^{+}\to M_{f,B}^{-}$ is any $K$-linear isomorphism. Next note that the natural map
$B_{f,{\mathop{\rm dR}\nolimits}}\to{\mathop{\rm Hom}\nolimits}_{K}(M_{f,{\mathop{\rm dR}\nolimits}},M_{f}(1)_{\mathop{\rm dR}\nolimits})\to{\mathop{\rm Hom}\nolimits}_{K}(\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}M_{f,{\mathop{\rm dR}\nolimits}},M_{f}(1)_{\mathop{\rm dR}\nolimits})$
factors through an isomorphism
$t_{B_{f}}\to{\mathop{\rm Hom}\nolimits}_{K}(\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}M_{f,{\mathop{\rm dR}\nolimits}},M_{f}(1)_{\mathop{\rm dR}\nolimits}).$
Using the isomorphism
${\mathop{\rm det}\nolimits}_{K}M_{f,{\mathop{\rm dR}\nolimits}}\to\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}M_{f,{\mathop{\rm dR}\nolimits}}\otimes_{K}(M_{f,{\mathop{\rm dR}\nolimits}}/\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}M_{f,{\mathop{\rm dR}\nolimits}})$
(with choice of sign again indicated by the ordering), we find that ${\mathop{\rm det}\nolimits}_{K}t_{B_{f}}$ is naturally isomorphic to
(50) ${\mathop{\rm Hom}\nolimits}_{K}(\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}M_{f,{\mathop{\rm dR}\nolimits}},M_{f,{\mathop{\rm dR}\nolimits}}/\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}M_{f,{\mathop{\rm dR}\nolimits}})\otimes{\mathbb{Q}}(2)_{\mathop{\rm dR}\nolimits}.$
We can therefore identify $\Delta_{\bf f}(B_{f})$ with
${\mathop{\rm Hom}\nolimits}_{K}(\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}M_{f,{\mathop{\rm dR}\nolimits}}\otimes{\mathbb{Q}}(2)_{B},(M_{f,{\mathop{\rm dR}\nolimits}}/\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}M_{f,{\mathop{\rm dR}\nolimits}})\otimes{\mathbb{Q}}(2)_{\mathop{\rm dR}\nolimits}),$
and we arrive at a canonical isomorphism
$\Delta_{\bf f}(A_{f})\otimes\Delta_{\bf f}({\mathbb{Q}}(2))\otimes A_{f,B}^{+}\xrightarrow{\sim}\Delta_{\bf f}(B_{f}).$
Fixing the basis $F_{\infty}$ of $A_{f,B}^{+}$ and the basis $\beta$ of $\Delta_{\bf f}({\mathbb{Q}}(2))$ which sends $(2\pi i)^{2}$ to $\iota^{-2}$ this defines an isomorphism of $K$-lines
(51) ${\rm tw}:\Delta_{\bf f}(A_{f})\rightarrow\Delta_{\bf f}(B_{f})$
so that ${\rm tw}(\phi)(x\otimes y)=\phi(x\otimes F_{\infty})\otimes\beta(y)$.
Lemma 8.4.
We have
$({\mathbb{R}}\otimes{\rm tw})(c^{+}(A_{f}))=-\frac{1}{2\pi^{2}}c^{+}(B_{f}).$
Proof.
Let $I^{\infty}_{M}:{\mathbb{C}}\otimes M_{f,{\mathop{\rm dR}\nolimits}}\cong{\mathbb{C}}\otimes M_{f,B}$ be the comparison isomorphism for $M_{f}$. Via the natural isomorphism ${\mathbb{C}}\otimes\text{End}_{K}(M_{f})_{?}\cong\text{End}_{{\mathbb{C}}\otimes K}({\mathbb{C}}\otimes M_{f,?})$ where ?=B?=B or $\mathop{\rm dR}\nolimits$}, $I^{\infty}_{M}$ induces the comparison isomorphism $I^{\infty}$ for both $\text{End}(M_{f})$ and $A_{f}$: $I^{\infty}(\phi)=I_{M}^{\infty}\circ\phi\circ(I_{M}^{\infty})^{-1}$. A similar formula holds for $c^{+}(A_{f})$.
Suppose now that $x$ is a non-zero element of ${\mathbb{R}}\otimes\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}M_{f,{\mathop{\rm dR}\nolimits}}$ (for example, take $x=g$), and write $I_{M}^{\infty}(x)=y^{+}+y^{-}$ with $y^{\pm}\in{\mathbb{C}}\otimes M_{f,B}^{\pm}$. Then
$c^{+}(A_{f})(x\otimes F_{\infty})=(I_{M}^{\infty})^{-1}(1\otimes F_{\infty})I_{M}^{\infty}(x)=(I^{\infty}_{M})^{-1}(y^{+}-y^{-})\mod{\mathbb{R}}\otimes\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}M_{f,{\mathop{\rm dR}\nolimits}}.$
Chasing through the above isomorphisms, one finds that $c^{+}(B_{f})$ is characterized by
$x\otimes c^{+}(B_{f})(x\otimes(2\pi i)^{2})\otimes\iota^{2}$ $=(2\pi i)^{2}(I^{\infty}_{M})^{-1}(y^{-})\wedge(I^{\infty}_{M})^{-1}(y^{+})$
$=x\otimes\frac{1}{2}(2\pi i)^{2}c^{+}(A_{f})(x\otimes F_{\infty}).$
Recall that $\Sigma_{e}(f)$ is the set of primes $p$ such that $L_{p}^{\mathop{\rm nv}\nolimits}(A_{f},s)=1$ but $L_{p}(A_{f},s)=(1+p^{-s})^{-1}$. We write $b_{{\mathop{\rm dR}\nolimits}}$ for the basis of $M_{\psi,{\mathop{\rm dR}\nolimits}}$ defined in §1.3, and set $\eta=0$ or 11 so $\eta\equiv k\bmod 2$. Note that by Proposition 5.5 of [De3], we have $\epsilon(M_{f}\otimes M_{\psi^{-1}})/\epsilon(M_{\psi^{-1}})\in K^{\times}$. The same proposition together with Lemma 6.1 gives $\epsilon(A_{f})\in K^{\times}$.
Theorem 8.5.
Let $b(A_{f})\in\Delta_{\bf f}(A_{f})$ be defined by the formula
$\langle f,b(A_{f})(f\otimes F_{\infty})\rangle=\frac{i^{k-\eta}((k-2)!)^{2}\epsilon(M_{f}\otimes M_{\psi^{-1}})}{2\epsilon(M_{\psi^{-1}})\epsilon(A_{f})}\prod_{p\in\Sigma_{e}(f)}(1+p^{-1})\cdot(b_{\mathop{\rm dR}\nolimits}\otimes\iota^{k-1}),$
and $b(B_{f})\in\Delta_{\bf f}(B_{f})$ by the formula
(52) $b(B_{f})=(1-k)\epsilon(A_{f}){\rm tw}(b(A_{f})).$
Then $L(A_{f},0)(1\otimes b(A_{f}))=c^{+}(A_{f})$ and $L(B_{f},0)(1\otimes b(B_{f}))=c^{+}(B_{f})$.
Proof.
If we show
$\langle f,c^{+}(A_{f})(f\otimes F_{\infty})\rangle=\frac{i^{k-\eta}(k-1)!(k-2)!\epsilon(M_{f}\otimes M_{\psi^{-1}})L^{\mathop{\rm nv}\nolimits}(A_{f},1)}{4\pi^{2}\epsilon(M_{\psi^{-1}})}\cdot(b_{\mathop{\rm dR}\nolimits}\otimes\iota^{k-1})$
in ${\mathbb{C}}\otimes M_{\psi}(1-k)_{\mathop{\rm dR}\nolimits}$, then the statement concerning $b(A_{f})$ is an immediate consequence of the functional equation (49). The identity $L(B_{f},0)(1\otimes b(B_{f}))=c^{+}(B_{f})$ then follows by applying $({\mathbb{R}}\otimes{\rm tw})$ to the identity $L(A_{f},0)(1\otimes b(A_{f}))=c^{+}(A_{f})$ and using (49) and Lemma 8.4.
As in §4.5 put $U=U_{0}(N)$, let $\sigma:U\rightarrow K^{\times}$ be the representation $\left(\begin{smallmatrix}a&b\\ c&d\end{smallmatrix}\right)\mapsto\psi^{-1}(a_{N})$, and choose a multiple $N^{\prime}$ of NN so that $U_{N^{\prime}}\subseteq U$ is sufficiently small. Put $w=\left(\begin{smallmatrix}0&-1\\ N&0\end{smallmatrix}\right)\in GL_{2}({\mathbb{A}}_{\bf f})$ and denote by $W=[UwU]_{\omega}:M(\sigma)_{N^{\prime}}\rightarrow M(\hat{\sigma}\otimes(\psi^{-1}\circ{\mathop{\rm det}\nolimits}))_{N^{\prime}}$ the isomorphism in Lemma 5.2. Note that $ww_{N}^{-1}\in U$ so we can work with $w$ instead of $w_{N}$. For any one-dimensional $K$-representation $\sigma$ of $U$ we shall view $M(\sigma)_{N^{\prime}}$ as a sub-PM-structure of $K\otimes M_{N^{\prime}}$. With $I^{\infty}$ denoting the comparison isomorphism for both $M_{f}$ and $K\otimes M_{N^{\prime}}$ we have
$\begin{array}[]{ll}\langle f,c^{+}(A_{f})(f\otimes F_{\infty})\rangle&=\langle f,(I^{\infty})^{-1}(1\otimes F_{\infty})I^{\infty}f\rangle\\ &=[f,(I^{\infty})^{-1}(1\otimes F_{\infty})I^{\infty}Wf]_{N^{\prime}}\\ &=[U:U_{N^{\prime}}]^{-1}(f,(I^{\infty})^{-1}(1\otimes F_{\infty})I^{\infty}Wf)_{N^{\prime}}.\end{array}$
We proceed with the computation of $Wf\in\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}M_{!}(\hat{\sigma}\otimes(\psi^{-1}\circ{\mathop{\rm det}\nolimits}))_{N^{\prime},{\mathop{\rm dR}\nolimits}}$. Note that the field $K_{f}$ generated by the Fourier coefficients of the newform $f$ is either totally real or a CM field and hence has a well defined automorphism $\rho$ induced by complex conjugation. It is known that the Fourier expansion $f^{\rho}(z)=\sum_{n=1}^{\infty}a_{n}^{\rho}e^{2\pi izn}$ is a newform of conductor NN and character $\psi^{-1}$ [Miy, 4.6.15(2)], hence represents an element of $\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}M_{!}(\hat{\sigma})_{N^{\prime},{\mathop{\rm dR}\nolimits}}$. Recall that $b_{\mathop{\rm dR}\nolimits}$ is the canonical basis of $M_{\psi,{\mathop{\rm dR}\nolimits}}\cong(H^{0}(X_{N^{\prime}})\otimes K_{\psi^{-1}})^{U}_{\mathop{\rm dR}\nolimits}$ (Lemma 4.6) and that $M_{\psi}\otimes_{K}M_{f^{\rho}}$ has a natural map into $M_{!}(\hat{\sigma}\otimes(\psi^{-1}\circ{\mathop{\rm det}\nolimits}))_{N^{\prime}}$ via the cup product on $X_{N^{\prime}}$ (see (26)).
Lemma 8.6.
We have
(53) $Wf=\psi(-1)\frac{i^{k-\eta}\epsilon(M_{f}\otimes M_{\psi^{-1}})}{N\epsilon(M_{\psi^{-1}})}b_{\mathop{\rm dR}\nolimits}\cup f^{\rho}.$
Proof.
We fix an embedding $\tau:K\rightarrow{\mathbb{C}}$ and compute the images of both sides in $S_{k}(U_{N^{\prime}})$ (we shall suppress $\tau$ in the notation and view all elements of $K$ as complex numbers via $\tau$). Let $\phi\in(S_{k}(U_{N^{\prime}})\otimes_{\mathbb{C}}{\mathbb{C}}_{\sigma})^{U}$ denote the element corresponding to $f$. We then have $\phi(xu)=\sigma^{-1}(u)\phi(x)$ for all $u\in U$ and $\beta(\phi)_{t}(z)=f(z)$ for all $t\in({\mathbb{Z}}/N^{\prime}{\mathbb{Z}})^{\times}$. Recall that the isomorphism $\beta:S_{k}(U_{N^{\prime}})\cong\bigoplus_{t\in({\mathbb{Z}}/N^{\prime}{\mathbb{Z}})^{\times}}S_{k}(\Gamma(N^{\prime}))$ was defined in (12) by .
$\beta(F)_{t}(\gamma(i)):=({\mathop{\rm det}\nolimits}\gamma)^{-1}j(\gamma,i)^{k}F(g_{t}\gamma(i))).$
for $\gamma\in GL_{2}({\mathbb{R}})^{+}$, $j(\left(\begin{smallmatrix}a&b\\ c&d\end{smallmatrix}\right),z)=cz+d$ and $g_{t}\equiv\left(\begin{smallmatrix}1&0\\ 0&t^{-1}\end{smallmatrix}\right)\mod N^{\prime}$. Note that since $g_{t}\in U$ we indeed have that
$\beta(\phi)_{t}(z)=$ $({\mathop{\rm det}\nolimits}\gamma)^{-1}j(\gamma,i)^{k}\phi(g_{t}\gamma(i))$
$=$ $({\mathop{\rm det}\nolimits}\gamma)^{-1}j(\gamma,i)^{k}\sigma^{-1}(g_{t})\phi(\gamma(i))$
$=$ $({\mathop{\rm det}\nolimits}\gamma)^{-1}j(\gamma,i)^{k}\phi(\gamma(i))$
$=$ $f(z)$
is independent of $t$.
Note that the element $e^{2\pi i/N^{\prime}}$ of $F={\mathbb{Q}}(e^{2\pi i/N^{\prime}})$ maps to $e^{-2\pi i/N^{\prime}}$ under the isomorphism $M_{F}\cong H^{0}(X_{N^{\prime}})$ of 4.5. Hence $b_{\mathop{\rm dR}\nolimits}=\sum_{a\in({\mathbb{Z}}/N{\mathbb{Z}})^{\times}}\psi(a)\otimes e^{2\pi ia/N}\in{\mathbb{C}}\otimes{\mathbb{Q}}(e^{2\pi i/N^{\prime}})$, when viewed as an element of $H^{0}_{\mathop{\rm dR}\nolimits}(X_{N^{\prime}})$, i.e. a locally constant function, is given by
$t\mapsto\sum_{a\in({\mathbb{Z}}/N{\mathbb{Z}})^{\times}}\psi(a)e^{-2\pi iat/N}$ $=\psi(-t)^{-1}\sum_{a\in({\mathbb{Z}}/N{\mathbb{Z}})^{\times}}\psi(a)e^{2\pi ia/N}$
$=\psi(-t)^{-1}G_{\psi}$
$=\psi(-t)^{-1}i^{\eta}\epsilon(M_{\psi^{-1}},\tau).$
If now $\phi^{\rho}\in S_{k}(U_{N^{\prime}})$ corresponds to $f^{\rho}$ then $\beta(\phi^{\rho})_{t}(z)=f^{\rho}(z)$ is again independent of $t$ and the right hand side of (53) is given by
(54) $t\mapsto\frac{i^{k}\epsilon(M_{f}\otimes M_{\psi^{-1}},\tau)}{N}\psi(t)^{-1}f^{\rho}.$
The perfect pairing $M_{f}\otimes_{K}M_{f}\rightarrow M_{\psi}(1-k)$ and the identity of Hecke eigenvalues [Miy, (4.6.17)] induce an isomorphism $M_{f}^{*}\cong M_{f}\otimes_{K}M_{\psi^{-1}}(k-1)\cong M_{f^{\rho}}(k-1)$ so that the functional equation for $\Lambda(M_{f}\otimes M_{\psi^{-1}},\tau,s)$ can be written
(55) $\Lambda(M_{f}\otimes M_{\psi^{-1}},\tau,s)=\epsilon(M_{f}\otimes M_{\psi^{-1}},\tau)N^{-s}\Lambda(M_{f^{\rho}}\otimes M_{\psi},\tau,k-s).$
Recall that the definition $(g|_{k}\gamma)(z)={\mathop{\rm det}\nolimits}(\gamma)^{k/2}j(\gamma,z)^{-k}g(\gamma(z))$ for $\gamma\in GL_{2}({\mathbb{R}})^{+}$ defines a right action of $GL_{2}({\mathbb{R}})^{+}$ on functions $g:{\mathfrak{H}}\rightarrow{\mathbb{C}}$. Put $W_{N}=\left(\begin{smallmatrix}0&-1\\ N&0\end{smallmatrix}\right)$. By [Miy, Th. 4.3.6] we have
$\Lambda(M_{f}\otimes M_{\psi^{-1}},\tau,s)=\Lambda(f,s)=i^{k}N^{-s+k/2}\Lambda(f|_{k}W_{N},k-s)$
which together with (55) yields
$f^{\rho}=\epsilon(M_{f}\otimes M_{\psi^{-1}},\tau)^{-1}i^{k}N^{k/2}f|_{k}W_{N}.$
Hence (54) becomes
(56) $t\mapsto(-1)^{k}N^{k/2-1}\psi(t)^{-1}f|_{k}W_{N}.$
Turning to the left hand side of (53) we have $(W\phi)(x):=\phi(xw)$ and $\phi(wh)=\phi(W_{\mathbb{Q}}W_{N}^{-1}h)=\phi(W_{N}^{-1}h)$ where $h\in GL_{2}({\mathbb{R}})^{+}$ and $W_{\mathbb{Q}}\in GL_{2}({\mathbb{Q}})$ is the matrix with image $w$ (resp. $W_{N}$) in $GL_{2}({\mathbb{A}}_{\bf f})$ (resp. $GL_{2}({\mathbb{R}})$). For $\gamma\in GL_{2}({\mathbb{R}})^{+}$ we have
${\mathop{\rm det}\nolimits}(h)^{-1}j(h,i)^{k}\phi(\gamma h)=$ ${\mathop{\rm det}\nolimits}(\gamma){\mathop{\rm det}\nolimits}(\gamma h)^{-1}j(\gamma,h(i))^{-k}j(\gamma h,i)^{k}\phi(\gamma h)$
$=$ ${\mathop{\rm det}\nolimits}(\gamma)j(\gamma,h(i))^{-k}f(\gamma h(i))$
$=$ ${\mathop{\rm det}\nolimits}(\gamma)^{1-k/2}(f|_{k}\gamma)(h(i)).$
Combining these equations we find that $W\phi$ corresponds to
$t\mapsto{\mathop{\rm det}\nolimits}(h)^{-1}j(h,i)^{k}(W\phi)(g_{t}h)=$ ${\mathop{\rm det}\nolimits}(h)^{-1}j(h,i)^{k}\phi(ww^{-1}g_{t}wh)$
$=$ ${\mathop{\rm det}\nolimits}(h)^{-1}j(h,i)^{k}\sigma^{-1}(w^{-1}g_{t}w)\phi(wh)$
$=$ ${\mathop{\rm det}\nolimits}(h)^{-1}j(h,i)^{k}\sigma^{-1}(w^{-1}g_{t}w)\phi(W_{N}^{-1}h)$
$=$ $\sigma^{-1}(w^{-1}g_{t}w){\mathop{\rm det}\nolimits}(W_{N}^{-1})^{1-k/2}(f|_{k}W_{N}^{-1})(h(i)).$
Since $f|_{k}W_{N}^{2}=(-1)^{k}f$ this last expression equals
(57) $\sigma^{-1}(w^{-1}g_{t}w)(-1)^{k}N^{k/2-1}(f|_{k}W_{N})(h(i)).$
For $g_{t}\equiv\left(\begin{smallmatrix}1&0\\ 0&t^{-1}\end{smallmatrix}\right)\mod N$ we have $w^{-1}g_{t}w\equiv\left(\begin{smallmatrix}t^{-1}&*\\ 0&1\end{smallmatrix}\right)\mod N$ and $\sigma^{-1}(w^{-1}g_{t}w)=\psi(t^{-1})=\psi(t)^{-1}$. So (56) and (57) agree which finishes the proof of the Lemma. ∎
The definition (28) of the pairing on $\sigma$-constructions shows that $(x,\alpha\cup y)_{N^{\prime}}=(x,y)\otimes_{K}\alpha$ where $\alpha\in M_{\psi}$ and $(x,y)$ is the $K$-linear extension of the ${\mathbb{Q}}(1-k)$-valued pairing on ${\mathcal{M}}_{N^{\prime}}$ defined in §2.4. Combining this with Lemma 53 the last term in (8.2) equals
(58) $U:U_{N^{\prime}}]^{-1}(f,(I^{\infty})^{-1}(1\otimes F_{\infty})I^{\infty}f^{\rho})\otimes_{K}\alpha_{{\mathop{\rm dR}\nolimits}}$
in ${\mathbb{C}}\otimes K(1-r)_{{\mathop{\rm dR}\nolimits}}\otimes_{K}M_{\psi,{\mathop{\rm dR}\nolimits}}$ where
$\alpha_{\mathop{\rm dR}\nolimits}$ $=\psi(-1)\frac{i^{k-\eta}\epsilon(M_{f}\otimes M_{\psi^{-1}})}{N\epsilon(M_{\psi^{-1}})}(I^{\infty})^{-1}(1\otimes F_{\infty})I^{\infty}b_{\mathop{\rm dR}\nolimits}$
$=\psi(-1)\frac{i^{k-\eta}\epsilon(M_{f}\otimes M_{\psi^{-1}})}{N\epsilon(M_{\psi^{-1}})}\psi(-1)^{-1}b_{\mathop{\rm dR}\nolimits}$
(with $I^{\infty}$ also denoting the comparison isomorphism for $M_{\psi}$). For any premotivic structure we have $(F_{\infty}\otimes F_{\infty})I^{\infty}=I^{\infty}(F_{\infty}\otimes 1)$ and we have $(F_{\infty}\otimes 1)(f^{\rho})=f^{\rho}$ since $f^{\rho}\in K\otimes M_{N^{\prime},{\mathop{\rm dR}\nolimits}}\subset{\mathbb{C}}\otimes K\otimes M_{N^{\prime},{\mathop{\rm dR}\nolimits}}$. Hence
$(I^{\infty})^{-1}(1\otimes F_{\infty})I^{\infty}f^{\rho}=(I^{\infty})^{-1}(F_{\infty}\otimes 1)I^{\infty}f^{\rho}.$
Under the natural isomorphism ${\mathbb{C}}\otimes K\otimes M_{N^{\prime},B}\cong({\mathbb{C}}\otimes M_{N^{\prime},B})^{{\bf I}_{K}}$ the action of $F_{\infty}\otimes 1\otimes 1$ on the left hand side gets transformed into the action sending $(x_{\tau})$ to $\tau\mapsto(F_{\infty,\tau}\otimes 1)(x_{\overline{\tau}})$ where $F_{\infty,\tau}$ is complex conjugation acting on $\mathbb{C}$} in the factor indexed by $\tau$. Hence the $\tau$-component of (58) equals
$[U:U_{N^{\prime}}]^{-1}(\tau(f),(I^{\infty})^{-1}(F_{\infty,\tau}\otimes 1)I^{\infty}\overline{\tau}(f^{\rho}))\otimes_{\mathbb{C}}\tau(\alpha_{{\mathop{\rm dR}\nolimits}})$
$=$ $[U:U_{N^{\prime}}]^{-1}(k-2)!(4\pi)^{k-1}\phi(N^{\prime})(\tau(f),\tau(f))_{\Gamma(N^{\prime})}\,\tau(\alpha_{\mathop{\rm dR}\nolimits})\otimes\iota^{k-1}$
where $\phi$ is Euler’s function and we have used Lemma $2.$ Therefore (58) equals
(59) $\begin{array}[]{l}[U:U_{N^{\prime}}]^{-1}(k-2)!(4\pi)^{k-1}\phi(N^{\prime})(f,f)_{\Gamma(N^{\prime})}\cdot\alpha_{\mathop{\rm dR}\nolimits}\otimes\iota^{k-1}\\ =\frac{[\bar{\Gamma}_{1}(N):\bar{\Gamma}(N^{\prime})]}{[U:U_{N^{\prime}}]}\phi(N^{\prime})(k-2)!(4\pi)^{k-1}(f,f)_{\Gamma_{1}(N)}\cdot\alpha_{\mathop{\rm dR}\nolimits}\otimes\iota^{k-1}\end{array}$
in ${\mathbb{C}}\otimes M_{\psi}(1-k)_{\mathop{\rm dR}\nolimits}$ where $\bar{\Gamma}_{1}(N):\bar{\Gamma}(N^{\prime})]$ is the degree of the covering $\Gamma(N^{\prime})\backslash{\mathfrak{H}}\rightarrow\Gamma_{1}(N)\backslash{\mathfrak{H}}$. Since the maps ${\mathop{\rm det}\nolimits}:U\rightarrow({\mathbb{Z}}/N^{\prime}{\mathbb{Z}})^{\times}$ and $\text{SL}_{2}({\mathbb{Z}})\rightarrow\text{SL}_{2}({\mathbb{Z}}/N^{\prime}{\mathbb{Z}})$ are surjective one finds
(60) $\begin{array}[]{ll}[U:U_{N^{\prime}}]&=\phi(N^{\prime})[\text{SL}_{2}({\mathbb{Z}})\cap U:\text{SL}_{2}({\mathbb{Z}})\cap U_{N^{\prime}}]\\ &=\phi(N^{\prime})[\Gamma_{0}(N):\Gamma(N^{\prime})]\\ &=\phi(N^{\prime})\phi(N)[\Gamma_{1}(N):\Gamma(N^{\prime})]\\ &=\phi(N^{\prime})\phi(N)\delta(N)[\bar{\Gamma}_{1}(N):\bar{\Gamma}(N^{\prime})]\end{array}$
where $\delta(N)=1$ if $N>2$ and $\delta(N)=2$ if $N\leq 2$ (note that $1\in\Gamma_{1}(N)$ iff $N\leq 2$ whereas $1\notin\Gamma(N^{\prime})$). Combining this with Lemma 8.7 below we find that (59) equals
$\frac{(k-2)!(4\pi)^{k-1}}{\phi(N)\delta(N)}\cdot\frac{(k-1)!\delta(N)N\phi(N)L^{\mathop{\rm nv}\nolimits}(A_{f},1)}{4^{k}\pi^{k+1}}\cdot\frac{i^{k-\eta}\epsilon(M_{f}\otimes M_{\psi^{-1}})}{N\epsilon(M_{\psi^{-1}})}\cdot b_{\mathop{\rm dR}\nolimits}\otimes\iota^{k-1}$
$=\frac{i^{k-\eta}(k-2)!(k-1)!L^{\mathop{\rm nv}\nolimits}(A_{f},1)\epsilon(M_{f}\otimes M_{\psi^{-1}})}{4\pi^{2}\epsilon(M_{\psi^{-1}})}\cdot b_{\mathop{\rm dR}\nolimits}\otimes\iota^{k-1}.$
This finishes the proof of Theorem 8.5. ∎
Lemma 8.7.
If $f$ is a newform of conductor NN, weight $k$ and with coefficients in the number field $K$ we have
$(\tau(f),\tau(f))_{\Gamma_{1}(N)}=\frac{(k-1)!\delta(N)N\phi(N)L^{\mathop{\rm nv}\nolimits}(A_{f},\tau,1)}{4^{k}\pi^{k+1}}$
for any embedding $\tau:K\rightarrow{\mathbb{C}}$ and $\delta(N)$ as in (60).
Proof.
We fix $\tau$ and write $f$ for $\tau(f)$ to ease notation. By Theorem 5.1 of [Hi1] (essentially a reformulation of a Theorem of Rankin and Shimura) we have
$L(k,f,\bar{\psi})=\frac{4^{k}\pi^{k+1}(f,f)_{\Gamma_{1}(N)}}{(k-1)!\delta(N)NN_{\psi}\phi(N/N_{\psi})}$
where $L(s,f,\bar{\psi})=\prod_{p}L_{p}(s,f,\bar{\psi})$,
$L_{p}(s,f,\bar{\psi})^{-1}=(1-\bar{\psi}(p)\alpha_{p}^{2}p^{-s})(1-\bar{\psi}(p)\alpha_{p}\beta_{p}p^{-s})(1-\bar{\psi}(p)\beta_{p}^{2}p^{-s})$
and $\alpha_{p},\beta_{p}$ are defined as in §5.5 for $p\nmid N$ and $\alpha_{p}+\beta_{p}=a_{p},\alpha_{p}\beta_{p}=0$ for $p|N$. Denote by $M_{p}$ the exact power of $p$ dividing an integer $M$. To show the Lemma it suffices to show that
(61) $L_{p}(k,f,\bar{\psi})\frac{\phi(N_{p}/N_{\psi,p})}{N_{p}/N_{\psi,p}}=L_{p}^{\mathop{\rm nv}\nolimits}(A_{f},\tau,1)\frac{\phi(N_{p})}{N_{p}}$
for all primes pp. If $p\nmid N$ this is immediate from §$5.$ If $N_{p}=p$ and $N_{\psi,p}=1$ we have $a_{p}^{2}=\psi(p)p^{k-2}$ by [Miy, Th. 4.6.17(2)] and $\pi_{p}(f)$ is special so that (61) holds true by (38). The only other case in which $a_{p}\neq 0$ is when $N_{p}=N_{\psi,p}$ [Miy, Th. 4.6.17]. In this case $\bar{\psi}(p)=0$ and hence $L_{p}(k,f,\bar{\psi})=1$ whereas $\pi_{p}(f)$ is principal series so that $L_{p}^{\mathop{\rm nv}\nolimits}(A_{f},\tau,1)=(1-p^{-1})^{-1}=N_{p}/\phi(N_{p})$ by (38). Finally, if $N_{p}>1$ and $a_{p}=0$ then $L_{p}(k,f,\bar{\psi})=L_{p}^{\mathop{\rm nv}\nolimits}(A_{f},\tau,1)=1$, $N_{p}/N_{\psi,p}>1$ and both sides in (61) equal $(1-p^{-1})$. ∎
Remark. In the following, we shall not need the full precision of Theorem 8.5 but only the fact that $i^{k-\eta}((k-2)!)^{2}\epsilon(M_{f}\otimes M_{\psi^{-1}})/2\epsilon(M_{\psi^{-1}})\epsilon(A_{f})$ is a unit in ${\mathcal{O}}={\mathcal{O}}_{K}[(Nk!)^{-1}]$. This in turn is a consequence of Lemma 8.8 below.
Lemma $8.$
Let $M$ be an object of $\operatorname{{\bf PM}}_{K}$ which is $L$-admissible everywhere and let $\tau:K\rightarrow{\mathbb{C}}$ be an embedding. Then $\epsilon(M,\tau)=\epsilon(M,\tau,0)$ is a unit in $\overline{{\mathbb{Z}}}[c(M)^{-1}]$ where $\overline{{\mathbb{Z}}}$ is the ring of algebraic integers.
Proof.
By definition $\epsilon(M,\tau)=\prod_{p}\epsilon(D_{{\mathop{\rm pst}\nolimits}}(M_{\lambda}|G_{p})\otimes_{K_{\lambda},\tau^{\prime}}{\mathbb{C}},\psi_{p},dx_{p})$ is a product over all places $p$ of $\mathbb{Q}$} where the additive characters $\psi_{p}$ and the Haar measures $dx_{p}$ are chosen as in [De3, 5.3] and $\tau^{\prime}:K_{\lambda}\rightarrow{\mathbb{C}}$ is any extension of $\tau$. The assumption that $M$ is $L$-admissible at $p$ implies that the isomorphism class of $D_{{\mathop{\rm pst}\nolimits}}(M_{\lambda}|G_{p})\otimes_{K_{\lambda},\tau^{\prime}}{\mathbb{C}}$ is independent of $\tau^{\prime}$. The definition of $\epsilon$ in [De2, (8.12)] and [De2, Th. 6.5 (a),(b)] show that
$\epsilon(D_{{\mathop{\rm pst}\nolimits}}(M_{\lambda}|G_{p})\otimes_{K_{\lambda},\tau^{\prime}}{\mathbb{C}},\psi_{p},dx_{p})=\tau^{\prime}\epsilon(D_{{\mathop{\rm pst}\nolimits}}(M_{\lambda}|G_{p}),\psi_{p},dx_{p})\in\tau^{\prime}(K_{\lambda}(\mu_{p^{\infty}})).$
Replacing $\tau^{\prime}$ by $\gamma\tau^{\prime}$, $\gamma\in\text{Aut}({\mathbb{C}}/K(\mu_{p^{\infty}}))$, and using the $L$-admissibility again, we deduce from this formula that $\epsilon(D_{{\mathop{\rm pst}\nolimits}}(M_{\lambda}|G_{p})\otimes_{K_{\lambda},\tau^{\prime}}{\mathbb{C}},\psi_{p},dx_{p})\in K(\mu_{p^{\infty}})$. The remark after [De2, (8.12.4)] shows that $\epsilon$ can be directly expressed in terms of the $\lambda$-adic representation $M_{\lambda}$ for $\lambda\nmid p$. Namely
$\epsilon(D_{{\mathop{\rm pst}\nolimits}}(M_{\lambda}|G_{p}),\psi_{p},dx_{p})=\epsilon_{0}((M_{\lambda}|W_{p})^{ss},\psi_{p},dx_{p}){\mathop{\rm det}\nolimits}(-\operatorname{{\mathop{\rm Frob}\nolimits}}|M_{\lambda}^{I_{p}})^{-1}$
where $\epsilon_{0}$ is introduced in [De2, §5] and $(M_{\lambda}|W_{p})^{ss}$ is the semisimplification of $M_{\lambda}$ as a representation of $W_{p}$. Now for any $\lambda\nmid p$ the $W_{p}$-representation $M_{\lambda}$ is the restriction of a continuous $G_{p}$-representation, hence carries a $W_{p}$-stable ${\mathcal{O}}_{\lambda}$-lattice. This implies, on the one hand, that ${\mathop{\rm det}\nolimits}(-\operatorname{{\mathop{\rm Frob}\nolimits}}|M_{\lambda}^{I_{p}})\in{\mathcal{O}}_{\lambda}^{\times}$ and on the other hand, via [De2, Th. 6.5(c)], that $\epsilon_{0}((M_{\lambda}|W_{p})^{ss},\psi_{p},dx_{p})\in{\mathcal{O}}_{\lambda}[\mu_{p^{\infty}}]^{\times}$. Noting that with our choice of $\psi_{p},dx_{p}$ the epsilon factor equals 1 (resp. a power of ii) for $p\nmid c(M)$ (resp. $p=\infty$) the Lemma follows. ∎
8.3. Bloch-Kato conjecture
We now recall the formulation of the $\lambda$-part of the Bloch-Kato conjecture. We assume that $M$ is a premotivic structure in $\operatorname{{\bf PM}}_{K}$ such that $M$ is critical, $L(M,0)\neq 0$ and Conjecture 8.3 holds. We assume that $\lambda$ is a prime of $K$ such that
(62) $H^{0}({\mathbb{Q}},M_{\lambda})\cong H^{1}_{\bf f}({\mathbb{Q}},M_{\lambda})\cong H^{1}_{\bf f}({\mathbb{Q}},M_{\lambda}^{D})\cong H^{0}({\mathbb{Q}},M_{\lambda}^{D})\cong 0.$
This is conjectured to hold for all $\lambda$ under our hypotheses on $M$ and it implies Conjecture 8.1. If $M=A_{f}$ or $A_{f}(1)$ and $\lambda$ is a prime in $S_{f}$ then (62) holds by Theorem 8.2.
Fontaine and Perrin-Riou ([F-P, II.4]) define an ${\mathcal{O}}_{\lambda}$-lattice $\delta_{{\bf f},\lambda}(M)$ in $K_{\lambda}\otimes_{K}\Delta_{\bf f}(M)$. They assume $K={\mathbb{Q}}$, denote their lattice $\Delta_{S}(T)$ (where $S$ is a finite set of primes and $T$ is a Galois-stable lattice in $M_{\lambda}$) and then prove it is independent of the choice of $S$ and $T$. One checks that the definition and independence argument carry over to arbitrary $K$ by taking determinants relative to ${\mathcal{O}}_{\lambda}$ and $K_{\lambda}$ instead of $\mathbb{Z}$$_{\ell}$ and $\mathbb{Q}$$_{\ell}$. The arguments of [F-P, II.5] carry over as well, giving another description of $\delta_{{\bf f},\lambda}(M)$ for which we need more notation. Choose a Galois stable lattice ${\mathcal{M}}_{\lambda}\subset M_{\lambda}$ and a free rank one ${\mathcal{O}}_{\lambda}$-module $\omega\subset K_{\lambda}\otimes_{K}{\mathop{\rm det}\nolimits}_{K}t_{M}$. We let $\theta({\mathcal{M}}_{\lambda})={\mathop{\rm det}\nolimits}_{{\mathcal{O}}_{\lambda}}{\mathcal{M}}_{\lambda}^{+}$, regarded as a lattice in $K_{\lambda}\otimes_{K}{\mathop{\rm det}\nolimits}_{K}M_{B}^{+}$ via the comparison isomorphism $I_{\lambda}^{B}$. We let ${\mathcal{M}}_{\lambda}^{D}={\mathop{\rm Hom}\nolimits}_{{\mathcal{O}}_{\lambda}}({\mathcal{M}}_{\lambda},{\mathcal{O}}_{\lambda}(1))\subset M_{\lambda}^{D}$. Under our hypotheses, the Tate-Shafarevich group $\mbox{\cyr X}({\mathcal{M}}_{\lambda})$ can be identified with $H^{1}_{\bf f}({\mathbb{Q}},M_{\lambda}/{\mathcal{M}}_{\lambda})$ (which we know is an ${\mathcal{O}}_{\lambda}$-module of finite length, as in the proof of Theorem 8.2). The same holds for ${\mathcal{M}}_{\lambda}^{D}$. Furthermore, by the main result of [Fl2] (also [F-P, II.5.4.2]), $\mbox{\cyr X}({\mathcal{M}}_{\lambda})$ and $\mbox{\cyr X}({\mathcal{M}}_{\lambda}^{D})$ have the same length. In fact, there is an ${\mathcal{O}}_{\lambda}$-linear isomorphism
(63) $\mbox{\cyr X}({\mathcal{M}}_{\lambda}^{D})\cong{\mathop{\rm Hom}\nolimits}_{{\mathbb{Z}}_{\ell}}(\mbox{\cyr X}({\mathcal{M}}_{\lambda}),{\mathbb{Q}}_{\ell}/{\mathbb{Z}}_{\ell}).$
Finally, the Tamagawa ideal of ${\mathcal{M}}_{\lambda}$ relative to $\omega$ is defined as
$\operatorname{{\mathop{\rm Tam}\nolimits}}_{\omega}^{0}({\mathcal{M}}_{\lambda})=\operatorname{{\mathop{\rm Tam}\nolimits}}_{\ell,\omega}^{0}({\mathcal{M}}_{\lambda})\cdot\operatorname{{\mathop{\rm Tam}\nolimits}}_{\infty}^{0}({\mathcal{M}}_{\lambda})\cdot\prod_{p\neq\ell}\operatorname{{\mathop{\rm Tam}\nolimits}}_{p}^{0}({\mathcal{M}}_{\lambda}),$
where the factors are defined as in I.4.1 (and III.5.3.3) of [F-P]. Recall that $\operatorname{{\mathop{\rm Tam}\nolimits}}_{p}^{0}({\mathcal{M}}_{\lambda})=1$ if ${\mathcal{M}}_{\lambda}$ is unramified at $p\neq\ell$ and that
$\operatorname{{\mathop{\rm Tam}\nolimits}}_{\infty}^{0}({\mathcal{M}}_{\lambda})=\operatorname{Fitt}_{{\mathcal{O}}_{\lambda}}H^{1}({\mathbb{R}},{\mathcal{M}}_{\lambda})={\mathcal{O}}_{\lambda}$
if $\ell$ is odd. The argument of [F-P, I.4.2.2] shows that if $p\neq\ell$, then
$\operatorname{{\mathop{\rm Tam}\nolimits}}_{p}^{0}({\mathcal{M}}_{\lambda})=\operatorname{Fitt}_{{\mathcal{O}}_{\lambda}}H^{1}(I_{p},{\mathcal{M}}_{\lambda})^{G_{{\mathbb{Q}}_{p}}}_{{\mathop{\rm tor}\nolimits}}$
from which it is not hard to deduce that
(64) $\operatorname{{\mathop{\rm Tam}\nolimits}}_{p}^{0}({\mathcal{M}}_{\lambda})=\operatorname{{\mathop{\rm Tam}\nolimits}}_{p}^{0}({\mathcal{M}}_{\lambda}^{D}).$
Viewing ${\mathop{\rm Hom}\nolimits}_{{\mathcal{O}}_{\lambda}}(\theta({\mathcal{M}}_{\lambda}),\omega)$ as a lattice in $K_{\lambda}\otimes_{K}\Delta_{\bf f}(M)$, we have by [F-P, Th. II.5.3.6]
(65) $\delta_{{\bf f},\lambda}(M)\!=\!\frac{\operatorname{Fitt}_{{\mathcal{O}}_{\lambda}}\!\!H^{0}({\mathbb{Q}},M_{\lambda}/{\mathcal{M}}_{\lambda})\cdot\operatorname{Fitt}_{{\mathcal{O}}_{\lambda}}\!\!H^{0}({\mathbb{Q}},M_{\lambda}^{D}/{\mathcal{M}}_{\lambda}^{D})}{\operatorname{Fitt}_{{\mathcal{O}}_{\lambda}}\mbox{\cyr X}({\mathcal{M}}_{\lambda}^{D})\cdot\operatorname{{\mathop{\rm Tam}\nolimits}}_{\omega}^{0}({\mathcal{M}}_{\lambda})}{\mathop{\rm Hom}\nolimits}_{{\mathcal{O}}_{\lambda}}(\theta({\mathcal{M}}_{\lambda}),\omega).$
The $\lambda$-part of the Bloch-Kato conjecture can then be formulated as follows:
Conjecture 8.9.
Let $M$ in $\operatorname{{\bf PM}}_{K}$ be critical, $b(M)$ as in Conjecture 8.3, $\lambda$ a place of $K$ such that (62) holds and $\delta_{{\bf f},\lambda}(M)$ as in (65). Then
$\delta_{{\bf f},\lambda}(M)=(1\otimes b(M)){\mathcal{O}}_{\lambda}.$
Theorem 8.10.
Let $f$ be a newform of weight $k\geq 2$ and $S_{f}$ the set of places defined in §7.4. Then Conjecture 8.9 holds for both $M=A_{f}$ and $M=A_{f}(1)$ if $\lambda\in S_{f}$.
Proof.
We first prove a Lemma which relates Conjecture 8.9 to the value $L^{\Sigma}(M,0)$ for a suitable finite set of primes $\Sigma$.
Lemma 8.11.
Suppose $M$, $b(M)$ and $\lambda$ are as in Conjecture 8.9 and $S$ is a set of places of $\mathbb{Q}$} containing $\ell$, $\infty$ and those where $M_{\lambda}$ is ramified. Assume $\Sigma:=S\setminus\{\ell,\infty\}$ is nonempty and $L_{p}(M,0)^{-1}\neq 0$ for all $p\in\Sigma$. Put $b^{\Sigma}(M)=\prod_{p\in\Sigma}L_{p}(M,0)b(M)$. Then Conjecture 8.9 is equivalent to
$\frac{\operatorname{Fitt}_{{\mathcal{O}}_{\lambda}}H^{0}({\mathbb{Q}},M_{\lambda}^{D}/{\mathcal{M}}_{\lambda}^{D})}{\operatorname{Fitt}_{{\mathcal{O}}_{\lambda}}H^{1}_{\Sigma}({\mathbb{Q}},M_{\lambda}^{D}/{\mathcal{M}}_{\lambda}^{D})\operatorname{{\mathop{\rm Tam}\nolimits}}_{\ell,\omega}^{0}({\mathcal{M}}_{\lambda})}{\mathop{\rm Hom}\nolimits}_{{\mathcal{O}}_{\lambda}}(\theta({\mathcal{M}}_{\lambda}),\omega)=(1\otimes b^{\Sigma}(M)){\mathcal{O}}_{\lambda}.$
where $H^{1}_{\Sigma}({\mathbb{Q}},M_{\lambda}^{D}/{\mathcal{M}}_{\lambda}^{D})$ was defined in §7.1.
Proof.
By [Fl1, Prop 1.4] there is a long exact sequence
(66) $0\rightarrow H^{1}_{\emptyset}({\mathbb{Q}},M_{\lambda}^{D}/{\mathcal{M}}_{\lambda}^{D})\rightarrow H^{1}(G_{S},M_{\lambda}^{D}/{\mathcal{M}}_{\lambda}^{D})\rightarrow\bigoplus_{p\in S}\frac{H^{1}({\mathbb{Q}}_{p},M_{\lambda}^{D}/{\mathcal{M}}_{\lambda}^{D})}{H^{1}_{f}({\mathbb{Q}}_{p},M_{\lambda}^{D}/{\mathcal{M}}_{\lambda}^{D})}\xrightarrow{\rho^{*}}\\ H^{1}_{f}({\mathbb{Q}},{\mathcal{M}}_{\lambda})^{*}\rightarrow H^{2}(G_{S},M_{\lambda}^{D}/{\mathcal{M}}_{\lambda}^{D})\rightarrow\bigoplus_{p\in S}H^{2}({\mathbb{Q}}_{p},M_{\lambda}^{D}/{\mathcal{M}}_{\lambda}^{D})\rightarrow H^{0}({\mathbb{Q}},{\mathcal{M}}_{\lambda})^{*}\rightarrow 0$
where $G_{S}$ is the Galois group of the maximal extension of $\mathbb{Q}$} unramified outside $S$ and denotes the Pontryagin dual. From (62) we know that the groups $H^{1}_{\emptyset}({\mathbb{Q}},M_{\lambda}^{D}/{\mathcal{M}}_{\lambda}^{D})$ and $H^{1}_{f}({\mathbb{Q}},{\mathcal{M}}_{\lambda})$ are finite. The map $\rho^{*}$ is Pontryagin dual to the restriction map
$H^{1}_{f}({\mathbb{Q}},{\mathcal{M}}_{\lambda})\xrightarrow{\rho}\bigoplus_{p\in S}H^{1}_{f}({\mathbb{Q}}_{p},{\mathcal{M}}_{\lambda}).$
Clearly $\rho$ is injective as $H^{1}_{f}({\mathbb{Q}},{\mathcal{M}}_{\lambda})\cong H^{0}({\mathbb{Q}},M_{\lambda}/{\mathcal{M}}_{\lambda})$ injects into $H^{0}({\mathbb{Q}}_{p},M_{\lambda}/{\mathcal{M}}_{\lambda})\cong H^{1}_{f}({\mathbb{Q}}_{p},{\mathcal{M}}_{\lambda})_{tors}$ for any $p$, for example some $p\in\Sigma$. This argument also shows that $\rho^{*}$ restricted to
$\bigoplus_{p\in\Sigma\cup\{\infty\}}\frac{H^{1}({\mathbb{Q}}_{p},M_{\lambda}^{D}/{\mathcal{M}}_{\lambda}^{D})}{H^{1}_{f}({\mathbb{Q}}_{p},M_{\lambda}^{D}/{\mathcal{M}}_{\lambda}^{D})}\cong\bigoplus_{p\in\Sigma\cup\{\infty\}}H^{1}_{f}({\mathbb{Q}}_{p},{\mathcal{M}}_{\lambda})^{*}$
is still surjective since the dual map is still injective. Hence we find an exact sequence
(67) $0\rightarrow H^{1}_{\emptyset}({\mathbb{Q}},M_{\lambda}^{D}/{\mathcal{M}}_{\lambda}^{D})\rightarrow H^{1}_{\Sigma}({\mathbb{Q}},M_{\lambda}^{D}/{\mathcal{M}}_{\lambda}^{D})\\ \rightarrow\bigoplus_{p\in\Sigma\cup\{\infty\}}H^{1}_{f}({\mathbb{Q}}_{p},{\mathcal{M}}_{\lambda})^{*}\xrightarrow{\rho^{*}}H^{0}({\mathbb{Q}},M_{\lambda}/{\mathcal{M}}_{\lambda})^{*}\rightarrow 0.$
By [F-P, Proof of I.4.2.2] we have for $p\in\Sigma$
$\operatorname{Fitt}_{{\mathcal{O}}_{\lambda}}H^{1}_{f}({\mathbb{Q}}_{p},{\mathcal{M}}_{\lambda})=L_{p}(M,0)^{-1}\operatorname{{\mathop{\rm Tam}\nolimits}}_{p}^{0}({\mathcal{M}}_{\lambda})$
since $L_{p}(M,0)^{-1}\neq 0$. In particular all terms in (67) are finite ${\mathcal{O}}_{\lambda}$-modules, and the statement of Lemma 8.11 follows easily by taking Fitting ideals in (67), together with (65). ∎
For a prime $\lambda\in S_{f}$, recall that ${\mathcal{A}}_{f,\lambda}={\mathop{\rm ad}\nolimits}^{0}_{{\mathcal{O}}_{\lambda}}{\mathcal{M}}_{f,\lambda}$. Put ${\mathcal{B}}_{f,\lambda}={\mathcal{A}}_{f,\lambda}(1)$. Identifying ${\mathop{\rm det}\nolimits}_{K}t_{A_{f}}$ with ${\mathop{\rm Hom}\nolimits}_{K}(\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}M_{f,{\mathop{\rm dR}\nolimits}},M_{f,{\mathop{\rm dR}\nolimits}}/\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}M_{f,{\mathop{\rm dR}\nolimits}})$, we let
$\omega_{A}=$ ${\mathcal{O}}_{\lambda}\otimes_{\mathcal{O}}{\mathop{\rm Hom}\nolimits}_{\mathcal{O}}(\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}{\mathcal{M}}_{f,{\mathop{\rm dR}\nolimits}},{\mathcal{M}}_{f,{\mathop{\rm dR}\nolimits}}/\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}{\mathcal{M}}_{f,{\mathop{\rm dR}\nolimits}})$
$\cong$ ${\mathop{\rm Hom}\nolimits}_{{\mathcal{O}}_{\lambda}}(\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}{\mathcal{M}}_{f,{\mathop{\rm dR}\nolimits}}\otimes_{\mathcal{O}}{\mathcal{O}}_{\lambda},{\mathcal{M}}_{f,{\mathop{\rm dR}\nolimits}}\otimes_{\mathcal{O}}{\mathcal{O}}_{\lambda}/\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}{\mathcal{M}}_{f,{\mathop{\rm dR}\nolimits}}\otimes_{\mathcal{O}}{\mathcal{O}}_{\lambda}).$
Similarly, identifying ${\mathop{\rm det}\nolimits}_{K}t_{B_{f}}$ with ${\mathop{\rm det}\nolimits}_{K}t_{A_{f}}\otimes{\mathbb{Q}}(2)_{\mathop{\rm dR}\nolimits}$ we define $\omega_{B}$ as $\omega_{A}\otimes\iota^{-2}$.
Proposition 8.12.
We have $\operatorname{{\mathop{\rm Tam}\nolimits}}_{\ell,\omega_{A}}^{0}({\mathcal{A}}_{f,\lambda})=\operatorname{{\mathop{\rm Tam}\nolimits}}_{\ell,\omega_{B}}^{0}({\mathcal{B}}_{f,\lambda})={\mathcal{O}}_{\lambda}$ for $\lambda\in S_{f}$.
Proof.
With the notation in §1.1, we further denote by MF the category of filtered $\phi$-modules as defined in [Fo1, 1.2.1]. We denote by ${\mathcal{O}}_{\lambda}$-$\mathcal{M}\mathcal{F}$}} (resp. $K_{\lambda}$-MF) the category of ${\mathcal{O}}_{\lambda}$- (resp. $K_{\lambda}$-) modules in the abelian category $\mathcal{M}\mathcal{F}$}} (resp. additive category MF) and by $\text{${\mathcal{O}}_{\lambda}$-${{\mathcal{M}\mathcal{F}}}$}^{a}$ (resp. $\text{$K_{\lambda}$-MF}^{a}$) the full subcategories with filtration restrictions as in §$1.$ Scalar extension -$\otimes_{{\mathbb{Z}}_{\ell}}{\mathbb{Q}}_{\ell}$ induces an exact functor $\text{${\mathcal{O}}_{\lambda}$-${{\mathcal{M}\mathcal{F}}}$}\rightarrow\text{$K_{\lambda}$-MF}$ where the notion of exactness in MF is defined in [Fo1, 1.2.3].
Now assume that ${\mathcal{D}}_{1},{\mathcal{D}}_{2}$ are torsion free objects of $\text{${\mathcal{O}}_{\lambda}$-${{\mathcal{M}\mathcal{F}}}$}^{a}$ for some $a$ and put $D_{i}={\mathcal{D}}_{i}\otimes_{{\mathbb{Z}}_{\ell}}{\mathbb{Q}}_{\ell}$. Set ${\mathcal{D}}={\mathop{\rm Hom}\nolimits}_{{\mathcal{O}}_{\lambda}}({\mathcal{D}}_{1},{\mathcal{D}}_{2})$ and $D={\mathop{\rm Hom}\nolimits}_{K_{\lambda}}(D_{1},D_{2})$ which are objects of $\mathcal{M}\mathcal{F}$}} and MF respectively. We also have $D\cong\mathcal{D}\otimes_{{\mathbb{Z}}_{\ell}}{\mathbb{Q}}_{\ell}$. An elementary computation shows that the first two rows in the following commutative diagram are exact
[…]
where $\pi$ is defined as follows. For $\eta\in{\mathcal{D}}$ define an extension $E_{\eta}$ of ${\mathcal{D}}_{1}$ by ${\mathcal{D}}_{2}$ in ${\mathcal{O}}_{\lambda}$-$\mathcal{M}\mathcal{F}$}} with underlying ${\mathcal{O}}_{\lambda}$-module ${\mathcal{D}}_{2}\oplus{\mathcal{D}}_{1}$, filtration
$\operatorname{{\mathop{\rm Fil}\nolimits}}^{i}E_{\eta}:=\operatorname{{\mathop{\rm Fil}\nolimits}}^{i}{\mathcal{D}}_{2}\oplus\operatorname{{\mathop{\rm Fil}\nolimits}}^{i}{\mathcal{D}}_{1}$
and Frobenius maps $\phi^{i}:\operatorname{{\mathop{\rm Fil}\nolimits}}^{i}E_{\eta}\rightarrow E_{\eta}$
(69) $\phi^{i}(x,y)=(\phi^{i}(x)+\eta\phi^{i}(y),\phi^{i}(y)).$
The same definitions for $\eta\in D$ lead to an extension in $K_{\lambda}$-MF. Then $\pi(\eta)$ is the class of the Yoneda extension $E_{\eta}$ in $\operatorname{{\mathop{\rm Ext}\nolimits}}^{1}$ (we shall identify $\operatorname{{\mathop{\rm Ext}\nolimits}}^{1}$ with the group of Yoneda extensions throughout).
To explain the remaining part of diagram (68) we first recall the notion of admissibility from [Fo1, 3.6.4]. A filtered $\phi$-module $D^{\prime}$ in MF is called admissible if the natural map $B_{cris}\otimes_{{\mathbb{Q}}_{\ell}}D^{\prime}\cong B_{cris}\otimes_{{\mathbb{Q}}_{\ell}}V(D^{\prime})$ is an isomorphism where $V(D^{\prime})$ is the $G_{\ell}$-representation
$V(D^{\prime})=\operatorname{{\mathop{\rm Fil}\nolimits}}^{0}(D^{\prime}\otimes B_{cris})^{\phi\otimes\phi=1}.$
The functor $D^{\prime}\rightarrow V(D^{\prime})$ is fully faithful and exact on the category of admissible filtered $\phi$-modules, and induces an equivalence of this category with the category ${\rm Rep}_{cris}(G_{\ell})$ of crystalline $K_{\lambda}[G_{l}]$-representations (see [Fo1, 3.6.5]). If $D^{\prime}={\mathcal{D}}^{\prime}\otimes_{{\mathbb{Z}}_{l}}{\mathbb{Q}}_{l}$ for some object $\mathcal{D}$$^{\prime}$ of ${{\mathcal{M}\mathcal{F}}}^{0}$ then $D^{\prime}$ is admissible by [F-L, Th. 8.4], and for such $D^{\prime}$ we have a natural isomorphism $V(D^{\prime})\cong{\mathbb{V}}({\mathcal{D}}^{\prime})\otimes_{{\mathbb{Z}}_{\ell}}{\mathbb{Q}}_{\ell}$ by (1). If $D^{\prime}={\mathcal{D}}^{\prime}\otimes_{{\mathbb{Z}}_{l}}{\mathbb{Q}}_{l}$ for some object $\mathcal{D}$$^{\prime}$ of ${{\mathcal{M}\mathcal{F}}}^{a}$ then we can extend the definition of $\mathbb{V}$} by ${\mathbb{V}}({\mathcal{D}}^{\prime})={\mathbb{V}}({\mathcal{D}}^{\prime}[-a])(a)$ (Tate twist) and we deduce again that $D^{\prime}$ is admissible. In particular $D_{1}$ and $D_{2}$ are admissible, and then $D$ is admissible by [Fo1, Prop. 3.4.3]. Putting $V:=V(D)$ and $V_{i}:=V(D_{i})$ we have an isomorphism of $G_{\ell}$-representations $V={\mathop{\rm Hom}\nolimits}_{K_{\lambda}}(V_{1},V_{2})$ by [Fo1, 3.6].
Coming back to diagram (68), the map $\iota$ is just the natural isomorphism
$D\xrightarrow{1\otimes-}B_{cris}\otimes_{{\mathbb{Q}}_{\ell}}D\cong B_{cris}\otimes_{{\mathbb{Q}}_{\ell}}V\leftarrow H^{0}({\mathbb{Q}}_{\ell},B_{cris}\otimes_{{\mathbb{Q}}_{\ell}}V)=:D(V)$
and $e$ is the boundary map in Galois cohomology induced from the short exact sequence of $G_{{\mathbb{Q}}_{\ell}}$-modules
(70) $0\rightarrow V(D)\rightarrow\operatorname{{\mathop{\rm Fil}\nolimits}}^{0}(B_{cris}\otimes_{{\mathbb{Q}}_{\ell}}D)\xrightarrow{1-\phi\otimes\phi}B_{cris}\otimes_{{\mathbb{Q}}_{\ell}}D\rightarrow 0$
as in the proof of [B-K, Lemma4.5(b)]. It is clear that $\text{$K_{\lambda}$-MF}^{a}$ is closed under extensions inside $K_{\lambda}$-MF hence we obtain a chain of isomorphisms
(71) $\theta^{i}:\operatorname{{\mathop{\rm Ext}\nolimits}}^{i}_{\text{$K_{\lambda}$-MF}}(D_{1},D_{2})\leftarrow\operatorname{{\mathop{\rm Ext}\nolimits}}^{i}_{\text{$K_{\lambda}$-MF}^{a}}(D_{1},D_{2})\xrightarrow{v^{i}}\\ \operatorname{{\mathop{\rm Ext}\nolimits}}^{i}_{{\rm Rep}_{cris}(G_{\ell})}(V_{1},V_{2})\xrightarrow{\Delta^{i}}\operatorname{{\mathop{\rm Ext}\nolimits}}^{i}_{{\rm Rep}_{cris}(G_{\ell})}(K_{\lambda},{\mathop{\rm Hom}\nolimits}_{K_{\lambda}}(V_{1},V_{2}))\rightarrow H^{i}_{f}({\mathbb{Q}}_{\ell},V)$
for $i=0,1$. Here $\Delta^{1}$ sends a Yoneda extension
$0\rightarrow V_{2}\rightarrow V_{3}\rightarrow V_{1}\rightarrow 0$
to the pull back to $K_{\lambda}\cdot 1_{V_{1}}\subseteq{\mathop{\rm Hom}\nolimits}_{K_{\lambda}}(V_{1},V_{1})$ of the induced extension
$0\rightarrow{\mathop{\rm Hom}\nolimits}_{K_{\lambda}}(V_{1},V_{2})\rightarrow{\mathop{\rm Hom}\nolimits}_{K_{\lambda}}(V_{1},V_{3})\rightarrow{\mathop{\rm Hom}\nolimits}_{K_{\lambda}}(V_{1},V_{1})\rightarrow 0.$
The maps $v^{i}$ (defined by applying $V$ to a Yoneda extension) are isomorphisms because $V$ is fully faithful and exact.
The three lower rows in (68) with the indicated maps form a commutative diagram, and all these rows are exact (see [B-K, Lemma 4.5(b)] for the two lower rows). We shall verify the identity $\theta^{1}\pi=e\iota$, all the others being straightforward. Consider the commutative diagram
[…]
where all unnamed arrows are natural projection or inclusion maps, the top row is (70) with $D$ replaced by $D_{2}$, and the action of $\phi$ on $D_{2}\oplus D_{1}$ is given by (69). For $\psi\in D$, the extension $e\iota(\psi)$ is the pullback of (70) under $K_{\lambda}(1\otimes\psi)\subset B_{cris}\otimes_{{\mathbb{Q}}_{\ell}}D$. To compute $(\Delta^{1})^{-1}e\iota(\psi)$ apply the exact functor ${\mathop{\rm Hom}\nolimits}_{K_{\lambda}}(V_{1},-)$ to diagram (72). Via the isomorphisms
${\mathop{\rm Hom}\nolimits}_{K_{\lambda}}(V_{1},B_{cris}\otimes_{{\mathbb{Q}}_{\ell}}D_{2})\cong$ ${\mathop{\rm Hom}\nolimits}_{B_{cris}\otimes K_{\lambda}}(B_{cris}\otimes_{{\mathbb{Q}}_{\ell}}V_{1},B_{cris}\otimes_{{\mathbb{Q}}_{\ell}}D_{2})$
$\cong$ ${\mathop{\rm Hom}\nolimits}_{B_{cris}\otimes K_{\lambda}}(B_{cris}\otimes_{{\mathbb{Q}}_{\ell}}D_{1},B_{cris}\otimes_{{\mathbb{Q}}_{\ell}}D_{2})$
$\cong$ $B_{cris}\otimes_{{\mathbb{Q}}_{\ell}}{\mathop{\rm Hom}\nolimits}_{K_{\lambda}}(D_{1},D_{2}),$
${\mathop{\rm Hom}\nolimits}_{K_{\lambda}}(V_{1},\operatorname{{\mathop{\rm Fil}\nolimits}}^{0}(B_{cris}\otimes_{{\mathbb{Q}}_{\ell}}D_{2}))\cong$ $\operatorname{{\mathop{\rm Fil}\nolimits}}^{0}(B_{cris}\otimes_{{\mathbb{Q}}_{\ell}}{\mathop{\rm Hom}\nolimits}_{K_{\lambda}}(D_{1},D_{2}))$
the first row becomes isomorphic to (70) and the image of
$1_{V_{1}}\in{\mathop{\rm Hom}\nolimits}_{K_{\lambda}}(V_{1},V_{1})={\mathop{\rm Hom}\nolimits}_{K_{\lambda}}(V_{1},\operatorname{{\mathop{\rm Fil}\nolimits}}^{0}(B_{cris}\otimes_{{\mathbb{Q}}_{\ell}}D_{1})^{\phi=1})$
in $B_{cris}\otimes_{{\mathbb{Q}}_{\ell}}D$ is $1\otimes\psi$. Hence $(\Delta^{1})^{-1}e\iota(\psi)$ is represented by the lower row in (72). But from the definition of $\pi$ it is immediate that the lower row in (72) is the image of $\pi(\psi)$ under the functor $V$. This gives the identity $\theta^{1}\pi=e\iota$.
Put $T_{i}={\mathbb{V}}({\mathcal{D}}_{i})$ for $i=1,2$ and $T={\mathop{\rm Hom}\nolimits}_{{\mathcal{O}}_{\lambda}}(T_{1},T_{2})$ with its natural $G_{\ell}$-action. Then, since $T_{1}$ is torsion free, $H^{1}({\mathbb{Q}}_{\ell},T)$ naturally identifies with the set of equivalence classes of extensions of ${\mathcal{O}}_{\lambda}[G_{\ell}]$-modules
(73) $0\rightarrow T_{2}\rightarrow T_{3}\rightarrow T_{1}\rightarrow 0.$
Since the functor $\mathbb{V}$} is exact on $\text{${\mathcal{O}}_{\lambda}$-${{\mathcal{M}\mathcal{F}}}$}^{a}$, and since $\text{${\mathcal{O}}_{\lambda}$-${{\mathcal{M}\mathcal{F}}}$}^{a}$ is closed under extensions in ${\mathcal{O}}_{\lambda}$-$\mathcal{M}\mathcal{F}$}}, we obtain maps
$\Theta^{i}:\operatorname{{\mathop{\rm Ext}\nolimits}}^{i}_{\text{${\mathcal{O}}_{\lambda}$-${{\mathcal{M}\mathcal{F}}}$}}({\mathcal{D}}_{1},{\mathcal{D}}_{2})\xleftarrow{\sim}\operatorname{{\mathop{\rm Ext}\nolimits}}^{i}_{\text{${\mathcal{O}}_{\lambda}$-${{\mathcal{M}\mathcal{F}}}$}^{a}}({\mathcal{D}}_{1},{\mathcal{D}}_{2})\rightarrow\operatorname{{\mathop{\rm Ext}\nolimits}}^{i}_{{\mathcal{O}}_{\lambda}[G_{\ell}]}(T_{1},T_{2})\cong H^{i}({\mathbb{Q}}_{\ell},T)$
analogous to the maps $\theta^{i}$. The faithfulness of $\mathbb{V}$} implies that $\Theta^{0}$ is injective and fullness of $\mathbb{V}$} implies that $\Theta^{0}$ is surjective and that $\Theta^{1}$ is injective. The image of $\Theta^{1}$ lies in the subgroup
$H^{1}_{f}({\mathbb{Q}}_{\ell},T):=\{[T_{3}]\in H^{1}({\mathbb{Q}}_{\ell},T)|[V_{3}]:=[T_{3}\otimes_{{\mathbb{Z}}_{\ell}}{\mathbb{Q}}_{\ell}]\in H^{1}_{f}({\mathbb{Q}}_{\ell},V)\}$
since ${\mathbb{V}}({\mathcal{D}}_{3})\otimes_{{\mathbb{Z}}_{\ell}}{\mathbb{Q}}_{\ell}\cong V({\mathcal{D}}_{3}\otimes_{{\mathbb{Z}}_{\ell}}{\mathbb{Q}}_{\ell})$ is a crystalline representation. Conversely, if (73) lies in $H^{1}_{f}({\mathbb{Q}}_{\ell},T)$, the $G_{\ell}$-module $T_{3}$ is a submodule of a crystalline representation $V_{3}$ so that $D(V_{3})$ lies in $\text{$K_{\lambda}$-MF}^{a}$ and hence $T_{3}$ lies in the essential image of the Fontaine Laffaille functor $\mathbb{V}$}, $T_{3}={\mathbb{V}}({\mathcal{D}}_{3})$, say. Since $\mathbb{V}$} is full the extension (73) is the image of a sequence $0\rightarrow{\mathcal{D}}_{2}\rightarrow{\mathcal{D}}_{3}\rightarrow{\mathcal{D}}_{1}\rightarrow 0$ in $\text{${\mathcal{O}}_{\lambda}$-${{\mathcal{M}\mathcal{F}}}$}^{a}$ and since $\mathbb{V}$} is fully faithful and exact this sequence is exact, hence represents an element of $\operatorname{{\mathop{\rm Ext}\nolimits}}^{1}_{\text{${\mathcal{O}}_{\lambda}$-${{\mathcal{M}\mathcal{F}}}$}}({\mathcal{D}}_{1},{\mathcal{D}}_{2})$. We conclude that
(74) $\Theta^{1}:\operatorname{{\mathop{\rm Ext}\nolimits}}^{1}_{\text{${\mathcal{O}}_{\lambda}$-${{\mathcal{M}\mathcal{F}}}$}}({\mathcal{D}}_{1},{\mathcal{D}}_{2})\cong H^{1}_{f}({\mathbb{Q}}_{\ell},T)$
is an isomorphism. It is clear that $\theta^{i}\epsilon^{i}=\tilde{\epsilon}^{i}\Theta^{i}$ where $\tilde{\epsilon}^{i}:H^{i}({\mathbb{Q}}_{\ell},T)\rightarrow H^{i}({\mathbb{Q}}_{\ell},V)$ are the natural maps. The last row in (68) induces an isomorphism
${\mathop{\rm det}\nolimits}_{K_{\lambda}}H^{0}({\mathbb{Q}}_{\ell},V)\otimes_{K_{\lambda}}{\mathop{\rm det}\nolimits}_{K_{\lambda}}^{-1}H^{1}_{f}({\mathbb{Q}}_{\ell},V)$ $\cong{\mathop{\rm det}\nolimits}_{K_{\lambda}}D\otimes_{K_{\lambda}}{\mathop{\rm det}\nolimits}_{K_{\lambda}}^{-1}D\otimes_{K_{\lambda}}{\mathop{\rm det}\nolimits}_{K_{\lambda}}^{-1}t_{V}$
$\cong{\mathop{\rm det}\nolimits}_{K_{\lambda}}^{-1}t_{V}$
and the Tamagawa ideal is defined in [F-P, I.4.1.1] so that
(75) ${\mathop{\rm det}\nolimits}_{{\mathcal{O}}_{\lambda}}H^{0}({\mathbb{Q}}_{\ell},T)\otimes_{{\mathcal{O}}_{\lambda}}{\mathop{\rm det}\nolimits}_{{\mathcal{O}}_{\lambda}}^{-1}H^{1}_{f}({\mathbb{Q}}_{\ell},T)\cong\operatorname{{\mathop{\rm Tam}\nolimits}}_{\ell,\omega}^{0}(T)\omega^{-1}.$
Using the fact that $\Theta^{0}$ is an isomorphism together with (74) and (68) one computes that the left hand side in (75) equals ${\mathop{\rm det}\nolimits}_{{\mathcal{O}}_{\lambda}}^{-1}{\mathcal{D}}/\operatorname{{\mathop{\rm Fil}\nolimits}}^{0}{\mathcal{D}}$ so that $\operatorname{{\mathop{\rm Tam}\nolimits}}_{\ell,\omega}^{0}(T)={\mathcal{O}}_{\lambda}$ if $\omega$ is a basis of ${\mathop{\rm det}\nolimits}_{{\mathcal{O}}_{\lambda}}{\mathcal{D}}/\operatorname{{\mathop{\rm Fil}\nolimits}}^{0}{\mathcal{D}}$.
These arguments apply to ${\mathcal{D}}_{1}={\mathcal{D}}_{2}={\mathcal{M}}_{f,{\lambda\mbox{\scriptsize-crys}}}$ which is an object of $\text{${\mathcal{O}}_{\lambda}$-${{\mathcal{M}\mathcal{F}}}$}^{0}$ if $\lambda\in S_{f}$, more specifically if $\ell\nmid N$ and $\ell>k$. We have $T_{1}=T_{2}={\mathcal{M}}_{f,\lambda}$ and $T={\mathcal{A}}_{f,\lambda}\oplus{\mathcal{O}}_{\lambda}$. Our choice of $\omega_{A}$ then ensures that $\operatorname{{\mathop{\rm Tam}\nolimits}}_{\ell,\omega_{A}}^{0}({\mathcal{A}}_{f,\lambda})={\mathcal{O}}_{\lambda}$. For $B_{f}=A_{f}(1)$ we can use the same argument as long as both ${\mathcal{D}}_{1}={\mathcal{M}}_{f,{\lambda\mbox{\scriptsize-crys}}}$ and ${\mathcal{D}}_{2}={\mathcal{M}}_{f,{\lambda\mbox{\scriptsize-crys}}}[1]$ are objects of $\text{${\mathcal{O}}_{\lambda}$-${{\mathcal{M}\mathcal{F}}}$}^{1}$. This is the case if $\ell>k+1$ or if $l=k+1$ and ${\mathcal{M}}_{f,{\lambda\mbox{\scriptsize-crys}}}$ has no nonzero quotient AA in ${\mathcal{O}}_{\lambda}$-$\mathcal{M}\mathcal{F}$}} with $\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}A=A$.
Lemma 8.13.
If $a_{\ell}\equiv 0\mod\lambda$ then ${\mathcal{M}}_{f,{\lambda\mbox{\scriptsize-crys}}}$ has no nonzero quotient AA in ${\mathcal{O}}_{\lambda}$-$\mathcal{M}\mathcal{F}$}} with $\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}A=A$.
Proof.
By [Scho2] we know that the characteristic polynomial of $\phi$ on ${\mathcal{M}}:={\mathcal{M}}_{f,{\lambda\mbox{\scriptsize-crys}}}$ is $X^{2}-a_{\ell}X+\psi(\ell)\ell^{k-1}$ hence $\bar{\phi}$ has characteristic polynomial $X^{2}$ on $\bar{{\mathcal{M}}}:={\mathcal{M}}\otimes_{\mathcal{O}}({\mathcal{O}}_{\lambda}/\lambda)$. Since
(76) $\bar{{\mathcal{M}}}=\bar{\phi}(\bar{{\mathcal{M}}})+\bar{\phi}^{k-1}(\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}\bar{{\mathcal{M}}})$
and $\dim_{{\mathcal{O}}_{\lambda}/\lambda}\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}\bar{{\mathcal{M}}}=1$ the map $\bar{\phi}$ is nonzero, hence conjugate to $\left(\begin{smallmatrix}0&1\\ 0&0\end{smallmatrix}\right)$. Since $\bar{\phi}=l^{k-1}\bar{\phi}^{k-1}=0$ on $\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}\bar{{\mathcal{M}}}$ and because of (76) we have $\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}\bar{{\mathcal{M}}}=\text{ker}(\bar{\phi})=\bar{\phi}(\bar{{\mathcal{M}}})$. It is now easy to see that $\bar{{\mathcal{M}}}$ is a simple object in ${\mathcal{O}}_{\lambda}$-$\mathcal{M}\mathcal{F}$}}: Any proper subobject $N\subset\bar{{\mathcal{M}}}$ is $\bar{\phi}$-stable, hence contained in $\text{ker}(\bar{\phi})=\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}\bar{{\mathcal{M}}}$ and we have $\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}N=N$. But again by (76) we find $\bar{\phi}^{k-1}(\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}\bar{{\mathcal{M}}})\not\subseteq\text{ker}(\bar{\phi})=\bar{\phi}(\bar{{\mathcal{M}}})$ so that $N=\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}N=0$. If AA is a nonzero quotient of $\mathcal{M}$} then $\bar{A}$ is a nonzero quotient of $\bar{{\mathcal{M}}}$ hence equal to $\bar{{\mathcal{M}}}$ and we find $\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}\bar{A}\neq\bar{A}$ and $\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}A\neq A$. ∎
It remains to prove Proposition 8.12 for ${\mathcal{B}}_{f,\lambda}$ in the ordinary case $a_{\ell}\not\equiv 0\mod\lambda$ (and $l=k+1$). We use the fact that $B_{f}\cong A_{f}^{*}(1)$ and appeal to the following conjecture, a slight generalisation (from $\mathbb{Z}$$_{\ell}$ to ${\mathcal{O}}_{\lambda}$) of conjecture $C_{EP}(V)$ of [PR2] (we also use a similar generalisation of [PR2, Prop. C.2.6]).
Let $V$ be a crystalline representation of $G_{\ell}$ over $K_{\lambda}$ and $T\subseteq V$ a $G_{\ell}$-stable ${\mathcal{O}}_{\lambda}$-lattice. Let $\omega$ (resp. $\omega^{*}$) be a lattice of
${\mathop{\rm det}\nolimits}_{K_{\lambda}}D(V)/\operatorname{{\mathop{\rm Fil}\nolimits}}^{0}D(V){\rm\ (resp.\ }{\mathop{\rm det}\nolimits}_{K_{\lambda}}D(V^{*}(1))/\operatorname{{\mathop{\rm Fil}\nolimits}}^{0}D(V^{*}(1)))$
so that we obtain a lattice $\omega\otimes\omega^{*,-1}$ of ${\mathop{\rm det}\nolimits}_{K_{\lambda}}D(V)$ via the exact sequence
(77) $0\rightarrow(D(V^{*}(1))/\operatorname{{\mathop{\rm Fil}\nolimits}}^{0}D(V^{*}(1)))^{*}\rightarrow D(V)\rightarrow D(V)/\operatorname{{\mathop{\rm Fil}\nolimits}}^{0}D(V)\rightarrow 0.$
Let $\eta(T,\omega,\omega^{*})\in B_{cris}\otimes_{{\mathbb{Q}}_{\ell}}K_{\lambda}$ be such that $t^{-t_{H}(V)}\otimes{\mathop{\rm det}\nolimits}_{{\mathcal{O}}_{\lambda}}T=\eta(T,\omega,\omega^{*})\omega\otimes\omega^{*,-1}$ under the comparison isomorphism $B_{cris}\otimes_{{\mathbb{Q}}_{\ell}}{\mathop{\rm det}\nolimits}_{K_{\lambda}}V\cong B_{cris}\otimes_{{\mathbb{Q}}_{\ell}}{\mathop{\rm det}\nolimits}_{K_{\lambda}}D(V)$. One shows that $\eta(T,\omega,\omega^{*})\in{\mathbb{Q}}_{\ell}^{ur}\otimes_{{\mathbb{Q}}_{\ell}}K_{\lambda}$ [PR2, Lemme C.2.8] and that in fact $\eta(T,\omega,\omega^{*})\in 1\otimes K_{\lambda}$ up to an element in $({\mathbb{Z}}_{\ell}^{ur}\otimes_{{\mathbb{Z}}_{\ell}}{\mathcal{O}}_{\lambda})^{\times}$.
Conjecture 8.14.
For $j\in{\mathbb{Z}}$, put $h_{j}(V)=\dim_{K_{\lambda}}\operatorname{{\mathop{\rm Fil}\nolimits}}^{j}D(V)/\operatorname{{\mathop{\rm Fil}\nolimits}}^{j+1}D(V)$, and put $\Gamma^{*}(j)=(j-1)!$ if $j>0$ and $\Gamma^{*}(j)=(-1)^{j}((-j)!)^{-1}$ if $j\leq 0$. Then
${\mathcal{O}}_{\lambda}\frac{\operatorname{{\mathop{\rm Tam}\nolimits}}_{\ell,\omega}^{0}(T)}{\operatorname{{\mathop{\rm Tam}\nolimits}}_{\ell,\omega^{*}}^{0}(T^{*}(1))}={\mathcal{O}}_{\lambda}\prod_{j}\Gamma^{*}(-j)^{-h_{j}(V)}\eta(T,\omega,\omega^{*})^{-1}$
Remark. One can show that upon taking the norm from $K_{\lambda}$ to $\mathbb{Q}$$_{\ell}$ all quantities in this formula transform into the corresponding quantities obtained by viewing $V$ as a representation over $\mathbb{Q}$$_{\ell}$ rather than $K_{\lambda}$. Since the norm map $K_{\lambda}^{\times}/{\mathcal{O}}_{\lambda}^{\times}\rightarrow{\mathbb{Q}}_{\ell}^{\times}/{\mathbb{Z}}_{\ell}^{\times}$ is injective it suffices to prove the conjecture for $K_{\lambda}={\mathbb{Q}}_{\ell}$.
We make Conjecture 8.14 more explicit for $V=A_{f,\lambda}$. In this case we have $h_{j}(V)=1$ for i=1k,0,k1i=1-k,0,k-1 and $h_{j}(V)=0$ otherwise so that $\prod_{j}\Gamma^{*}(-j)^{-h_{j}(V)}=(-1)^{k-1}(k-1)$. For $\lambda\in S_{f}$ Lemma 6.1 shows that the isomorphism $B_{cris}\otimes_{{\mathbb{Q}}_{\ell}}{\mathop{\rm det}\nolimits}_{K_{\lambda}}V\cong B_{cris}\otimes_{{\mathbb{Q}}_{\ell}}{\mathop{\rm det}\nolimits}_{K_{\lambda}}D(V)$ is induced by the functor $\mathbb{V}$} for the unit object in ${\mathcal{P}M}^{S}_{K}$, hence sends ${\mathop{\rm det}\nolimits}_{{\mathcal{O}}_{\lambda}}{\mathcal{A}}_{f,\lambda}$ to ${\mathop{\rm det}\nolimits}_{{\mathcal{O}}_{\lambda}}{\mathcal{A}}_{f,{\mathop{\rm dR}\nolimits}}\otimes_{\mathcal{O}}{\mathcal{O}}_{\lambda}$. The computation of $t_{B_{f}}$ before (50) works with $M_{f,{\mathop{\rm dR}\nolimits}}$ replaced by ${\mathcal{M}}_{f,{\mathop{\rm dR}\nolimits}}\otimes_{\mathcal{O}}{\mathcal{O}}_{\lambda}$ and the pairing (36) on $A_{f}$ gives a perfect pairing $({\mathcal{A}}_{f,{\mathop{\rm dR}\nolimits}}\otimes_{\mathcal{O}}{\mathcal{O}}_{\lambda})\otimes({\mathcal{A}}_{f,{\mathop{\rm dR}\nolimits}}\otimes_{\mathcal{O}}{\mathcal{O}}_{\lambda})\rightarrow{\mathcal{O}}_{\lambda}$. Hence we find that $\omega_{A}\otimes\omega_{B}^{-1}$ is a basis of ${\mathop{\rm det}\nolimits}_{{\mathcal{O}}_{\lambda}}{\mathcal{A}}_{f,{\mathop{\rm dR}\nolimits}}\otimes_{\mathcal{O}}{\mathcal{O}}_{\lambda}$ via the exact sequence (77). We conclude that $\eta({\mathcal{A}}_{f,\lambda},\omega_{A},\omega_{B})=1$ and that Conjecture 8.14 reduces to the assertion
${\mathcal{O}}_{\lambda}\operatorname{{\mathop{\rm Tam}\nolimits}}_{\ell,\omega_{A}}^{0}({\mathcal{A}}_{f,\lambda})={\mathcal{O}}_{\lambda}\operatorname{{\mathop{\rm Tam}\nolimits}}_{\ell,\omega_{B}}^{0}({\mathcal{A}}_{f,\lambda}(1)).$
Moreover, we know from the first part of the proof that the left hand side equals ${\mathcal{O}}_{\lambda}$ if $\lambda\in S_{f}$. Now Conjecture 8.14 is shown in [PR1] for $K_{\lambda}={\mathbb{Q}}_{\ell}$ and $V$ an ordinary representation of $G_{{\mathbb{Q}}_{\ell}}$ (combine Proposition 4.2.5, Theorem 3.5.4 of loc. cit.) under the assumption of another conjecture Rec(V) which has meanwhile been proved in [Col]. Ordinarity of $A_{f,\lambda}$ is implied by ordinarity of $M_{f,\lambda}$ which in turn is implied by $a_{\ell}\not\equiv 0\mod\lambda$. This finishes the proof of Proposition 8.12. ∎
We shall first prove Theorem 8.10 for $M=B_{f}=A_{f}(1)$ in which case Lemma 8.11 applies. Fix a prime $\lambda\in S_{f}$ and let $\Sigma$ be the set of primes dividing NN if $N>1$ or put $\Sigma=\{p\}$ for some prime $\lambda\nmid p$ if $N=1$. The isomorphism $\gamma:M_{f}\rightarrow M_{f}^{\Sigma}$ of Proposition 6.3 satisfies
$\gamma^{t}=\gamma^{-1}\phi\prod_{p\in\Sigma}L_{p}^{{\mathop{\rm nv}\nolimits}}(B_{f},0)^{-1}$
by Proposition 6.5 where $\phi=\prod_{\delta_{p}=1}(-a_{p})\prod_{\delta_{p}=2}\psi(p)p^{k-1}\in{\mathcal{O}}_{\lambda}^{\times}$ (it is well known that $a_{p}^{2}=\psi(p)p^{k-1}$ or $a_{p}^{2}=\psi(p)p^{k-2}$ if $\delta_{p}=1$ [Miy, 4.6.17]). Moreover, $\gamma$ induces an isomorphism
$B_{f}={\mathop{\rm Hom}\nolimits}_{K}(M_{f},M_{f}(1))\rightarrow B_{f}^{\Sigma}:={\mathop{\rm Hom}\nolimits}_{K}(M_{f}^{\Sigma},M_{f}^{\Sigma}(1))$
and an isomorphism $\gamma:\Delta_{\bf f}(B_{f})\rightarrow\Delta_{\bf f}(B_{f}^{\Sigma})$ so that
$\gamma(b)(x\otimes(2\pi i)^{2})\otimes\iota^{2}=\gamma_{{\mathop{\rm dR}\nolimits}}b(\gamma_{{\mathop{\rm dR}\nolimits}}^{-1}(x)\otimes(2\pi i)^{2})\otimes\iota^{2}$
for $b\in\Delta_{\bf f}(B_{f})$ and $x\in M_{f,{\mathop{\rm dR}\nolimits}}^{\Sigma}$. For such $b$ and $x$ we have
$\langle x,\gamma(b)(x\otimes(2\pi i)^{2})\otimes\iota^{2}\rangle^{\Sigma}=$ $\langle\gamma_{{\mathop{\rm dR}\nolimits}}^{-1}(x),\gamma_{{\mathop{\rm dR}\nolimits}}^{t}\gamma(b)(x\otimes(2\pi i)^{2})\otimes\iota^{2}\rangle$
$=$ $\langle\gamma_{{\mathop{\rm dR}\nolimits}}^{-1}(x),\gamma_{{\mathop{\rm dR}\nolimits}}^{-1}\gamma(b)(x\otimes(2\pi i)^{2})\otimes\iota^{2}\rangle\phi\prod_{p\in\Sigma}L_{p}^{{\mathop{\rm nv}\nolimits}}(B_{f},0)^{-1}$
(78) $=$ $\langle\gamma_{{\mathop{\rm dR}\nolimits}}^{-1}(x),b(\gamma_{{\mathop{\rm dR}\nolimits}}^{-1}(x)\otimes(2\pi i)^{2})\otimes\iota^{2}\rangle\phi\prod_{p\in\Sigma}L_{p}^{{\mathop{\rm nv}\nolimits}}(B_{f},0)^{-1}.$
Recall that $\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}{\mathcal{M}}_{f,{\mathop{\rm dR}\nolimits}}^{\Sigma}={\mathcal{O}}\cdot f^{\Sigma}={\mathcal{O}}\cdot\gamma(f)$ by Propositions 6.6 and 6.3 where ${\mathcal{O}}=\bigcap_{\lambda\in S_{f}}K\cap{\mathcal{O}}_{\lambda}$. Note that if $b^{\prime}$ is an ${\mathcal{O}}_{\lambda}$-basis for
${\mathop{\rm Hom}\nolimits}_{{\mathcal{O}}_{\lambda}}(\theta({\mathcal{B}}_{f,\lambda}),\omega_{B})$
$\cong{\mathop{\rm Hom}\nolimits}_{{\mathcal{O}}_{\lambda}}(\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}{\mathcal{M}}_{f,{\mathop{\rm dR}\nolimits}}^{\Sigma}\otimes_{\mathcal{O}}\theta({\mathcal{B}}_{f,\lambda}),({\mathcal{M}}_{f,{\mathop{\rm dR}\nolimits}}^{\Sigma}/\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}{\mathcal{M}}_{f,{\mathop{\rm dR}\nolimits}}^{\Sigma})\otimes_{\mathcal{O}}{\mathcal{O}}_{\lambda}\otimes\iota^{-2})$
then
${\mathcal{O}}_{\lambda}\cdot b^{\prime}(f^{\Sigma}\otimes(2\pi i)^{2})\otimes\iota^{2}=({\mathcal{M}}_{f,{\mathop{\rm dR}\nolimits}}^{\Sigma}/\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}{\mathcal{M}}_{f,{\mathop{\rm dR}\nolimits}}^{\Sigma})\otimes_{\mathcal{O}}{\mathcal{O}}_{\lambda}$
and hence
${\mathcal{O}}_{\lambda}\cdot\langle f^{\Sigma},b^{\prime}(f^{\Sigma}\otimes(2\pi i)^{2})\otimes\iota^{2}\rangle^{\Sigma}={\mathcal{O}}_{\lambda}\eta^{\Sigma}_{f}{\mathcal{M}}_{\psi}(1-k)_{{\mathop{\rm dR}\nolimits}}$
where $\eta^{\Sigma}_{f}$ was defined before Proposition $6.$ On the other hand by (78), Theorem 8.5 and the remark after the proof of Theorem 8.5 we have for $\lambda\in S_{f}$
${\mathcal{O}}_{\lambda}\cdot\langle f^{\Sigma},\gamma b^{\Sigma}(B_{f})(f^{\Sigma}\otimes(2\pi i)^{2})\otimes\iota^{2}\rangle^{\Sigma}$
$={\mathcal{O}}_{\lambda}\cdot\langle f,b^{\Sigma}(B_{f})(f\otimes(2\pi i)^{2})\otimes\iota^{2}\rangle\prod_{p\in\Sigma}L_{p}^{{\mathop{\rm nv}\nolimits}}(B_{f},0)^{-1}$
$={\mathcal{O}}_{\lambda}\cdot\langle f,b(B_{f})(f\otimes(2\pi i)^{2})\otimes\iota^{2}\rangle\prod_{p\in\Sigma_{e}(f)}L_{p}(B_{f},0)$
$={\mathcal{O}}_{\lambda}(b_{\mathop{\rm dR}\nolimits}\otimes\iota^{k-1})={\mathcal{O}}_{\lambda}{\mathcal{M}}_{\psi}(1-k)_{\mathop{\rm dR}\nolimits}.$
Applying Lemma 8.11 we must therefore show
$\frac{\operatorname{Fitt}_{{\mathcal{O}}_{\lambda}}H^{0}({\mathbb{Q}},B_{f,\lambda}^{D}/{\mathcal{B}}_{f,\lambda}^{D})}{\operatorname{Fitt}_{{\mathcal{O}}_{\lambda}}H^{1}_{\Sigma}({\mathbb{Q}},B_{f,\lambda}^{D}/{\mathcal{B}}_{f,\lambda}^{D})\operatorname{{\mathop{\rm Tam}\nolimits}}_{\ell,\omega_{B}}^{0}({\mathcal{B}}_{f,\lambda})}\,\eta_{f}^{\Sigma}={\mathcal{O}}_{\lambda}.$
Using Proposition 8.12, the fact that ${\mathcal{A}}_{f}={\mathcal{B}}_{f}^{D}$ and the vanishing of $H^{0}({\mathbb{Q}},A_{f,\lambda}/{\mathcal{A}}_{f,\lambda})$ for $\lambda\in S_{f}$ this reduces to the identity
$\operatorname{Fitt}_{{\mathcal{O}}_{\lambda}}H^{1}_{\Sigma}({\mathbb{Q}},A_{f,\lambda}/{\mathcal{A}}_{f,\lambda})={\mathcal{O}}_{\lambda}\eta^{\Sigma}_{f}$
which is Theorem 7.15.
By (63), (64) and Lemma 8.12 the factor in front of ${\mathop{\rm Hom}\nolimits}(\theta({\mathcal{M}}_{\lambda}),\omega)$ in (65) is the same for $M=A_{f}$ and $M=B_{f}$. The isomorphism $\rm tw$} defined in (51) maps ${\mathop{\rm Hom}\nolimits}_{{\mathcal{O}}_{\lambda}}(\theta({\mathcal{A}}_{f,\lambda}),\omega_{A})$ to ${\mathop{\rm Hom}\nolimits}_{{\mathcal{O}}_{\lambda}}(\theta({\mathcal{B}}_{f,\lambda}),\omega_{B})$, hence $\delta_{{\bf f},\lambda}(A_{f})$ to $\delta_{{\bf f},\lambda}(B_{f})$. Theorem 8.10 for $M=A_{f}$ therefore follows from Theorem 8.10 for $M=B_{f}$, together with Theorem 8.5 and the fact that $(1-k)\epsilon(A_{f})$ is a unit in ${\mathcal{O}}_{\lambda}$. ∎

### v2

@@
 Abstract.
-Let $f$ be a newform of weight $k\geq 2$, level NN with coefficients in a number field $K$, and AA the adjoint motive of the motive $M$ associated to $f$.
-We carefully discuss the construction of the realisations of $M$ and AA, as well as natural integral structures in these realisations.
+Let $f$ be a newform of weight $k\geq 2$, level $N$ with coefficients in a number field $K$, and $A$ the adjoint motive of the motive $M$ associated to $f$.
+We carefully discuss the construction of the realisations of $M$ and $A$, as well as natural integral structures in these realisations.
 We then use the method of Taylor and Wiles to verify the $\lambda$-part of the Tamagawa number conjecture of Bloch and Kato for $L(A,0)$ and $L(A,1)$.
@@
 The Taylor-Wiles construction yields a formula relating the size of a certain module measuring congruences between modular forms to that of a certain Galois cohomology group.
-This was carried out in [Wi] and [T-W] in the context of modular forms of weight 22, where it was used to prove results in the direction of the Fontaine-Mazur conjecture [F-M].
+This was carried out in [Wi] and [T-W] in the context of modular forms of weight $2$, where it was used to prove results in the direction of the Fontaine-Mazur conjecture [F-M].
 While it was no surprise that the method could be generalized to higher weight modular forms and that the resulting formula would be related to the Bloch-Kato conjecture, there remained many technical details to verify in order to accomplish this.
@@
 This was already done up to a rational multiple by conjectures of Deligne and Beilinson; the additional precision of the Bloch-Kato conjecture can be thought of as a generalized class number formula, where ideal class groups are replaced by groups defined using Galois cohomology.
-Dirichlet’s class number formula amounts to the conjecture for the Dedekind zeta function for a number field at $s=0$ or 11.
-The conjecture is also known up to a power of 22 for Dirichlet $L$-functions (including the Riemann zeta function) at any integer ([Ma-W], [B-K], [H-K]).
+Dirichlet’s class number formula amounts to the conjecture for the Dedekind zeta function for a number field at $s=0$ or $1$.
+The conjecture is also known up to a power of $2$ for Dirichlet $L$-functions (including the Riemann zeta function) at any integer ([Ma-W], [B-K], [H-K]).
 It is known up to an explicit set of bad primes for the $L$-function of a CM elliptic curve at $s=1$ if the order of vanishing is $\leq 1$ ([C-W], [Ru1], [Ko]).
 There are also partial results for $L$-functions of other modular forms at the central critical value ([G-Z], [Kato3], [Ko-L], [N], [Z]) and for values of certain Hecke $L$-functions ([Hn], [Gu2], [Ki]).
-Here we consider the adjoint $L$-function of a modular form of weight $k\geq 2$ at $s=0$ and 11.
+Here we consider the adjoint $L$-function of a modular form of weight $k\geq 2$ at $s=0$ and $1$.
 Special values of the $L$-function associated to the adjoint of a modular form, and more generally, twists of its symmetric square, have been studied by many mathematicians.
@@
   • a weight filtration.
-Suppose that $f$ is a newform of weight $k\geq 2$ and level NN.
+Suppose that $f$ is a newform of weight $k\geq 2$ and level $N$.
 Much of the paper is devoted to the construction of the motivic structure $A_{f}$ for which we prove the conjecture.
@@
 Let us briefly recall here how the construction works.
-We start with the modular curve $X_{N}$ parametrizing elliptic curves with level NN structure.
+We start with the modular curve $X_{N}$ parametrizing elliptic curves with level $N$ structure.
 Then one takes the Betti, de Rham and $\ell$-adic cohomology of $X_{N}$ with coefficients in a sheaf defined as the (k2)(k-2)-nd symmetric power of the relative cohomology of the universal elliptic curve over $X_{N}$.
@@
 Theorem 0.1.
-(=Theorem 8.10) Let $f$ be a newform of weight $k\geq 2$ and level NN with coefficients in $K$.
+(=Theorem 8.10) Let $f$ be a newform of weight $k\geq 2$ and level $N$ with coefficients in $K$.
 If $\lambda$ is not in $T_{f}$, then the $\lambda$-part of the Bloch-Kato conjecture holds for $A_{f}$ and $A_{f}(1)$.
@@
 Theorem 0.2.
-(=Theorem 7.15) Let $f$ be a newform of weight $k\geq 2$ and level NN with coefficients in $K$.
-Suppose $\Sigma$ is a finite set of rational primes containing those dividing NN.
+(=Theorem 7.15) Let $f$ be a newform of weight $k\geq 2$ and level $N$ with coefficients in $K$.
+Suppose $\Sigma$ is a finite set of rational primes containing those dividing $N$.
 Suppose that $\lambda$ is a prime of $K$ which is not in $T_{f}$ and does not divide any prime in $\Sigma$.
@@
 1. Generalities and examples of premotivic structures
-$1.$ Galois representations
+1.1. Galois representations
 For a field $F$, $\bar{F}$ will denote an algebraic closure, and $G_{F}=\operatorname{{\mathop{\rm Gal}\nolimits}}(\bar{F}/F)$.
@@
 Let $I_{F}$ denote the inertia subgroup of $G_{F}$, $W_{F}$ the Weil subgroup, $\nu^{-1}({\mathbb{Z}})$ and $^{\prime}W_{F}$ the Weil-Deligne group of $F$ (see [De2, §8]).
-Recall that if $K$ is a field of characteristic zero, then to give a representation of $^{\prime}W_{F}$ with coefficients in $K$ is equivalent to giving a finite-dimensional $K$-vector space $V$ (with the discrete topology), a continuous homomorphism $\rho:W_{F}\to\operatorname{\operatorname{{\mathop{\rm Aut}\nolimits}}}_{K}V$ and a nilpotent endomorphism NN of $V$ satisfying $\rho(g)N=p^{-\nu(g)}N\rho(g)$.
+Recall that if $K$ is a field of characteristic zero, then to give a representation of $^{\prime}W_{F}$ with coefficients in $K$ is equivalent to giving a finite-dimensional $K$-vector space $V$ (with the discrete topology), a continuous homomorphism $\rho:W_{F}\to\operatorname{\operatorname{{\mathop{\rm Aut}\nolimits}}}_{K}V$ and a nilpotent endomorphism $N$ of $V$ satisfying $\rho(g)N=p^{-\nu(g)}N\rho(g)$.
 If $F={\mathbb{Q}}_{p}$, we write simply $G_{p}$, $I_{p}$, $W_{p}$ and $^{\prime}W_{p}$ for these groups and identify $I_{p}\subset W_{p}\subset G_{p}$ with their images in $G_{\mathbb{Q}}$.
@@
 Similarly $D_{\mbox{\scriptsize crys}}(V)=(B_{\mbox{\scriptsize crys}}\otimes_{{\mathbb{Q}}_{\ell}}V)^{G_{F}}$ is a filtered free $F_{0}\otimes_{{\mathbb{Q}}_{\ell}}K_{\lambda}$-module of finite rank with a $(\phi_{\ell}\otimes 1)$-semilinear endomorphism $\phi$, and $V$ is called crystalline if $\dim_{F_{0}}D_{\mbox{\scriptsize crys}}(V)=\dim_{{\mathbb{Q}}_{\ell}}V$.
-Similarly $D_{\rm st}(V)=(B_{\rm st}\otimes_{{\mathbb{Q}}_{\ell}}V)^{G_{F}}$ is a filtered free $F_{0}\otimes_{{\mathbb{Q}}_{\ell}}K_{\lambda}$-module of finite rank with a $(\phi_{\ell}\otimes 1)$-semilinear endomorphism $\phi$ and an endomorphism NN satisfying $N\phi=\ell\phi N$, and $V$ is called semistable if $\dim_{F_{0}}D_{\rm st}(V)=\dim_{{\mathbb{Q}}_{\ell}}V$.
+Similarly $D_{\rm st}(V)=(B_{\rm st}\otimes_{{\mathbb{Q}}_{\ell}}V)^{G_{F}}$ is a filtered free $F_{0}\otimes_{{\mathbb{Q}}_{\ell}}K_{\lambda}$-module of finite rank with a $(\phi_{\ell}\otimes 1)$-semilinear endomorphism $\phi$ and an endomorphism $N$ satisfying $N\phi=\ell\phi N$, and $V$ is called semistable if $\dim_{F_{0}}D_{\rm st}(V)=\dim_{{\mathbb{Q}}_{\ell}}V$.
 One defines $D_{\mathop{\rm pst}\nolimits}(V)$ as the direct limit of $(B_{\rm st}\otimes_{{\mathbb{Q}}_{\ell}}V)^{G_{F^{\prime}}}$ where $F^{\prime}$ runs over finite extensions of $F$ in $\bar{{\mathbb{Q}}}_{\ell}$; this is a filtered free ${\mathbb{Q}}_{\ell}^{\rm ur}\otimes_{{\mathbb{Q}}_{\ell}}K_{\lambda}$-module and one checks that $(\rho,N)$ with $\rho(g)=g\phi^{\nu(g)}$ makes $D_{\mathop{\rm pst}\nolimits}(V)$ a $K_{\lambda}$-rational representation of $^{\prime}W_{F}$, and $V$ is called potentially semistable if $\dim_{{\mathbb{Q}}_{\ell}^{\rm ur}}D_{\mathop{\rm pst}\nolimits}(V)=\dim_{{\mathbb{Q}}_{\ell}}V$.
@@
 We let ${{\mathcal{M}\mathcal{F}}}({\mathcal{O}}_{F})$ denote the category whose objects are finitely generated ${\mathcal{O}}_{F}$-modules equipped with
-  • a decreasing filtration such that $\operatorname{{\mathop{\rm Fil}\nolimits}}^{a}A=A$ and $\operatorname{{\mathop{\rm Fil}\nolimits}}^{b}A=0$ for some $a,b\in{\mathbb{Z}}$, and for each $i\in{\mathbb{Z}}$, $\operatorname{{\mathop{\rm Fil}\nolimits}}^{i}A$ is a direct summand of AA;
+  • a decreasing filtration such that $\operatorname{{\mathop{\rm Fil}\nolimits}}^{a}A=A$ and $\operatorname{{\mathop{\rm Fil}\nolimits}}^{b}A=0$ for some $a,b\in{\mathbb{Z}}$, and for each $i\in{\mathbb{Z}}$, $\operatorname{{\mathop{\rm Fil}\nolimits}}^{i}A$ is a direct summand of $A$;
   • $\phi_{\ell}$-semilinear maps $\phi^{i}:\operatorname{{\mathop{\rm Fil}\nolimits}}^{i}A\to A$ for $i\in{\mathbb{Z}}$ satisfying $\phi^{i}|_{\operatorname{{\mathop{\rm Fil}\nolimits}}^{i+1}A}=\ell\phi^{i+1}$ and $A=\sum{\mathop{\rm Im}\nolimits}\phi^{i}$.
 It follows from [F-L, 1.8] that ${{\mathcal{M}\mathcal{F}}}({\mathcal{O}}_{F})$ is an abelian category.
-Let ${{\mathcal{M}\mathcal{F}}}^{a}({\mathcal{O}}_{F})$ denote the full subcategory of objects AA satisfying $\operatorname{{\mathop{\rm Fil}\nolimits}}^{a}A=A$ and $\operatorname{{\mathop{\rm Fil}\nolimits}}^{a+\ell}A=0$ and having no non-trivial quotients $A^{\prime}$ such that $\operatorname{{\mathop{\rm Fil}\nolimits}}^{a+\ell-1}A^{\prime}=A^{\prime}$, and let ${{\mathcal{M}\mathcal{F}}}^{a}_{\mathop{\rm tor}\nolimits}({\mathcal{O}}_{F})$ denote the full subcategory of ${{\mathcal{M}\mathcal{F}}}^{a}({\mathcal{O}}_{F})$ consisting of objects of finite length.
+Let ${{\mathcal{M}\mathcal{F}}}^{a}({\mathcal{O}}_{F})$ denote the full subcategory of objects $A$ satisfying $\operatorname{{\mathop{\rm Fil}\nolimits}}^{a}A=A$ and $\operatorname{{\mathop{\rm Fil}\nolimits}}^{a+\ell}A=0$ and having no non-trivial quotients $A^{\prime}$ such that $\operatorname{{\mathop{\rm Fil}\nolimits}}^{a+\ell-1}A^{\prime}=A^{\prime}$, and let ${{\mathcal{M}\mathcal{F}}}^{a}_{\mathop{\rm tor}\nolimits}({\mathcal{O}}_{F})$ denote the full subcategory of ${{\mathcal{M}\mathcal{F}}}^{a}({\mathcal{O}}_{F})$ consisting of objects of finite length.
 So ${{\mathcal{M}\mathcal{F}}}^{0}_{\mathop{\rm tor}\nolimits}({\mathcal{O}}_{F})$ is the category denoted $\underline{MF}_{\mathop{\rm tor}\nolimits}^{f,\ell^{\prime}}$ in [F-L], and it follows from [F-L, 6.1] that ${{\mathcal{M}\mathcal{F}}}^{a}({\mathcal{O}}_{F})$ and ${{\mathcal{M}\mathcal{F}}}^{a}_{\mathop{\rm tor}\nolimits}({\mathcal{O}}_{F})$ are abelian categories, stable under taking subobjects, quotients, direct products and extensions in ${{\mathcal{M}\mathcal{F}}}({\mathcal{O}}_{F})$.
@@
 We can regard $\mathbb{V}$} as a functor from ${\mathcal{O}}_{\lambda}$-${{\mathcal{M}\mathcal{F}}}^{0}$ to the category of ${\mathcal{O}}_{\lambda}[G_{\ell}]$-modules.
-If AA and $A^{\prime}$ are objects of ${\mathcal{O}}_{\lambda}$-${{\mathcal{M}\mathcal{F}}}^{0}$ such that $A\otimes_{{\mathcal{O}}_{\lambda}}A^{\prime}$ defines an object of ${\mathcal{O}}_{\lambda}$-${{\mathcal{M}\mathcal{F}}}^{0}$, then there is a canonical isomorphism
+If $A$ and $A^{\prime}$ are objects of ${\mathcal{O}}_{\lambda}$-${{\mathcal{M}\mathcal{F}}}^{0}$ such that $A\otimes_{{\mathcal{O}}_{\lambda}}A^{\prime}$ defines an object of ${\mathcal{O}}_{\lambda}$-${{\mathcal{M}\mathcal{F}}}^{0}$, then there is a canonical isomorphism
 ${\mathbb{V}}(A\otimes_{{\mathcal{O}}_{\lambda}}A^{\prime})\cong{\mathbb{V}}(A)\otimes_{{\mathcal{O}}_{\lambda}}{\mathbb{V}}(A^{\prime}).$
@@
 If $K\subseteq K^{\prime}$, we let $S^{K^{\prime}}$ denote the set of primes in $S_{\bf f}(K^{\prime})$ lying over those in $S$, then $K^{\prime}\otimes_{K}\cdot$ defines a functor from $\operatorname{{\bf PM}}_{K}^{S}$ to $\operatorname{{\bf PM}}_{K^{\prime}}^{S^{K^{\prime}}}$.
-Definition $1.$
+Definition 1.1.
 An object $M$ of $\operatorname{{\bf PM}}_{K}^{S}$
@@
   • is $L$-admissible at $p$ if $M_{\lambda}$ is geometric for every $\lambda\not\in S$ dividing $p$ and the representations $D_{{\mathop{\rm pst}\nolimits}}(M_{\lambda}|G_{p})^{{\mathop{\rm ss}\nolimits}}$, for $\lambda\not\in S$, form a compatible system of $K$-rational representations of the Deligne-Weil group $^{\prime}W_{p}$ (see [De2, §8]);
-  • is $L$-admissible everywhere if it is $L$-admissible at $p$ for all primes pp.
+  • is $L$-admissible everywhere if it is $L$-admissible at $p$ for all primes $p$.
 If $M$ is $L$-admissible at $p$, then the local factor associated to $D_{{\mathop{\rm pst}\nolimits}}(M_{\lambda}|G_{p})^{{\mathop{\rm ss}\nolimits}}$ is of the form $P(p^{-s})^{-1}$ for some polynomial $P(u)\in K[u]$ independent of $\lambda$ not in $S$.
@@
 We say that $\mathcal{M}$} has good reduction at $p$, is $L$-admissible at $p$ or is $L$-admissible everywhere according to whether the same is true for $\mathbb{Q}$$\otimes{\mathcal{M}}$.
-Note that if $\mathcal{M}$} is $L$-admissible at $p$ and $p$ is not invertible in $\mathcal{O}$}, then $\mathcal{M}$} necessarily has good reduction at pp.
+Note that if $\mathcal{M}$} is $L$-admissible at $p$ and $p$ is not invertible in $\mathcal{O}$}, then $\mathcal{M}$} necessarily has good reduction at $p$.
 For objects $\mathcal{M}$} and $\mathcal{M}$$^{\prime}$ of ${\mathcal{P}M}_{K}^{S}$, we can form ${\mathcal{M}}\otimes_{{\mathcal{O}}_{K}}{\mathcal{M}}^{\prime}$ in ${\mathcal{P}M}_{K}^{S}$ provided ${\mathcal{M}}_{{\lambda\mbox{\scriptsize-crys}}}\otimes_{{\mathcal{O}}_{\lambda}}{\mathcal{M}}^{\prime}_{{\lambda\mbox{\scriptsize-crys}}}$ defines an object of ${{\mathcal{M}\mathcal{F}}}^{0}_{\lambda}$ for all $\lambda\not\in S$.
@@
 Note that if $\mathcal{M}$} is an object of ${\mathcal{P}M}_{K}^{S}$, then $\operatorname{{\mathop{\rm End}\nolimits}}{\mathcal{M}}$ is a finitely generated ${\mathcal{O}}_{K}$-module.
-If II is an ${\mathcal{O}}_{K}$-submodule of $\operatorname{{\mathop{\rm End}\nolimits}}{\mathcal{M}}$, then we define an object ${\mathcal{M}}[I]$ of ${\mathcal{P}M}_{K}^{S}$ as the kernel of
+If $I$ is an ${\mathcal{O}}_{K}$-submodule of $\operatorname{{\mathop{\rm End}\nolimits}}{\mathcal{M}}$, then we define an object ${\mathcal{M}}[I]$ of ${\mathcal{P}M}_{K}^{S}$ as the kernel of
 $(x_{1},\ldots,x_{r}):{\mathcal{M}}\to{\mathcal{M}}^{r}$
-where $x_{1},\ldots,x_{r}$ generate II.
+where $x_{1},\ldots,x_{r}$ generate $I$.
 This is independent of the choice of generators.
-This applies in particular when II is the image in $\operatorname{{\mathop{\rm End}\nolimits}}{\mathcal{M}}$ of an ideal in a commutative ${\mathcal{O}}_{K}$-algebra $R$ mapping to $\operatorname{{\mathop{\rm End}\nolimits}}{\mathcal{M}}$, or the augmentation ideal in ${\mathcal{O}}_{K}[G]$ where $G$ is a group acting on $\mathcal{M}$}.
+This applies in particular when $I$ is the image in $\operatorname{{\mathop{\rm End}\nolimits}}{\mathcal{M}}$ of an ideal in a commutative ${\mathcal{O}}_{K}$-algebra $R$ mapping to $\operatorname{{\mathop{\rm End}\nolimits}}{\mathcal{M}}$, or the augmentation ideal in ${\mathcal{O}}_{K}[G]$ where $G$ is a group acting on $\mathcal{M}$}.
 In the latter case, we write ${\mathcal{M}}^{G}$ instead of ${\mathcal{M}}[I]$.
@@
 We regard $\psi$ also as a character of $\mathbb{A}$$^{\times}$ and $G_{\mathbb{Q}}$ via the isomorphisms $\hat{{\mathbb{Z}}}^{\times}\cong{\mathbb{A}}^{\times}/{\mathbb{R}}_{>0}^{\times}{\mathbb{Q}}^{\times}\cong G_{\mathbb{Q}}^{\mathop{\rm ab}\nolimits}$, where the first isomorphism is induced by the natural inclusion $\hat{{\mathbb{Z}}}^{\times}\to{\mathbb{A}}^{\times}$ and the second is given by class field theory.
-(Our convention is that a uniformizer in ${\mathbb{Q}}_{p}^{\times}$ maps to $\operatorname{{\mathop{\rm Frob}\nolimits}}_{p}$ in the Galois group of any abelian extension of $\mathbb{Q}$} unramified at pp.)
+(Our convention is that a uniformizer in ${\mathbb{Q}}_{p}^{\times}$ maps to $\operatorname{{\mathop{\rm Frob}\nolimits}}_{p}$ in the Galois group of any abelian extension of $\mathbb{Q}$} unramified at $p$.)
 Suppose we are given an embedding of a number field $F\to\bar{{\mathbb{Q}}}$ such that $\psi$ is trivial on the image of $G_{F}$ in $G_{\mathbb{Q}}$.
@@
 One can check that the construction is independent of the choice of $F$ and embedding $F\to\bar{{\mathbb{Q}}}$ as above.
-To describe $M_{\psi}$ explicitly, we choose $F={\mathbb{Q}}(e^{2\pi i/N})\subset{\mathbb{C}}$ where $\psi$ has conductor NN.
+To describe $M_{\psi}$ explicitly, we choose $F={\mathbb{Q}}(e^{2\pi i/N})\subset{\mathbb{C}}$ where $\psi$ has conductor $N$.
 We let $\tau_{0}:F\to\bar{{\mathbb{Q}}}$ denote the embedding compatible with our fixed $\bar{{\mathbb{Q}}}\to{\mathbb{C}}$, and we regard $\psi$ as a Dirichlet character via the canonical isomorphism $({\mathbb{Z}}/N{\mathbb{Z}})^{\times}\cong G$.
-Let $S$ denote the set of primes in $K$ lying over those dividing NN and define an object ${\mathcal{M}}_{\psi}$ of ${\mathcal{P}M}_{K}^{S}$ by $({\mathcal{V}}\otimes{\mathcal{M}}_{F})^{G}$ where ${\mathcal{V}}={\mathcal{O}}_{K}^{S}$ with $G$ acting by $\psi$.
+Let $S$ denote the set of primes in $K$ lying over those dividing $N$ and define an object ${\mathcal{M}}_{\psi}$ of ${\mathcal{P}M}_{K}^{S}$ by $({\mathcal{V}}\otimes{\mathcal{M}}_{F})^{G}$ where ${\mathcal{V}}={\mathcal{O}}_{K}^{S}$ with $G$ acting by $\psi$.
 We then have:
@@
 2. Premotivic structures for modular forms
-In this section we review the construction of premotivic structures associated to the space of modular forms of weight $k$ and level NN.
+In this section we review the construction of premotivic structures associated to the space of modular forms of weight $k$ and level $N$.
 More precisely, if $k\geq 2$ and $N\geq 3$, we construct objects of ${\mathcal{P}M}_{\mathbb{Q}}^{S}$ whose de Rham realization contains the space of such forms, where $S=S_{N}$ is the set of primes dividing $Nk!$.
-2.1. Level NN modular curves
-Let $k$ and NN be integers with $k\geq 2$ and $N\geq 3$.
-Let $T=\operatorname{{\mathop{\rm Spec}\nolimits}}{\mathbb{Z}}[1/Nk!]$, and consider the functor which associates to a $T$-scheme $T^{\prime}$ the set of isomorphism classes of generalized elliptic curves over $T^{\prime}$ with level NN structure [D-R, IV.6.6].
+2.1. Level $N$ modular curves
+Let $k$ and $N$ be integers with $k\geq 2$ and $N\geq 3$.
+Let $T=\operatorname{{\mathop{\rm Spec}\nolimits}}{\mathbb{Z}}[1/Nk!]$, and consider the functor which associates to a $T$-scheme $T^{\prime}$ the set of isomorphism classes of generalized elliptic curves over $T^{\prime}$ with level $N$ structure [D-R, IV.6.6].
 By [D-R, IV.6.7], the functor is represented by a smooth, proper curve over $T$.
-We denote this curve by $\bar{X}$, and we let $\bar{s}:\bar{E}\to\bar{X}$ denote the universal generalized elliptic curve with level NN structure.
+We denote this curve by $\bar{X}$, and we let $\bar{s}:\bar{E}\to\bar{X}$ denote the universal generalized elliptic curve with level $N$ structure.
 We let $X$ denote the open subscheme of $\bar{X}$ over which $\bar{E}$ is smooth.
@@
 More precisely, for $\alpha>0$, we let $D_{\alpha}$ denote the disk in $D[q^{1/N}]$ defined by $|q|<e^{-2\pi\alpha}$ and let $D_{\alpha}^{0}$ denote the corresponding punctured disk.
-Then over a punctured neighborhood around each cusp in $\bar{X}_{N}$, we have an isomorphism between $E_{N}$ and the NN-sided analytic Tate curve $E_{q}^{\mathop{\rm an}\nolimits}$ over $D_{1}^{0}$, and we define $\bar{E}_{N}$ by gluing NN-sided analytic Tate curves over $D_{1}$ to $E_{N}$.
-For each $a$, we define a level NN-structure on $E_{N,t}$ by the pair of sections $X_{N,t}\to E_{N,t}$ induced by $\tau\mapsto(\tau,\tau/N)$ and $\tau\mapsto(\tau,t/N)$.
-The resulting level NN-structure on $E_{N}$ in fact extends to $\bar{E}_{N}$.
+Then over a punctured neighborhood around each cusp in $\bar{X}_{N}$, we have an isomorphism between $E_{N}$ and the $N$-sided analytic Tate curve $E_{q}^{\mathop{\rm an}\nolimits}$ over $D_{1}^{0}$, and we define $\bar{E}_{N}$ by gluing $N$-sided analytic Tate curves over $D_{1}$ to $E_{N}$.
+For each $a$, we define a level $N$-structure on $E_{N,t}$ by the pair of sections $X_{N,t}\to E_{N,t}$ induced by $\tau\mapsto(\tau,\tau/N)$ and $\tau\mapsto(\tau,t/N)$.
+The resulting level $N$-structure on $E_{N}$ in fact extends to $\bar{E}_{N}$.
 Using for example [Ha, App.
@@
 One can check the resulting morphism $\bar{X}_{N}^{\mathop{\rm alg}\nolimits}\to\bar{X}_{\mathbb{C}}$ induces a bijection $X_{N}\to X({\mathbb{C}})$, and since both are smooth, proper algebraic curves over $\mathbb{C}$}, the map is an isomorphism.
-The analytification of the universal generalized elliptic curve with level NN-structure is therefore isomorphic to $\bar{E}_{N}\to\bar{X}_{N}$ with the level NN-structure defined above.
-$2.$ Realizations of level NN modular forms
+The analytification of the universal generalized elliptic curve with level $N$-structure is therefore isomorphic to $\bar{E}_{N}\to\bar{X}_{N}$ with the level $N$-structure defined above.
+2.2. Realizations of level $N$ modular forms
 To construct the Betti realization, define ${\mathcal{F}}_{B}$ as the locally constant sheaf $R^{1}s^{\mathop{\rm an}\nolimits}_{*}{\mathbb{Z}}$ on $X^{\mathop{\rm an}\nolimits}$.
@@
 Proof.
-Over a neighborhood of each cusp in $\bar{X}^{\mathop{\rm an}\nolimits}$, $\bar{E}^{\mathop{\rm an}\nolimits}$ is isomorphic to the NN-sided analytic Tate curve $E_{q}^{\mathop{\rm an}\nolimits}$ over $D_{1}$.
+Over a neighborhood of each cusp in $\bar{X}^{\mathop{\rm an}\nolimits}$, $\bar{E}^{\mathop{\rm an}\nolimits}$ is isomorphic to the $N$-sided analytic Tate curve $E_{q}^{\mathop{\rm an}\nolimits}$ over $D_{1}$.
 We claim that there is an isomorphism between the pull-back of ${\mathcal{F}}_{{\mathop{\rm dR}\nolimits}}^{k,{\mathop{\rm an}\nolimits}}$ to $D_{1}$ and ${\mathcal{O}}_{D_{1}}^{k-1}$ such that the corresponding connection on ${\mathcal{O}}_{D_{1}}^{k-1}$ is given by $\nabla(f)=(A(t)f+tf^{\prime})\dfrac{dt}{t}$ where $A(t)\in M_{k-1}({\mathcal{O}}_{D_{1}})$ is such that $A(0)$ is nilpotent.
@@
 ${\mathbb{V}}({\mathcal{A}})={\mathbb{D}}({\mathcal{A}})^{*}={\mathop{\rm Hom}\nolimits}({\mathbb{D}}({\mathcal{A}}),{\mathbb{Q}}_{\ell}/{\mathbb{Z}}_{\ell}).$
-In the case of $Y=\bar{Y}=\operatorname{{\mathop{\rm Spec}\nolimits}}{\mathcal{O}}_{F}$ where $F$ is an unramified extension of $\mathbb{Q}$$_{\ell}$, Faltings’ category can be identified with the full subcategory of ${{\mathcal{M}\mathcal{F}}}^{0}_{\mathop{\rm tor}\nolimits}({\mathcal{O}}_{F})$ whose objects AA satisfy $\operatorname{{\mathop{\rm Fil}\nolimits}}^{a+1}A=0$.
+In the case of $Y=\bar{Y}=\operatorname{{\mathop{\rm Spec}\nolimits}}{\mathcal{O}}_{F}$ where $F$ is an unramified extension of $\mathbb{Q}$$_{\ell}$, Faltings’ category can be identified with the full subcategory of ${{\mathcal{M}\mathcal{F}}}^{0}_{\mathop{\rm tor}\nolimits}({\mathcal{O}}_{F})$ whose objects $A$ satisfy $\operatorname{{\mathop{\rm Fil}\nolimits}}^{a+1}A=0$.
 Furthermore, identifying the category of étale sheaves on $\operatorname{{\mathop{\rm Spec}\nolimits}}F$ with that of continuous $G_{F}$-modules, one finds that the functor $\mathbb{D}$} agrees with $\underline{U}_{S}$ (see [Fo3, 3.8]).
@@
 where ${\mathcal{M}}_{c,{\ell\mbox{\scriptsize-crys}}}$ has underlying filtered module ${\mathcal{M}}_{c,{\mathop{\rm dR}\nolimits}}$.
-2.3. Level NN modular forms
+2.3. Level $N$ modular forms
 We recall the relation between $M_{{\mathop{\rm dR}\nolimits}}$ and the space of modular forms of weight $k$ with respect to $\Gamma(N)$.
@@
 For $g\in{\mathbb{C}}\otimes\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}M_{c,{\mathop{\rm dR}\nolimits}}$ and $t\in({\mathbb{Z}}/N{\mathbb{Z}})^{\times}$, write $g_{t}$ for the image of $g$ in the corresponding component of $S_{k}(\Gamma(N))$, and similarly for $M_{\mathop{\rm dR}\nolimits}$ and $M_{k}(\Gamma(N))$.
-Lemma $2.$
+Lemma 2.2.
 If $g\in{\mathbb{C}}\otimes\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}M_{c,{\mathop{\rm dR}\nolimits}}$ and $h\in{\mathbb{C}}\otimes\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}M_{\mathop{\rm dR}\nolimits}$, then we have
@@
 $\left(\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}V\right)\,\cap\,\left((I^{\infty})^{-1}(F_{\infty}\otimes 1)I^{\infty}\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}V\right)=0,$
-where $V=\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}({\mathbb{C}}\otimes M_{c,{\mathop{\rm dR}\nolimits}})\cong\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}W_{k-1}({\mathbb{C}}\otimes M_{\mathop{\rm dR}\nolimits})$, and this is immediate from Lemma $2.$
+where $V=\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}({\mathbb{C}}\otimes M_{c,{\mathop{\rm dR}\nolimits}})\cong\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}W_{k-1}({\mathbb{C}}\otimes M_{\mathop{\rm dR}\nolimits})$, and this is immediate from Lemma 2.2.
 We can now regard $\mathcal{M}$} and ${\mathcal{M}}_{c}$ as objects of ${\mathcal{P}M}_{\mathbb{Q}}^{S}$, and $M={\mathbb{Q}}\otimes{\mathcal{M}}$ and $M_{c}={\mathbb{Q}}\otimes{\mathcal{M}}_{c}$ as objects of $\operatorname{{\bf PM}}_{\mathbb{Q}}^{S}$, where $S$ contains the set of primes dividing $Nk!$.
@@
 which holds if $k=2$ or $\ell$ does not divide N(k2)!N(k-2)!. ∎
-We now summarize main properties of the premotivic structures ${\mathcal{M}}_{N}$ and $M_{N}$ for level NN modular forms.
+We now summarize main properties of the premotivic structures ${\mathcal{M}}_{N}$ and $M_{N}$ for level $N$ modular forms.
 Theorem 2.4.
@@
 Suppose now that $(h,N,N^{\prime})$ is an admissible triple.
-Let $\bar{E}/\bar{X}$ (respectively, $\bar{E}^{\prime}/\bar{X}^{\prime}$) denote the universal generalized elliptic curve with level NN (respectively $N^{\prime}$) structure.
+Let $\bar{E}/\bar{X}$ (respectively, $\bar{E}^{\prime}/\bar{X}^{\prime}$) denote the universal generalized elliptic curve with level $N$ (respectively $N^{\prime}$) structure.
 […]
 $({\mathbb{Z}}/N{\mathbb{Z}})^{2}\rightarrow({\mathbb{Z}}/N^{\prime}{\mathbb{Z}})^{2}/(({\mathbb{Z}}/N^{\prime}{\mathbb{Z}})^{2}(N^{\prime}h^{-1})$
-and so gives rise to a level NN-structure on $E^{\prime}/G$.
-Letting $(\bar{E}^{\prime}/G)_{\rm cont}$ denote the contraction of $\bar{E}^{\prime}/G$ whose cuspidal fibers are NN-gons [D-R, IV.1], we obtain a level NN structure on $(\bar{E}^{\prime}/G)_{\rm cont}\to\bar{X}^{\prime}$.
-By the universal property of $\bar{E}_{/\bar{X}}$, this defines a map $\chi_{h}=\chi_{h,N,N^{\prime}}:\bar{X}^{\prime}\to\bar{X}$ such that there is an isomorphism $(\bar{E}^{\prime}/G)_{\rm cont}\to\bar{E}\times_{\bar{X}}\bar{X}^{\prime}$ of generalized elliptic curves with level NN-structures.
+and so gives rise to a level $N$-structure on $E^{\prime}/G$.
+Letting $(\bar{E}^{\prime}/G)_{\rm cont}$ denote the contraction of $\bar{E}^{\prime}/G$ whose cuspidal fibers are $N$-gons [D-R, IV.1], we obtain a level $N$ structure on $(\bar{E}^{\prime}/G)_{\rm cont}\to\bar{X}^{\prime}$.
+By the universal property of $\bar{E}_{/\bar{X}}$, this defines a map $\chi_{h}=\chi_{h,N,N^{\prime}}:\bar{X}^{\prime}\to\bar{X}$ such that there is an isomorphism $(\bar{E}^{\prime}/G)_{\rm cont}\to\bar{E}\times_{\bar{X}}\bar{X}^{\prime}$ of generalized elliptic curves with level $N$-structures.
 We let $\varepsilon_{h}=\varepsilon_{h,N,N^{\prime}}$ denote the composite of the natural map $\bar{E}^{\prime}\to(\bar{E}^{\prime}/G)_{\rm cont}$ with this isomorphism, so we obtain a commutative diagram
@@
 $\tilde{\gamma}_{\tau^{\prime}}:E_{\tau^{\prime}}\to E_{\tau}$
-induces an isomorphism $E_{\tau^{\prime}}/G_{\tau^{\prime}}\cong E_{\tau}$ compatible with level NN-structures.
+induces an isomorphism $E_{\tau^{\prime}}/G_{\tau^{\prime}}\cong E_{\tau}$ compatible with level $N$-structures.
 Recall that the level $N^{\prime}$-structure on $E_{\tau^{\prime}}$ is defined by
@@
 $G_{\tau^{\prime}}={\mathbb{Z}}^{2}\gamma{\mbox{\scriptsize$\left(\begin{array}[]{c}{{\tau^{\prime}}}\\ {{1}}\end{array}\right)$}}/\Lambda_{\tau^{\prime}}.$
-Furthermore, since multiplication by $u$ is the identity on $({\mathbb{Z}}/N{\mathbb{Z}})^{2}$, the level NN-structure on $E_{\tau^{\prime}}/G_{\tau^{\prime}}$ is given by
+Furthermore, since multiplication by $u$ is the identity on $({\mathbb{Z}}/N{\mathbb{Z}})^{2}$, the level $N$-structure on $E_{\tau^{\prime}}/G_{\tau^{\prime}}$ is given by
 $(m,n)g_{\kappa(t^{\prime})}\bmod N\hat{{\mathbb{Z}}}^{2}\ \mapsto\ (m,n)N^{-1}\gamma{\mbox{\scriptsize$\left(\begin{array}[]{c}{{\tau}}\\ {{1}}\end{array}\right)$}}\bmod{\mathbb{Z}}^{2}\gamma{\mbox{\scriptsize$\left(\begin{array}[]{c}{{\tau^{\prime}}}\\ {{1}}\end{array}\right)$}},$
 and the lemma follows. ∎
-Lemma $3.$
+Lemma 3.3.
 Suppose that $(h,N,N^{\prime})$ and $(h^{\prime},N^{\prime},N^{\prime\prime})$ are admissible triples.
@@
 So we can take $\gamma^{\prime\prime}=\gamma\gamma^{\prime}$ and apply Lemma 3.2 to conclude the desired equalities. ∎
-$3.$ Action on the premotivic structures
+3.3. Action on the premotivic structures
 Suppose that $(h,N,N^{\prime})$ is an admissible triple.
 Let $S$ denote the set of primes dividing $Nk!$ and $S^{\prime}$ the set of primes dividing $N^{\prime}k!$.
-We let $\mathcal{M}$}, ${\mathcal{M}}_{c}$ and $\mathcal{M}$$_{!}$ (respectively, $\mathcal{M}$$^{\prime}$, ${\mathcal{M}}^{\prime}_{c}$ and $\mathcal{M}$$^{\prime}_{!}$) denote the objects of ${\mathcal{P}M}_{\mathbb{Q}}^{S}$ (respectively, ${\mathcal{P}M}_{\mathbb{Q}}^{S^{\prime}}$) structures defined in §2.5 associated to modular forms of weight $k$ and level NN (respectively, level $N^{\prime}$).
+We let $\mathcal{M}$}, ${\mathcal{M}}_{c}$ and $\mathcal{M}$$_{!}$ (respectively, $\mathcal{M}$$^{\prime}$, ${\mathcal{M}}^{\prime}_{c}$ and $\mathcal{M}$$^{\prime}_{!}$) denote the objects of ${\mathcal{P}M}_{\mathbb{Q}}^{S}$ (respectively, ${\mathcal{P}M}_{\mathbb{Q}}^{S^{\prime}}$) structures defined in §2.5 associated to modular forms of weight $k$ and level $N$ (respectively, level $N^{\prime}$).
 We shall define compatible morphisms ${\mathcal{M}}_{c}^{S^{\prime}}\to{\mathcal{M}}_{c}^{\prime}$ and ${\mathcal{M}}^{S^{\prime}}\to{\mathcal{M}}^{\prime}$, giving also ${\mathcal{M}}_{!}^{S^{\prime}}\to{\mathcal{M}}_{!}^{\prime}$.
@@
 $\begin{array}[]{ccccccccc}0&\to&\varepsilon_{h}^{-1}\omega^{\bullet}_{\bar{E}/\bar{X}}\otimes_{\varepsilon_{h}^{-1}{\mathcal{O}}_{\bar{E}}}\varepsilon_{h}^{-1}\bar{s}^{*}\omega^{1}_{\bar{X}/T}&\to&\varepsilon_{h}^{-1}\omega^{\bullet}_{\bar{E}/T}&\to&\varepsilon_{h}^{-1}\omega^{\bullet}_{\bar{E}/\bar{X}}&\to&0\\ &&\downarrow&&\downarrow&&\downarrow&&\\ 0&\to&\omega^{\bullet}_{\bar{E}^{\prime}/\bar{X}^{\prime}}\otimes_{{\mathcal{O}}_{\bar{E}^{\prime}}}\bar{s}^{\prime,*}\omega^{1}_{\bar{X}^{\prime}/T^{\prime}}&\to&\omega^{\bullet}_{\bar{E}^{\prime}/T^{\prime}}&\to&\omega^{\bullet}_{\bar{E}^{\prime}/\bar{X}^{\prime}}&\to&0,\end{array}$
-where the top row is obtained by applying $\varepsilon_{h}^{-1}$ to (3), the bottom row is (3) with $N^{\prime}$ instead of NN, and the vertical arrows are given by the canonical $\varepsilon_{h}^{-1}{\mathcal{O}}_{\bar{E}}$-linear maps in each degree.
+where the top row is obtained by applying $\varepsilon_{h}^{-1}$ to (3), the bottom row is (3) with $N^{\prime}$ instead of $N$, and the vertical arrows are given by the canonical $\varepsilon_{h}^{-1}{\mathcal{O}}_{\bar{E}}$-linear maps in each degree.
 Applying ${\bf R}^{\bullet}\bar{s}^{\prime}_{*}$ and using the natural transformation $\chi_{h}^{-1}\bar{s}_{*}\to\bar{s}^{\prime}_{*}\varepsilon_{h}^{-1}$ gives a commutative diagram
@@
 Theorem 3.6.
-Fix positive integers $k\geq 2$, and NN, $N^{\prime}\geq 3$.
+Fix positive integers $k\geq 2$, and $N$, $N^{\prime}\geq 3$.
 Let $g$ be an element in $\operatorname{{\mathop{\rm GL}\nolimits}}_{2}({\mathbb{A}}_{\bf f})$ such that $g^{-1}U_{N^{\prime}}g\subseteq U_{N}$.
@@
 Combining (14) and (15) we see that if $(\beta^{-1}\circ\alpha)(\theta)=\phi$, then $(\beta^{-1}\circ\alpha)(h\theta)=h\phi$.
-Further let $r\in{\mathbb{Z}}$. Then the action of rIrI on both spaces are multiplication by $r^{k-2}$. This proves the proposition. ∎
+Further let $r\in{\mathbb{Z}}$. Then the action of $rI$ on both spaces are multiplication by $r^{k-2}$. This proves the proposition. ∎
 3.5. Compatibility with the pairings
@@
 Then we have $g]:{\mathcal{M}}_{N}\to{\mathcal{M}}_{N^{\prime}}$ and $g]:{\mathcal{M}}_{M}\to{\mathcal{M}}_{M^{\prime}}$.
-Let II be the identity matrix in $\operatorname{{\mathop{\rm GL}\nolimits}}_{2}({\mathbb{A}}_{\bf f})$.
+Let $I$ be the identity matrix in $\operatorname{{\mathop{\rm GL}\nolimits}}_{2}({\mathbb{A}}_{\bf f})$.
 From $N|M$ we get $I^{-1}U_{M}I\subset U_{N}$ and we have
@@
 ∎
-4. The premotivic structure for forms of level NN and character $\psi$
+4. The premotivic structure for forms of level $N$ and character $\psi$
 4.1. $\sigma$-constructions
@@
 The proof for (b) is the same. ∎
-Because of the lemma, we will suppress NN in the notations by using ${\mathcal{M}}(\sigma)_{c}$, ${\mathcal{M}}(\sigma)_{!}$ and $M(\sigma)$ when there is no danger of confusion.
+Because of the lemma, we will suppress $N$ in the notations by using ${\mathcal{M}}(\sigma)_{c}$, ${\mathcal{M}}(\sigma)_{!}$ and $M(\sigma)$ when there is no danger of confusion.
 4.2. Action of double cosets
@@
   • $S_{N^{\prime}}^{K}\subset S$.
-To see that this is possible, first choose NN divisible by 44 so that $S_{N}^{K}\subset S$ and the second condition holds, then choose $N^{\prime}$ so that the third condition holds, and then for each $\ell\not\in S$, replace $N^{\prime}$ by $N^{\prime}/\ell^{r}$ where $\ell^{r}||N^{\prime}$.
+To see that this is possible, first choose $N$ divisible by $4$ so that $S_{N}^{K}\subset S$ and the second condition holds, then choose $N^{\prime}$ so that the third condition holds, and then for each $\ell\not\in S$, replace $N^{\prime}$ by $N^{\prime}/\ell^{r}$ where $\ell^{r}||N^{\prime}$.
 The condition that $S_{N}^{K}\subset S$ and $g_{\ell}\in\operatorname{{\mathop{\rm GL}\nolimits}}_{2}({\mathbb{Z}}_{\ell})$ ensures that we still have $g^{-1}U_{N^{\prime}}g\subset U_{N}.$ Note that $N|N^{\prime}$ so $N^{\prime}\geq 3$, $U_{N^{\prime}}\subset U_{N}\subset\ker\sigma\cap\ker\sigma^{\prime}$ and $S_{N}^{K}\subset S_{N^{\prime}}^{K}\subset S$.
@@
 We will give a formula for the composition $U^{\prime\prime}g^{\prime}U^{\prime}]_{\tau^{\prime},N^{\prime},N^{\prime\prime},\sharp}\circ[U^{\prime}gU]_{\tau,N,N^{\prime},\sharp}$ for $\sharp\in\{\phi,c,!\}$.
-To simplify the notation, we will suppress NN, $N^{\prime}$ and $\sharp$ when there is no danger of confusion.
+To simplify the notation, we will suppress $N$, $N^{\prime}$ and $\sharp$ when there is no danger of confusion.
 We first recall the definition of the product formula for double cosets [Sh1, §3.1],[Miy, §2.7].
@@
 Thus $U^{\prime\prime}g^{\prime}u_{k}gU]_{\tau_{k}}$ is well-defined.
-Proposition $4.$
+Proposition 4.4.
 $U^{\prime\prime}g^{\prime}U^{\prime}]_{\tau^{\prime}}\circ[U^{\prime}gU]_{\tau}=\sum_{k}[U^{\prime\prime}g^{\prime}u_{k}gU]_{\tau_{k}}.$
@@
 This proves the proposition. ∎
-Using Lemma 3.9 and Proposition 4.4, we conclude that $U^{\prime}gU]$ is independent of the choice of NN and $N^{\prime}$.
+Using Lemma 3.9 and Proposition 4.4, we conclude that $U^{\prime}gU]$ is independent of the choice of $N$ and $N^{\prime}$.
 More precisely, if $U^{\prime}gU]_{\tau,M,M^{\prime}}$ is also defined, then $U^{\prime}gU]_{\tau,N,N^{\prime}}^{S}=[U^{\prime}gU]_{\tau,M,M^{\prime}}^{S}$ when $S\supset S_{\operatorname{{\mathop{\rm lcm}\nolimits}}(N^{\prime},M^{\prime})}$.
-$4.$ Pairing on $\sigma$-constructions
+4.4. Pairing on $\sigma$-constructions
 In this section we will construct perfect pairings between certain $\sigma$-constructions and compute the adjoints of double coset actions with respect to the pairings.
@@
 Lemma 4.5.
-For NN, $N^{\prime}$ and $g$ as above, we have the commutative diagram
+For $N$, $N^{\prime}$ and $g$ as above, we have the commutative diagram
 […]
@@
 Lemma 4.6.
-Suppose NN is such that $N\geq 3$, $U_{N}\subset U$, the conductor of $\psi$ divides NN. Then the morphism ${\mathcal{M}}_{F}\to{\mathcal{H}}$ induces an isomorphism
+Suppose $N$ is such that $N\geq 3$, $U_{N}\subset U$, the conductor of $\psi$ divides $N$. Then the morphism ${\mathcal{M}}_{F}\to{\mathcal{H}}$ induces an isomorphism
 ${\mathcal{M}}_{\psi}^{S_{N}}\to({\mathcal{O}}_{\psi}\otimes{\mathcal{H}})^{U}$
@@
 Given a representation $\sigma:U\to\operatorname{\operatorname{{\mathop{\rm Aut}\nolimits}}}_{{\mathcal{O}}_{K}}{\mathcal{V}}$, let $\hat{\sigma}\otimes(\psi^{-1}\circ{\mathop{\rm det}\nolimits})$ denote the representation defined by the action of $U$ on ${\mathop{\rm Hom}\nolimits}_{{\mathcal{O}}_{K}}({\mathcal{V}},{\mathcal{O}}_{\psi})$.
-For $N\geq 3$ such that $U_{N}\subset\ker\sigma$ and the conductor of $\psi$ divides NN, Lemma (4.7) and the duality homomorphism $\delta^{L}_{N}:{\mathcal{M}}_{N,c}\to{\mathop{\rm Hom}\nolimits}({\mathcal{M}}_{N},{\mathcal{T}}^{S_{N},\otimes(k-1)})$ from §2.5 give a morphism
+For $N\geq 3$ such that $U_{N}\subset\ker\sigma$ and the conductor of $\psi$ divides $N$, Lemma (4.7) and the duality homomorphism $\delta^{L}_{N}:{\mathcal{M}}_{N,c}\to{\mathop{\rm Hom}\nolimits}({\mathcal{M}}_{N},{\mathcal{T}}^{S_{N},\otimes(k-1)})$ from §2.5 give a morphism
 (27) $\begin{array}[]{l}{\mathop{\rm Hom}\nolimits}_{{\mathcal{O}}_{K}}({\mathcal{V}},{\mathcal{O}}_{\psi^{-1}})\otimes{\mathcal{M}}_{N,c}\cong{\mathop{\rm Hom}\nolimits}_{{\mathcal{O}}_{K}}({\mathcal{V}},{\mathcal{M}}_{N,c}\otimes{\mathcal{M}}_{\psi}^{S_{N}^{K}})\\ \to{\mathop{\rm Hom}\nolimits}_{{\mathcal{O}}_{K}}({\mathcal{V}},{\mathop{\rm Hom}\nolimits}_{{\mathcal{O}}_{K}}({\mathcal{M}}_{N},{\mathcal{T}}^{S_{N},\otimes(k-1)}\otimes{\mathcal{M}}_{\psi}^{S_{N}^{K}}))\\ \cong{\mathop{\rm Hom}\nolimits}_{{\mathcal{O}}_{K}}({\mathcal{V}}\otimes{\mathcal{M}}_{N},{\mathcal{M}}_{\psi}(1-k)),\end{array}$
@@
 so $\bar{\delta}^{L}_{N}$ is an isomorphism. The compatibility with change of levels follows from applying Proposition 3.8 and (25) with $g=I$. The proof for $\bar{\delta}^{L}_{N,!}$ is the same. ∎
-We will therefore omit the subscript NN in the notation for the morphisms $\bar{\delta}^{L}$.
-(Note however that $\delta^{L}_{N}$ depends on NN, and $\delta^{L}_{N,!}$ depends on NN only in the normalization.)
+We will therefore omit the subscript $N$ in the notation for the morphisms $\bar{\delta}^{L}$.
+(Note however that $\delta^{L}_{N}$ depends on $N$, and $\delta^{L}_{N,!}$ depends on $N$ only in the normalization.)
 […]
@@
 We only need to prove the assertion for Betti realizations, for then $\delta^{L}_{N}=[U:U_{N}]\tilde{\delta}^{L}$ for a unique morphism $\tilde{\delta}^{L}$ with the desired property.
-To do this, we give an alternate description of the Betti realizations and pairings which is independent of NN.
+To do this, we give an alternate description of the Betti realizations and pairings which is independent of $N$.
 and define a perfect pairing on the new Betti realizations.
@@
 Case 3. In this case $\tau:V\to V^{\prime}$ is a morphism of $U$-modules and the formula follows directly from the definitions. ∎
-4.5. Premotivic structure of level NN and character $\psi$
+4.5. Premotivic structure of level $N$ and character $\psi$
 […]
@@
 for all ${\mbox{\scriptsize$\left(\begin{array}[]{cc}{{a}}&{{b}}\\ {{c}}&{{d}}\end{array}\right)$}}\in U_{0}(N)$.
-It is straightforward to check that this is precisely the space $M_{k}(N,\psi)$ of classical modular forms of weight $k$, level NN and character $\psi$, i.e., the set of forms $f\in M_{k}(\Gamma(M))$ such that
+It is straightforward to check that this is precisely the space $M_{k}(N,\psi)$ of classical modular forms of weight $k$, level $N$ and character $\psi$, i.e., the set of forms $f\in M_{k}(\Gamma(M))$ such that
 $f\left(\frac{a\tau+b}{c\tau+d}\right)=\psi(d)(c\tau+d)^{k}f(\tau)$
@@
 Using the notations in §4.5 and letting $S$ be a superset of $S_{N}^{K}$, we now consider the action of Hecke operators on ${\mathcal{M}}(N,\psi)_{M,!}$ as an object in ${\mathcal{P}M}^{S}_{K}$.
-Fix a rational prime pp.
+Fix a rational prime $p$.
 By Lemma 4.2, $U{\mbox{\scriptsize$\left(\begin{array}[]{cc}{{p}}&{{0}}\\ {{0}}&{{1}}\end{array}\right)$}}_{p}U]_{\psi(p_{p})^{-1}}$ defines an endomorphism of ${\mathcal{M}}(N,\psi)_{M,!}$ regarded as an object in ${\mathcal{P}M}^{S_{pN}}$.
@@
 $U(-w)U]_{\omega^{\prime}}\circ[UwU]_{\omega}=[U(NI)U]_{1}=N^{k-2}.$
-Since NN is invertible in $\mathcal{O}$}, $UwU]_{\omega}$ is an isomorphism in ${\mathcal{P}M}^{S_{N}^{K}}_{K}$ and
+Since $N$ is invertible in $\mathcal{O}$}, $UwU]_{\omega}$ is an isomorphism in ${\mathcal{P}M}^{S_{N}^{K}}_{K}$ and
 $UwU]_{\omega}^{-1}=N^{2-k}[U(-w)U]_{\omega^{\prime}}=[Uw^{-1}U]_{\omega^{t}\otimes\psi({\mathop{\rm det}\nolimits}w)}.$
@@
 Then ${\mathbb{T}}\subset\operatorname{{\mathop{\rm End}\nolimits}}_{K}(\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}{\mathcal{M}}(N,\psi)_{M,!,{\mathop{\rm dR}\nolimits}})$ and ${\mathbb{T}}^{\prime}\subset\operatorname{{\mathop{\rm End}\nolimits}}_{K}(\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}M(N,\psi)_{M,{\mathop{\rm dR}\nolimits}})$ .
-Lemma $5.$
+Lemma 5.5.
   1. (a)
@@
 We next show that $\mathbb{T}$} acts on ${\mathcal{M}}(N,\psi)_{M,!}$ as an object in ${\mathcal{P}M}^{S_{N}^{K}}_{K}$.
-Fix a finite prime pp.
+Fix a finite prime $p$.
 Recall that ${\mathcal{M}}(N,\psi)_{M,!}$ is in ${\mathcal{P}M}^{S_{N}^{K}}_{K}$ and, by Lemma 4.2, $T_{p}:{\mathcal{M}}(N,\psi)_{M,!}\to{\mathcal{M}}(N,\psi)_{pM,!}$ is defined in ${\mathcal{P}M}^{S_{pM}^{K}}_{K}={\mathcal{P}M}^{S_{pN}^{K}}_{K}$.
@@
 ∎
-$5.$ The $L$-function
+5.5. The $L$-function
 Suppose that $f(\tau)=\sum a_{n}e^{2\pi in\tau}$ is a newform of weight $k$, conductor $N_{f}$ and character $\psi_{f}$. Associated to $f$ is the $L$-function with Euler product factorization:
@@
 Then we have $U^{\prime}_{1}=U^{\prime}$.
-So in the double coset decomposition $U^{\prime}=\coprod_{k}U^{\prime}_{2}u_{k}U^{\prime}_{1}$ there is only one $u_{k}$ which we can take II.
+So in the double coset decomposition $U^{\prime}=\coprod_{k}U^{\prime}_{2}u_{k}U^{\prime}_{1}$ there is only one $u_{k}$ which we can take $I$.
 Further $W_{1}=W_{1}^{\prime}=U^{\prime\prime}$ and $\tau_{1}=1$.
@@
 For this we first display the following elementary relations.
-Fix a prime number pp.
+Fix a prime number $p$.
 For each integer $i\geq 0$, let $V_{i}=\{{\mbox{\scriptsize$\left(\begin{array}[]{cc}{{a}}&{{b}}\\ {{c}}&{{d}}\end{array}\right)$}}\in\operatorname{{\mathop{\rm GL}\nolimits}}_{2}({\mathbb{Z}}_{p})|c\equiv 0\mod p^{i}\}.$ Let $A_{t}={\mbox{\scriptsize$\left(\begin{array}[]{cc}{{1}}&{{0}}\\ {{0}}&{{p^{t}}}\end{array}\right)$}},t=0,1,2$.
@@
 $\gamma=\sum_{i=1}^{r}\sum_{e_{i}=0}^{\delta_{p_{i}}}\gamma_{\prod_{i}p_{i}^{e_{i}}}^{\Sigma_{0},\Sigma_{r}}=\sum_{i=1}^{r}\sum_{e_{i}=0}^{\delta_{p_{i}}}\circ_{i=1}^{r}\gamma_{p_{i}^{e_{i}}}^{\Sigma_{i-1},\Sigma_{i}}=\circ_{i=0}^{r}(\sum_{e_{i}=0}^{\delta_{p_{i}}}\gamma_{p_{i}^{e_{i}}}^{\Sigma_{i-1},\Sigma_{i}}).$
-Thus to prove Proposition 6.3, we only need to prove it in the case when $\Sigma$ contains one prime pp.
+Thus to prove Proposition 6.3, we only need to prove it in the case when $\Sigma$ contains one prime $p$.
 Let $\Sigma_{0}=\phi$ and $\Sigma_{1}=\Sigma$.
-We consider three cases depending on $\delta_{p}=0,1$ or 22.
+We consider three cases depending on $\delta_{p}=0,1$ or $2$.
 Case 1: When $\delta_{p}=0$, then $N^{\Sigma}=N$ and $\gamma$ is just the identity map. We have also seen that $a_{p}=0$. Thus $a_{pn}=0$. So $f=f^{\Sigma}$ and the proposition is proved in this case.
@@
 $=\gamma\circ T_{q}.$
-When $q=p$, by (41) and (42), the double coset product of the local factor at primes other than $p$ involves II and is obviously zero. The local factor at $p$ is
+When $q=p$, by (41) and (42), the double coset product of the local factor at primes other than $p$ involves $I$ and is obviously zero. The local factor at $p$ is
 $[V_{i+1}BV_{i+1}]([V_{i+1}A_{0}V_{i}][V_{i}B_{0}V_{i}]-p^{1-k}[V_{i+1}A_{1}V_{i}][V_{i+1}BV_{i}])$
@@
 First $\epsilon_{1}\circ\phi_{1}=[U{{}^{\Sigma}}IU]_{1}$ acts on ${\mathcal{M}}(N,\psi)_{M}\subseteq{\mathcal{M}}_{M}\otimes{\mathcal{V}}$ by $I\otimes 1$, where $M\geq 3$ is chosen such that $N|M$ and $S_{M}=S_{N}$.
-The action corresponding to II on the classical forms is II.
+The action corresponding to $I$ on the classical forms is $I$.
 Since $f$ has rational coefficients, its image in $\oplus_{t\in({\mathbb{Z}}/M{\mathbb{Z}})^{\times}}M(M,\psi)$ is of the form $(f_{0}(z))_{t}$ with $f_{0}$ independent of $t$.
@@
 When $q\neq p$, the relation $T_{q}\gamma=\gamma T_{q}$ is proved in the same way as in Case 2.
-When $q=p$, we only need to consider the local factor of $T_{p}\gamma$ at the prime pp.
+When $q=p$, we only need to consider the local factor of $T_{p}\gamma$ at the prime $p$.
 The double coset product of the local factor is
@@
 Proof.
-As in the proof of Proposition 6.3, we can again specialize to the case when $\Sigma$ has a single prime pp. Then we only need to consider the place at pp.
+As in the proof of Proposition 6.3, we can again specialize to the case when $\Sigma$ has a single prime $p$. Then we only need to consider the place at $p$.
 When $\delta_{p}=0$, $\gamma$ is the identity map and the formula is clear.
@@
 Note that $\eta_{f}^{\Sigma}$ is well-behaved under extension of scalars ${\mathcal{O}}^{\prime}\otimes_{\mathcal{O}}\cdot$ if $K\subset K^{\prime}$ and $S^{\prime}$ is contained in the set of primes dividing those in $S$.
-Proposition $6.$
+Proposition 6.6.
 If ${\mathcal{M}}_{f,\lambda}/\lambda$ is an irreducible $({\mathcal{O}}/\lambda)[G_{\mathbb{Q}}]$-module for every $\lambda$ in $S$, then the map $M_{f}\to M_{f}^{\Sigma}$ induces an isomorphism ${\mathcal{M}}_{f}\to{\mathcal{M}}_{f}^{\Sigma}$ in ${\mathcal{P}M}^{S}$.
@@
 $\varepsilon:{\mathcal{O}}_{F}^{\times}\to({\mathcal{O}}_{F}/p^{c_{p}/2})^{\times}\to{\mathbb{C}}^{\times},$
-where $V_{p}^{\prime}$ denotes the set of matrices in $\operatorname{{\mathop{\rm GL}\nolimits}}_{2}({\mathbb{Z}}_{p})$ congruent to II mod $p^{c_{p}/2}$, and $\Theta(\varepsilon)$ is defined in §3.2 of [C-D-T].
+where $V_{p}^{\prime}$ denotes the set of matrices in $\operatorname{{\mathop{\rm GL}\nolimits}}_{2}({\mathbb{Z}}_{p})$ congruent to $I$ mod $p^{c_{p}/2}$, and $\Theta(\varepsilon)$ is defined in §3.2 of [C-D-T].
 Our hypotheses ensure that $\varepsilon/\varepsilon\circ\operatorname{{\mathop{\rm Frob}\nolimits}}_{p}$ has conductor $(p{\mathcal{O}}_{F})^{c_{p}/2}$.
@@
 where ${\mathcal{V}}^{\Sigma}_{p,\Sigma_{0}}={\mathcal{V}}_{p}^{0}$ if $p\in\Sigma_{0}-\Sigma$ and ${\mathcal{V}}_{p}^{1}$ (restricted to $U_{p}$) otherwise.
-Here the tensor product is over all primes $p$, but the action is trivial for all but finitely many pp.
+Here the tensor product is over all primes $p$, but the action is trivial for all but finitely many $p$.
 Note that if $\Sigma_{0}\subset\Sigma$, then $U^{\Sigma}_{\Sigma_{0}}=U^{\Sigma}$ and $\sigma^{\Sigma}_{\Sigma_{0}}=\sigma(N,\psi)$.
@@
 ∎
-It follows from the lemma and the definitions that $\gamma_{\Sigma_{0}}^{\Sigma,\Sigma\cup\Sigma_{0}}$ is injective and commutes with the action of the Hecke operators $T_{p}$ for all primes pp.
+It follows from the lemma and the definitions that $\gamma_{\Sigma_{0}}^{\Sigma,\Sigma\cup\Sigma_{0}}$ is injective and commutes with the action of the Hecke operators $T_{p}$ for all primes $p$.
 In particular, for $p\in S$, the operator $T_{p}$ can be written as a polynomial in $T_{q}$ for $q\neq p$, and the same arguments as in Proposition 5.6 show that $\mathbb{T}$} acts on ${\mathcal{M}}(\sigma_{\Sigma_{0}}^{\Sigma})_{!}$ as an object of ${\mathcal{P}M}^{S}_{K}$.
@@
 $A=(K/{\mathcal{O}})\otimes_{\mathcal{O}}{\mathop{\rm Hom}\nolimits}_{\mathcal{O}}(L_{1},L_{2}).$
-Below we shall define an $\mathcal{O}$}-submodule $H^{1}_{w}(G_{p},A)\subset H^{1}(G_{p},A)$ for each prime pp. For each finite set of primes $\Sigma\subset\Sigma_{0}$, we define
+Below we shall define an $\mathcal{O}$}-submodule $H^{1}_{w}(G_{p},A)\subset H^{1}(G_{p},A)$ for each prime $p$. For each finite set of primes $\Sigma\subset\Sigma_{0}$, we define
 $H^{1}_{\Sigma}(G_{\mathbb{Q}},A)\subset H^{1}(G_{\mathbb{Q}},A)$
@@
 Using the stability of $\mathcal{F}\mathcal{R}$ under direct sums, subobjects and quotients, one checks that $H^{1}_{w}(G_{\ell},A_{n})$ is an $\mathcal{O}$}-submodule, and furthermore that it is the preimage of $H^{1}_{w}(G_{\ell},A_{n+1})$ under the natural map.
-Now define $H^{1}_{w}(G_{p},A)=\displaystyle{\lim_{\longrightarrow}}\,H^{1}_{w}(G_{p},A_{n})$ for every prime pp.
+Now define $H^{1}_{w}(G_{p},A)=\displaystyle{\lim_{\longrightarrow}}\,H^{1}_{w}(G_{p},A_{n})$ for every prime $p$.
 Recall that Bloch and Kato define a divisible $\mathcal{O}$}-submodule $H^{1}_{\bf f}(G_{p},A)\subset H^{1}(G_{p},A)$ for each $p$ (see [B-K] or [F-P]). Their Selmer group
 $H^{1}_{\bf f}(G_{\mathbb{Q}},A)\subset H^{1}(G_{\mathbb{Q}},A)$
-is then the $\mathcal{O}$}-submodule of elements with restrictions in $H^{1}_{\bf f}(G_{p},A)$ for all primes pp. If $p\neq\ell$, it is clear that
+is then the $\mathcal{O}$}-submodule of elements with restrictions in $H^{1}_{\bf f}(G_{p},A)$ for all primes $p$. If $p\neq\ell$, it is clear that
 $H^{1}_{\bf f}(G_{p},A)\subset H^{1}_{w}(G_{p},A)$
@@
 whose construction we leave as an exercise (see also diagram (68) for a similar computation). ∎
-Now we specialize to the case where $\rho_{1}=\rho_{2}$, each having dimension 22 over $K$ with $\dim_{K}\operatorname{{\mathop{\rm Fil}\nolimits}}^{1}D_{\mbox{\scriptsize crys}}(V_{i})$ =1=1 (so also $\ell>2$).
+Now we specialize to the case where $\rho_{1}=\rho_{2}$, each having dimension $2$ over $K$ with $\dim_{K}\operatorname{{\mathop{\rm Fil}\nolimits}}^{1}D_{\mbox{\scriptsize crys}}(V_{i})$ =1=1 (so also $\ell>2$).
 We assume also that $L_{1}=L_{2}$, and let
@@
 where ${\mathop{\rm ad}\nolimits}^{0}_{\mathcal{O}}L_{1}$ denotes the kernel of the trace map $\operatorname{{\mathop{\rm End}\nolimits}}_{\mathcal{O}}L\to{\mathcal{O}}$.
-We let $A^{0}_{n}=A^{0}[\lambda^{n+1}]$, identify $A^{0}$ with a submodule of AA, and define $H^{1}_{w}(G_{p},A^{0}_{n})$, $H^{1}_{w}(G_{p},A^{0})$ and $H^{1}_{w}(G_{\mathbb{Q}},A^{0})$ as the preimage of the corresponding group for AA.
+We let $A^{0}_{n}=A^{0}[\lambda^{n+1}]$, identify $A^{0}$ with a submodule of $A$, and define $H^{1}_{w}(G_{p},A^{0}_{n})$, $H^{1}_{w}(G_{p},A^{0})$ and $H^{1}_{w}(G_{\mathbb{Q}},A^{0})$ as the preimage of the corresponding group for $A$.
 From the proposition above, we get
@@
 We let $\mathcal{C}$} denote the category of complete local Noetherian $\mathcal{O}$}-algebras.
-Recall that if AA is an object of $\mathcal{C}$} with maximal ideal $\mathfrak{m}$}, then an AA-deformation of $V_{0}$ is an isomorphism class of free AA-modules $M$ endowed with continuous $AG_{\mathbb{Q}}$-action such that $M/{\mathfrak{m}}M$ is $(A/{\mathfrak{m}})G_{\mathbb{Q}}$-isomorphic to $(A/{\mathfrak{m}})\otimes_{\kappa}V_{0}$.
+Recall that if $A$ is an object of $\mathcal{C}$} with maximal ideal $\mathfrak{m}$}, then an $A$-deformation of $V_{0}$ is an isomorphism class of free $A$-modules $M$ endowed with continuous $AG_{\mathbb{Q}}$-action such that $M/{\mathfrak{m}}M$ is $(A/{\mathfrak{m}})G_{\mathbb{Q}}$-isomorphic to $(A/{\mathfrak{m}})\otimes_{\kappa}V_{0}$.
 Suppose that $\Sigma$ is a finite subset of $\Sigma_{0}$. We say that $M$ is of type $\Sigma$ if the following hold:
@@
   • the $AG_{\mathbb{Q}}$-module $\wedge^{2}_{A}M$ is isomorphic to $A\otimes_{{\mathcal{O}}}{\mathcal{O}}_{\psi}(1-k)$.
-Consider the functor on $\mathcal{C}$} which associates to AA the set of AA-deformations of $\rho_{0}$ of type $\Sigma$.
+Consider the functor on $\mathcal{C}$} which associates to $A$ the set of $A$-deformations of $\rho_{0}$ of type $\Sigma$.
 By the results of Mazur and Ramakrishna, this functor is representable by an object of $\mathcal{C}$}.
@@
 Thus ${\mathcal{M}}_{!}^{\Sigma}$ can be canonically identified with a quotient (also a direct summand) of ${\mathcal{M}}(\sigma^{\Sigma})_{!,\lambda_{0}}$.
-Lemma $7.$
+Lemma 7.7.
 The composites of the natural maps
@@
 Let ${\mathcal{V}}_{0}={\mathcal{V}}_{0}^{\Sigma}={\mathcal{V}}_{0}^{\Sigma^{\prime}}$, $V^{\prime\prime}=\alpha_{p}^{-1}V^{\prime}\alpha_{p}$ and $A={\mathcal{V}}_{0}\otimes_{{\mathcal{O}}_{0}}{\mathcal{M}}_{!}(V^{\prime\prime})$.
-Then $V^{\prime\prime}$ is normal in both $U^{\Sigma}$ and $\alpha_{p}^{-1}U^{\Sigma}\alpha_{p}$, and both groups act on AA, where the action of $\alpha_{p}^{-1}U^{\Sigma}\alpha_{p}$ on ${\mathcal{V}}_{0}$ is defined by first conjugating by $\alpha_{p}$.
+Then $V^{\prime\prime}$ is normal in both $U^{\Sigma}$ and $\alpha_{p}^{-1}U^{\Sigma}\alpha_{p}$, and both groups act on $A$, where the action of $\alpha_{p}^{-1}U^{\Sigma}\alpha_{p}$ on ${\mathcal{V}}_{0}$ is defined by first conjugating by $\alpha_{p}$.
 Note that the actions agree on
@@
 is injective.
-To prove this, suppose there were a nonzero vector in AA fixed by both $\alpha_{p}^{-1}U^{\Sigma}\alpha_{p}$ and $A^{U^{\Sigma}}$.
+To prove this, suppose there were a nonzero vector in $A$ fixed by both $\alpha_{p}^{-1}U^{\Sigma}\alpha_{p}$ and $A^{U^{\Sigma}}$.
 Then there would be a nonzero vector in ${\mathcal{M}}_{!}(V^{\prime\prime})$ on which $\left(\begin{array}[]{cc}{{a}}&{{b}}\\ {{c}}&{{d}}\end{array}\right)$ acts via $\psi_{p}^{-1}(d)$ for all matrices in $U_{p}^{\mathop{\rm min}\nolimits}$ or $\alpha_{p}^{-1}U_{p}^{\mathop{\rm min}\nolimits}\alpha_{p}$.
@@
 (48) $0\to C^{U^{\Sigma}}\times C^{U^{\Sigma}}\to D^{\tilde{U}}$.
-Now let $A_{c}$, $B_{c}$, $C_{c}$ and $D_{c}$ denote the $\tilde{{\mathbb{T}}}-$modules defined in the same way as AA, $B$, $C$ and $D$, but using ${\mathcal{M}}_{c}$ instead of $\mathcal{M}$$_{!}$.
+Now let $A_{c}$, $B_{c}$, $C_{c}$ and $D_{c}$ denote the $\tilde{{\mathbb{T}}}-$modules defined in the same way as $A$, $B$, $C$ and $D$, but using ${\mathcal{M}}_{c}$ instead of $\mathcal{M}$$_{!}$.
 Then the natural maps $C_{c}\to A_{c}^{V}$, $D_{c}\to B_{c}^{\tilde{V}}$ and $D_{c}\to A_{c}^{\tilde{V}}$ are isomorphisms.
@@
 Writing $\tilde{\ }$ for Teichmuller liftings, we conclude that $\rho\otimes_{K}\tilde{\psi}^{\prime}\psi$ is modular, where $\psi$ is a character of $\ell$-power order such that $\chi_{\ell}^{1-k}\psi^{2}{\mathop{\rm det}\nolimits}\rho$ has order not divisible by $\ell$.
-Suppose that $K_{0}$ is a number field with ring of integers ${\mathcal{O}}_{0}$, and $f=\sum a_{n}e^{2\pi inz}$ is a newform of weight $k\geq 2$ and level NN with coefficients $a_{n}$ in $K_{0}$.
+Suppose that $K_{0}$ is a number field with ring of integers ${\mathcal{O}}_{0}$, and $f=\sum a_{n}e^{2\pi inz}$ is a newform of weight $k\geq 2$ and level $N$ with coefficients $a_{n}$ in $K_{0}$.
 For primes $\lambda$ of $K_{0}$, we write $\bar{\rho}_{f,\lambda}$ for the representation of $G_{\mathbb{Q}}$ on
@@
 Theorem 7.15.
-Let $f$ be a newform of weight $k\geq 2$ and level NN with coefficients in $K_{0}$.
+Let $f$ be a newform of weight $k\geq 2$ and level $N$ with coefficients in $K_{0}$.
 Suppose that $\lambda$ is a prime in $S_{f}$ (as defined above), and let ${\mathcal{O}}={\mathcal{O}}_{0,\lambda}$ and $K=K_{0,\lambda}$.
@@
 Recall that $\Sigma_{e}(f)$ is the set of primes $p$ such that $L_{p}^{\mathop{\rm nv}\nolimits}(A_{f},s)=1$ but $L_{p}(A_{f},s)=(1+p^{-s})^{-1}$.
-We write $b_{{\mathop{\rm dR}\nolimits}}$ for the basis of $M_{\psi,{\mathop{\rm dR}\nolimits}}$ defined in §1.3, and set $\eta=0$ or 11 so $\eta\equiv k\bmod 2$.
+We write $b_{{\mathop{\rm dR}\nolimits}}$ for the basis of $M_{\psi,{\mathop{\rm dR}\nolimits}}$ defined in §1.3, and set $\eta=0$ or $1$ so $\eta\equiv k\bmod 2$.
 Note that by Proposition 5.5 of [De3], we have $\epsilon(M_{f}\otimes M_{\psi^{-1}})/\epsilon(M_{\psi^{-1}})\in K^{\times}$.
@@
 The identity $L(B_{f},0)(1\otimes b(B_{f}))=c^{+}(B_{f})$ then follows by applying $({\mathbb{R}}\otimes{\rm tw})$ to the identity $L(A_{f},0)(1\otimes b(A_{f}))=c^{+}(A_{f})$ and using (49) and Lemma 8.4.
-As in §4.5 put $U=U_{0}(N)$, let $\sigma:U\rightarrow K^{\times}$ be the representation $\left(\begin{smallmatrix}a&b\\ c&d\end{smallmatrix}\right)\mapsto\psi^{-1}(a_{N})$, and choose a multiple $N^{\prime}$ of NN so that $U_{N^{\prime}}\subseteq U$ is sufficiently small.
+As in §4.5 put $U=U_{0}(N)$, let $\sigma:U\rightarrow K^{\times}$ be the representation $\left(\begin{smallmatrix}a&b\\ c&d\end{smallmatrix}\right)\mapsto\psi^{-1}(a_{N})$, and choose a multiple $N^{\prime}$ of $N$ so that $U_{N^{\prime}}\subseteq U$ is sufficiently small.
 Put $w=\left(\begin{smallmatrix}0&-1\\ N&0\end{smallmatrix}\right)\in GL_{2}({\mathbb{A}}_{\bf f})$ and denote by $W=[UwU]_{\omega}:M(\sigma)_{N^{\prime}}\rightarrow M(\hat{\sigma}\otimes(\psi^{-1}\circ{\mathop{\rm det}\nolimits}))_{N^{\prime}}$ the isomorphism in Lemma 5.2.
@@
 Note that the field $K_{f}$ generated by the Fourier coefficients of the newform $f$ is either totally real or a CM field and hence has a well defined automorphism $\rho$ induced by complex conjugation.
-It is known that the Fourier expansion $f^{\rho}(z)=\sum_{n=1}^{\infty}a_{n}^{\rho}e^{2\pi izn}$ is a newform of conductor NN and character $\psi^{-1}$ [Miy, 4.6.15(2)], hence represents an element of $\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}M_{!}(\hat{\sigma})_{N^{\prime},{\mathop{\rm dR}\nolimits}}$.
+It is known that the Fourier expansion $f^{\rho}(z)=\sum_{n=1}^{\infty}a_{n}^{\rho}e^{2\pi izn}$ is a newform of conductor $N$ and character $\psi^{-1}$ [Miy, 4.6.15(2)], hence represents an element of $\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}M_{!}(\hat{\sigma})_{N^{\prime},{\mathop{\rm dR}\nolimits}}$.
 Recall that $b_{\mathop{\rm dR}\nolimits}$ is the canonical basis of $M_{\psi,{\mathop{\rm dR}\nolimits}}\cong(H^{0}(X_{N^{\prime}})\otimes K_{\psi^{-1}})^{U}_{\mathop{\rm dR}\nolimits}$ (Lemma 4.6) and that $M_{\psi}\otimes_{K}M_{f^{\rho}}$ has a natural map into $M_{!}(\hat{\sigma}\otimes(\psi^{-1}\circ{\mathop{\rm det}\nolimits}))_{N^{\prime}}$ via the cup product on $X_{N^{\prime}}$ (see (26)).
@@
 $=$ $[U:U_{N^{\prime}}]^{-1}(k-2)!(4\pi)^{k-1}\phi(N^{\prime})(\tau(f),\tau(f))_{\Gamma(N^{\prime})}\,\tau(\alpha_{\mathop{\rm dR}\nolimits})\otimes\iota^{k-1}$
-where $\phi$ is Euler’s function and we have used Lemma $2.$ Therefore (58) equals
+where $\phi$ is Euler’s function and we have used Lemma 2.2. Therefore (58) equals
 (59) $\begin{array}[]{l}[U:U_{N^{\prime}}]^{-1}(k-2)!(4\pi)^{k-1}\phi(N^{\prime})(f,f)_{\Gamma(N^{\prime})}\cdot\alpha_{\mathop{\rm dR}\nolimits}\otimes\iota^{k-1}\\ =\frac{[\bar{\Gamma}_{1}(N):\bar{\Gamma}(N^{\prime})]}{[U:U_{N^{\prime}}]}\phi(N^{\prime})(k-2)!(4\pi)^{k-1}(f,f)_{\Gamma_{1}(N)}\cdot\alpha_{\mathop{\rm dR}\nolimits}\otimes\iota^{k-1}\end{array}$
@@
 Lemma 8.7.
-If $f$ is a newform of conductor NN, weight $k$ and with coefficients in the number field $K$ we have
+If $f$ is a newform of conductor $N$, weight $k$ and with coefficients in the number field $K$ we have
 $(\tau(f),\tau(f))_{\Gamma_{1}(N)}=\frac{(k-1)!\delta(N)N\phi(N)L^{\mathop{\rm nv}\nolimits}(A_{f},\tau,1)}{4^{k}\pi^{k+1}}$
@@
 (61) $L_{p}(k,f,\bar{\psi})\frac{\phi(N_{p}/N_{\psi,p})}{N_{p}/N_{\psi,p}}=L_{p}^{\mathop{\rm nv}\nolimits}(A_{f},\tau,1)\frac{\phi(N_{p})}{N_{p}}$
-for all primes pp.
-If $p\nmid N$ this is immediate from §$5.$ If $N_{p}=p$ and $N_{\psi,p}=1$ we have $a_{p}^{2}=\psi(p)p^{k-2}$ by [Miy, Th.
+for all primes $p$.
+If $p\nmid N$ this is immediate from §5.5.
+If $N_{p}=p$ and $N_{\psi,p}=1$ we have $a_{p}^{2}=\psi(p)p^{k-2}$ by [Miy, Th.
 4.6.17(2)] and $\pi_{p}(f)$ is special so that (61) holds true by (38).
@@
 This in turn is a consequence of Lemma 8.8 below.
-Lemma $8.$
+Lemma 8.8.
 Let $M$ be an object of $\operatorname{{\bf PM}}_{K}$ which is $L$-admissible everywhere and let $\tau:K\rightarrow{\mathbb{C}}$ be an embedding.
@@
 Noting that with our choice of $\psi_{p},dx_{p}$ the epsilon factor equals 1 (resp.
-a power of ii) for $p\nmid c(M)$ (resp.
+a power of $i$) for $p\nmid c(M)$ (resp.
 $p=\infty$) the Lemma follows.
@@
 additive category MF) and by $\text{${\mathcal{O}}_{\lambda}$-${{\mathcal{M}\mathcal{F}}}$}^{a}$ (resp.
-$\text{$K_{\lambda}$-MF}^{a}$) the full subcategories with filtration restrictions as in §$1.$ Scalar extension -$\otimes_{{\mathbb{Z}}_{\ell}}{\mathbb{Q}}_{\ell}$ induces an exact functor $\text{${\mathcal{O}}_{\lambda}$-${{\mathcal{M}\mathcal{F}}}$}\rightarrow\text{$K_{\lambda}$-MF}$ where the notion of exactness in MF is defined in [Fo1, 1.2.3].
+$\text{$K_{\lambda}$-MF}^{a}$) the full subcategories with filtration restrictions as in §1.1.
+Scalar extension -$\otimes_{{\mathbb{Z}}_{\ell}}{\mathbb{Q}}_{\ell}$ induces an exact functor $\text{${\mathcal{O}}_{\lambda}$-${{\mathcal{M}\mathcal{F}}}$}\rightarrow\text{$K_{\lambda}$-MF}$ where the notion of exactness in MF is defined in [Fo1, 1.2.3].
 Now assume that ${\mathcal{D}}_{1},{\mathcal{D}}_{2}$ are torsion free objects of $\text{${\mathcal{O}}_{\lambda}$-${{\mathcal{M}\mathcal{F}}}$}^{a}$ for some $a$ and put $D_{i}={\mathcal{D}}_{i}\otimes_{{\mathbb{Z}}_{\ell}}{\mathbb{Q}}_{\ell}$.
@@
 For $B_{f}=A_{f}(1)$ we can use the same argument as long as both ${\mathcal{D}}_{1}={\mathcal{M}}_{f,{\lambda\mbox{\scriptsize-crys}}}$ and ${\mathcal{D}}_{2}={\mathcal{M}}_{f,{\lambda\mbox{\scriptsize-crys}}}[1]$ are objects of $\text{${\mathcal{O}}_{\lambda}$-${{\mathcal{M}\mathcal{F}}}$}^{1}$.
-This is the case if $\ell>k+1$ or if $l=k+1$ and ${\mathcal{M}}_{f,{\lambda\mbox{\scriptsize-crys}}}$ has no nonzero quotient AA in ${\mathcal{O}}_{\lambda}$-$\mathcal{M}\mathcal{F}$}} with $\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}A=A$.
+This is the case if $\ell>k+1$ or if $l=k+1$ and ${\mathcal{M}}_{f,{\lambda\mbox{\scriptsize-crys}}}$ has no nonzero quotient $A$ in ${\mathcal{O}}_{\lambda}$-$\mathcal{M}\mathcal{F}$}} with $\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}A=A$.
 Lemma 8.13.
-If $a_{\ell}\equiv 0\mod\lambda$ then ${\mathcal{M}}_{f,{\lambda\mbox{\scriptsize-crys}}}$ has no nonzero quotient AA in ${\mathcal{O}}_{\lambda}$-$\mathcal{M}\mathcal{F}$}} with $\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}A=A$.
+If $a_{\ell}\equiv 0\mod\lambda$ then ${\mathcal{M}}_{f,{\lambda\mbox{\scriptsize-crys}}}$ has no nonzero quotient $A$ in ${\mathcal{O}}_{\lambda}$-$\mathcal{M}\mathcal{F}$}} with $\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}A=A$.
 Proof.
@@
 But again by (76) we find $\bar{\phi}^{k-1}(\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}\bar{{\mathcal{M}}})\not\subseteq\text{ker}(\bar{\phi})=\bar{\phi}(\bar{{\mathcal{M}}})$ so that $N=\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}N=0$.
-If AA is a nonzero quotient of $\mathcal{M}$} then $\bar{A}$ is a nonzero quotient of $\bar{{\mathcal{M}}}$ hence equal to $\bar{{\mathcal{M}}}$ and we find $\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}\bar{A}\neq\bar{A}$ and $\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}A\neq A$.
+If $A$ is a nonzero quotient of $\mathcal{M}$} then $\bar{A}$ is a nonzero quotient of $\bar{{\mathcal{M}}}$ hence equal to $\bar{{\mathcal{M}}}$ and we find $\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}\bar{A}\neq\bar{A}$ and $\operatorname{{\mathop{\rm Fil}\nolimits}}^{k-1}A\neq A$.
 ∎
@@
 We shall first prove Theorem 8.10 for $M=B_{f}=A_{f}(1)$ in which case Lemma 8.11 applies.
-Fix a prime $\lambda\in S_{f}$ and let $\Sigma$ be the set of primes dividing NN if $N>1$ or put $\Sigma=\{p\}$ for some prime $\lambda\nmid p$ if $N=1$.
+Fix a prime $\lambda\in S_{f}$ and let $\Sigma$ be the set of primes dividing $N$ if $N>1$ or put $\Sigma=\{p\}$ for some prime $\lambda\nmid p$ if $N=1$.
 The isomorphism $\gamma:M_{f}\rightarrow M_{f}^{\Sigma}$ of Proposition 6.3 satisfies
@@
 ${\mathcal{O}}_{\lambda}\cdot\langle f^{\Sigma},b^{\prime}(f^{\Sigma}\otimes(2\pi i)^{2})\otimes\iota^{2}\rangle^{\Sigma}={\mathcal{O}}_{\lambda}\eta^{\Sigma}_{f}{\mathcal{M}}_{\psi}(1-k)_{{\mathop{\rm dR}\nolimits}}$
-where $\eta^{\Sigma}_{f}$ was defined before Proposition $6.$ On the other hand by (78), Theorem 8.5 and the remark after the proof of Theorem 8.5 we have for $\lambda\in S_{f}$
+where $\eta^{\Sigma}_{f}$ was defined before Proposition 6.6. On the other hand by (78), Theorem 8.5 and the remark after the proof of Theorem 8.5 we have for $\lambda\in S_{f}$
 ${\mathcal{O}}_{\lambda}\cdot\langle f^{\Sigma},\gamma b^{\Sigma}(B_{f})(f^{\Sigma}\otimes(2\pi i)^{2})\otimes\iota^{2}\rangle^{\Sigma}$